A Complete Bibliography of Publications in the Journal of Systems and Software (2000–2009)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

25 May 2021 Version 3.00

Title word cross-reference

2 [Aba08]. 3 [BMOKAM09, BGG09]. k [BRTT08, SSCL08]. n [LKJL01]. R [SC00]. t [LYX09].

-fault [LYX09]. -medoid [BRTT08]. -NN [SSCL08]. -tree [SC00]. -way [LKJL01].

.NET [BS03].

10th [DGV08]. 1471 [KvV06]. 15504 [EG00, EB00, EJ01, JH01]. 1st [CBVD07].

2 [AACL02, CT00]. 2004 [LC06b]. 2007 [GH08, HLM^{+}09]. 2008 [Sai09]. 24-h [JJ06].

3-Disjoint [CLC03]. 3/layer [DGV^{+}07].

60 [Ano02e]. 64 [LKH^{+}08].

7 [DK08].

9001 [JH01]. 95 [RW00]. 9D [LLT^{+}09]. 9D-SPA [LLT^{+}09].

abbreviated [ONR02]. abstraction [SD02]. Academia [Wey01]. academic [Fra07]. Academics [Gla00b]. Accelerating [BRTT08]. acceleration [EHKH04]. acceptance [UN09]. access [CET^{+}08, Cho04a, Cho04b, CHL05, CC05, CHY^{+}05, HH05, HY03, HCC05, JW06, LNC01, LCCH02, LY01, Oi08, SC07, WL05, KFS^{+}02]. Accessing [LNY06, LHLY05, LO04]. accounting [TDL^{+}02]. accuracy [KPME02, KPME05, LMYMGT08]. accurate [ED04]. Achieving [FMP09]. ACODF [TTWY04]. acquisition [CS01, RR09]. across [IBP03, LT09, MGEB03]. action [CC07, Moy00]. action-based [CC07]. Actions [CGP^{+}09]. Active [KPG^{+}07, WOH08, AJCM08, DPSU06, KRC00, LLL00, WZJ01, PK02a]. Active/Standby [PK02a]. activities [MG04, SSA08]. activity [CCC05]. actually [SLS08]. ad [BMES04, WF07, YZ05, YSK09, ZMN05]. Ada [RW00]. Adaptable [CS04, HK09, RS06, SK04]. adaptation [BBC05, DPSU06, FdSdP08, HKY01, INS00, OC04, Rad04]. adapting [CELS07]. Adaptive [CGHL07, CL08, CKL09, DGV^{+}07, GL05, HYC04, LU06, LWHS05, LG05a, PSK05, PSH06, TSSD09, BLM^{+}08, BFV04, CJHB08, CSGL05, CKL08, HC04a, KSAOK04, KCB05, KD05, LT07, LT09, LCF^{+}06, LGL08, MPST06, MV05, MK06, MHC00, Rav03, RH06, SMG08, SG06, TC06, VA08, YHZ^{+}09, ZC06, ZL07, CH05]. adaptors [AMNT08]. adapts [EK00]. adding [CLL05]. adjoint [ULN06]. adjusted [CH07b]. adopt [SNJ^{+}07]. adopting [MKK09]. adoption [AW07, QHS08, Tan00, TW08a, UN09]. adult [CHZY03]. Advanced [CY04, ZS05a]. advantages [CDS07, PW09]. affect [Kel09]. affecting [RH03, ZP00]. affinity [Kar00]. after [JBA08, PTRW04]. against [SCH05, SC09]. agency [ML03]. agenda [WD07]. Agent [AM04, CL04b, BHAM09, BWM06, CPT05, CC08b, CET^{+}08, CLC08a, Cho05, CMNA^{+}09, GTA09, JZL07, LH04, LT09, LSH09, MV05, MV06, OKS08, PLCC09, RMC05, SCdS^{+}06, TKA^{+}02, WHN^{+}01, YGH^{+}08]. agent-aided [CPT05]. Agent-based [AM04, CC08b, Cho05, LH04, SCdS^{+}06]. agent-oriented [OKS08]. agents [CFN07, GMB^{+}09, GHK05, GDH05, HWH^{+}03, MHW01, WGC02, WBW^{+}06, ZK04a]. aggregation [BLM^{+}08]. Agile [NRG08, AL05, CC08c, GR05, HF08, HDGZ06, MKK09, PW09, QHS08]. agilization [TBD^{+}08]. aging [JX07]. agreement [LLY07, LKH09, NLKW05, SCH05, Tse07, WY04]. aided [CPT05]. Ajanta [TKA^{+}02]. AJAX [MvD08]. al [SCL07, WYL06]. algorithm [CDC09, CHC01, CCSC07, CLH07, CTL08, DXPY03, JE02b, JFC08, KKK08, KK07b, LC07, LLLZ06a, LLLZ06b, LCCH02, LK04, Loo05, OW04, OOD09, Özm09, PK01a, TTC04, WGW^{+}09, WHYT06, YSK06, Yoo09]. algorithms [AN01, BRMA^{+}09, CBZ00, CWC04, CD07, GBL08, GLJ00, Gho01, HSR01, IWF07, JJP02, MMM00, MBB01, PH06, Rav03, SMDM05, SM06b]. alignment [BCV06, CBVD07]. allocating [IJC03]. allocation [Aba06, Aba08, BMOKAM09, BHAM09, CLSa01, DXPY03, GP05, HCDJ08, HC01b, HL06a, KMOS09, LLL06, LCLL07, MC01, MV05, MK06, SK03, SOC^{+}03, YYWW07, vdSJK^{+}07]. allocator [HC06]. ALMA [BLBvV04, LBvVB02]. alone [DF00]. alternatives [DC09]. among [CLW05, HGK^{+}06, WSJK08]. Amorphous [HBD03]. AN/BSY [AACL02]. AN/BSY-2 [AACL02]. analogy [CH07b, JIS03, LXG09]. analogy-based [CH07b]. Analysing [GW01, JMSS07, dL04, LTK^{+}06]. Analysis [BNW^{+}08, CUY09, DHKV06, HH08b, KSAOK04, LG03, WYL06, Aml00, BH02, BH03, BRC09, BDMK03, BBS00, BLBvV04, BHL00, BLOS06, CC08a, CCC05, CCW02a, CH09, CPL^{+}04, CGW08, CZH^{+}08, DCAC09, DH09, DDGR09, FAB^{+}07, FL05, FRR09, Gok09, GPML06, GAWW07, HPT07, HH07, jHjW08, HY00, Hua05b, HZCD05, IBP03, JMP07, JX07, JCYC04, Kar04b, KBK06, KG09, KKP06, KPS08, KSS03, KDEK04, KSH09, LKH^{+}08, LYC04, LGW09, LKJL01, LDL07, LSaC04, LHC^{+}05, LH06, Lop03, LHSK06, DPS03, MYZC06, MT07, MLB09, Mil00b, MR00a, ML08, Mur08, PG05, Par00, PK02a, PC04, PS00, RK00, RH02, RH03, SKZ^{+}04, SNBH08, SW09, SS04, SM08, Sut00, SPZ06, TNJH07, TSA08, TBD^{+}08, UN09, WG05, WPP^{+}09, XNP07, YS02, ZP00, ZSP01, ZJDB02, dBvV08, LBvVB02, ADZ^{+}09]. analyst [MG04]. analytical [FCSM09, MA09, Mil00a, ZM06]. analyzable [DGL^{+}08]. analyze [MR00b, PSMB01]. Analyzing [CC02a, CWC04, CBKK08, HYS^{+}04, Lop03]. angle [HDLK00]. animation [LW07, PH07]. annealing [PH06, TVA04]. anomalies [LM03, WW09]. Anomalous [HWM01]. Anomaly [Rad04, CCKM09, DLW08]. Anomaly-free [Rad04]. anonymous [CHL^{+}08, FHHL09]. anti [CHY^{+}05]. anti-spyware [CHY^{+}05]. anticipation [TSSD09]. Antoniou [LZ07]. API [CPiLH09, HS03, SL07]. APIs [TDK^{+}07]. APIs-based [TDK^{+}07]. applets [HWM01]. Application [Alz08, Lop03, MR01, SK07, TQ05, ASS07, Aml00, CPT05, CH07a, Cho05, hCSW^{+}04, DPSU06, ELH00, FAB^{+}07, FIGCLN^{+}02, Hus01, JE02a, Kel09, LORB03, LS04, LGW09, LP05, MMTL06, MR00b, OC04, PTRW04, Pon03, RDD02, SCdS^{+}06, SRDLCP09, Tan04, WK00, YWT07]. Application-specific [SK07]. Applications [EC04, IT03, LZN04, AP09, ALT^{+}09, AMHJ09, BBG^{+}04, Boz00, CELS07, CCCT06, CJZ04, CLGL05, CZL07, CJ09, CC05, CBKK08, DGV^{+}07, DBO05, DCH02, DK01, FL09, GD04, HL01, HH08b, HKW00, KDS^{+}08, KCS01, LLY07, LG05b, LGL08, LT08, MV05, MV06, SUSO04, SHS^{+}07, SFSE05, TL09a, VA08, WDCL08, YS04, ZCT^{+}09, Zhu03, Zhu04c]. applied [LNPAGD^{+}06, PB00]. Applying [CDS02, Gon08, Mil00a, SLC00, TPRW04, BKB^{+}07]. Approach [CW09, ÅCF^{+}07, AM04, BM00a, BDGR01, BHN02, BBC05, CCW^{+}01, CPT05, CFFT08, CELS07, CCHT09, CCW02b, CC03, CC07, CCKM09, CC09b, CJT01, ESW06, EBB09, FdOdL04, GM02, GPHS08, GMS07, GSB^{+}07, HTK00, HK09, HCC08, HZCD05, HWML04, KY08, KGT02, KMOS09, LMvV09, LNC01, Lee07, LNY06, LW07, Lut00, MLB09, MM00a, MDMC06, MR00b, Mur08, Mus03, Nae01, NRG08, ÖKT09, PCC02, PTBP08, PSG^{+}09, RT07, RW00, SM09, SL03, ST07, TVA04, TK00, TTWY04, TL07, WDN05, YR09, ZERO00, ZSM04, Zhu03, KLRW01]. Approaches [LCY00, Bat08, DA07, HKN^{+}07, JZ05, LS05b, MH04, Rey07, SH07]. Approximating [BMES04]. APRIL [PH07]. Arabic [Mus03]. arbitrary [CCW02b, NXS00]. arbitrary-rate [NXS00]. Arches [DSSL09]. Archetypal [RRC07]. Architecting [FB04, dLGR06, FM08]. architects [Kru08]. Architectural [YWLG02, dBvV08, AAAC07, BGG^{+}06, GPML06, HYS^{+}04, JBA08, KKLP09, MvD08, PTBP08, RLvV06, SMR09, WDS09, Wil03, WSQM05, XZAR06, ZKL^{+}09]. Architecture [BLBvV04, LH04, WPC06, ANH07, AG08, BKZ^{+}06, BL09, BJ03, BNW^{+}08, BL03, CDS02, CLL05, CJZ04, CS04, CMS04, CBS00, DHL06, DK01, EK00, HKN^{+}07, IWF07, JAvdV09, JHSB09, KDS^{+}08, KBK06, KPS^{+}04, KLY03, KPT09, KKLP09, KKK08, LRvV03, LC07, LGL08, Lop03, LICA09, LG03, MK08, MKNS06, MDR06, PWCC01, PNL07, RRD06, RS06, SNBH08, SMHMA08, ST07, SSM^{+}09, SC09, TBGH06, TJH07, TNJH07, TSA08, WT01, ZMAV08, dBvV09, dBvV03, AJCM08, LBvVB02]. Architecture-based [WPC06, MDR06, ST07]. architecture-centric [SNBH08].

Architecture-level [BLBvV04, LBvVB02]. architectures [BGH03, CCG01, CS01, DGP02, Del08, ELK06, FdSBR06, GWvD08, IT03, JE02b, KPS08, LCM^{+}04, PNM04, RSP03, SO03, SG06, SM07, TDL^{+}02, UZ09, YHZ^{+}09, CFFT08]. ARDIN [CG03]. area [HYC04, LY09]. array [HY00, HY01]. arrays [Çam00a, CCW02a]. art [PW09, CWP09]. article [ML08]. artifact [WW09]. artifacts [RGBM06]. artificial [dBTdSS08]. ASM [ZM06]. ASM-based [ZM06]. aspect [ADZ^{+}09, LVMM07, VP07, KCS08]. aspect-oriented [LVMM07, VP07, KCS08]. aspects [OC04, Wij03, WPP^{+}09]. assembling [AMNT08]. assembly [HMSW03, PTBP08]. assertions [SM00]. Assessing [JZ07, DPS03, NR04, FN00, KPS^{+}04]. Assessment [KB07, SZZ06, SP08, AS00, BW01, Bud00, CJHB08, EJ01, Gla99, Gla00c, Gla00d, GC01, GC02, GC03, GC05, HCN00, KPS^{+}04, LSV^{+}06, LHC^{+}05, MSA08, ONZ09, PIGÖ08, REF^{+}07, SKW06, SPSM03, TCG06, WTG^{+}08, WTG^{+}09, ZSP01]. assigned [WWB09]. Assigning [JJ06]. assignment [AS01, CY00, LLL00]. assisting [NWZ05a]. association [LcLsW06, YHHR03]. associative [Sta03]. assumptions [DI01a, RLvV06, dlRT06]. assurance [SM00, ZE03]. ASWEC [GH08]. asynchronous [GLJ00, Gho01, LR04, Rav03]. ATF [CH05]. ATLAS [CL04a]. Atomic [CGP^{+}09, MK00]. attack [GJ08, WYL06]. attacks [GMB^{+}09, SKZ^{+}04, SCH05]. attractiveness [AADAD02]. attribute [PK01b, WZG09, YHZ^{+}09]. attributes [BL03, CGSGR06, Wij03]. auction [CHL^{+}08, LLL06]. audio [HHL06]. audit [WZG09]. audits [McD02, dBvV08]. augment [SW09]. augmented [GHK05]. authenticated [CLC08b, WH02, YC09]. authenticating [Lin01]. authentication [CJT01, CJ03, LT04, Lin07, TM06, WHHT08, YCYW07, YS04]. author [FMSG08, Gla00a]. authoring [BBG^{+}04]. authorization [Lin07]. authorizations [LWL04]. authors [SM06b]. authorship [DS04]. automata [KH06, WKH09, CR06]. Automated [BLOS06, GML05, HWHM02, MB06, TBD^{+}08, Yeu00, ZAO08, HCWN05, LW07, LT08, PG05, CSSW05]. Automatic [CCCT06, DBO05, Phi06, SA08, YLC06, ZL06, dlRT06, CCHT09, KGMI06, LHP^{+}09, LHP^{+}10, PTBP08, RjHHK08, TH02, VA08, WBW^{+}06, ZC08, ZS05b]. automation [ACDG02, DL06]. automotive [GD04, SP08]. Autonomic [NKJT09, BDK08, WDCL08]. Autonomous [BHAM09, MHW01, GHK05]. Availability [Aba08, PK02a]. Availability-based [Aba08]. AVDL [RS06]. average [MM01b]. Avoiding [O’B08, Shu03]. aware [AKP04, CDEV08, DPMD07, GBL08, HLYL06, HZ07, KPTV09, KK07b, MA09, PSH06, PS09, RT07, YGH^{+}08]. Axiomatic [TDT08]. axiomatization [LORB03].

back [CE08, PJ09]. back-off [PJ09]. back-propagation [CE08]. background [KM04]. backtracking [CC01, YZ08]. backup [MK08]. backward [CPL^{+}04]. bad [LS07]. balance [CHL04]. balancing [Boz00, CBZ00, MCCC03, RwJK01, WGW^{+}09, ZK09]. bandwidth [MV05, MK06]. Barry [Fra07, Vau07]. Based [CSSW05, ÅCF^{+}07, Aba08, AZGvG09, AZW07, AA07, AKP04, AM04, Aml00, ACGS^{+}08, AMNT08, BM05, Bai05, BHXN05, BRMA^{+}09, BW01, CC09a, CDI07, CGL^{+}04, CC04, CCCT06, CCY^{+}09, CL06b, CC07, CWP09, CBG09, CK00a, CLGL05, CZL07, CC08b, CLC08a, CKL09, CHZY03, CCC06, CLG08, CJT01, CH07b, CPL^{+}04, Cho04a, Cho05, CC05, CE08, CNLV07, CL02, DCAC09, DHL06, DPSU06, DACY07, DK01, DZW^{+}09, EBGR01, EHKH04, FY04, FHHL09, FdOdL04, FCL^{+}00, FL09, GMR08, GML05, GBL08, GM02, GR05, Gok09, GMS07, GJ07, GHKR04, Gru07, GJ08, GAWW07, HRL09, HSM^{+}07, HYC02, HH06, HH08b, HDLK00, HC01b, HH05, HZCD05, HTH09, HZ07, HWML04, JR09, JW06, JHSB09, JTM04, Jun00, KBM05, KBD09, KK07a, KDS^{+}08, KSAOK04, KKP06, KYPW06, KY08, KJ04, KTK01, KSS03, KEK04, KLGH07, LWS^{+}03, LHLY05, LJB05, LK01, LNC01, LH04]. based [LS04, LPR04, LKW^{+}09, LS05a, LZL^{+}06, LXG09, LCCH02, LP05, LCLL08, LWL04, LS05b, LZ06, LDL07, LGL08, LC08, LLV^{+}09, LHC^{+}05, Lok06, LWL09, MYZC06, MKL^{+}00, MV05, MV06, MLB09, MvD08, MGI07, MSB^{+}02, MDR06, NLKW05, NG08, NGM08, OÓ08, OW04, Oi08, ONR02, Özm09, ONZ09, PG05, PKL03, PSK05, PDC01, PTBP08, PILO06, Phi04, PW03, PLP04, RH02, RRW00, RSP03, Rey07, RDD02, SM09, SZ06, SRGL08, SCdS^{+}06, Sha09, ST07, mSgFtL05, SL02, SM06b, SSM^{+}09, SS04, SM08, SZPMK04, SM03, SH07, TJH07, dBTdSS08, TL09a, TDK^{+}07, VMJS06, Wal05, WCH03, WPC06, WC07, WGY^{+}08, WDCL08, WKH09, WG05, WQ06, WLC08, XYS07, YY04, YGH^{+}08, YS04, YLC08, YL09, YSK06, YC08b, ZEY04, ZC08, ZM06, Zhu00, ZS05b]. batch [dSSJV08]. Bayesian [Bai05, BHXN05, TNJH07]. BBN [FY04]. BBN-based [FY04]. be [ED04]. BEACH [Tan04]. beat [Gla00k]. before [ZP06]. behavior [CLSa01, CGW08, Oi08, RRC07]. Behavioral [HFC^{+}01, CMT02, Gla00k, HZCD05, KZDX09]. behaviors [CCCT06, MM00b]. behaviour [OFWP07, Phi04]. behavioural [HCWN05]. behaviours [HCWN05, dL04]. behind [Gla00n]. belief [BG09, TNJH07]. belief-theoretic [BG09]. believing [Gla00m]. benchmark [ZBLG07].

Benchmarking [NG08]. benefit [NGC02]. benefit-oriented [NGC02]. Best [GH08, Sai02, VE03, KK07b]. Better [JTM04]. between [BMES04, BBS00, BGH^{+}08, BFPAGS^{+}08, BWDP00, CGMPAP08, GMS07, JH01, MR00a, PW09, WGH00, dBvV09]. Beyond [RGBM06]. Bi [FL05]. Bi-directional [FL05]. Bidder [CHL^{+}08]. Bidder-anonymous [CHL^{+}08]. binary [CY00, CPiLH09]. binding [CDEV08]. biometric [UN09]. birthmark [CPiLH09]. blind [CZL07, HH08b, HC04b, JL04, SHT05, WYL06, ZC05]. block [Gok09, HOR01, LKH^{+}08, WQ06, WLC08]. blocking [KW00, Shu03]. blue [Gla00n]. board [Ano02h, Ano02i, Ano02j, Ano02k, Ano02l, Ano02m, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03l, Ano03m, Ano03n, Ano03o, Ano03p, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano04m, Ano04n, Ano04o, Ano04p, Ano04q, Ano05h, Ano05i, Ano05j, Ano05k, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p, Ano05q, Ano04f, Ano04g]. Boehm [Fra07, Vau07]. Book [LC06b]. Boolean [YLC06]. both [HZCD05, YYL^{+}06]. Botswana [UN09]. bottlenecks [HRN^{+}01]. bound [MC01, SSO05]. boundaries [KRHŽ05]. bounding [HDLK00]. Branch [SC00, MC01, PG04, SLC00]. branch-and-bound [MC01]. Brazilian [LdSBA^{+}08]. Breadth [LC00]. Breadth-first [LC00]. bribery [CW09]. bribery-free [CW09]. Bridging [MGEB03]. broadcast [CBK02, DY03, LK04, MK00, PJ09, PLF05, SC07, SC08, WHYT06]. broadcast-and-select [WHYT06]. Broadcasting [KM04, CK00b, LCY00]. browsing [KY09, LZL^{+}06]. browsing-based [LZL^{+}06]. Brute [ZK04a]. BSP [TW07]. BSY-2 [AACL02]. buffer [CSGL05]. buffers [SLC00]. bug [CCHW09]. build [HFRHS09]. Building [DSSL09, WHC07, CJZ04, GRRX01, LLY07, SRGL08, SL01, XYS07]. built [Wal05]. bundle [CZH^{+}08]. bursty [GAWW07]. bus [KBM05]. Business [ACDG02, CBVD07, ABCT06, LC09, LCL04, LWC06, Pfl00, PNL07, Rey07, SL03, TK00, WW09, ZMAV08]. bye [Gla00f, Gla02]. Bytecode [GK08, CY04]. Byzantine [BDK08, Zha09].

C [CLSa01, CL04a, CC05, CN00, WK00]. C/ATLAS [CL04a]. C4.5 [SSCL08]. Cache [Kar00, CE08, DPMD07, ED06, JFC08]. caching [AKP04, CLG08, HL06b, PKL03, SM06a, TYH04]. CAE [WL09]. calculating [BS09]. calibration [LHP^{+}09, LHP^{+}10]. Call [Ano02a, Ano02b]. Cambridge [LZ07]. CAmkES [KLGH07]. Can [SSCL08, HH08a, LRvV03, KBM05, LJB05]. CAN-based [LJB05]. CAN-bus [KBM05]. capabilities [LH08, TDL^{+}02]. capability [EB00, JH01]. capacity [PK02b, WLT^{+}09]. capstone [RR09]. Capture [PTRW04, SL03, TR00]. capture-recapture [TR00]. car [PG05]. card [BNvdH05]. cards [BPM06, BBC^{+}08]. Cartesian [WDS09]. Case [AN01, CL04a, WKH09, AAAC07, ASS07, Aml00, AACL02, BFPAGS^{+}08, BCF^{+}05, CCCT06, CW02, CCC06, CP07, DSB05, Del08, DZRH04, DF00, DJW08, ED04, EG00, EBGR01, EBB09, FAB^{+}07, FCL^{+}00, FLA^{+}01, Fra04, FWA09, GR05, Gur01, HF08, JG08, KK06, LC08, Lok06, NRG08, PW09, PB04, PSG^{+}09, RRD06, RRW00, RGBM06, THGL07, VAS^{+}04, Yeu00, BT03, LL04]. case-based [EBGR01]. cases [CKL08, DJW08, LWN03, YLC06, ZQZ^{+}06]. cash [FHHL09]. catalog [PTK00]. catalogues [PB00]. catastrophes [SC09]. Categorical [SA06]. categorization [KGMI06]. causal [HYC04, JJP02, JFC08]. causally [CN04]. cause [LM03]. causes [vGB02]. CBT [PKR01]. cellular [DST^{+}04, KMOS09, LLKL04]. center [WDCL08]. centered [KPS^{+}04]. centers [MH04]. centred [KK06]. centric [LD00, PTM08, Pon03, SNBH08, WDN05]. certain [SC01]. certificate [GMR08]. certificate-based [GMR08]. certificateless [HRL09]. certificates [ZSM05, RMC05]. certification [JH01, VP00]. certified [BCW05, CWH00, LL06, Sha07, WH03, WL09, YYL^{+}06]. chain [JJP02, PP04]. Challenge [CJ09]. Challenges [CL02, Rey07]. Chang [ZC05]. change [AD07, BM00b, CC09b, HKvVvdV07, Hua05b, JLC04, KL07, Shy03, TNJH07, WK00, ZLCY06]. change-point [Hua05b, Shy03, ZLCY06]. change-prone [KL07]. changeability [AS00]. changes [EK00, LRvV03]. Changing [LL04, MM00b]. channel [KMOS09]. channels [LG05a]. character [MM01b, WCLL09]. Characteristics [JFG07, BGG^{+}06, CBKK08, MMTL06, RSS00, SRDLCP09]. Characterization [CT08, HFC^{+}01, KT03]. Characterizing [CLB05, GWDE07, KL07, TNAA01]. CHC [BH09]. CHC-cluster [BH09]. checker [BDDG04]. checkers [FWA09]. Checking [Wal05, CT09, CCP05]. checklists [KLMZ08]. Checkpointing [DCH02]. checkpoints [Lea08]. Chidamber [Gur01]. Chief [Car08]. Chinese [WCLL09]. chip [CGL^{+}04, ELK06]. Chord [LZ06]. churn [AD07, HM00]. cipher [LKH^{+}08]. circuits [CGL^{+}04]. circular [CZH^{+}08]. class [BvD06, CBKK08, GAWW07, HA03, KLMC06, LKL02, LWN03, LS07, MM00b, PG04, Rad04, SM03]. class-based [SM03]. classes [CP07, EMM01, KL07, SL08, SPSM03]. Classification [DZW^{+}09, LPS02, CCCT06, CCHT09, CP09, FMSG08, KSH05, MT07]. classifier [JE02a]. classifiers [EBGR01, PS05]. Client [MSA08, CCDD00, CPL^{+}04, HC04a, NGC02, Pon05, YS04]. client-based [CPL^{+}04]. client-server [CCDD00]. Client-side [MSA08]. client/server [CPL^{+}04]. clients [FHT07]. closeness [WGH00]. cluster [AKP04, ABW07, BH09, CLG08, MKMS05, MB06, PK02a, WZJ01]. cluster-based [AKP04]. clustered [CDC09, WWC00]. clustering [ACGS^{+}08, CBK02, HLMB07, HWML04, KCB05, LZN04, LXZS06, MB06, MK06, SMDM05, TTWY04, Zhu04d, Zhu06]. clusters [CBKK08, SHS^{+}07]. CMMI [Rei00, SNJ^{+}07, YYL^{+}06]. Co [LC06b, XYS07, ZS01]. co-operation [ZS01]. co-verification [XYS07]. coarse [ZPEL01]. coarse-grained [ZPEL01]. Cobol [JPK00]. COCOA [MGI07]. COCOMO [Fai07, Sai07]. Code [DHKV06, LSC04, AD07, BHN02, BGH^{+}08, BFV04, CHL04, DDGR09, FMSG08, HM00, HJ00, LK09, Lea08, LC07, Phi06, PUPT03, RGBM06, SMR09, SHW09, THGL07, WG05, ZQZ^{+}06, ZCT^{+}09, WGC02]. Code-on-Demand [WGC02]. coding [CCH09, LWL09]. coercion [CW09]. coercion-free [CW09]. cognitive [Kuo00, ST01, ZS01]. coherent [CN04]. collaboration [CSNS05, Tan04]. collaborative [BG09, LL09, LNC01, LNPAGD^{+}06, NRG08, RR00, SG01]. collection [Fra04, LSaC01, SvV08, SK07]. collector [KCS01]. Collocation [VP07]. color [CC04, HH06, TW07]. color-complexity [CC04]. color-spatial [CC04]. com [Sha01, IT03]. COM/DCOM [IT03]. coma [Sha01]. combination [DLW08]. combinational [SH07]. combinatorial [YZ08]. combined [SCdS^{+}06]. combiner [LL06]. Combining [MS03, ED06, LC08, MØHB08]. come [Mea09]. Comments [LZ07, CJT04]. commerce [CCF^{+}04, SL02, WGC02, YC09]. commercial [CW02, SPSM03]. commit [QL03]. Common [KJLK07, FAB^{+}07].

Communication

[PH07, ZS01, ZK09, AM04, CLC08a,

CNLV07, DSC^{+}08, ELK06, HKW00, IBP03, Lai02, LLY07, LUS^{+}00, MHW01, OS09, Rav03, RwJK01, Tse07, YZ05, ZH05]. communication-efficient [Tse07]. communications [AACL02, HYC04, WF07]. community [JR09]. community-driven [JR09]. companies [KJLK07]. company [MDFG08]. Comparative [BMOKAM09, BGG^{+}06, CGP^{+}09, DZ05, EFG^{+}08, GRRX01, GR05, LO04, SUSO04, SSCL08]. compare [HBVG08]. Comparing [EBGR01, MA08, SPZ06]. Comparison [JRB^{+}06, KT03, KLMC06, LASE00, LMYMGT08, LICA09, MBB01, Mül05, NLSK04, OD05, PCV^{+}08, PW09, SM06b, WBP^{+}03, YL06, YSC^{+}06, ZPEL01]. comparisons [MM01b, Tho06]. compatibility [FK01, RFZ08]. compendium [CTY01]. competing [CLW05]. competitive [HPT07]. complete [BG06, HLWC04]. completely [DGJ^{+}03]. complex [DZRH04, Gho01]. complexity [AHGSS05, CC04, CG05, JPK00, KT03, KRHŽ05]. compliance [Kim07a]. compliant [LLK05]. Component [CSSW05, DPSU06, HTH09, TDT08, XYS07, ÅCF^{+}07, AMNT08, BKR09, Ber03, BBC05, BWM06, CGL^{+}04, CLGL05, CL02, DL06, DGP02, DGL^{+}08, FBM09, Fra04, GMS07, GDH05, Gru07, GJ08, HZ07, KBH07, KLGH07, LS04, LZL^{+}06, LZXS09, MYZC06, MvD08, PDC01, PTBP08, PKR01, Rad04, RSP03, SDG^{+}07, SPZ06, Wil03, Zhu00, Zhu06, ZS05b, dL04, HTH09]. component- [MvD08]. Component-Based [CSSW05, HTH09, XYS07, ÅCF^{+}07, AMNT08, CLGL05, CL02, GMS07, Gru07, GJ08, HZ07, MYZC06, PDC01, PTBP08, RSP03, ZS05b]. Component-Interface [HTH09]. component-level [DL06]. Component-Oriented [TDT08]. componentized [SRGL08]. components [BTV06, CCD^{+}04, DACY07, EBGR01, GS07, HH07, HGK^{+}06, KBK06, KBH07, Sch03, VP00, WGH00, WDN05]. Composing [DACY07, WDN05]. composite [Çam00b, CDEV08]. composition [CPT05, FL09, JZL07, KDS^{+}08, KBH07, KKK08, KSH09, PW03, dBvV03, MGI07]. composition-based [FL09]. compositionality [Sch03]. comprehensible [VMB^{+}08]. Comprehension [KLT07, DRW00, SKW06]. comprehensive [CELS07, VK08]. compressed [LZG07, WC02]. compression [BGG09, LSC04, TC06, WCH03, WW00]. compromise [RFZ08]. computation [Alz08, CLC08b, CLC08a, RMC05, TH05]. computation-efficient [CLC08b]. computational [CL04b]. computations [ULN06]. Computer [CPT05, Kar04a, AACL02, Gla00i, Kar04b, LNC01, RGV04, RRC07, Sta02, WZG09, Zel09]. computerized [JJP02]. Computing [PP04, SPDT06, ALT^{+}09, ANH07, BCF04, DHL06, DB06, DPMD07, GL05, HC01b, HL06b, KR08, KK07b, LCY00, LLKL04, LK04, MKMS05, MGI07, MPG^{+}08, PK01a, Tan04, WT01, ZERO00, FdSdP08]. concept [MM01b, ONR02, Par00, Xia00]. concept-based [ONR02]. concepts [BDMK03, BGH^{+}08, JE02a, MH04, ZPEL01]. conceptual [BG09, CT09, DGJ^{+}03, GPHS07, PDC01]. concerning [Mül05]. concerns [KPS^{+}04]. concise [HWHM02]. Concurrency [DY03, Jun00, KMS04, MMCB00, SY02, Shu03]. concurrent [AMNT08, CD05, FRR09]. condition [JLYK09]. conditioner [DDF^{+}05]. Conference [LP07, DGV08, Sai09, LKH09, Tse07, VE03]. conference-key [LKH09, Tse07]. confidence [JTM04]. configuration [SDG^{+}07]. Conflict [LL00, ZWX^{+}08, HGK^{+}06]. conflicts [jHjW08, LKL02]. conformance [KYP^{+}03]. congestion [GAWW07]. connected [Aba06, Aba08, SK03]. connection [JE02a, LJB05]. connection-oriented [LJB05]. connectionist [TN05]. connectivity [BMES04]. consequences [ST01]. considerations [PK01b]. consistencies [JFC08]. consistency [HC01a, WSJK08]. consistent [BG09, CN04]. consonance [KJ01]. constant [BCF^{+}05]. constrained [KP07, LKL05, SK01, ZWX^{+}08]. Constraint [VMJS06, CBG09, GM02, HCDJ08]. constraint-based [CBG09, GM02]. constraints [LcLsW06, LYC04, MK08, ZKL^{+}09]. Constructing [WZG09, YZ08, dBvV08]. constructs [BBS00]. ConSUS [DDF^{+}05]. contained [LY01]. Content [AKP04, CLG08, FdSdP08, KY08, LK01, SL01, TR00]. Content-aware [AKP04]. content-based [CLG08, KY08, LK01]. content-oriented [SL01]. contention [MA09]. Contents [Ano01c, Ano01d, Ano01e, Ano01a, Ano01b, Ano02e, Ano02f, Ano02g, Ano02c, Ano02d, Ano03a, Ano03b, Ano03c, Ano03d, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano05e, Ano05f, Ano05g, Ano05a, Ano05b, Ano05c, Ano05d]. Context [KPTV09, CELS07, CMNA^{+}09, DPMD07, KK07b, Kri06, MPG^{+}08, RT07, SW05]. Context-aware [KPTV09, DPMD07, KK07b, RT07]. contiguous [BMOKAM09, SK03]. continue [KWT^{+}00]. continuous [LU06]. contracting [AG08, LGW09]. contracts [BS03]. contribution [RSM00]. contributions [VM07]. Control [LVMM07, AAAC07, CDS02, CCW02b, CLH07, CSGL05, Cho04a, Cho04b, Cho05, CHL05, CC05, CC06, CHY^{+}05, CFN07, DY03, DZRH04, Fer00, GWvD08, GAWW07, HSM^{+}07, HYC02, HC04a, JMP07, JE02b, JW06, Jun00, KRC00, KMS04, LNC01, LZG07, LY01, MV09, MDMC06, MH04, NKJT09, PTM08, SP08, SY02, Shu03, ULN06, WCLK07]. control-based [HSM^{+}07]. control-theoretic [MDMC06]. controllability [HYC02]. controlled [DSA^{+}04, Mül05, PUPT03]. controller [LCF^{+}06]. Controlling [HY03, ELH00, WL05]. Controversy [Ano01f]. COnversation [MGI07]. COnversation-based [MGI07]. conversion [CGMPAP08]. Convertible [WH02]. cooperated [TCSC04]. cooperative [AKP04, Dar02, FRR09, RDD02]. coordinated [MHW01, CGP^{+}09]. coordination [CJKC09, JF04, mJKME01, PNL07]. coordinator [LSH09]. Coping [Moy00]. COPS [Dar02]. copy [LC02, WLC07]. copyright [CWP09]. CoRAL [AT09]. CORBA [CLCY04, LJB05, RDD02]. Corba-based [RDD02]. core [LK09]. Corner [Car02, Gla02, Wey01, Wyn01, Car04, Ano01f, Ano01g]. correct [PTBP08]. correction [DBO05, LH06]. correctly [AMNT08]. correctness [BGH03, DACY07]. correlated [GAWW07]. correlation [LP05, LGL08]. Corrigendum [Gla00d, LHP^{+}10]. COSMIC [CGMPAP08, KBM05]. Cost [Hua05a, ACGS^{+}08, CMC04, CGSGR06, HL06a, KSS03, KRCK08, LP00, LXG09, MMC05, MA08, NR04, PUPT03, SA06, Wes02, ZS01, ZK09]. cost-effectiveness [NR04]. cost-estimation [CGSGR06]. Cost-reliability-optimal [Hua05a]. costs [EL07]. COTS [CCD^{+}04, MSB^{+}02]. COTS-based [MSB^{+}02]. counting [HOR01, OR00]. Coupling [Fer00, WK00, FAB^{+}07, GS07, Xia00]. courses [VM07]. COVAMOF [SD08]. coverage [CFN07, Gok09, TH05, YL06].

CPLD [KK07a]. CPLD-s [KK07a]. CR [LLL06]. CR-CSFQ [LLL06]. crawling [YWLG02]. created [KP07]. crew [GH04]. crises [Gla00j]. Crisis [Gla00l]. crisscross [CCP05]. criteria [LVMM07, VMJS06, YL06]. criterion [PG04]. critical [CC08c, CGW08, DGV^{+}07, GD04, LJS05, LM03, MM01b, SS04]. Cronus [ACT^{+}08]. Cross [MDFG08, NLSK04, PCV^{+}08]. Cross-company [MDFG08]. cross-cultural [PCV^{+}08]. cross-national [NLSK04]. Cryptanalysis [LLCL08, Sha05, TM06, WL05, ZC05]. Cryptographic [LKJL01, HY03]. cryptography [YC09]. cryptosystem [DHL06, JW06, LL06]. cryptosystems [CHC01, EHKH04]. CSEET08 [Sai09]. CSFQ [LLL06]. CSLF [LLL06]. CSP [Yeu00]. CUDL [KNYS09]. cultural [PCV^{+}08]. culture [TW08a]. cumulative [BS09]. curricula [KBBW05]. curriculum [BM05, BT05, Cow05, Wen03]. curve [EHKH04, JW06, LWHS05, YC09]. customer [Cha06, LCL04, Lin01, LS05b]. customer-oriented [LCL04]. customized [CCF^{+}04, ZBLG07]. CVM [DSC^{+}08]. cybernetics [BCDM06, XSS06]. cycle [mJKME01, LD00]. cycle-time [mJKME01]. cycles [SHS^{+}07]. cyclic [PK01a].

D [Aba08, BRC09, BMOKAM09, BGG09]. Data [CET^{+}08, HY00, VM00, AAAC07, BRMA^{+}09, BNW^{+}08, CC02a, CCY^{+}09, CD00, CY00, CCW02b, CL06a, CWC04, CLB05, CTL08, CK00b, CBK02, Har04, HCS09, HL00a, HC01a, HY01, HL06b, KNYS09, KUK07, KRC00, LCY00, LKL02, LKL04, LLC^{+}09, LVMM07, LWC06, LWL09, LO04, MCCC03, MPST06, MK08, MLC09, MR00b, Özm09, PSH06, Sal02, SHS^{+}07, SA06, SA08, SS07, SSCL08, TVA04, TTWY04, TW07, TC06, TL07, VK08, WDCL08, WZG09, WQ06, WLT^{+}09, WDN05, YHHR03, ZM06]. data-centric [WDN05]. Database [DK08, PK01b, AV02, CM05, HLWC04, HDLK00, JR09, Jun00, KRK00, KRP02, KLC02, LKL02, LK01, LPJP09, LY01, MDFG08, NG08, NGM08, PS09, PQLN04, SVMAM04, YLC08, ZHS01]. databases [AJCM08, BH09, HL09, HLL01a, KYPW06, LLL00, LL00, LLT^{+}09, LLKL04, TTWY04, ÜDUG04, YC08a]. dataflow [CD07]. Db4XML [SVMAM04]. DBMS [LKW^{+}09]. DC [YL06]. DCOM [Dar02, DZ05, IT03]. DDH [NLKW05]. DDH-based [NLKW05]. DDO [LKH^{+}08]. DDO-64 [LKH^{+}08]. De-motivators [BH03]. deadline [LLL00]. Deadlock [IT03, KSAOK04]. Deadlock-free [IT03]. debtor [SLS08]. debugging [Alz08, BBS00, Shy03, WSQM05, WQ06]. decay [AS00]. Decentralized [AS01, AMNT08, CHL05]. decision [ABG02, BFV04, GLJ00, Gho01, KWT^{+}00, KLMZ08, mJKME01, UZ09, ZKL^{+}09]. decision-making [GLJ00, Gho01, KLMZ08]. decisions [BL09, JBA08, YL06]. Declarative [Lok06]. Decomposing [CCDD00]. decomposition [KK07a, WHG01, YGH^{+}08]. decompressing [LCLL08]. decryption [SWH^{+}09]. Defect [CC07, ZR04, Bif03, EE08, KT03, LASE00, SPSM03]. defect-prone [EE08, SPSM03]. defects [Bif03, JMP07, KJLK07, LPS02]. defining [KvV06]. definitions [CK02b]. delay [KTK01]. Delegated [WHG01]. delegation [SM09]. delegation-based [SM09]. delivery [KD05]. delta [AD07, HM00]. demand [HH05, NXS00, PLF05, WW00, WGC02]. denial [SKZ^{+}04]. density [HWML04]. density-based [HWML04]. Dependability [VP00, BGG^{+}06, DB06, LC09, XZAR06]. dependable [CGP^{+}09, GRRX01, dLGR06]. dependence [BGH^{+}08, CCW02a, CCW02b, HY00, HY01]. dependences [PC01]. dependencies [DCAC09]. Dependency [HTH09, LSC04, WQ06, YR09, ZKL^{+}09]. dependent [FS05, LU06, LH08, TSSD09]. Deployed [GDH05]. deployment [PDC01, SDG^{+}07, ZP06]. depth [PUPT03]. derivation [DSB05]. derive [FCL^{+}00]. derived [HKN^{+}07]. Deriving [AJCM08]. Description [OKS08, FIGCLN^{+}02, LZXS09, RS06, SMG08]. descriptions [CP07]. Design [ALT^{+}09, BL09, CLG08, DDGR09, HRL09, KRP02, LWS^{+}03, LJB05, LLK05, LKW^{+}09, PWCC01, SKZ^{+}04, TDT08, TKA^{+}02, YY04, Zha09, vGB02, AA07, AL05, ACDF01, Bat08, BM07, BWDP00, CLX^{+}04, CGL^{+}04, CH07a, CL04a, CCC06, DI05, DSRS03, DSA^{+}04, ED04, EMM01, FBM09, FIGCLN^{+}02, FHT07, GD04, Gla00h, Gla00i, GPM08, GMS07, HALS08, HKN^{+}07, HL00a, HCC08, HC04b, Hus01, JBA08, JMSS07, KBK06, KK06, KCS08, KP07, LASE00, LRvV03, LH04, LT09, LSH09, LZG07, LY09, LSaC04, MLB09, MR00a, NWZ05a, PK01b, RDD02, SNBH08, Spi01, SDG^{+}07, SPSM03, SC09, TA02, TBGH06, TJH07, TNJH07, WZJ01, Wij03, WSQM05, YWLG02, ZM06, Zhu04c, KY09]. designated [CC09a, KBD09]. Designing [Ber03, GH02, LCLL08, PB04, SZ06, SVMAM04, SD02, CCG^{+}07, CGP^{+}09, CW09, ZMAV08]. designs [PG05, SK02]. desires [HKvVvdV07]. detect [FW00]. Detecting [Sta03, WCH03, WW09]. Detection [WC02, CCP05, DBO05, HWM01, HWHM02, HWH^{+}03, HZ07, JZ07, LASE00, LH06, TR00, WBW^{+}06, WZG09, WJT09, WLC07, WHC07, ZWX^{+}08]. determinants [VEM^{+}01]. determined [ZWX^{+}08]. Determining [Kel09, SvV08]. developer [SHW09]. Developing [BM05, JHSB09, LK09, SG06, CCF^{+}04, O’B08, SPZ06, REF^{+}07]. Development [RO09, ÅCF^{+}07, AW07, BG09, BM00a, BDGR01, CH09, CBS00, CL02, DZ00, EB00, FLA^{+}01, GML05, GR05, GPHS07, GJ07, HDGZ06, Har00, HH08a, HHW01, HMC01, JPKP04, JJ06, JK00, JTM04, Jør04, KWT^{+}00, KKLP09, KPME02, KPME05, KRCK08, LS04, LCL04, LK02, LMYMGT08, MR01, McB08, MMTL06, MKK09, MSB^{+}02, NSL^{+}07, NER01, OAZ08, PW09, PLP04, RGBM06, RDD02, RS00, RMO^{+}08, SCdS^{+}06, SFJ04, ST01, She02, Sta09, SJK07, TDT08, TK00, Wes02, ZE03, ZGH^{+}07, ZS01, dOZR^{+}04, DL06]. device [BBG^{+}04]. devices [LKW^{+}09, LKL05, PSG^{+}09, SFJ04, VA08, ZK04a]. DHARMA [MMM00]. diagnosability [LORB03]. Diagnosis [SK02, JC02, LORB03, LYX09, WY04]. diagram [YLC08]. diagrams [BCV06, KZDX09]. dialogue [LHLY05]. dialogue-based [LHLY05]. dictate [HKvVvdV07]. difference [CL06a, WLT^{+}09]. differences [BBS00, JKD02]. differencing [WWTH08]. different [Mül07]. Differential [LGW09]. differentiated [TYH04]. difficulties [KLT07, She02]. digested [LHLY05]. Digital [Lin01, CWH00, HRL09, HL00b, HYJL04, SRGL08, TCC02, YKC^{+}05, CDS07]. dimensional [Aba06, CCW02a, HLW08, LcLsW06, LO04]. DIPS [MC04]. Direct [CBZ00]. directed [KS04, CCHW09]. direction [CCW02a]. directional [FL05]. disadvantage [CDS07]. disciplined [RMO^{+}08]. disclosure [CLH07]. Discovering [CD05, MV09, KV05]. Discovery [SMR09, DMQ07, HWHM02, LK09, MPST06, SSM^{+}09, ZS05a, ZMN05, dBvV08, MPG^{+}08]. Discrete [GAWW07, HRN^{+}01, KDEK04]. Discrete-time [GAWW07]. Disjoint [CLC03]. disk [CCSC01, CCSC07, KEK04, LKL05, TSSD09, VM00]. disk-based [KEK04]. disk-scheduling [CCSC07]. dissemination [HL06b, PSH06]. distance [AM04, LKL04, WG05]. distances [CCW02b, CH07b]. Distinguishing [LUS^{+}00]. Distributed [HCS04, Loo05, MKM^{+}06, TDK^{+}07, AZW07, AACL02, Ati00, AMNT08, BKZ^{+}06, BLL02, CN04, CLX^{+}04, CJZ04, CET^{+}08, DGL^{+}08, ESW06, FL09, GBL08, GTA09, GLJ00, Gho01, GD04, HSM^{+}07, HC01b, JE02b, JLYK09, KMSMD08, Kar01, KUK07, KW00, KPG^{+}07, KMOS09, LLL00, LNC01, LPJP09, LR04, LUS^{+}00, LNPAGD^{+}06, LH01b, MEH05, MHW01, QL03, Rav03, SM09, SO03, SM00, SCdO02, SC07, SOC^{+}03, SK04, TMD07, USLC01, WT01, WBW^{+}06, WCLK07, WFWL09, WKH09, YY04, YYWW07, ZS01]. Distributing [CKL08, WZJ01]. distribution [CBZ00, CKL09, CLG08, WHHT08, YS04, ZK04b]. diversity [BFLP09, YS02]. DL [HRL09]. DL-based [HRL09]. DMMX [CSaLG02]. Do [Mül07, Kru08, PCV^{+}08, PVSG05, SNJ^{+}07, CPT05]. Document [LLH08, CDS07, CK02b, KY09, LL09, WHG01, ZSM04, ZL06]. documentation [Ber03, HS03, JAvdV09]. Documenting [JBA08]. documents [BHL00, CH07a, LASE00, PWLH06, TH02]. DOM [KY09]. Domain [Sut00, dOZR^{+}04, BRC09, BGH03, BKB^{+}07, CL06b, Del08, FBM09, FCL^{+}00, FLA^{+}01, Fra04, JOZ03, KG09, KKP06, KPS08, SL03, Spi01, ZGH^{+}07]. Domain-oriented [dOZR^{+}04]. domain-polymorph [FBM09]. domain-specific [Spi01, ZGH^{+}07]. dominance [MC01]. domino [LLLZ06a, LLLZ06b]. Dot [Sha01]. Dot-com [Sha01]. down [HWML04, WCLL09]. Downloadable [HCKY08]. DPE [CHL05]. DPE/PAC [CHL05]. DR [HCKY08]. DR-TCP [HCKY08]. DRAMA [KPS08]. DRE [LBS^{+}07, SDG^{+}07]. driven [ABCT06, BKR09, Boz00, CCHW09, CCC06, DI05, GWvD08, GMS07, HRN^{+}01, JR09, KKLP09, Mus03, Özm09, PLCC09, Phi05, Phi06, PQLN04, PSG^{+}09, TKM03, dBvV03, AJCM08, DL06]. drivers [BCB09]. DSEA [LLLZ06a, LLLZ06b]. DSM [INS00]. DSP [LC05, LC07, PNM04]. DTA [Rav03]. Dual [WY04, HCC05]. due [JLC04]. duplication [HTK00]. durations [ZWX^{+}08]. duty [LWL04]. DWT [CWP09]. DWT-based [CWP09]. Dynamic [CSaLG02, DTV09, INS00, OC04, QK08, SSM^{+}09, SM03, YC08b, ADZ^{+}09, BRC09, BRMA^{+}09, BGH03, Boz00, CDI07, CD00, CBG09, DI01a, FL09, GP05, GDH05, GWDE07, HSM^{+}07, HLL01a, KYP^{+}03, KBH07, KDEK04, KPG^{+}07, KMOS09, LBS^{+}07, LLY07, LLKL04, LSaC04, MMM00, MHW01, MLC09, MKM^{+}06, MM06, OD05, RwJK01, SM09, SLS08, SA08, SA05, SH07, WHYT06, WG05, Zhu06]. Dynamically [KPT09]. dynamics [KLRW01, RRT01, SG01]. DYNAMOS [RT07].

e-business [LC09, ZMAV08]. e-commerce [WGC02]. e-contracting [AG08]. e-science [SZZ06, ZL06]. each [LY01]. earlier [Ber02]. Early [Gru07, OAZ08, WJT09, ED04, JZ07, LKB06, MMC05, PCC02]. EASE06 [KB07]. EASY [MPG^{+}08]. eBizBench [LCL04]. ECEM [LP05]. eco [SZZ06]. eco-security [SZZ06]. economic [CC03]. economics [AW07]. edge [WLT^{+}09]. EDI [LH01b, LTC01]. editing [RDD02]. Editor [Ano01g, Car02, Car04, Car08, CDW07, Gla02, DGV08, Wey01, Wyn01]. Editor-in-Chief [Car08]. Editorial [Ano02h, Ano02i, Ano02j, Ano02k, Ano02l, Ano02m, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03l, Ano03m, Ano03n, Ano03o, Ano03p, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano04m, Ano04n, Ano04o, Ano04p, Ano04q, Ano05h, Ano05i, Ano05j, Ano05k, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p, Ano05q, Bae05, Bae06, BJM02, Car08, CF08, CSSW03, CHS^{+}07, EST06, HJP07, KMR05, LW02, MN08, PGTS08, RW01, TW08b, WM09, vV08, HLM^{+}09, Ano04f, Ano04g]. Editors [MW08]. educating [SJ05]. Education [Sai09, BT05, CHZY03, Fai07, FCL^{+}00, GSB^{+}07, Haz02, Let00, Mea09, PKR01, SW05]. effect [HCN00, WW00, XNP07]. Effective [LLL06, WQ06, IWF07, KPS^{+}04, LC05, LC07, NWZ05a, PC02, PC01, WZG09, ZG07, ZK09]. Effectiveness [CCL01, CKM06, ELH00, JK00, NR04, SL08]. Effects [HCWN05, Kri06, AW07, CGW08, FCSM09, HMC01, Hus01]. efficacy [MMTL06]. Efficiency [SKKL07, FMP09, Hua05a, MK06, YTH04, ZS05a]. Efficient [GH04, HPT07, HL06b, LLK04, Lee07, MPST06, OFWP07, PWLH06, SM00, TH05, ULN06, WL09, Aba06, AM04, BHAM09, CD00, CLH07, CLC08b, CLG08, CTL08, CBK02, DA07, FS06, HL09, HC04b, JW06, JC02, JLYK09, KPSK09, KMOS09, LWHS05, LC07, MC04, MLC09, MM06, PHN08, PJ09, SC08, SOC^{+}03, Tse07, TL07, TL09b, USLC01, ÜDUG04, YWLG02, YC09, YC08a, YSK06, YC08b, MPG^{+}08]. efficiently [IJC03]. effort [CH07b, GJ07, HBVG08, Hua05b, JIS03, JTM04, Jør04, LH08, LMYMGT08, MS03, MDFG08, RSS00]. EIS [Sal02]. electronic [Ber03, CW09, FHHL09, PTK00, SL02]. element [NG08]. element-based [NG08]. elements [HLWC04]. ElGamal [CWH00]. ElGamal-like [CWH00]. Eliciting [ASS07, CP07, DB06]. elimination [CCH09, Özm09]. elliptic [EHKH04, JW06, YC09]. elusive [SKZ^{+}04]. email [CP09]. embedded [BRMA^{+}09, CC03, CS04, CG05, Del08, HLC^{+}09, JHSB09, KCS01, KP07, KLGH07, LNY06, MYZC06, PB04, SO03, SP08, WCLK07, WDN05, XYS07]. Embedding [Cho04a, KC09, WLC08]. EMBOT [ZEY04]. Empirical [AW07, Har00, MBB01, RK00, SKW06, UN09, ACS07, AL05, AS00, BKZ^{+}06, BVN07, BBS00, BGH^{+}08, BvD06, BT03, CH09, CN00, CGSGR06, CGMPAP08, DSRS03, EJ01, HH07, IS03a, JPK00, JH01, KPME02, KPME05, KT03, LS07, LJS05, LTC01, LWC06, DPS03, MSA08, MM00a, MR00b, Mur08, NWZ05b, OOD09, OD05, PLM07, RGV04, SSA08, SC01, Tan00, VK08, VHF02, YR09]. employee [LC09]. Employing [Deu01, CDS02]. enabled [LPJP09, SDG^{+}07]. Enabling [HMSW03, JZL07, SKKL07]. enactment [GPHS08]. encoding [HL09, MLC09]. encompassing [LD00]. encryption [CHC01, GMR08, LLLZ06a, LLLZ06b, LLCL08, SWH^{+}09, WH02]. End [Gla00e, KD05, LS05a, WCLK07]. end-to-end [KD05, WCLK07]. Energy [TL07, TL09b, JLYK09, PJ09]. Energy-efficient [TL09b, JLYK09, PJ09]. enforce [AAAC07]. Engaging [JR09]. engine [CHL05, HKW00, SVMAM04]. Engineering [BL03, BB08, CBVD07, GR05, KB07, Sai09, VE03, AAAC07, ADZ^{+}09, AA07, Ale05, BM05, BNvdH05, Ber02, BDA^{+}02, BKB^{+}07, Bud00, BT05, BM00b, CC08a, CSNS05, CDZ07, Cow05, DA07, DJW08, DD01, EC04, Fai07, FCSM09, Gla99, Gla00c, Gla00d, GC02, GC03, GC05, GPM08, GSB^{+}07, HF08, Haz02, HAHH06, HFRHS09, JR09, KPTV09, Kim07a, Kim07b, KBBW05, Mea09, Mil00a, ML08, MR00b, PILO06, Phi06, PH07, PKB09, RR00, Sai02, SW05, SCdS^{+}06, SG01, TKM03, TL09a, TCG06, VM07, VHF02, VEM^{+}01, WTG^{+}08, WTG^{+}09, vDB05, CSSW05, GC01]. Engineering-based [GR05]. engineers [JFG07, Let00]. engines [CCF^{+}04]. England [LZ07]. English [CHL^{+}08]. enhance [FLA^{+}01]. enhanced [MC01, PK02c, ZEY04, ZSM05]. enhancements [OS09]. Enhancing [LH08, PTK00, ZS05a, ZSP01]. Enriching [JAvdV09]. Ensuring [ABW07]. enterprise [CCG01, CG03, LBS^{+}07, LK02, NKJT09, SL02, TSPH06, PNL07]. entity [SZ06, YLC08]. entity-life [SZ06]. entity-relationship [YLC08]. entropy [LZL^{+}06, Özm09, SS04]. entropy-based [Özm09]. environment [CPL^{+}04, DK01, HK09, HC04a, HLYL06, KKP06, LCL04, LPJP09, LNY06, NLKW05, PILO06, SZZ06, SOC^{+}03, TA02, TMB02, VA08, ZR04, dOZR^{+}04]. environmental [HCWN05, ZSP01, ZLCY06]. environments [ADZ^{+}09, CELS07, CL04a, DI05, DSSL09, DY03, DTV09, DPMD07, HL06b, HCC05, KGT02, LLK04, LSZ^{+}07, LLH08, MC04, MGI07, MPG^{+}08, NKJT09, PJ09, RT07, SCdO02, SC08, Tan04, YC09, ZMN05]. epidemic [MK08]. equation [SM08]. Era [Gla00e, Gla00g]. erosion [vGB02]. ERP [NGC02, WSJK08, WOH08]. ERP-client [NGC02]. Erratum [KPME05, LLLZ06a]. Error [TC06, LP00, LS07, MT07, SL08, TBD^{+}08]. error-prone [SL08]. errors [BG06, HCS09, Wes02]. essential [KBK06]. Establishing [BVN07]. establishment [XSS06]. estimate [BPM06]. estimates [ELH00, GJ07, MØHB08]. Estimating [EG00, SeMC02, LP00, MM01b]. Estimation [BHL00, ABG02, ACGS^{+}08, Bif03, CH07b, CGSGR06, HLW08, JIS03, Jør04, KPME02, KPME05, KPG^{+}07, KRCK08, LXG09, MMC05, SSCM^{+}04, SA06, SH07, THP^{+}06, THGL07, OOD09]. estimations [TR00]. estimators [TR00]. ETCS [ZH05]. ETOOD [TA02]. evaluate [AP09, BM00b]. Evaluating [BGH03, BS09, Bif03, CCG^{+}07, CdOBT07, ABG02, Bat08, HCC08, KV05, LCLL08, MMM00, SM07, YR09]. Evaluation [Bud00, HJ00, AZGvG09, BKZ^{+}06, BMOKAM09, BM00a, BNW^{+}08, BM07, BGG^{+}06, BT03, CJ05, DZW^{+}09, EJ01, Fug03, FL09, GLJ00, GPML06, HHW01, HLWC04, LH04, LPS02, LZG07, Lop03, MK06, MM00a, Nae01, NsL00, OS09, ONR02, ÖKT09, PWLH06, PTRW04, PB00, PG04, QHS08, SM06a, SS04, SSCL08, SK02, TK00, TDK^{+}07, TMD07, TMB02, VK08, WHB01, YWLG02, ZH05, KB07]. evaluations [SUSO04]. evenly [CKL08]. event [DPSU06, HRN^{+}01, KBM05, KDEK04, LP05, LGL08, PLCC09, SFSE05]. event-based [DPSU06, KBM05]. event-driven [PLCC09]. event-triggered [SFSE05]. eventual [BDK08]. every [GSB^{+}07]. evidence [JR09, Wen03, Wes02, NSL^{+}07]. evidence-based [JR09]. Evolution [AK08, AD07, AN01, AL05, ABCT06, BM00b, CT08, DD01, FL09, GPM08, HM00, Har00, KLRW01, Kel09, KBH07, KP07, LS07, LM03, PLM07, SM09, SL08, ZR04]. Evolutionary [WWB09, BCB09, Sal02, SA08, TN05]. Evolving [PG05, PTBP08]. exact [LHSK06]. examination [MR00a]. Examining [FMSG08]. example [KLRW01, LK09, Vau07]. examples [HS03]. Exception [CCHW09, FdSBR06, FRR09, GRRX01, JCYC04]. exchange [CLC08b, YC09, ZSM04]. exclusion [KTK01]. executable [HS03, SM00]. executables [CPiLH09]. execute [CLW05, SHS^{+}07]. Execution [CZH^{+}08, CBZ00, LU06, SOC^{+}03, WQ06]. exogenous [BCB09]. expansion [CL06a, WLT^{+}09]. Experience [Fra07, TNAA01, TL09a, ADZ^{+}09, ACDG02, CCF^{+}04, CP07, FM08, LG03, McD02, OR00, DB06, LNY06]. experienced [Moy00]. Experiences [LBvVB02, SN07, BT03, VJB06, LNPAGD^{+}06]. experiment [BS09, DSA^{+}04, PUPT03, SHW02]. Experimental [AD07, HCN00, YS02, ZPEL01, BNvdH05, CJHB08, CCCT06, LASE00, MMTL06, PG04, RSS00, SK02, Zel09]. experimentation [HJ00]. Experiments [JG08, AP09, CGP^{+}05, Mil00a, Mil04, Mül05, SKW06]. expert [BHB^{+}05, GJ07, Jør04, MØHB08, THGL07]. explicitly [GJ08]. Exploiting [BFPAGS^{+}08, CFN07, SHS^{+}07, HH00]. exploration [GD04]. explorative [KLT07]. Exploratory [ZSP01, ONR02, PVSG05, PV06, SNJ^{+}07, Tan00, ZGH^{+}07]. Exploring [BWDP00, DC09, HRN^{+}01]. expressed [BGH^{+}08]. Expressing [BNR09]. expressions [CK02a, PC02, PWLH06]. extended [LLK05, MDFG08, LKJL01]. Extending [HL09, JF04, Lut00]. extensibility [KFS^{+}02]. extensible [CLL05, CC03, KLMC06]. extension [CG03, KCS08]. extensions [CSaLG02, JSBR09]. extent [EG00]. external [ABG02, GMB^{+}09, PS09]. extract [IWF07, TH02]. extracted [WPP^{+}09]. Extracting [YLC08]. Extraction [DS04, EKV05]. extranet [DK01]. extreme [HBM05, TW08a, SJ05]. Eye [LSZ^{+}07].

F [FLA^{+}01]. facilitate [LT09]. Facilitating [KCS08, ZMN05, WSJK08]. facilities [PK01b]. facility [WHN^{+}01]. fact [JBA08]. factor [HMC01, MM01b, PUPT03, Tan00]. factors [ACS07, CH09, CC08c, Gla00k, HFC^{+}01, Kel09, MKK09, RH02, RH03, WSJK08, ZP00, ZSP01]. failure [BHXN05, CCCT06, CGW08, DMQ07, JX07, TSA08, WGW^{+}09, ZP06, dL04]. Fair [FHHL09, JL04, SA05, LLL06, ZSM04]. families [DSB05]. family [AP09, CGP^{+}05, Del08, Lut00]. far [Mea09]. fast [LK01, PS09]. faster [LHSK06]. FasTLInC [GM02]. Fault [CC01, LY09, AZGvG09, AT09, BFLP09, CJZ04, CT00, GK08, GH02, Gon08, HTK00, Lea08, LKH09, LGW09, LCH^{+}04, Lin07, LYX09, LH06, MR00b, NJ07, RW00, SSO05, SS04, TR00, THGL07, Tse07, VMB^{+}08, WY04, WKH09, Zha09]. fault-proneness [Gon08, MR00b]. fault-tolerance [GH02, Lea08]. Fault-tolerant [CC01, LY09, AT09, CJZ04, CT00, HTK00, LKH09, Lin07, Tse07]. faults [DBO05, JLC04, Sta03]. faulty [EMM01]. FBCM [KMKY07]. FC [WCLK07]. FC-ORB [WCLK07]. FDB [KNYS09]. FDDI [CCL01]. FDDI-M [CCL01]. FEA [LL07]. FEA-M [LL07]. Fears [HKvVvdV07]. Feasibility [PC04, BRC09]. Feature [GPML06, ESW06, LXG09, TBD^{+}08, WDS09, WG05]. feature-based [WG05]. features [CC04, CP09, FMSG08, PHN08, RS00, WBP^{+}03, WGH00, FdOdL04]. Feedback [HSM^{+}07, CGHL07, KMSMD08, KCB05, KY08, YL09]. field [HAHH06, ZP06]. File [ZK04b, CLG08, CT00, KFS^{+}02, MCC02]. files [HH05]. Filling [GMS07, LWHS05]. filter [PCC02]. Filtered [WDS09]. filtering [HCC05, KY08, LL09]. Final [Gla02]. financial [Aml00, LHLY05]. finding [HFC^{+}01]. findings [Sal02]. Fine [FAB^{+}07]. Fine-grain [FAB^{+}07]. fingerprints [DS04]. finite [HM09]. FIPA [CMNA^{+}09]. fire [WJT09]. first [CCDD00, Gla00g, Gla00i, LC00]. fit [WSJK08]. five [HKN^{+}07, IBP03]. fixed [CGHL07, LHSK06]. fixed-memory [CGHL07]. fixed-priority [LHSK06]. fixing [CCHW09]. flags [WWB09]. flash [BH09, LKW^{+}09, PS09]. Flash-aware [PS09]. flash-based [LKW^{+}09]. Flattening [WDS09]. Flexible [LSH09, ZL04, Cho04b, DA07, Har04, KBH07]. FLOSS [BCB09]. flow [AAAC07, CCW02b, Cho04b, Cho05, CC05, CC06, Fer00, FdSBR06, FRR09, HKY01, HC04a, LL09, LVMM07, LZG07, SKKL07, ULN06, ZG07]. FMF [RH06]. FNDS [LHLY05]. folder [LLH08]. folding [TCSC04]. force [ZK04a]. forecasting [JJP02, PH06]. forensics [CDS07]. Formal [CW02, LNPAGD^{+}06, RDD02, BBC05, FIGCLN^{+}02, FIBRGCLN05, GHKR04, HRŽ06, JE02a, KSS03, Wal05, WW09, YKC^{+}05, ZAO08]. formalism [Ale05]. formalisms [KEK04]. Formally [HYS^{+}04, Phi04]. formative [PB00]. formats [CF07]. formed [BM07]. formidable [Rei00]. formulation [GP05]. forward [Tse07]. foundation [GPHS07, PDC01]. fourth [DHKV06]. FPA [KRHŽ05]. FPGA [EHKH04]. fractal [WCH03]. fragmentation [SeMC02]. fragments [Zhu04d]. frame [GLJ00]. framework [AZW07, AS00, BG09, CDEV08, hCSW^{+}04, CL04b, DH09, DSSL09, DB06, FdOdL04, FLA^{+}01, FL09, Gru07, HALS08, HLMB07, HCWN05, HZ07, JCC05, KPS08, LBS^{+}07, Lop03, NWZ05a, OAdLC07, DNAM05, QHS08, RMC05, SRGL08, SCdS^{+}06, SK02, SL07, Tan04, TSPH06, WHB01, ZC08, Zha09, CH05]. Frameworks [CGP^{+}09, FCL^{+}00]. Frank [LZ07]. free [Aba06, CW09, IT03, LL00, Rad04, SSA08, WCH03]. free-list [Aba06]. free/open [SSA08]. FreeBSD [YSC^{+}06]. frequency [BPM06, HH05]. frequency-hopping [BPM06]. frequent [LLT^{+}09, SPDT06]. friends [CN00]. frustrated [Gla00a]. Full [CMNA^{+}09, LKH^{+}08]. full-round [LKH^{+}08]. fully [KSAOK04]. Function [OR00, HOR01, SHW09, WWTH08, WWB09, ZLCY06, AHGSS05]. function-assigned [WWB09]. Functional [TT09, AP09, CGMPAP08, HRŽ06, LC08, MLB09, Nae01, XZAR06]. functionalities [CFFT08]. functionality [PLF05]. functions [KPT09]. Fundamental [BDA^{+}02]. fundamentals [Aml00]. fusion [HF08]. future [MKNS06, PSK05, TDL^{+}02, Wen03]. Fuzzy [Zhu04a, ACGS^{+}08, EL07, LMYMGT08, CWP09].

Galois [JE02a]. game [BNvdH05]. gamma [CC01, CLC03]. gap [GMS07]. gaps [CJKC09, JKWL09, O’B08]. garbage [KCS01, LSaC01, SK07]. gathering [CLCY04]. general [BJ03, CCW02b, HKN^{+}07]. Generalized [SM06b]. Generating [LWN03]. generation [CS04, FWA09, GH04, JR09, LU06, LC07, LC08, Phi05, Phi06, PQLN04, SA08, SZPMK04, THP^{+}06, VA08, WBW^{+}06, YLC06, ZAO08, ZBLG07, ZL06, dlRT06, RR09]. generic [CHY^{+}05, Gru07]. generics [RFZ08]. Genetic [OW04, PS05, BRMA^{+}09, DXPY03, GBL08, JJP02, Yoo09]. Genetic-algorithm-based [OW04]. genomes [HLWC04]. geographic [KPSK09]. geometric [CJT01]. Gibbs [BT05]. global [CCSC07, SC09]. GM [SOC^{+}03]. GM-WTA [SOC^{+}03]. GMPLS [WGY^{+}08]. go [Mea09]. Goal [KKP06, BCV06, CCCT06]. goal-oriented [CCCT06]. goals [CCHW09, OW04]. Gompertz [OOD09]. Good [Gla02, Gla00f, MM01b]. Good-bye [Gla02, Gla00f]. grades [TYH04]. graduate [VM07]. grafting [SC00]. grain [FAB^{+}07]. grained [ZPEL01]. Granular [KK07b, PS05]. granularity [INS00, Jun00]. graph [CLX^{+}04, KZDX09, LL00, SM06b]. graph-based [SM06b]. graph-oriented [CLX^{+}04]. graphically [Wal05]. graphs [QK08]. gray [HH06]. gray-level [HH06]. Grid [LPJP09, Zhu04b, CL04b, LK09, Zhu06, ZG07, ZK09, LC06b, NKJT09, SZZ06, ZL06]. Grids [FMP09]. Grigoris [LZ07]. grounded [CO08]. Grounding [OHS01]. groundwater [LHP^{+}09, LHP^{+}10]. Group [CCSC01, AS01, CJT04, CNLV07, HYC04, HDLK00, KPG^{+}07, LL06, LLY07, NLKW05, RDD02, Sha05, WF07, WHHT08, XY02, YZ05]. group-oriented [LL06, WHHT08]. grouping [WHYT06]. groups [HBM05]. groupware [BKZ^{+}06]. growth [Hua05b, LHC^{+}05, ZLCY06]. GSM [FIGCLN^{+}02]. GSR [CCSC07]. guessing [SCH05]. Guest [Bae06, BJM02, CSSW03, CHS^{+}07, LW02, RW01, CDW07, DGV08, MW08, Wey01, Wyn01]. guide [PIGÖ08, dBvV08]. guidelines [SN07]. Guo [LLLZ06a, LLLZ06b].

h [JJ06]. H. [LC06b, Zha08]. halftone [CCP05]. handheld [PSG^{+}09]. Handling [JOZ03, LH01a, CCHW09, GRRX01, MPST06]. handoff [HLYL06]. hands [FIBRGCLN05]. hands-on [FIBRGCLN05]. HANet [JCC05]. hard [KCS01, PC04, SY02]. hardware [CGL^{+}04, EHKH04, Gla00i, KPT09, Oi08, SP08, TCSC04, XYS07]. hardware-translation [Oi08]. hardware/software [CGL^{+}04, XYS07]. Harmelen [LZ07]. harmful [SJ05]. hash [TMB02]. haves [HKvVvdV07]. HDWT [CCY^{+}09]. HDWT-based [CCY^{+}09]. heap [LSaC01]. helpful [SJ05]. Heterogeneous [GHKR04, KLC02, FBM09, JZL07, Kar01, PWLH06, SKKL07]. Heuristic [AAM00, PCC02, DSRS03, DSA^{+}04, MHW01, SMDM05]. heuristics [FLA^{+}01]. Hiding [RwJK01, CCY^{+}09, CL06a, CL06b, HCS09, LWL09, TW07, WCLL09, WLT^{+}09]. Hierarchical [WWC00, vdSJK^{+}07, BS09, JW06, KBH07, LLKL04, TYH04, WF07]. hierarchically [YR09]. hierarchies [BS09, HY03, Lee07, WL05]. hierarchy [LY01]. High [QL03, BGG09, CD07, CT00, CTL08, DB06, EBGR01, ELK06, FMSG08, HCS09, KC09, KT03, LO04, NsL00, PLCC09, Phi06, SMG08, SP08, SVMAM04, WWTH08, WLT^{+}09, HA03]. high-dimensional [LO04]. high-integrity [SP08]. high-level [CD07, FMSG08, Phi06, SMG08]. high-performance [CT00, SVMAM04]. high-quality [BGG09]. high-speed [ELK06, NsL00]. highly [RS06, WDS09]. HIPaG [JLYK09]. histogram [HLW08]. History [Boz00]. History-driven [Boz00]. hoc [BMES04, WF07, YZ05, YSK09, ZMN05]. holistic [CC09b, WSJK08]. home [vdSJK^{+}07]. hopping [BPM06]. hospital [ÖKT09]. host [CL06a]. hostile [HWM01]. Hotswapping [LC06a]. Hsu [BCW05]. HTML [RDD02]. Huang [ZC05]. human [HH08a, KK06, MV09]. human-centred [KK06]. human-related [HH08a]. Hwang [WL05]. Hybrid [DI01b, LS05b, BDGR01, CJ03, HC06, KH06, MR01, MR00b, TM06, YYWW07]. HyMIS [MK08]. hypercubes [KM04]. hypermedia [SL01]. hypermesh [LYX09]. Hyppocrates [BDDG04].

I&C [KSS03]. I/O [LP05]. I/O-intensive [LP05]. iconic [YC08a, YL09]. ICPS [LP07]. ID [CZL07, HH08b]. ID-based [CZL07, HH08b]. identification [DS04, HH06, SPSM03]. identifier [CDP05]. Identifying [BCB09, KL07, MKK09, She02, ZQZ^{+}06, SL08, TNAA01]. Identity [WC07, CC09a, KBD09, Sha09]. identity-based [KBD09, Sha09]. idle [SHS^{+}07]. IDRS [HL00b]. IEC [EG00, EB00, EJ01, JH01]. IEC61850 [PW03]. IEEE [Sai09, CMNA^{+}09, KvV06]. IEEE-FIPA [CMNA^{+}09]. IFPUG [CGMPAP08]. ignorance [Ber02]. Illustrating [ST01]. Image [CC04, CC02b, CHC01, HH06, KC09, KLC02, KCB05, KY08, LWS^{+}03, LK01, LLT^{+}09, LLCL08, Lin00, LT04, LWL09, PHN08, mSgFtL05, WLC07, YCYW07, YC08a, YL09, Zhu04d]. images [CL06a, CCP05, FWTC05, HCS09, LC02, TCC02, TW07, WCH03, WC02, Zhu04d]. Impact [VM07, BLOS06, CH09, CC09b, CBS00, DGP02, JMSS07, LJS05, RRD06, RSS00, Tan00, TNJH07, TMD07, YS02]. impacts [CG05, Ebe07]. Impartial [CJ05]. imperfect [Shy03]. imperfect-debugging [Shy03]. Implementation [HCC05, JE02b, ALT^{+}09, BBC^{+}08, CLX^{+}04, CH07a, CLG08, DGJ^{+}03, HYJL04, KRP02, KY09, KSH09, KLMC06, LWS^{+}03, LLK05, LKW^{+}09, LK02, NWZ05a, NWZ05b, NGM08, PLF05, SC00, SJK07, WZJ01, WSJK08, WOH08, YY04, YYL^{+}06, Zha09]. implementations [LL07]. implemented [LCH^{+}04]. Implementing [CMS04, CGP^{+}09, RH02]. implements [JFC08]. importance [Ber02, RGBM06]. important [MKK09]. imprecise [ANH07]. improper [LL07]. improve [Lea08, MK00, SSCL08, SKW06, TPRW04]. Improved [GMR08, KRHŽ05, LL06, LL07, LWL09]. Improvement [BH02, SCL07, Sha09, BH03, BHB^{+}05, LPJP09, LKB06, MT07, MM01a, MM00a, NWZ05a, NWZ05b, PK02a, PIGÖ08, QHS08, RH02, WHB01, XSS06]. Improvements [YCYW07, Hua05a, SJK07]. Improving [CHL04, FRR09, Mil00b, Pon05, SLS08, SK03, SK01, HLMB07, JMP07, MKNS06, RR09, VJB06]. imputation [SA06, SS07, SSCL08, VK08]. in-home [vdSJK^{+}07]. in-network [BLM^{+}08, JLYK09]. Incentive [FK01]. including [Aml00]. incomplete [XNP07]. inconsistencies [SK02]. Inconsistency [GJ07, NER01]. incorporating [Hua05a]. Increasing [BFLP09]. Incremental [CT09, VAS^{+}04, jHjW08, MM00a, MC04, PW09, PLP04]. independence [Mil02]. independent [DGJ^{+}03, DNAM05, SRDLCP09]. index [BH09, CL06b, CK00b, HLL01a, HL06b, LWL09, PSK05, SC08]. index-domain [CL06b]. indexes [HWML04]. indexing [LK01, RVM06, SC07, YC08a]. Indian [IS03a]. indicator [CCH09]. indicator-elimination [CCH09]. individual [RSS00]. Industrial [SD08, Wey01, ADZ^{+}09, AZW07, ASS07, Fra07, FS01, HF08, HDGZ06, HKN^{+}07, Kim07a, Kim07b, KGT02, LW02, DPS03, PW09, PKB09, THGL07, TL09a, VHF02]. industry [CCG^{+}07, CSNS05, EBB09, IS03a, JZ05, LdSBA^{+}08, Wes02]. industry/university [CSNS05]. inexact [Zhu03]. influence [BT05, Fai07, KLMZ08, TW08a, Vau07]. influences [Sai07]. influential [HFC^{+}01]. informal [Wyn01]. Information [KJ04, ML03, BWM06, CLCY04, CL06b, CK00a, Cho04b, Cho05, CC05, CLW05, CC06, CBK02, DHJ05, Fra04, HL02, HFRHS09, Kim07a, KJ01, KJLK07, LK01, LW02, LK02, LZL^{+}06, LWC06, MCC02, MMTL06, ÖKT09, ONZ09, PWLH06, PB00, PNL07, SKKL07, WCLL09, Wen03]. infrastructure [CL04a, LLV^{+}09]. infrastructures [DST^{+}04]. inheritance [HCN00, Lee07, Phi04, PUPT03, TB00]. inheritor [SLS08]. injection [GK08]. innovation [CDZ07, PKB09]. innovative [ANH07, CMS04]. input [LT08]. insight [MB06]. insights [CTY01]. inspection [DRW00, KS04, LD00]. inspections [BFV04, ELH00, PTRW04, TPRW04]. inspectors [Mil02]. instability [AL05]. instantiation [FdOdL04, OAdLC07]. Institutionalization [ACS07]. institutions [Gla99, Gla00c, Gla00d, GC01, GC02, GC03, GC05, TCG06, WTG^{+}08, WTG^{+}09]. instrumentation [DH09, Özm09, TLW07]. integrated [CLCY04, DI05, KLY03, LNC01, LK02, Lok06, PKR01]. Integrating [Ale05, BW01, LL09, SNBH08, MLB09]. integration [BG09, CCG01, CG03, DPSU06, GML05, GD04, LH06, RRW00, SD02, UZ09, WD07, Yeu00]. Integrity [WGC02, CT09, SP08, ZKL^{+}09]. intelligent [BFPAGS^{+}08, CHZY03, CG05, LKB06]. intensive [LP05, O’B08]. inter [CH05, Cho05, HCC05, LKL02, SL02, WQ06]. inter-application [Cho05]. inter-block [WQ06]. inter-class [LKL02]. inter-enterprise [SL02]. inter-stream [CH05]. Interaction [IWF07, BJK06, dL04]. interactions [CD05]. interactive [CFFT08, HYC02, HL00b, HKW00, JF04, SMHMA08]. interception [FIGCLN^{+}02]. interconnection [CGL^{+}04, CC01, CLC03, LYX09, RS00]. interface [AA07, CGL^{+}04, CH07a, MV09, THP^{+}06, HTH09]. interfaces [HYC02, SFJ04]. Interfacing [HSR01]. interference [BPM06]. interleaving [LCLL08]. internal [SeMC02]. International [CBVD07, LP07]. Internet [CJ09, DK01, HL00b, HLT09, KD05, LWS^{+}03, LCL04, MHC00, PTM08, SL02, SC09]. Internet-based [LWS^{+}03]. interoperability [CMNA^{+}09, DGP02]. interpolation [FWTC05]. interpretation [ML03]. Interprocedural [XNP07, MM06]. Interprocess [AACL02, IBP03]. interval [LLC^{+}09, NG08, YC08b]. interval-based [NG08, YC08b]. intervals [JTM04, TSSD09]. interviews [HJ00]. interworking [SKKL07]. Intranet [Tan00]. intraprocedural [ULN06]. introducing [DL06, HCWN05]. Introduction [BCDM06, Cha09, KB07, LK02, DGV08, TDL^{+}02, CDW07]. Intrusion [HZ07, CNLV07, HWM01, HWHM02, HWH^{+}03, LCLL07, SC09, WBW^{+}06, WZG09, WHC07]. intrusion-tolerant [CNLV07]. invalid [CJT04]. inventory [CDS02]. Investigating [BM00a, CO08, MMC05, RRD06]. investigation [ACS07, BVN07, CN00, DSRS03, DSA^{+}04, KLRW01, KWT^{+}00, KBBW05, LJS05, MKL^{+}00, RSS00, dBTdSS08]. IP [HHL06, Lin07]. IPv6 [HLYL06, LY09]. IRIS [Çam00b]. IS/software [Moy00]. ISO [EG00, EB00, EJ01, JH01, LCM^{+}04, YYL^{+}06]. ISO-certified [YYL^{+}06]. ISO/IEC [EG00, EB00, EJ01, JH01]. Issue [CUY09, CGA08, GH08, HLM^{+}09, LP07]. Issues [FWA09, MSB^{+}02, PW09]. IT-based [Rey07]. item [MCCC03, MM01b]. itemsets [CTL08, SPDT06]. iteration [WCH03]. iteration-free [WCH03]. iterative [BBS00, JPKP04, JE02b].

J [LHP^{+}10]. J2EE [ZP05]. JAIN [TDK^{+}07]. Java [CY04, CYH04, CDP05, DS04, GK08, HWM01, IS03b, JCYC04, MKM^{+}06, Oi08, RFZ08, SeMC02, TB00, TCSC04]. Job [Kar01]. jobs [ZK09]. join [JLYK09, LWHS05, LCCH02]. Joint [KCS01]. Journal [Gla00d, KPME05, LLLZ06a, Gla00a]. journey [BDA^{+}02]. JPEG [WC02]. JSD [Yeu00]. judgement [JTM04]. judgment [GJ07]. judgment-based [GJ07]. just [YY04]. just-in-time [YY04]. justification [ÖKT09]. JXTA [AMHJ09].

Kanji [Kuo00]. keep [RFZ08]. Kemerer [Gur01]. Kendra [MHC00]. kernel [CC03, CHY^{+}05, LC06a]. kernels [YSC^{+}06]. Key [RH02, ACS07, CLC08b, EHKH04, JW06, LLY07, LKH09, LKJL01, NLKW05, NJ07, SCH05, Tse07, WF07, WHHT08, YC09, YS04, ZSM05]. key-management [JW06]. keys [BCW05, CWH00, HY03, WH03]. kleptomania [Sta02]. Knowledge [KPS^{+}04, LC06b, Zhu04b, CSNS05, CDZ07, FM08, LL09, LLH08, MMTL06, RR09, RO09, SSA08, TL09a, Zhu06, ZG07, dBvV08, SZZ06, ZL06]. knowledge-based [TL09a]. known [HWW01, YTH04]. Korea [NSL^{+}07]. Korean [KJLK07].

label [LLL06]. labeling [MLC09, YC08b]. labels [MB06]. Lagrange [FWTC05]. landscape [GW01]. language [BFLP09, DDGR09, KNYS09, KRK00, MBM^{+}09, OAdLC07, ONR02, DNAM05, RS06, SMG08, SCdS^{+}06, TL09a, Wal05, YS02, ZMAV08, ZGH^{+}07]. language-independent [DNAM05]. languages [Ayr04, PCDG02, Spi01, Wen03]. large [AM04, BMES04, CJ03, CFN07, CCD^{+}04, Deu01, HBM05, HY03, JLC04, KY09, KGT02, KL07, KPG^{+}07, LK01, DPS03, MPST06, MHW01, PWLH06, TM06, TTWY04, TTC04, WL05, YSK09]. large-scale [BMES04, Deu01, JLC04, KL07, KPG^{+}07, PWLH06]. largeness [KEK04]. latency [LS05a]. latent [dBvV08]. lattice [JE02a]. laws [DZRH04, PLCC09]. layer [CH05, DGV^{+}07, WGY^{+}08]. layer-2 [DGV^{+}07]. layer-3 [DGV^{+}07]. layer-3/layer-2 [DGV^{+}07]. layered [IWF07, LBS^{+}07, OFWP07, PTM08]. layering [SMR09]. layers [SMR09]. Leadership [Vau07]. leaf [PHN08]. leakage [CLW05]. learn [Pfl00]. learning [Gon08, LNPAGD^{+}06, MKL^{+}00, ZHS01, BJK06]. least [SSO05]. left [WCLL09]. left-right [WCLL09]. legacy [ADZ^{+}09, BHN02, BT05, CCDD00, CFFT08, Lea08, LG05b, SCdO02, WBP^{+}03, YLC08]. legitimate [Lin01]. Lessons [BKB^{+}07]. letter [Gla00b, Gla00a]. Level [HA03, BLBvV04, CD07, CHY^{+}05, DL06, DHL06, FMSG08, HH06, HH00, JC02, LC05, PLCC09, Phi06, SMG08, TL09b, LBvVB02]. levels [JH01]. leveraging [LZXS09]. lexical [BHL00]. libraries [Ber03, SRGL08, SPZ06]. library [LLY07]. life [LD00, SZ06]. lifecycle [MGEB03]. lifetime [LS05b]. light [DDF^{+}05, PIGÖ08]. light-weight [DDF^{+}05]. Lightweight [CM05, DCAC09, HWH^{+}03]. like [CWH00]. likelihood [JZ07]. Lin [CC02b]. line [ACS07, AD07, AK08, Çam00b, HF08, KDS^{+}08, LG03, NRG08, TN05, ZR04, ZM06]. linear [HY01, PWLH06, dSSJV08, TMB02]. lines [EBB09, FL05, KG09, KPS08, LDL07, MR00a, SHW09]. lingual [RMC05]. Linguistic [Sta02]. link [WY04, WGW^{+}09]. linkability [WYL06]. Linking [BJ03]. links [Zhu04d]. Linux [FAB^{+}07, LC06a, YSC^{+}06]. List [Çam00a, Aba06, BG06, CHY^{+}05]. lists [CC05]. literature [BKB^{+}07, EFG^{+}08, ML08]. LMR [Rav03]. load [Boz00, CBZ00, MCCC03, RwJK01, TH02, WGW^{+}09, ZK09]. load/extract [TH02]. Local [Oi08, FLA^{+}01]. Locality [ZG00, YR09]. localization [AZGvG09]. locally [TC06]. locating [WBP^{+}03]. Location [HLYL06, LLKL04, ZS05a, ESW06, LU06, LPR04, PSK05, PSH06]. Location-aware [HLYL06, PSH06]. location-based [LPR04, PSK05]. location-dependent [LU06]. locking [CM05, Jun00]. locking-based [Jun00]. locks [HPT07]. log [CPL^{+}04]. logging [CPL^{+}04]. Logic [KK07a, EL07, IS03b, dSSJV08, TL09a, ZC06]. logic-based [TL09a]. logical [YL06]. logics [BNR09]. login [CJT01]. logistic [SA06]. logs [LZXS09]. London [LZ07]. long [Kel09]. long-term [Kel09]. loop [HH00, SP08, WWC00]. loop-level [HH00]. loops [CCW02b, KMSMD08, LC05, LC07, PC01]. loss [ED06, HHL06]. lossless [LWL09]. lossy [BGG09]. LRU [ED06].

M [CCL01, LL07]. MAC [NsL00]. MacGuffin [LGW09]. machine [DSC^{+}08, Gon08, GWvD08, MKL^{+}00, MV09, Oi08]. machinery [PNL07]. machines [DCH02, EE08, HM09, PH06, WWC00]. mail [LCLL08]. Maintainability [YSC^{+}06, CH09, DSRS03, DSA^{+}04, HCN00, LJS05, ZL07]. Maintaining [AV02, BVN07, Har04, KLT07]. Maintenance [DGV08, FS05, Gla00g, Gla00j, HLW08, KPME02, KPME05, KP07, DPS03, MR00b, NGC02, OÓ08, PUPT03, PNL07, RSS00]. maintenance-first [Gla00g]. major [Bif03, LWHS05]. make [Mül07]. Making [GSB^{+}07, NER01, BFV04, CC03, GLJ00, Gho01, KLMZ08]. malfunctions [ZHS01]. Management [MM01a, AZW07, Ban08, dOBWT04, BH09, CDI07, Cha06, CD00, CSaLG02, CBG09, CM05, CDZ07, DST^{+}04, DTV09, DLB04, Ebe07, FY04, FN00, GTA09, HSM^{+}07, HRN^{+}01, HMC01, HTH09, JW06, JJP02, JKWL09, KMSMD08, KRC00, Kim07a, Kim07b, KJLK07, LMvV09, LBS^{+}07, LP00, LNC01, LLKL04, LRS^{+}07, LKB06, LSaC04, LWC06, MS03, ML03, McB08, McD02, MMTL06, MDMC06, PSK05, PKR01, Pfl00, PK01b, Sai07, SM03, TLW07, dBTdSS08, WF07, WDCL08, WFWL09, WOH08]. manager [LP05, SHW09]. managers [ABG02, Moy00, PV06]. Managing [CC06, EBB09, PNL07, Rey07, ZKL^{+}09, CWC04, CC09b]. Manipulation [DHKV06]. mapping [BM00b, MKNS06]. mappings [Phi05]. maps [PSMB01]. market [HRN^{+}01, LLL06, ZG07]. market-driven [HRN^{+}01]. Markov [BHXN05, PP04]. MARKS [ALT^{+}09]. Massachusetts [LZ07]. massive [CZH^{+}08]. match [CL06b, KK07b, LC00]. matching [HDLK00, LCCH02, LZG07, MMTL06, Mus03, Zhu03]. matching-based [LCCH02]. MATLAB [HSR01, ZC08]. MATLAB/Simulink [ZC08]. matrix [HSR01, HTH09]. matters [SvV08]. maturity [NWZ05b, RH02, SJK07]. maturity-based [RH02]. maximize [PC01]. maximizing [TYH04, YYWW07]. MB [VA08]. MB-UID [VA08]. MBASE [VM07]. MC [YL06]. MC/DC [YL06]. MDA [ZBLG07]. MDABench [ZBLG07]. me [CPT05]. mean [JIS03, KTK01]. measure [RO09]. Measurement [CGA08, LS05a, AP09, BW01, CLB05, HALS08, HRŽ06, LORB03, NSL^{+}07, SMR09, VK08, Xia00]. Measurement-based [LS05a, BW01]. measurements [Gok09]. measures [BFLP09, BWDP00, CKM06, EB00, Fer00, HH08a, PSMB01, WK00]. Measuring [BTV06, LC09, SvV08, DHJ05, LH08, LCM^{+}04]. mechanism [CL08, CSGL05, GAWW07, HC01a, PC01, SKKL07, WGY^{+}08]. mechanisms [AK08, AJCM08, CJKC09, GRRX01, IBP03, Lin00, McB08]. mechatronic [DZRH04, SFSE05]. media [CDC09, SM03]. mediator [CDS02]. mediators [BJK06]. medical [BGG09, LWS^{+}03]. medoid [BRTT08]. melody [RH06]. memory [BH09, CGHL07, CN04, CD00, CSaLG02, CBG09, CHL04, HC06, JFC08, Kar00, LUS^{+}00, LSaC04, SeMC02, THP^{+}06, USLC01]. mental [KV05]. mesh [Aba06, Aba08, BMOKAM09, CCHT09, SK03]. mesh-connected [Aba06, Aba08, SK03]. Message [MW08, HYC04]. meta [Mil00a, SMDM05, ZGH^{+}07]. meta-analytical [Mil00a]. meta-heuristic [SMDM05]. meta-tool [ZGH^{+}07]. Metadata [DI05, CDS07, WFWL09]. Metadata-driven [DI05]. metamodel [OHS01]. metamodels [HFRHS09, TT09]. metasearch [LDN04]. method [CCY^{+}09, CCH09, CD07, CBK02, FN00, Har04, HFRHS09, JC02, KYP^{+}03, KPSK09, KMKY07, KvV06, KSS03, KRHŽ05, LTK^{+}06, LC05, LC08, Lop03, MLC09, MM06, PJ09, SNBH08, SC00, SOC^{+}03, SK04, SS07, Tho06, WWTH08, WJT09, YKC^{+}05, Zhu04c]. methodologies [GR05, GPHS07, KLMC06, MMTL06]. Methodology [Cha06, FL09, Kim07b, BRMA^{+}09, CCC06, GPHS08, LS04, LK02, MYZC06, NGM08, ONZ09, SL01]. methods [BDMK03, FIBRGCLN05, GPML06, HALS08, HRŽ06, HJ00, HH06, QHS08, SUSO04, SPZ06, WBP^{+}03, Yeu00]. metric [AL05, CC08b, Hus01, NJ07]. metrics [Aml00, CGP^{+}05, CKL09, DLW08, EMM01, FN00, GS07, Gur01, Har04, JKD02, JPK00, KJ04, MMC05, SvV08, SL08, SC01, SPSM03, TQ05, WG05, ZG00]. MHS [WFWL09]. micro [HLL01b]. micro-payment [HLL01b]. microaggregation [CLH07]. microkernel [KLGH07]. microkernel-based [KLGH07]. middleware [ALT^{+}09, DGV^{+}07, DLB04, KBM05, LGL08, SDG^{+}07, TDK^{+}07, TMD07, VP07, WCLK07, YZ05]. middleware-based [LGL08]. midwest [Wen03]. Migratable [MKMS05]. migrating [CCDD00, CFFT08, HL01]. migration [CLC08a, GWvD08, UZ09]. millennium [MG04]. miner [WHC07]. minimize [LUS^{+}00]. Minimizing [KTK01, PK01a]. Minimum [LKL04]. Mining [LcLsW06, LLT^{+}09, LLC^{+}09, VMB^{+}08, YHHR03, CTL08, DZW^{+}09, KLNS07, LL09, LPR04, SHS^{+}07, TTWY04, TL07, TL09b]. MIP [GP05]. mirrored [VM00]. missing [DZ00, SA06, VK08]. mission [DGV^{+}07, LJS05]. mission-critical [DGV^{+}07, LJS05]. mistakes [Mül07]. mitigating [SKZ^{+}04]. mixed [CCSC01, LKL02]. MLC [LKW^{+}09]. MMDB [DK08]. mobile [BHAM09, BMES04, CLC08a, CJ03, CMNA^{+}09, DPMD07, FIGCLN^{+}02, GMB^{+}09, GTA09, HLYL06, HLT09, HL06b, LCY00, LLK04, LKW^{+}09, LNY06, LLKL04, LRS^{+}07, LK04, Lin07, MK08, NLKW05, PSH06, PJ09, PS09, RT07, TM06, TKA^{+}02, VA08, WGC02, WBW^{+}06, WF07, WHN^{+}01, YC09, YZ05, YSK09, ZK04a, LY09]. mobile-commerce [YC09]. mobility [CMS04, HLT09]. mode [CGW08]. Model [AA07, CDI07, Gok09, GWvD08, Phi05, PLP04, BHXN05, BKR09, BHB^{+}05, CCC05, CC09a, CGL^{+}04, CELS07, CLB05, CMC04, Cho04a, Cho04b, Cho05, CC05, CC06, DLW08, DGL^{+}08, DGJ^{+}03, EJ01, FBM09, FWA09, GMR08, GMS07, HAHH06, HKN^{+}07, HK09, JPKP04, JJ06, JHSB09, KRK00, KBH07, KLGH07, LZ06, LT08, LH01b, MMCB00, MR01, MA09, NR04, NWZ05b, NGM08, OOD09, PLCC09, PG05, PK02b, Phi06, PW03, RRT01, Shy03, SM08, Tan04, Tan00, TJH07, TN05, TCSC04, TMB02, UZ09, WW09, YYL^{+}06, ZE03, ZLCY06, ZyCkP01, Zhu03, Zhu04a, Zhu04c, AJCM08, FdOdL04, MYZC06, Zha08]. Model-based [AA07, CDI07, Gok09, PG05]. Model-driven [GWvD08, BKR09, GMS07, AJCM08]. modeled [GJ08]. Modeling [CS01, CUY09, EL07, HA03, HLC^{+}09, HYJL04, JX07, JLC04, PLM07, BCV06, CCW^{+}01, CW02, DB06, FCSM09, GH04, Iso01, JOZ03, JZ05, Kar04b, KPS08, KMKY07, KSS03, KEK04, KDEK04, LH04, LSH09, LDL07, LHC^{+}05, Nae01, OD05, PS05, Phi04, RK00, SZ06, SRDLCP09, TBD^{+}08, WPC06, WKH09, Xia00, YWT07, ZH05]. Modelling [CBG09, ELK06, GPHS07, RW01, CFN07, Cow05, DI01b, KLRW01, PH07, PSG^{+}09, RRW00, SG01, Wal05]. models [ABG02, ACGS^{+}08, BG09, Bif03, BLOS06, CGP^{+}05, DA07, GBL08, GMS07, HBVG08, HFC^{+}01, Hua05a, Hua05b, IWF07, JHSB09, JZ07, KSH05, KH06, KV05, Lin01, LH08, LHP^{+}09, LHP^{+}10, LMYMGT08, MGEB03, MDFG08, MA08, NG08, OFWP07, PS00, PP04, Sai07, SFJ04, SKW06, SH07, SPSM03, THP^{+}06, THGL07, VMB^{+}08, Wal05, WPP^{+}09, Zel09, ZKL^{+}09]. modern [BM00a]. Modifiability [LBvVB02, BLBvV04]. Modification [AHGSS05, HCS09]. modified [CJT01, THGL07]. Modify [KFS^{+}02]. Modify-on-Access [KFS^{+}02]. modular [DXPY03, HL06a, dlRT06]. Modularized [HL00a]. Module [RS00, LHC^{+}05, MR00b, PKR01]. modules [BT05, EE08, KT03, LC06a, TNAA01, XNP07]. modulus [WWTH08]. Monitoring [LCF^{+}06, OAZ08, ZS05b]. monitors [HL00a]. morphology [Mus03]. morphology-driven [Mus03]. motion [ZEY04]. motion-based [ZEY04]. Motivators [BH02, BH03]. movement [TL09b]. moving [LPR04, LSZ^{+}07, RVM06]. MPEG [DK08, DK08]. MPEG-7 [DK08, DK08]. MPI [DCH02]. MPSoC [JHSB09]. MRP [LPJP09, MM00b]. MRRL [ED06]. MUDABlue [KGMI06]. Multi [LKL05, SFJ04, WGY^{+}08, BBG^{+}04, BWM06, CCW02a, CET^{+}08, DCH02, FMP09, FWTC05, GAWW07, HLW08, JZL07, Jun00, LBS^{+}07, LcLsW06, LSH09, NXS00, PLCC09, RMC05, Sha09, TL09b, WC07, WDCL08]. multi-agent [BWM06, CET^{+}08, JZL07, LSH09, PLCC09]. multi-class [GAWW07]. multi-device [BBG^{+}04]. Multi-devices [SFJ04]. multi-dimensional [CCW02a, HLW08, LcLsW06]. Multi-disk [LKL05]. multi-granularity [Jun00]. Multi-layer [WGY^{+}08]. multi-layered [LBS^{+}07]. multi-level [TL09b]. multi-lingual [RMC05]. multi-organizational [FMP09]. multi-processor [DCH02]. multi-secret [FWTC05]. multi-server [NXS00]. multi-signature [Sha09, WC07]. multi-tier [WDCL08]. multicast [LT07, TTC04]. multicasting [WGW^{+}09]. multicomputers [Aba06, Aba08, BMOKAM09, RwJK01]. multidimensional [HWML04]. multigranularity [CM05]. Multimedia [DK08, HLT09, BRMA^{+}09, CCCT06, CCSC01, CH05, DLB04, GL05, HKW00, HKY01, HL02, HLYL06, LTK^{+}06, LG05b, MV05, MV06, PK02b, TVA04, TTC04, YY04]. multinomial [SA06]. multiobjective [Yoo09]. multiple [BFV04, CCF^{+}04, CK02a, Deu01, HKY01, KMSMD08, KC09, Lee07, LSV^{+}06, Loo05, OW04, PC02, DM07, SFJ04, TB00, TCC02]. multiple-stream [HKY01]. multipoint [CBK02]. multiprocessing [Kar00]. multiprocessor [HTK00, LC05, LYX09, MMM00, PWCC01]. multiprocessors [PK01a, SA05]. multiscape [LY01]. multisignature [CWH00, WHG01]. multiuser [LNPAGD^{+}06]. multivariate [ZL07]. MUMCUT [YLC06, YL06]. MUS-T [VAS^{+}04]. MUSEMBLE [RjHHK08]. music [RjHHK08]. must [HKvVvdV07]. Mutation [HM09, SW09, HLM^{+}09]. mutual [KTK01].

nAIT [DH09]. NAND [LKW^{+}09]. NASA [DB06]. NAT [CJ09]. national [NLSK04]. native [HL09, SVMAM04]. Natural [BFLP09]. nature [FS05]. nearest [LSZ^{+}07]. Need [CBVD07]. negotiation [LR04]. negotiators [HCWN05]. nesC [DH09]. nested [LC05, MMCB00, PC01, TMB02]. net [CCC06, JR09, KDEK04, LKJL01, RR09]. NetBSD [YSC^{+}06]. nets [dSSJV08, BM07, CR06, HA03, Phi06]. network [AN01, AADAD02, Bai05, BHXN05, BDMK03, BLM^{+}08, CCW^{+}01, CLL05, CC01, CJ03, CE08, DGV^{+}07, DST^{+}04, FS06, GTA09, HLYL06, HCC05, JLYK09, JCC05, LCLL07, LLV^{+}09, NsL00, OS09, PNL07, SSM^{+}09, SH07, SC09, TTC04, WHC07, ZHS01, ZK04b, ZCT^{+}09]. networking [DJW08, HC04a, KPT09, LK09, Zhu06]. networks [AM04, BPM06, BMES04, Boz00, BLM^{+}08, CLC03, CC08b, CFN07, DGV^{+}07, HC01b, HHL06, HLT09, JLYK09, Kar04b, KPSK09, KMOS09, KV05, KRCK08, LT09, Lin07, LY09, MBM^{+}09, MK08, MHW01, Rav03, TM06, TNJH07, TQ05, TPN^{+}09, dBTdSS08, TL07, TL09b, WF07, WGY^{+}08, WGW^{+}09, WHYT06, YZ05, YSK09]. Neural [SH07, CCW^{+}01, CE08, KRCK08, TQ05, dBTdSS08]. Neural-network-based [SH07]. neuro [LCF^{+}06]. neuro-adaptive [LCF^{+}06]. news [CT08, LHLY05]. Newsmonger [MK00]. NN [SSCL08]. no [ED06]. no-state-loss [ED06]. node [MK08, NJ07]. nodes [BMES04]. noisy [VK08]. Non [HY01, KW00, Gla00j, KMOS09, LC07, MLB09, PC01, DM07, XZAR06, ZWX^{+}08]. Non-blocking [KW00]. non-crises [Gla00j]. non-determined [ZWX^{+}08]. non-functional [MLB09, XZAR06]. Non-linear [HY01]. non-orthogonal [LC07]. non-perfect [DM07]. non-real-time [CCSC01, KMOS09]. non-uniform [PC01]. nonclairvoyant [ZK09]. noncontiguous [Aba08]. nonrepudiable [HWW01, YTH04]. nonuniform [SC08]. Norm [BT05]. Notable [Spi01]. note [DD01]. Novel [WLC07, KBD09, RjHHK08, mSgFtL05, SSM^{+}09, TTWY04, TTC04, TW07]. NPP [KSS03, YS02]. NT [AS01, LCH^{+}04]. NT-SwiFT [LCH^{+}04]. nuclear [YKC^{+}05]. NUMA [WWC00]. number [MM01b].

O-intensive [LP05]. obfuscation [CY04, CDP05, KJ04]. Object [ACDF01, CGL^{+}04, Iso01, BWDP00, CL04a, Cho04a, DSRS03, DSA^{+}04, DHL06, EKV05, EMM01, FN00, FCL^{+}00, FS05, GRRX01, HCN00, Jun00, KMSMD08, KLT07, LSZ^{+}07, LS07, LJS05, MLB09, PSMB01, Phi04, Phi05, Pon03, Pon05, QK08, RS00, SNBH08, ST01, She02, SK02, SC01, SPSM03, TA02, TQ05, TK00, TMD07, TH02, TL07, TL09b, WT01, XNP07, ZEY04, ZL07, GHKR04]. Object-based [CGL^{+}04]. Object-oriented [ACDF01, Iso01, BWDP00, CL04a, Cho04a, DSRS03, DSA^{+}04, DHL06, EMM01, FN00, FCL^{+}00, FS05, GRRX01, HCN00, Jun00, KLT07, LS07, LJS05, MLB09, PSMB01, Phi04, RS00, SNBH08, ST01, She02, SK02, SC01, SPSM03, TA02, TQ05, XNP07, ZL07]. object-relational [Phi05, TH02]. Object-Z [GHKR04]. objects [HL02, IS03b, Pon06, RVM06, SM09]. observations [IS03a]. observe [ZHS01]. observers [JL04]. occluded [ZERO00]. OCL [CT09]. OCL2Trigger [AJCM08]. ODCHP [PC01]. ODMG [LLK05]. ODMG-compliant [LLK05]. Odyssey [BWM06]. Odyssey-Search [BWM06]. off [ELK06, PJ09]. off-chip [ELK06]. offs [Bat08]. OLAP [PKL03]. OML [OHS01, OD05, ZPEL01]. On-line [TN05, Çam00b, ZM06]. One [LYX09, JZ07, KMS04, OR00]. one-part [JZ07]. One-step [LYX09]. ones [Gla00k]. online [LCF^{+}06, NKJT09, PTK00, TH05]. ontologies [FdSdP08]. ontology [OHS01]. OO [BDGR01, CBKK08]. OODBMS [HLMB07, LLK05]. Open [CF07, Fug03, AW07, CLL05, DST^{+}04, KL07, MMCB00, MSB^{+}02, PLCC09, SLS08, SM08, SSA08, TDK^{+}07, YSC^{+}06, ZE03, KGMI06]. open-source [CLL05, KL07, YSC^{+}06]. OpenBSD [YSC^{+}06]. OpenVPN [LLV^{+}09]. OPERA [CLL05]. operand [LSC04]. operating [IBP03, PLM07, WW00, YSC^{+}06]. operation [ZS01, ZH05]. Operational [LM03, Bai05]. operations [TCSC04]. operator [BLM^{+}08]. Opportunistic [HLMB07]. opportunities [SFJ04]. optical [LYX09, WGY^{+}08]. Optimal [CY00, DXPY03, HL06a, CSGL05, Hua05a, JE02b, PK01a, WDS09]. optimisation [PG05]. optimistic [JFG07]. Optimization [BRMA^{+}09, CDC09, CHL04, CK02a, PCC02, PK02c, ÜDUG04, YYWW07, ZCT^{+}09]. optimizations [VP07]. optimize [MS03]. Optimizing [HYC02, HLL01a, CCSC07]. options [ÖKT09, WOH08]. oracles [CCHT09]. ORB [WCLK07]. order [CCH09, LWHS05]. ordered [HY03, WL05]. ordering [HYC04, KLMC06]. organization [JH01, LZ06, DPS03]. organizational [ACS07, ABG02, BCV06, FMP09, RSS00, TW08a]. organizations [CLW05, SNJ^{+}07, YYL^{+}06]. organizing [BNR09, PSMB01]. orientation [ADZ^{+}09]. Oriented [CFFT08, TDT08, ACDF01, BWDP00, CLX^{+}04, CCCT06, CL06a, CL04a, Cho04a, DSRS03, DSA^{+}04, DST^{+}04, DHL06, DTV09, EMM01, FN00, FCL^{+}00, FS05, GRRX01, HCN00, Iso01, Jun00, KLT07, KSH05, KKK08, LJB05, LC00, LCL04, LL06, LVMM07, LS07, LJS05, MLB09, Mur08, NGC02, OKS08, PSMB01, Phi04, RS00, SNBH08, ST01, She02, SK02, SC01, SPSM03, SL01, TA02, TQ05, TMD07, UZ09, VP07, WHHT08, XNP07, ZL07, Zhu00, dOZR^{+}04, KCS08]. original [CL06a]. orthogonal [LC07]. orthogonality [RFZ08]. OSA [TDK^{+}07]. OSA/Parlay [TDK^{+}07]. other [Gla00j, RGBM06, YL06]. Outgoing [Car08]. Outsourcing [Gla00e, BVN07]. Over-confidence [JTM04]. Overcoming [CDP05]. overhead [MA09]. overheads [RwJK01]. overlay [SSM^{+}09]. overloaded [MA09]. overriding [Rad04]. ownership [CL08, HH06, Lin01].

P [BRC09]. P2P [LZ06]. PAC [CHL05]. Pacific [LW02]. Packaging [VJB06]. packet [ELK06, HHL06, HCC05]. page [CN04, LCCH02, LSaC01]. page-coherent [CN04]. pages [Pon06]. paging [HH05, WW00]. pair [CCG^{+}07, Mül05]. pairing [BBC^{+}08]. pairings [Sha07]. pairs [Mül07]. PAL [KK07a]. PAL-based [KK07a]. Palladio [BKR09]. panacea [Mil04]. panel [BHB^{+}05]. paper [Gla00a]. papers [Ano02a, Ano02b, DHKV06, GH08, Sai09, VE03]. para [BG09]. para-consistent [BG09]. paradigm [DGJ^{+}03, EL07, LJS05]. Parallel [MEH05, CLX^{+}04, HBVG08, HSR01, JE02b, MCC02, OFWP07, SK03, SPDT06, WT01]. parallelism [HH00, PC01]. parallelize [CCW02b]. parallelizing [LC05]. Parameter [PQLN04]. parameterized [MHW01]. parametric [NGM08]. parametrized [JFC08]. Parlay [TDK^{+}07]. parsing [DDGR09]. PART [HLM^{+}09, JZ07]. partial [CC02a, SPDT06]. partially [CZL07, HH08b, HC04b, HY03, WL05, ZC05]. particle [YYWW07]. partitioning [JC02, KPT09, LO04, LZN04, MCC02, SK04]. party [CLC08b, SCH05, YC09]. passive [KPG^{+}07]. Password [YS04, BDDG04]. Password-based [YS04]. past [MKNS06, RVM06]. past-time [RVM06]. PAT [CCHT09]. path [CK02a, HL09, LZG07, Mur08, PC02, PWLH06]. path-oriented [Mur08]. Pathfinder [KV05]. pattern [CCHT09, DDGR09, DACY07, HZCD05, Hus01, KPS^{+}04, KLNS07, LPR04, WHC07, XZAR06, ZMAV08]. pattern-based [DACY07]. Patterns [HGK^{+}06, SO03, AA07, ACDF01, BJ03, BNR09, DJW08, HCC08, KCS08, KP07, LLT^{+}09, OKS08, PB04, SMHMA08, SL03, SC07, Spi01, TL09b]. payload [KC09]. payment [HLL01b]. PCs [SHS^{+}07]. peer [BGG^{+}06, Lok06, Loo05, MK08, Mül05, SM06a, SSM^{+}09, ZK04b]. peer-to-peer [BGG^{+}06, Lok06, Loo05, MK08, SM06a, ZK04b]. Perception [JKD02, CJKC09, JKWL09, KWT^{+}00, KLMZ08]. Perceptions [RSM00, LL04, SHW09]. perfect [DM07]. Performance [BM07, CLGL05, CUY09, HLWC04, Hua05b, IBP03, Kar04b, NsL00, PLF05, SM06a, TMB02, WPP^{+}09, ABW07, BJK06, BKR09, BBS00, CDI07, CT00, DI05, Del08, ED04, GLJ00, Gok09, GMS07, GAWW07, HH07, HLMB07, IWF07, JKWL09, JRB^{+}06, KDEK04, LTK^{+}06, LJB05, LS05a, LSaC01, MK06, MK00, NLSK04, NSL^{+}07, OS09, OFWP07, Pon05, PSG^{+}09, QL03, SO03, SK03, ST07, SVMAM04, SK01, SSCL08, SJK07, SDG^{+}07, TDK^{+}07, TMD07, WW00]. performance-driven [PSG^{+}09]. performance/reliability [GMS07]. performances [CCG^{+}07]. Period [BRC09, PK01a]. periodic [PC04]. persistent [IJC03]. personal [LLH08]. personalization [BFPAGS^{+}08, KUK07]. perspective [CO08, Haz02, mJKME01, Kim07a, Kuo00, LC09, Vau07, WOH08]. perspectives [LW02, LSV^{+}06]. Pervasive [LP07, ALT^{+}09, CELS07, CJ09, JZL07, MPG^{+}08, MGI07]. Petri [CR06, HA03, LKJL01, CCC06, KDEK04, Phi06, dSSJV08]. petri-net-based [CCC06]. Petri-Nets [Phi06]. phase [CK02a, HL06a, MDMC06, Mil00b]. phrases [CP09]. physical [ZG00]. picture [LC00, LY01]. piece [DZ00]. pilot [HBVG08]. Pinned [HH05]. pipeline [LCF08]. pixel [WWTH08]. pixel-value [WWTH08]. placement [HL02, MCCC03, TVA04, VM00]. planned [NR04]. planning [AN01, KKK08, MØHB08, NRG08, PK02b, PIGÖ08, ZHS01]. plant [YKC^{+}05]. Plate [Car02]. Platform [SRDLCP09, GTA09, JHSB09, KPT09, Zhu04a]. Platform-independent [SRDLCP09]. platforms [CCDD00, GD04]. playback [NXS00]. playing [BPM06]. please [CPT05]. point [Hua05b, HCC05, OR00, Shy03, ZLCY06, AHGSS05]. pointer [MC04]. pointers [EKV05]. points [SvV08, SHW09]. poker [MØHB08]. policies [Aba08, ZK09]. policy [Çam00b, HC04a, Hua05a, KLRW01, MBM^{+}09, TYH04, YY04]. Polyhedral [THP^{+}06]. polymorph [FBM09]. pool [GPHS08]. portable [RW00]. Portal [CC03]. portfolio [CdOBT07]. portfolios [KGT02]. post [LS07, SL08]. post-release [LS07, SL08]. Posting [MCC02]. potential [HMC01]. Pounamu [ZGH^{+}07]. power [ED04]. power/performance [ED04]. powerful [Ayr04]. Practical [CP07, LT09, SPSM03, AZGvG09, CCF^{+}04, DB06, HH00, HZCD05, PSG^{+}09]. practice [CJ05, CO08, Fer00, Gla00m, QHS08]. practices [DD01, HDGZ06, MKK09, Sai02, VHF02]. practitioner [BH02, BH03, Haz02, KLMZ08, PIGÖ08]. practitioners [PCV^{+}08, PVSG05]. precise [CCW02a]. predicate [Sta03]. predict [LRvV03, MR00b]. predictable [HMSW03]. Predicting [EE08, ZL07, EBGR01, TL09b]. prediction [Bai05, BHXN05, BKR09, BFLP09, CLGL05, EMM01, Gon08, Gru07, GJ08, HCS09, JTM04, KR08, LS05a, MKL^{+}00, MA08, RSP03, Sch03, SRDLCP09, TQ05, TN05, THGL07, dBTdSS08, VMB^{+}08, WHB01, WLC08, WLT^{+}09, ZP06]. predictions [JFG07, MS03]. Predictive [LMYMGT08, HWHM02, LH08]. predictors [Gla00k]. preferences [LS05b]. prefetching [Pon03, Pon06]. prescription [MM01a]. present [MKNS06]. presentations [CH05, HKY01, YY04, YWT07]. Press [LZ07]. Preventing [CLW05]. prevention [CC07, CCKM09, LCLL07]. Primer [LZ07, AV04, AV08]. principles [BM00a, BDA^{+}02]. Priorities [Let00, BS09]. prioritised [HLMB07]. prioritization [JG08, ZCT^{+}09]. Priority [LLL00, LSV^{+}06, BRC09, BCF^{+}05, HC01b, LHSK06]. priority-based [HC01b]. privacy [Cho04a, CHL^{+}08, ZSM05]. privacy-enhanced [ZSM05]. private [KUK07]. Proactive [HLW08, LR04, BDDG04]. probabilistic [HM09]. probability [LS07]. problem [BCV06, CH09, CJT04, Dar02, DSSL09, HCDJ08, KEK04, PW03, TNAA01, Wij03, Zhu00]. problem-oriented [Zhu00]. problem-prone [TNAA01]. problem-solving [DSSL09]. problems [GH04, JE02b, KRHŽ05, LL07, vGB02]. procedure [AP09, SD02]. procedures [Mil00a]. Process [BH02, CGA08, HF08, LCF08, RW01, AK08, AL05, ACDG02, BKZ^{+}06, BH03, BM05, BHB^{+}05, BKB^{+}07, BM00b, CGP^{+}05, CCC05, CS01, CHL05, CO08, CGSGR06, DCAC09, DA07, DHJ05, DI01a, DI01b, DZW^{+}09, EB00, FCSM09, GW01, HL01, HAHH06, HHW01, HFC^{+}01, HFRHS09, JPKP04, JMP07, JH01, KRHŽ05, LPJP09, LSV^{+}06, MR01, MM01a, NWZ05a, NWZ05b, PIGÖ08, PH07, QK08, RK00, RH02, REF^{+}07, SL08, SJK07, WW09, WHB01, Wyn01, XSS06, ZyCkP01]. process-related [CGSGR06]. Processes [MSB^{+}02, BNvdH05, CC07, CBS00, DI01a, GR05, HH08a, HRN^{+}01, KLRW01, LH06, DPS03, PS00, RH03]. processing [BLM^{+}08, CK02b, HL09, KRP02, KW00, LWHS05, LCCH02, MLC09, PJ09, SK01, YC08b, ZM06]. processor [Aba08, Çam00a, CHL04, DCH02, HSR01, SK03]. processor-in-memory [CHL04]. processors [ELK06, LCLL07, TCSC04]. Producing [SHW02]. Product [CGA08, DSB05, ACS07, AD07, AK08, BW01, CHL05, Del08, Ebe07, EBB09, FL05, GPHS08, HF08, KDS^{+}08, KG09, KPS08, LS05b, LDL07, Lut00, LG03, NRG08, ZR04, dBvV08]. product-line [KDS^{+}08]. production [HK09, ZKL^{+}09]. productivity [FS01]. products [KL07]. profile [Bai05, CK00a, NLSK04, TR00]. profiles [GJ08]. Program [LXZS06, Alz08, CH07a, DDF^{+}05, DS04, HBD03, Kri06, LNY06, OR00, PCDG02, RSS00, WGH00, WQ06, ZG00, ZC06, ZCT^{+}09]. programmer [Mül07]. programmers [Mül07]. programming [Ayr04, CLX^{+}04, DBO05, FMSG08, HBM05, HCDJ08, HBVG08, Lok06, Mül05, TW08a, TKA^{+}02, Wen03, KCS08, SJ05]. programs [CCDD00, CCHT09, CLSa01, CDP05, EKV05, FS05, JPK00, LVMM07, LMYMGT08, MKM^{+}06, SeMC02]. progress [DHJ05]. progresses [LW02]. project [BM05, dOBWT04, BJK06, BDGR01, CdOBT07, DB06, FY04, JK00, JKWL09, KWT^{+}00, LPS02, LXG09, LKB06, DPS03, MS03, ML03, MR01, McB08, McD02, Moy00, MH04, NLSK04, NSL^{+}07, PCV^{+}08, PKR01, PVSG05, PV06, RRT01, Sai07, SS07, SSCL08, Sta09, SJK07, dBTdSS08]. projects [Ban08, BCB09, CC08c, FN00, FHT07, HH07, MØHB08, RR09, SSA08, SM07]. Prolog [Lok06]. prone [EE08, KL07, SL08, SPSM03, TNAA01]. proneness [Gon08, MR00b]. propagation [CE08]. propensity [KWT^{+}00]. properties [BGH03, CKM06, IS03b, Wil03]. property [CCH09]. Proportional [CTY01]. proposal [DF00]. protect [Cho04a]. Protecting [GMB^{+}09]. protection [CL08, KUK07, LC02, WGC02, WGW^{+}09, YKC^{+}05]. Proteus [USLC01]. protocol [CN04, CCL01, CLC08b, CJ03, FS06, FIBRGCLN05, HCC05, KYP^{+}03, KMS04, Lai02, LTK^{+}06, LT07, LJB05, LKH09, OS09, QL03, TM06, Tse07, WHHT08, YC09, ZS05a]. protocols [CLL05, CFN07, MK00, Shu03, SCH05, YS04]. Prototyping [BJK06, CLCY04, DZRH04, HCS04, WKL04]. provider [DST^{+}04, JKD02]. Providing [Cho04b, Lin07, KBH07, TYH04]. provision [TDK^{+}07]. provisioning [KUK07, RT07]. Proxy [RMC05, CE08, DK01, HWW01, HC04a, HLYL06, LT09, LCLL08, SCL07, Sha09, SHT05, WC07, WYL06, WL09, YTH04]. proxy-based [DK01]. pruning [PC02]. public [BCW05, CWH00, CHL^{+}08, EHKH04, LC02, WH03, ZSM05, ZMN05]. public-key [ZSM05]. publications [SM06b]. publish [YSK06, YSK09]. publish/subscribe [YSK06, YSK09]. publisher [SO03]. publisher/subscriber [SO03]. Publishing [LC06b, CCC05]. purchase [LS05b]. purposes [LH01a]. push [MvD08]. push-based [MvD08].

QoS [CDEV08, DGV^{+}07, DLB04, KD05, LR04, MYZC06, MGI07, MPG^{+}08, NKJT09, SDG^{+}07]. QoS-aware [CDEV08]. QoS-enabled [SDG^{+}07]. qualitative [RH03]. Quality [KKLP09, KKK08, TKM03, YHZ^{+}09, ZE03, dBvV03, ABG02, BGG09, BL03, BWDP00, CC08b, FMP09, FS01, Gru07, HCS09, HK09, HCC08, JMP07, KSH05, KC09, LKB06, LCM^{+}04, Mil00b, OAZ08, PS05, SJK07, TQ05, WWTH08, Wij03]. Quality-driven [KKLP09, TKM03, dBvV03]. Quality-of-service [KKK08]. Quantifying [ST07, WGH00]. Quantitative [GJ08, PS00, SS04, HCC08, LSaC04, RH03]. quantities [KLNS07]. quantization [CL06b]. queries [BG06, CMC04, CBK02, CK02a, LU06, LKL04, PSK05, ÜDUG04]. query [BLM^{+}08, CK02b, DCAC09, HL09, KRK00, KRP02, LZXS09, MLC09, ONR02, PC02, PCC02, PK02c, PKL03, RjHHK08, ÜDUG04, YC08b, RH06]. query-based [DCAC09]. queue [BCF^{+}05, SM03]. Queueing [BDMK03]. queues [KMOS09]. quite [BG06]. quorum [KTK01]. quorum-based [KTK01].

R [HLL01a, LZ06]. R-Chord [LZ06].

R-tree [HLL01a]. radio

[BPM06, BMES04, HL00b]. RAID [LKL05, Tho06]. RAID-0 [LKL05]. random [CKL08, CKL09, PJ09]. Ranking [GS07, Çam00a, SM06b]. Rapid [DZRH04, GD04, WKL04, KSH09]. rare [YHHR03]. rate [AD07, CSGL05, NXS00, PTM08, ZP06]. rate-control [CSGL05]. rating [KRHŽ05]. ratio [JZ07]. rationale [BL09, LICA09, TBGH06, TJH07, BB08]. rationale-based [TJH07]. RCES [LLCL08]. RCES/RSES [LLCL08]. RDL [OAdLC07]. re [AAAC07, CDEV08, HC04a, TKM03]. re-binding [CDEV08]. re-engineering [AAAC07, TKM03]. re-transmission [HC04a]. reactive [CJZ04, HLW08, KSH09, OAZ08, SD02, ZAO08]. Reading [MR00a, LASE00, dBvV08]. Real [LK04, dSSJV08, Yoo09, AV02, Ati00, BCK00, BNR09, Çam00b, CBZ00, CCSC01, CCSC07, CCKM09, CG05, Del08, DY03, DZRH04, DGL^{+}08, GBL08, GP05, Gho01, GWDE07, HA03, HSM^{+}07, HCDJ08, HLC^{+}09, HHL06, Iso01, JE02a, KBM05, KMSMD08, KCS01, KLY03, KMS04, KMOS09, LLL00, LKL02, LL00, LR04, LRS^{+}07, LJS05, LLV^{+}09, LHP^{+}09, LHP^{+}10, MMM00, MEH05, Nae01, NsL00, OW04, OAZ08, ÖKT09, PC04, QL03, Rav03, SUSO04, SSO05, SLS08, SO03, SY02, Shu03, TLW07, THP^{+}06, TL09b, WCLK07, WDN05, ZAO08, ZH05, LJB05, WOH08]. Real-time [LK04, dSSJV08, Yoo09, AV02, Ati00, BCK00, BNR09, Çam00b, CCSC01, CCSC07, CCKM09, CG05, Del08, DY03, DZRH04, DGL^{+}08, GBL08, GP05, GWDE07, HA03, HSM^{+}07, HCDJ08, HLC^{+}09, HHL06, KBM05, KMSMD08, KCS01, KLY03, KMS04, KMOS09, LLL00, LKL02, LL00, LR04, LRS^{+}07, LLV^{+}09, LHP^{+}09, LHP^{+}10, MMM00, MEH05, Nae01, NsL00, OW04, OAZ08, PC04, QL03, Rav03, SUSO04, SSO05, SLS08, SO03, SY02, Shu03, TLW07, THP^{+}06, TL09b, WCLK07, WDN05, ZAO08, ZH05, LJB05]. real-time/non-real-time [CCSC01]. real-world [Gho01, Iso01, LJS05]. reality [GHK05, SSCM^{+}04]. really [Kru08, PCV^{+}08, PVSG05]. reasoning [ANH07, BFPAGS^{+}08, EBGR01, TJH07]. reassessment [Ban08]. Rebalanced [SWH^{+}09]. Rebalanced-RSA [SWH^{+}09]. REBNITA’05 [CBVD07]. recapture [PTRW04, TR00]. receiver [PTM08]. receiver-centric [PTM08]. recognition [ZERO00]. recommendation [LL09, LS05b, LLH08]. recommender [BFPAGS^{+}08]. Reconciling [SMHMA08]. reconfigurable [CWC04, DHL06, HCKY08, KPT09, USLC01]. reconfiguration [CDI07]. recoverer [BDK08]. Recovering [RLvV06, JBA08]. recovery [ACDF01, DDGR09, HZCD05, HHL06, KSAOK04, LKJL01, LT08, LZN04, MMCB00]. rectangular [KH06]. recurring [Boz00]. Recursive [WHHT08, BBS00]. RED [GAWW07]. redistributing [SUSO04]. reduce [FW00]. Reducing [CJKC09]. reduction [DLW08, KSS03]. redundant [Alz08]. REDUP [HHL06]. Reengineering [SCdO02, UZ09, ACDG02, DGV08]. refactoring [OÓ08]. refactorings [CCHW09]. reference [AG08, BGH03, Ber03, CCHT09, GLJ00, SL02]. referencer [PTK00]. References [CCG01, HY00]. refinement [PCC02]. Refining [LZXS09]. reflection [YC08a]. Reflections [FHT07, Sai07]. reflective [Haz02]. reformulation [RjHHK08]. region [BRC09, HL09, KY08]. register [LSC04]. regression [BFV04, JIS03, MBB01, MDR06, SA06, mSgFtL05, ZL07]. regular [CK02a, PC02]. rejuvenation [PK02a]. Rekeying [HLT09]. related [CGSGR06, HH08a]. relation [JKWL09, LC08, MC01]. relation-based [LC08]. relational [AJCM08, LLC^{+}09, Phi05, TH02]. relations [JE02a, SZPMK04, ZKL^{+}09]. relationship [BGH^{+}08, Cha06, JH01, LLK05, YLC08]. relationships [BVN07, BWDP00, CC06, CGSGR06, YL09]. relative [YHHR03]. release [Hua05a, LS07, SL08]. Relevance [KCB05, KY08, YL09]. relevant [JG08]. Reliability [LHC^{+}05, RSP03, Bai05, CCW^{+}01, CJHB08, CJ05, FRR09, GMS07, Hua05a, Hua05b, HL06a, JZ05, JZ07, KR08, LH08, LH06, MT07, OOD09, PH06, PP04, ST07, Shy03, SH07, TSA08, Tho06, TN05, WPC06, YYWW07, ZP00, ZSP01, ZLCY06]. Reliable [SFSE05, HKY01, JCC05, LT07, MK06, SHW02]. Remarks [BCW05]. remote [CJT01, Shu03]. renaming [CDP05]. renovation [DNAM05]. repeatability [CC02a]. repercussions [FM08]. repetitive [HLWC04]. replacement [CE08, LSaC01]. replay [GMB^{+}09]. replicated [CY00, CWC04, MSA08, OFWP07, SKZ^{+}04, VM00]. replication [CK00b, MK08, WZJ01]. report [ADZ^{+}09, FIBRGCLN05, LG03, McD02]. repositories [CCD^{+}04, KGMI06, TH02, VMB^{+}08]. repository [Har04, Zhu00]. representation [CCK02, CL04a, Gur01, HRŽ06, LC00, LLT^{+}09, OAdLC07, WCLL09]. request [CLG08]. requests [JLC04, LKL05]. Requirement [XSS06, CCK02, CJKC09, KSS03, KV05]. Requirements [CBVD07, HKvVvdV07, ASS07, BHB^{+}05, BS09, Ber02, BCV06, BHL00, CMT02, DB06, EK00, EBB09, FM08, FCSM09, Gla00k, HRN^{+}01, JOZ03, JKWL09, KKP06, KPS08, KMKY07, LSV^{+}06, LM03, MLB09, Moy00, PILO06, Rey07, SZPMK04, SG01, SPZ06, TL09a, XZAR06, YKC^{+}05, ZJDB02, dBvV09]. requirements-uncertainty [Moy00]. resampling [MA08]. Reschedulable [CCSC01]. Reschedulable-Group-SCAN [CCSC01]. Research [RGV04, Wey01, CC08a, PTRW04, PKB09, SFJ04, Tan00, WD07]. researchers [VEM^{+}01]. resequencing [Kar00]. residual [LWL09]. resilient [TC06]. resolution [KPSK09, ZWX^{+}08]. resolving [KRHŽ05, LKL02, Lin01]. Resource [KMSMD08, KSH05, LYC04, LRS^{+}07, LCLL07, Zha08, ZCT^{+}09, Zhu04c, AM04, BHAM09, DXPY03, GP05, HSM^{+}07, jHjW08, HC01b, HL06a, KP07, LK09, LBS^{+}07, LSH09, LZ06, MA09, MK06, SRDLCP09, TLW07, WDCL08, Zhu04a, vdSJK^{+}07]. resource-constrained [KP07]. Resource-oriented [KSH05]. resourceful [GH02]. resources [ZWX^{+}08, Zhu06]. respectable [NER01]. response [DMQ07]. restoration [RW00, WC02]. restoring [CL06a, WCH03]. restrictive [CZL07, HH08b]. restructuring [Lee07, LZN04, LXZS06]. Results [DL06, PKL03, PKB09, DM07]. retailing [CDS02]. retrieval [BWM06, CC04, CK00a, Fra04, HDLK00, KCB05, KYPW06, KY08, LC00, LK01, LZL^{+}06, MCC02, Par00, PWLH06, PHN08, PB00, Pon05, RH06, RjHHK08, YL09, ZL04]. Retrieve [Zhu04d]. retrieving [YY04]. retrospect [REF^{+}07]. retrospective [LPS02]. reusability [GS07]. reusable [DF00, Fra04, LK09, SPZ06]. reuse [BHN02, CBS00, FK01, FS01, LdSBA^{+}08, Lut00, NR04, Sut00, Zhu06]. Revealing [Wil03]. revenue [TYH04]. reversal [ULN06]. reverse [ADZ^{+}09, BM00b, vDB05]. reversibility [KC09]. Reversible [CL06b, HCS09, CCY^{+}09, WCLL09, WLT^{+}09]. review [BKB^{+}07, EFG^{+}08, HJ00, Jør04, KG09, LC06b, Mül05, SN07]. reviewing [Wyn01]. revisitation [Ber02]. revisited [AAM00, Iso01]. Revocation [ZSM05]. revolution [Gla00i]. rewarding [FHHL09]. rework [DLW08]. right [WCLL09]. rightful [CL08, Lin01]. rights [HYJL04]. Rim [LW02]. ring [Rav03]. ripples [WK00]. Risk [Aml00, Ban08, EBGR01, FY04, FW00, GJ08, HH07, HMC01, KWT^{+}00, KLMZ08, NSL^{+}07, Pfl00, Sai07, Aml00]. Risk-based [Aml00, GJ08]. risks [dOBWT04, CdOBT07, EL07, JK00]. Risky [Pfl00]. RMI [JRB^{+}06]. Roam [hCSW^{+}04]. robotics [BNSG05, LRS^{+}07]. Robust [BNSG05, Lin00, TR00, CWP09, GP05, WCLK07, vdSJK^{+}07]. robustness [FMP09, GWDE07]. Role [GPM08, LWL04, Cho04a, CC06, Cow05, DRW00, LNC01]. Role-based [LWL04, Cho04a, LNC01]. roles [KLMZ08]. rotating [WLC07]. rotation [YC08a]. Round [LSZ^{+}07, CLC08b, LKH^{+}08]. round- [CLC08b]. Round-Eye [LSZ^{+}07]. router [CLL05]. routines [DF00]. routing [AN01, AM04, BHAM09, KSAOK04, KRC00, KPSK09, MHW01, TTC04, WGY^{+}08, YSK06]. routinized [IS03a]. row [LWHS05]. RSA [SWH^{+}09]. RSES [LLCL08]. RTCOM [DGL^{+}08]. Rule [SZPMK04, GH04, Zhu00]. Rule-based [SZPMK04, Zhu00]. rules [DPSU06, HWHM02, LcLsW06, LLC^{+}09, YHHR03, ZKL^{+}09]. run [HH00, JZL07, MM00b, SM00]. run-time [HH00, JZL07, MM00b, SM00]. running [DZW^{+}09]. Runtime [BS03, CLX^{+}04, USLC01, YHZ^{+}09].

s [KK07a, ALT^{+}09, HL01]. S-MARKS [ALT^{+}09]. safe [JTM04, Lut00, MMCB00]. Safety [LDL07, FL05, GD04, KPS^{+}04, LM03, PG05, SS04]. safety-critical [GD04, LM03]. sailing [Gla00n]. SAM [HYS^{+}04]. sampled [ED06]. samples [Par00]. sampling [CTY01, HH06, TPRW04]. SAVE [ÅCF^{+}07]. scalability [ZS05a]. scalable [AM04, PWCC01, SM03, YSK06, YSK09]. scale [BMES04, Deu01, JLC04, KL07, KPG^{+}07, PWLH06, TTC04]. SCAM [DHKV06]. SCAN [CCSC01]. scanning [LCLL08]. SCARAB [CMS04]. scarlet [Gla00b]. scenario [BW01, KKP06, LdSBA^{+}08, PILO06]. scenario-based [BW01]. scenarios [BJ03, RRD06, TSA08, WPP^{+}09]. schedulability [LHSK06, SLS08]. schedule [YY04]. schedule-based [YY04]. Scheduling [LG05b, BLL02, BNSG05, Çam00b, CCSC01, CCSC07, CCKM09, GH04, HTK00, HYC04, Kar01, KCS01, LL00, LC05, LCF08, LKL05, LHSK06, MMM00, OW04, PK01a, dSSJV08, SA05, TSSD09, TSPH06, WWC00, Yoo09, ZK09]. schemas [CT09, DZW^{+}09]. scheme [Aba06, BCW05, CC09a, CCSC01, CL06a, CL06b, CWP09, CJT01, CK00b, CHL^{+}08, CW09, CE08, CDZ07, FWTC05, HWW01, HH06, HC04b, HHL06, HLL01b, JW06, KBD09, KC09, LLCL08, LWL09, MV05, MV06, MK06, PTM08, SKZ^{+}04, Sha05, SCL07, Sha07, Sha09, mSgFtL05, TW07, TH02, WZJ01, WL05, WF07, WCLL09, WHG01, WH02, WH03, WL09, WLT^{+}09, XY02, YTH04, YC08b, ZC05]. schemes [CWH00, GPM08, HKY01, KTK01, KM04, LU06, LZG07, NsL00, PSH06, SHT05, VM00, WYL06, OS09]. scholars [Gla99, Gla00c, Gla00d, GC01, GC02, GC03, GC05, TCG06, WTG^{+}08, WTG^{+}09]. science [CC02a, RGV04, SZZ06, Sta02, ZL06]. Scientific [LC06b, Kel09]. scoped [LMvV09]. scores [Hus01]. SCTL [VAS^{+}04]. SCTL/MUS [VAS^{+}04]. SCTL/MUS-T [VAS^{+}04]. SDL [WSQM05]. SEAL [LLY07]. seamless [hCSW^{+}04]. Search [BWM06, CCH09, OÓ08, AAM00, FLA^{+}01, LC00, YZ08, ZK04a, ZC08]. Search-based [OÓ08, ZC08]. Search-order [CCH09]. searches [PTK00]. Searching [TPN^{+}09, Mus03]. secrecy [Tse07]. Secret [LT04, FWTC05]. secrets [DM07, TCC02]. section [BCDM06, KB07, LW02]. Secure [KMS04, RMC05, SCH05, ALT^{+}09, CC09a, FS06, HLT09, KLGH07, LLY07, WF07, YZ05, ZMN05]. securing [CH07a]. Security [CC02b, LKH^{+}08, LKH09, LL07, AV02, AMHJ09, DK01, EFG^{+}08, GPM08, GJ08, Kim07b, KJLK07, LLLZ06a, LLLZ06b, MBM^{+}09, OS09, PNL07, RRC07, ST07, SZZ06, SHT05, VHF02, WPP^{+}09, JRB^{+}06]. security-engineering [VHF02]. seek [CCSC07]. seek-optimizing [CCSC07]. Seeking [KJ01]. segment [WGW^{+}09]. Segmented [ACGS^{+}08, CGSGR06]. SEI [BT05]. SEKE’01 [VE03]. select [WHYT06]. Selected [DHKV06, Sai09]. Selecting [CCD^{+}04, DF00, WDS09]. selection [KLC02, LXG09, Loo05, MBB01, MK08, MSA08, SM00, VJB06]. selectivity [HLW08]. Self [Sha07, BCW05, BDK08, CWH00, HPT07, LL06, LY01, PSMB01, WH03, WL09]. Self-certified [Sha07, BCW05, CWH00, LL06, WH03, WL09]. self-contained [LY01]. self-organizing [PSMB01]. self-stabilizing [BDK08]. self-tuning [HPT07]. Semantic [BG06, LZ07, Pon06, Zhu06, MPG^{+}08, BFPAGS^{+}08, LZ06, O’B08, Zhu04d, dBvV08, AV04, AV08, DJW08, LICA09]. semantic-based [LZ06]. semantics [GHKR04, KNYS09, KZDX09, LLK05, ZC06, ZL06]. semi [VA08]. semi-automatic [VA08]. senior [ABG02, FHT07]. sensitivity [BRC09, LHC^{+}05]. sensor [BLM^{+}08, CFN07, FS06, JLYK09, KPSK09, LHP^{+}09, LHP^{+}10, MBM^{+}09, TL07, TL09b, ZCT^{+}09]. separability [XY02]. separate [Deu01]. separated [PCC02]. separation [CCF^{+}04, LWL04]. sequence [CZH^{+}08, HDLK00]. sequences [LZL^{+}06]. sequencing [LCCH02]. sequential [JFC08, KLNS07]. Serfs [Sri07]. serialization [LL00]. series [KYPW06, LKL04]. server [ABW07, BHAM09, CCDD00, CPL^{+}04, HL01, HC04a, NXS00, OFWP07, SKZ^{+}04, YS04]. servers [AKP04, CDC09, HH05, MA09, OFWP07, SM03, TYH04]. Service [DST^{+}04, MPG^{+}08, AT09, CT00, DMQ07, DGV^{+}07, DTV09, FMP09, GML05, KPTV09, KDS^{+}08, KUK07, KKK08, LPR04, LT09, LGL08, MGI07, PK02b, PNM04, RT07, SKZ^{+}04, TYH04, TDK^{+}07, TDL^{+}02, UZ09, YZ05, YGH^{+}08, ZMN05, ZG07, CFFT08, SSM^{+}09, ZS05a]. service-based [GML05, KDS^{+}08, YGH^{+}08]. Services [LP07, CDEV08, CLL05, CC08b, CMS04, FdSdP08, JCC05, JRB^{+}06, JSBR09, KSH09, LNPAGD^{+}06, MSA08, PSH06, PNL07, SRGL08, TSPH06, WZJ01, ZP05, Zha09, MPST06, ZL04]. services-based [SRGL08]. session [HLT09]. set [SW09, SKW06]. sets [MPST06, SS07, SSCL08, WDS09]. Setting [Lea08, CW02]. settings [Fra07]. several [JE02b, YL06]. Shape [KYPW06, HDLK00, LK01]. Shape-based [KYPW06]. shapes [ZERO00]. shared [CN04, Kar00, LUS^{+}00, USLC01, WDCL08]. shared-memory [Kar00]. Sharing [TCC02, AAAC07, FWTC05, INS00, LSH09, LUS^{+}00, LT04, LLH08, DM07, SSA08, WHYT06, YCYW07]. shift [Sta03]. Shortcut [Tho06]. Shortening [LZL^{+}06]. shorter [ED06]. Should [SW09, ED04]. side [CL06b, MSA08, XNP07]. side-effect [XNP07]. side-match [CL06b]. SigDAQ [PK02c]. sighting [Ber02]. signal [LLLZ06a, LLLZ06b]. signature [BCW05, CC09a, CWH00, CJT04, HWW01, HC04b, KBD09, LH01a, Sha05, SCL07, Sha07, Sha09, SHT05, WC07, WH03, WYL06, XY02, YTH04, ZC05]. signatures [CZL07, HRL09, HH08b, JL04]. signer [CJT04]. signer-verified [CJT04]. signers [HWW01, YTH04]. significance [FMSG08, Mil04, SK02]. significant [YHHR03]. similar [TPN^{+}09]. Similarity [HDLK00, CH07b, KCB05, dBvV09]. simple [Ayr04, CCW02b, HL01, HLL01b]. simplification [CCHT09]. simplified [RRT01]. simulated [PH06, TVA04]. Simulating [GHK05, TB00, BMES04, CS01]. Simulation [RW01, BGG^{+}06, CBZ00, CFN07, DI01b, ED04, ED06, ELK06, FCSM09, GW01, HRN^{+}01, HFC^{+}01, HMC01, mJKME01, LK09, LLV^{+}09, MR01, NKJT09, PWCC01, PKR01, RK00, SLC00, SP08, SG01, ZK04b]. simulations [CET^{+}08]. simulator [DI01a, LSaC04]. Simulink [ZC08]. Singapore [LC06b]. Single [Sta09, ABW07, MDFG08, WGW^{+}09]. single-company [MDFG08]. single-link [WGW^{+}09]. site [CT08, Pon06]. situation [YGH^{+}08]. situation-aware [YGH^{+}08]. Situational [ANH07]. Six [SM07]. size [AP09, CGMPAP08, HRŽ06, HH06, JH01, KPG^{+}07, MCCC03, MMC05]. skewed [SC07]. skies [Gla00n]. skills [CSNS05, MG04]. SLA [LS05a]. slack [SUSO04]. SLAs [DTV09]. slices [JG08, WQ06]. slicing [HBD03, Kri06, MKM^{+}06, MM06]. slower [Pon06]. SM [Lop03]. small [LMYMGT08, SS07, SSCL08]. Smart [WSQM05, LLL06, WHN^{+}01, BBC^{+}08]. Smart-Cards [BBC^{+}08]. SmartTutor [CHZY03]. smells [LS07]. SMIL [CH05]. SMIL2.0 [YWT07]. smooth [Gla00n]. SMP [HL01]. SMPCkpt [DCH02]. SMS4 [LGW09]. snapshot [KMS04]. Snort [WHC07]. SOAP [DZ05]. SoC [KPT09]. social [DJW08, Wyn01]. sockets [MKMS05]. SOCKS [OS09]. soft [KMSMD08, KR08, SLS08, ZERO00]. Software [BH02, BHXN05, BL03, CUY09, CG05, CSSW05, CGA08, Del08, FM08, Gla00j, Gla00k, Gla00l, GC01, HM00, JK00, JIS03, KLY03, KR08, KPME05, KB07, KV05, KRCK08, LP00, LLLZ06a, LHP^{+}10, DGV08, LdSBA^{+}08, McD02, Mea09, ML08, MDR06, MH04, NSL^{+}07, PH06, Par00, PV06, PKB09, RW01, Sai09, SJK07, TSA08, VE03, ZP06, ZLCY06, ZJDB02, vDB05, AZW07, ACS07, AW07, AD07, AK08, Ale05, AL05, Aml00, ABG02, ACGS^{+}08, Ati00, BKZ^{+}06, BVN07, BL09, BH03, BM05, Bai05, BNvdH05, Ban08, dOBWT04, BJ03, BCDM06, Ber03, BTV06, BFLP09, BDA^{+}02, BKB^{+}07, BWDP00, BW01, BDK08, Bud00, BT05, BM00b, CCW^{+}01, CC02a, CGHL07, CC08a, CJHB08, CGP^{+}05, CSNS05, CGL^{+}04, CCCT06, CJ05, CC07, CH09, CC09b, CJKC09, CLB05, CH07b, CS01, CC08c, CBS00, CO08, CDZ07, CdOBT07]. software [CN00, Cow05, CGSGR06, CGMPAP08, DI05, DXPY03, DLW08, DZ00, DSB05, DBO05, Deu01, DJW08, DS04, DI01a, DI01b, DD01, DRW00, Ebe07, ELH00, EB00, EBGR01, EE08, EC04, EL07, EK00, Fai07, FY04, FCSM09, FK01, FdSBR06, FW00, Fug03, GML05, GRRX01, GR05, GK08, Gla99, Gla00c, Gla00d, Gla00g, Gla00i, Gla00m, GC02, GC03, GC05, Gon08, GPHS07, GJ07, GSB^{+}07, GDH05, GS07, HALS08, HH07, HF08, HDGZ06, Har00, Har04, Haz02, HH08a, HYS^{+}04, HHW01, HGK^{+}06, HKN^{+}07, HFC^{+}01, HMC01, Hua05a, Hua05b, HL06a, HTH09, IS03a, IT03, JPKP04, JJ06, JAvdV09, JR09, JHSB09, JZ05, JZ07, JKD02, JX07, JTM04, Jør04, JFG07, JMSS07, KLRW01, KBK06, KWT^{+}00, KLMZ08, KS04, Kel09, KPS^{+}04, KSAOK04, KSH05, KBH07, KPS08, KPT09, KJ04, KBBW05]. software [KMKY07, KSS03, KCS08, Kru08, LD00, Let00, LXG09, LUS^{+}00, LCH^{+}04, LJS05, LH08, LKB06, LSV^{+}06, LDL07, LHC^{+}05, LH06, LMYMGT08, DPS03, LZN04, LG03, MYZC06, MS03, MB06, MR01, McB08, MV09, MGEB03, MKNS06, Mil00a, Mil02, Mil04, MDMC06, MKK09, MØHB08, MR00b, MSB^{+}02, Moy00, Mur08, NLSK04, NR04, NJ07, NWZ05a, NWZ05b, NER01, O’B08, OÓ08, OOD09, Özm09, PK02a, PDC01, PSMB01, PS05, DNAM05, PCDG02, PCV^{+}08, PTRW04, PIGÖ08, PKR01, PB00, PVSG05, PNM04, RK00, RRD06, RH02, RH03, RSM00, RSS00, RSP03, RR00, RGBM06, RW00, RMO^{+}08, RRT01, RO09, Sai02, SW05, Sai07, SNBH08, SA06, ST07, SL08, SMDM05, Shy03, SHW02, SW09, SA08, SS04, SM08, SS07, SSCL08, SSA08, SH07, SC01, Sut00, SM07, TA02, TKM03, Tan04, TPRW04]. software [TT09, TQ05, TNAA01, TN05, TCSC04, dBTdSS08, TK00, TL09a, TCG06, VM07, VK08, VMB^{+}08, VEM^{+}01, VP00, WPC06, WWB09, Wes02, WBP^{+}03, WK00, WHB01, WSQM05, WTG^{+}08, WTG^{+}09, WDN05, XYS07, XNP07, YS02, YKC^{+}05, YR09, ZP00, ZSP01, ZL07, ZS05b, dBvV08, dBvV03, dOZR^{+}04, dlRT06, BB08]. software-based [AZW07, KSAOK04]. software-first [Gla00i]. software/hardware [TCSC04]. solo [Mül07]. solution [CHY^{+}05, Wij03]. solutions [CJ09, KG09]. Solving [CJT04, HCDJ08, Dar02, DSSL09, KEK04]. Some [IS03a, JZ05, Wyn01, CTY01, MKK09, SHT05, WYL06]. sorting [MM01b, PS09]. sorts [Gla00f]. source [AW07, BGH^{+}08, CF07, CLL05, DH09, DDGR09, FMSG08, Fug03, KL07, RGBM06, SMR09, SHW09, SM08, SSA08, YSC^{+}06, ZQZ^{+}06, ZE03, DHKV06, KGMI06]. SPA [LLT^{+}09]. Space [Zha08, LWHS05, LO04, THP^{+}06, Zhu04a, Zhu04c]. spaces [LO04]. Spatial [LY01, CC04, HLL01a, LC00, LWHS05, Lin00, PCC02, TPN^{+}09, YL09]. spatio [CMC04, ÜDUG04]. spatio-temporal [CMC04, ÜDUG04]. spatiotemporal [KRK00, KRP02]. Special [CUY09, CGA08, GH08, LW02, LP07, BCDM06, HLM^{+}09, KB07]. specific [CCW02a, JHSB09, SK07, Spi01, ZGH^{+}07]. Specification [FdSBR06, LKJL01, YGH^{+}08, BNR09, DLB04, GPHS08, GHKR04, HZ07, LW07, LNPAGD^{+}06, PLCC09, SCdS^{+}06, VAS^{+}04, YS02, YKC^{+}05]. specification-based [HZ07]. specifications [EBB09, HCS04, HYS^{+}04, jHjW08, LYC04, Nae01, WW09, YLC06, ZAO08, dlRT06]. specified [HCWN05]. specify [ZC06]. Specifying [BCK00, DHJ05, KZDX09]. Spectral [SMDM05]. spectrum [AZGvG09, BPM06]. spectrum-based [AZGvG09]. speed [ELK06, NsL00]. speeding [SWH^{+}09]. SPICE [REF^{+}07]. spin [HPT07]. spin-locks [HPT07]. spiral [Sai07]. splines [BFV04, ZL07]. sponsored [FHT07]. spread [BPM06]. Springer [Zha08]. spyware [CHY^{+}05]. SQL [BG06]. SQUIRE [KLNS07]. SSL [JRB^{+}06]. stabilizing [BDK08]. stack [SLC00, TCSC04]. stage [CCC05, ED04, KK07a]. stage-activity [CCC05]. stages [BCB09]. staggered [PLF05]. stakeholder [BM00a, JKWL09]. stand [DF00]. stand-alone [DF00]. standard [AHGSS05, CC09a, GMR08, KRHŽ05, REF^{+}07]. standards [CF07, CBS00, EG00, LCM^{+}04]. Standby [PK02a]. star [AADAD02, WHYT06]. starting [SvV08]. State [RW00, DHJ05, ED06, HM09, LDL07, PW09, SZ06]. state-based [LDL07, SZ06]. statecharts [GHKR04]. statement [TH05]. statements [HH06]. Static [EKV05, WG05, CPiLH09, PS00, SL07, Zhu06]. station [HL00b]. Statistical [Mil04, THGL07, CLH07, CKM06, ED04]. statistics [EC04]. Std [KvV06]. steganographic [KC09, WWTH08]. steganography [CDS07, LT04, YCYW07]. stego [KC09]. stego-image [KC09]. step [CCDD00, LYX09]. Stepping [Car02]. stepwise [CDC09]. stereotypes [SKW06]. still [LC02]. Stochastic [HMC01, HM09, KEK04, Shy03, BM07]. stock [KMS04]. storage [HLL01a, IJC03, MCCC03, MK08]. storefronts [CCF^{+}04]. strategic [BCV06, SM08, UZ09]. Strategies [LKL02, BMOKAM09, BFPAGS^{+}08, KLT07, mJKME01, LO04, NWZ05a, SJK07, TL07]. strategy [CTY01, HC01b, HL02, KC09, MC04, PCC02, ÜDUG04, WFWL09, YC08a, YLC06, KMKY07]. stream [CH05, HKY01, LCLL08]. stream-based [LCLL08]. Streaming [KFS^{+}02, KD05, CDC09, CSGL05, HHL06, LG05a, LT09, vdSJK^{+}07]. streams [CTL08, PTM08]. stress [GBL08]. string [MM01b, Mus03]. strong [KBD09, XY02]. Structural [CR06, HL09, HZCD05, LVMM07, LC08, SM08, VMJS06]. structure [CD00, DPMD07, HLL01a, LGW09, LHC^{+}05, TMB02]. structured [YR09]. structures [BRMA^{+}09, ZG00]. student [GSB^{+}07, SM07]. students [HBM05, FHT07]. studies [Del08, Har00, Jør04, KK06]. study [AAAC07, AW07, AN01, ASS07, AL05, Aml00, AACL02, BKZ^{+}06, BGH^{+}08, BFPAGS^{+}08, BvD06, BT03, CJHB08, CGP^{+}09, CCCT06, CLSa01, CW02, CL04a, CC08c, CGSGR06, CGMPAP08, DZ05, DSB05, DZRH04, DF00, DJW08, ED04, EBB09, FAB^{+}07, FCL^{+}00, FLA^{+}01, FS01, Fra04, GRRX01, GR05, Gur01, HF08, HAHH06, HBVG08, JPK00, JH01, KLT07, mJKME01, KPME02, KPME05, LS07, LXG09, LSaC01, LTC01, LWC06, LO04, MDFG08, MMTL06, NWZ05b, NRG08, PLM07, PB04, PLF05, PVSG05, PV06, PSG^{+}09, RRD06, RGV04, RGBM06, Sal02, SKKL07, SSA08, SNJ^{+}07, SC01, THGL07, TDK^{+}07, VHF02, VAS^{+}04, WKH09, Yeu00, ZK04b]. style [MvD08]. styles [BGG^{+}06, Wil03]. subjective [ELH00]. submesh [Aba06]. subscribe [YSK06, YSK09]. subscriber [SO03]. subscription [YSK06]. success [CC08c, Gla00k, JKD02, MKK09, PCV^{+}08, PVSG05, PV06, PKB09, RH02, WSJK08, WHB01]. successful [JZ05, SM08]. suite [CMT02, Gur01]. suites [YZ08, ZAO08]. summary [HL09]. SUMMITrak [BDGR01]. Sun [WYL06]. supervisory [GWvD08]. supply [JJP02]. support [AK08, Ati00, BKZ^{+}06, BBG^{+}04, BHL00, BFV04, CL04b, CDZ07, DLB04, EE08, GML05, HH08a, HK09, JZL07, JSBR09, KSH09, LL09, LWL04, Lut00, MGI07, MPG^{+}08, NXS00, PH06, PH07, QHS08, RR09, RT07, RDD02, mSgFtL05, SPDT06, Wen03, YHHR03, ZHS01, ZP05]. supported [FIBRGCLN05, LNC01]. Supporting [dOBWT04, LDN04, WT01, CCL01, CMS04, HYC04, HCC05, KLY03, KBH07, RW00]. sure [JTM04]. surrounders [LSZ^{+}07]. survey [AMHJ09, CC08c, FB04, KPTV09, Lai02, LD00, TBGH06]. survivable [WGY^{+}08, WGW^{+}09]. survival [HCWN05]. SW [BBC^{+}08]. swarm [YYWW07]. SwiFT [LCH^{+}04]. swizzling [MC04]. symbolic [LC00]. symmetric [DCH02]. Synchronization [HKY01, YWT07, CH05, FS06, MV06]. synchronized [SG06]. synchronous [CCL01, PK01a, Tan04]. synergies [BFPAGS^{+}08]. synergy [Zhu06]. Synthesis [AMNT08, CCC06, CD07, KK07a, SD02, YGH^{+}08, ZCT^{+}09]. synthetic [PQLN04]. Syst [LHP^{+}10]. System [Ati00, DK08, KLRW01, SG01, AV02, AHGSS05, AL05, AACL02, ABW07, Ayr04, BBG^{+}04, BWM06, CFFT08, CLX^{+}04, CDI07, CGL^{+}04, CC02b, CC04, CCSC01, CLCY04, CK00a, CJZ04, CHZY03, CCC06, CD05, CNLV07, DB06, FBM09, GH02, GH04, HWM01, HC01a, HYC02, JC02, JJP02, JKD02, JLC04, Kar00, KUK07, KGMI06, KFS^{+}02, KRP02, KGT02, KW00, KLMC06, KJLK07, LWS^{+}03, LHLY05, LH04, LSZ^{+}07, LS07, LKJL01, Lin00, LKB06, LHP^{+}09, LHP^{+}10, MHC00, MV09, MDMC06, NXS00, ÖKT09, PLP04, PP04, RH06, RjHHK08, SMHMA08, SK03, SL02, SVMAM04, TYH04, TKA^{+}02, USLC01, VP07, WBW^{+}06, WKH09, WKL04, WHC07, WW00, YWLG02, YYWW07, YSK09, ZHS01, ZSM04]. system-level [JC02]. system-on-a-chip [CGL^{+}04]. Systematic [Bat08, TDT08, BKB^{+}07, FK01, KG09, SL03, SN07]. Systems [KPME05, LLLZ06a, MG04, ÅCF^{+}07, AZW07, Ati00, AMNT08, ABW07, BCK00, BRC09, BRMA^{+}09, BFPAGS^{+}08, BWDP00, CGP^{+}09, CET^{+}08, CLC08a, CM05, Cho04a, CHL04, CBK02, CS04, CGW08, CG05, DMQ07, DXPY03, Del08, DST^{+}04, DZRH04, Deu01, DGL^{+}08, ESW06, FIGCLN^{+}02, FRR09, GBL08, GTA09, GP05, Gho01, Gla99, Gla00c, Gla00d, GC02, GC03, GC05, GMS07, Gru07, GJ08, GWDE07, HCN00, HTK00, HA03, HSM^{+}07, HL00a, HLC^{+}09, HDLK00, HL02, HL06a, HFRHS09, HZ07, IBP03, INS00, JZL07, Jun00, KMSMD08, KLT07, Kar01, Kar04a, Kar04b, KLY03, KMS04, Kim07a, Kim07b, KJ01, KK07b, KSS03, KDEK04, KPG^{+}07, KP07, KLGH07, LCY00, LKL02, LBS^{+}07, LW02, LK02, LH01a, LLKL04, LR04, LSH09, LUS^{+}00, LCLL07, LWL04, LYX09, Lok06, Loo05, LWC06, ML03, MKL^{+}00, MMM00, MEH05, MMTL06, MR00a, MM00b, Nae01]. systems [O’B08, OFWP07, OAZ08, ONZ09, PLCC09, PK02a, PK02b, PKL03, PS09, PTBP08, PLM07, Phi04, Phi06, PH07, PB04, PLF05, PK01b, RRD06, Rav03, SMG08, SSCM^{+}04, SZ06, SUSO04, SSO05, SLS08, SM00, SG06, SK03, SCdO02, ST01, SM06a, SP08, SY02, SFSE05, SKKL07, dSSJV08, SK04, SA05, SD02, SDG^{+}07, SPSM03, SL01, TLW07, TT09, THP^{+}06, TNAA01, dBTdSS08, TCG06, TMD07, VM00, Wen03, WTG^{+}08, WTG^{+}09, XYS07, YGH^{+}08, YKC^{+}05, YSK06, YSC^{+}06, YR09, ZM06, ZAO08, ZS05b, dLGR06, GC01].

t [LNPAGD^{+}06, VAS^{+}04]. t-learning [LNPAGD^{+}06]. table [YLC08]. table-based [YLC08]. tables [JLYK09]. tacit [RO09]. Tactics [MLB09, KKLP09]. TAIC [HLM^{+}09]. Taking [CDS07, vV08]. Talk [Gla00l]. tamper [CCP05]. tampered [WCH03, WC02]. target [SLC00]. Task [KS04, YYWW07, CCKM09, FS05, HTK00, LWL04, MC01, OW04, SOC^{+}03, SA05, TA02, Yoo09, ZJDB02]. task-based [LWL04]. task-dependent [FS05]. Task-directed [KS04]. tasks [Çam00b, JJ06, KCS01, LRS^{+}07, PK01a, PC04]. taxonomy [NGC02]. TCI [BDGR01]. TCP [HCKY08]. Teaching [HBM05, RMO^{+}08, BNvdH05, Fra07]. team [BNSG05, RO09, ZS01]. team-robotics [BNSG05]. teams [RSM00, RO09]. technical [Gla00k]. technique [CCP05, CPL^{+}04, CK02a, HOR01, HH00, KRC00, KEK04, KDEK04, LK01, LC02, LC05, MK00, PC02, PK02c, VJB06, WLC08, YL09]. Techniques [IJC03, CY04, FIGCLN^{+}02, Fra07, KR08, KPG^{+}07, LASE00, LCF^{+}06, LZN04, LXZS06, MS03, MPST06, MA08, PG05, PWLH06, PB00, SD02, TLW07, ZCT^{+}09]. technologies [BM00b, LICA09]. Technology [ABCT06, ACDG02, DJW08, LL04, Mil04, PKB09, UN09, Zel09, Kim07a]. Technology-driven [ABCT06]. telecommunication [JLC04, TNAA01]. tell [CPT05]. templates [OKS08]. Temporal [IS03b, LPR04, BNR09, CMC04, CTL08, KRC00, LCY00, LLC^{+}09, NG08, NGM08, O’B08, ÜDUG04, ZC06]. term [Kel09]. terminal [CMS04]. terminals [FIGCLN^{+}02]. terminology [BDMK03]. terms [DHJ05]. test [CKL08, DL06, DI01a, FWA09, HY01, JG08, KYP^{+}03, LWN03, LC08, MBB01, MDMC06, SW09, SA08, YZ08, YLC06, ZAO08, DL06]. Test-Driven [DL06]. testability [BvD06, SS04]. testers [SW09]. Testing [FW00, LT08, Aml00, CGHL07, CJHB08, CCCT06, CCHT09, CBG09, CKM06, CKL08, CKL09, DXPY03, EL07, FIBRGCLN05, GBL08, GDH05, HM09, Hua05a, Hua05b, HL06a, KGT02, Lai02, LVMM07, LH08, LC08, Mil04, MDR06, Mur08, Phi05, PG04, PLP04, RRW00, SA08, Sta03, VJB06, VMJS06, WWB09, ZC08]. testing-effort [Hua05b, LH08]. testing-resource [DXPY03]. tests [JZ07]. text [Mus03]. texts [MR00a]. TFRP [CLH07]. their [BT05]. theoretic [BG09, MDMC06]. theoretical [CGMPAP08]. theories [Moy00]. theories-of-action [Moy00]. theory [CO08, KJ04, ML03, PTRW04]. theory-based [KJ04]. think [PCV^{+}08, PVSG05]. thoughts [Wyn01]. thread [CD05]. Three [CH05, CLC08b, CDZ07, LO04, SCH05, YC09]. Three-layer [CH05]. three-party [CLC08b, SCH05, YC09]. three-tier [CDZ07]. Threshold [WH03, BCW05, HWW01, JL04, SCL07, YTH04]. throughout [BM05]. tier [CDZ07, WDCL08]. time [AV02, Ati00, BCK00, BNR09, BCF^{+}05, Çam00b, CCSC01, CCSC07, CCKM09, CGW08, CG05, Del08, DY03, DZRH04, DGL^{+}08, FS06, GBL08, GP05, GWDE07, GAWW07, HA03, HSM^{+}07, HCDJ08, HLC^{+}09, HH00, HHL06, JZL07, KBM05, KMSMD08, KCS01, mJKME01, KLY03, KMS04, KYPW06, KMOS09, LLL00, LKL02, LRvV03, LL00, LKL04, LR04, LRS^{+}07, LK04, LLV^{+}09, LKL05, LHP^{+}09, LHP^{+}10, MMM00, MEH05, MM00b, Nae01, NsL00, OW04, OAZ08, Özm09, PC04, QL03, RVM06, Rav03, SUSO04, SSO05, SLS08, SO03, SM00, SY02, Shu03, dSSJV08, SK01, TLW07, THP^{+}06, TL09b, WCLK07, WDN05, YY04, Yoo09, ZAO08, ZH05, CR06, LJB05]. time-constrained [LKL05, SK01]. time-critical [CGW08]. time-driven [Özm09]. time-series [KYPW06]. time/non [CCSC01]. Timeboxing [JPKP04]. Timed [CGW08, LT07, LKJL01, DZW^{+}09, NsL00, WKH09, ZyCkP01, CR06]. timed-token [NsL00]. timeliness [AV02]. Timeslot [WHYT06]. Timeslot-sharing [WHYT06]. timestamping [NG08]. timing [BCK00, Nae01]. TOFF [CT00]. TOFF-2 [CT00]. token [NsL00, Rav03]. token-ring [Rav03]. tolerance [GH02, Lea08, LCH^{+}04, RW00, SSO05, SC09, WLC07, Zha09]. tolerant [AT09, CC01, CJZ04, CT00, CNLV07, HTK00, LKH09, Lin07, LY09, Tse07, WKH09]. Tool [KSH09, AN01, BT03, CMT02, FN00, HHW01, KPS^{+}04, MMM00, MM00a, PNL07, WD07, YZ08, ZGH^{+}07]. Tooling [BBG^{+}04]. tools [ED04, RS00]. top [HWML04]. top-down [HWML04]. topics [CC08a]. TOS [ZPEL01]. TPR [CMC04]. TPR-tree [CMC04]. trace [CZH^{+}08, EKV05]. traceability [LMvV09, SZPMK04, TJH07]. traces [IWF07, LZG07]. tracing [GM02]. tracker [ZEY04]. tracking [LSZ^{+}07, TL07, TL09b]. trade [Bat08]. trade-offs [Bat08]. tradeoff [Lop03, YHZ^{+}09]. Trading [SWH^{+}09, KMS04]. traditional [SSCM^{+}04]. Traffic [GBL08, CCL01, GAWW07, KM04, KMOS09, NsL00]. Traffic-aware [GBL08]. Training [Sai09, Fai07, Let00, PKR01, SW05]. trajectories [TPN^{+}09]. transaction [CM05, KW00, LL00, MMCB00, PJ09, TMB02]. transactions [LLL00, LLK04]. transcoding [LG05a]. transcription [RjHHK08]. transfer [ACDG02]. transform [BGG09, LWHS05]. transformation [Rey07]. transformations [KZDX09]. transition [DC09]. Transitioning [Wey01]. translation [CR06, Oi08, Yeu00]. transmission [HKY01, HC04a, NsL00]. transparent [AT09, Lin00]. transport [LHP^{+}09, LHP^{+}10]. Traps [CYH04]. traversal [CJ09]. Treating [BLL02]. tree [CMC04, HLL01a, IWF07, KY09, NJ07, SC00, SS04, TW07]. trees [CGW08, SPDT06]. triggered [LLL00, SFSE05]. trip [GH04]. true [KSAOK04]. trust [AZW07, BVN07]. trustworthiness [LNY06]. Trustworthy [Sch03, XYS07]. Tseng [LKH09]. Tukutuku [MDFG08]. tunable [HC06]. tuning [Del08, HPT07]. tuplespaces [JF04]. Turning [PKB09]. tutoring [CHZY03]. TVIS [HKW00]. Two [Del08, Mül05, Aba06, BS09, CK02a, DHL06, Gur01, HBVG08, KK07a, KL07, LC05, KK06]. two-dimensional [Aba06]. two-level [DHL06, LC05]. two-stage [KK07a]. type [Ayr04, CK02b]. Types [RR00, LUS^{+}00, ML08].

Ubiquitous

[BCF04, CJ09, HLT09, Tan04, FdSdP08]. UDDI [JSBR09]. UFN [LGW09]. UID [VA08]. ultimately [JCC05]. UM-RTCOM [DGL^{+}08]. UML [BM07, BLOS06, CT09, FLA^{+}01, GBL08, JHSB09, KZDX09, KSS03, LASE00, OD05, PSG^{+}09, SKW06, WPP^{+}09, ZPEL01]. UML-based [JHSB09]. UML-F [FLA^{+}01]. UML/OCL [CT09]. unanticipated [SM09]. uncaching [MC04]. uncaught [JCYC04]. Uncertainty [NLSK04, BCK00, BLL02, JKWL09, Moy00]. unconstrained [HH06]. undergraduate [Ale05]. undergraduates [SJ05]. understand [AD07]. Understanding [DMQ07, FCSM09, Gho01, MKNS06, SHW09, SSA08, Bat08, KV05, Zhu04d]. unified [YYL^{+}06, ZSM04]. uniform [CCW02b, LC05, LC07, PC01]. units [CGMPAP08]. universal [CC09a, Har04]. university [CSNS05, Wen03]. UNIX [IBP03]. unreliable [PK02b]. untestable [LNY06]. up-down [WCLL09]. update [LU06, McD02, YC08b, Zel09]. updating [MLC09, KNYS09]. upgrade [CSNS05]. upon [WLC08]. upper [KRHŽ05, SSO05]. urgency [HLC^{+}09]. US$1bn [Rey07]. Usability [PKL03, BJ03, BTV06, FB04, HAHH06, JMSS07, ONZ09, RRD06, SMHMA08]. Usability-based [PKL03]. usage [PP04, RRW00, SRDLCP09]. usage-based [RRW00]. Use [CN00, HK09, AD07, CELS07, CCC06, CP07, EG00, HA03, RRW00, SW09, ZQZ^{+}06, DJW08]. use-case-driven [CCC06]. user [AA07, AS01, CK00a, CMS04, JKD02, LZXS09, SFJ04, YS04]. Using [ADZ^{+}09, BPM06, BHB^{+}05, BFV04, BM00b, BB08, CP09, DJW08, FCL^{+}00, FLA^{+}01, FdSdP08, MØHB08, PSMB01, RR09, TNJH07, ZC06, AJCM08, BM05, BCW05, BCV06, BH09, CCC05, CGP^{+}09, CDS07, CWH00, CPL^{+}04, CMC04, CK02b, Dar02, DPSU06, DCH02, EMM01, EE08, EL07, FWTC05, FCSM09, FWA09, GBL08, Gok09, GDH05, GS07, HPT07, HTK00, HYS^{+}04, HCS09, HFC^{+}01, HY03, HWML04, JG08, JJP02, JZ07, KMSMD08, KC09, KRC00, KCB05, KKLP09, KV05, KRCK08, KP07, LL00, LLL06, LZL^{+}06, LDL07, LXZS06, MB06, MK08, MDFG08, NKJT09, PG05, Par00, PK02a, PWLH06, PJ09, PB04, PP04, Rav03, SMHMA08, SP08, dSSJV08, SPDT06, SN07, SKW06, SH07, SPSM03, TSA08, TQ05, TN05, TW07, TDK^{+}07, Wal05, WCLL09, WH03, YC09, YYWW07, YLC06, YHHR03, ZK04a, ZK04b, ZL07, ZBLG07]. utility [CTL08]. utilization [CSGL05, HLL01a, SM08, WCLK07]. Utilizing [PHN08]. UWIS [ONZ09].

validate [BHB^{+}05, CGP^{+}05]. validated [HCS04]. Validating [BCV06, EB00, Zel09]. validation [DI05, FIBRGCLN05, GDH05, Gur01, KMKY07, LLK04, LW07, LT08, LHP^{+}09, LHP^{+}10, OOD09, SD08, ZJDB02]. validity [JZ07]. value [LS05b, LWL09, VK08, WWTH08]. Variable [MCCC03, Oi08]. Variable-size [MCCC03]. variant [JOZ03]. variants [WL09]. VCR [PLF05]. vector [AM04, CL06b, EE08, PH06, mSgFtL05]. vectors [LWN03]. vehicular [ÅCF^{+}07]. venation [PHN08]. verifiability [CHL^{+}08]. verifiable [LC02]. Verification [CD07, KH06, TLW07, BS03, CWP09, DC09, GHKR04, HALS08, HA03, HLC^{+}09, LT07, LS05a, SL07, XYS07, dlRT06]. verified [CJT04]. verifier [CC09a, KBD09]. verifying [BCK00, DACY07]. version [HTH09, JSBR09, LNC01]. Versioning [SY02]. versus [UZ09]. vertical [SK04]. very [KY09]. vessels [WJT09]. via [BNR09, GD04, Shu03, ZJDB02]. video [BGG09, HH05, KD05, LG05a, LT09, NXS00, PTM08, PLF05, THP^{+}06, TYH04, ÜDUG04, WJT09, vdSJK^{+}07]. video-on-demand [NXS00, PLF05]. videoconferencing [HYC02]. view [HC01a, RS06, WSJK08]. viewing [LWS^{+}03]. Viewpoint [Gur01, XSS06]. Viewpoints [PNM04, FCL^{+}00, KvV06]. views [BH02, BH03, CZH^{+}08, Deu01, JKD02]. violations [SMR09]. Virtual [SSCM^{+}04, ZG07, CG03, DSC^{+}08, GD04, HSR01, Oi08]. Virtualization [WDCL08]. Virtualization-based [WDCL08]. virus [LCLL08]. visibility [VEM^{+}01]. ViSta [CMT02]. Visual [CCK02, DDGR09, KDS^{+}08, ZGH^{+}07]. Visualization [LICA09, CMT02, KLMC06]. voice [RjHHK08]. void [KPSK09]. volatility [FCSM09]. Volume [Ano01c, Ano01e, Ano02f, Ano02g, Ano03a, Ano03b, Ano03c, Ano03d, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano05e, Ano05f, Ano05g, Ano01d, Ano02e]. vote [CY00]. voting [BS09, CW09]. VQ [LWL09]. VQ-index [LWL09]. vs [MDFG08, SSCM^{+}04, THGL07, TDK^{+}07]. vulnerabilities [MV09].

wallet [JL04]. warehouse [HL00a]. warehouses [ZM06]. warehousing [HC01a]. warmup [ED06]. warning [LKB06]. watermark [CL08]. watermarking [CC02b, Lin00, Lin01, mSgFtL05]. wavelet [BGG09, KRCK08]. way [LKJL01]. ways [BS09]. WDM [WHYT06]. weak [PG04]. weak-branch [PG04]. weakness [LKH09]. weaving [WPP^{+}09]. Web [LZ07, Pon03, Zha08, AP09, AT09, AKP04, ASS07, AV04, AV08, BM05, CT08, CDEV08, CCC05, CHZY03, CLG08, CE08, DA07, DJW08, FCL^{+}00, FdSdP08, HYC02, JR09, JRB^{+}06, JSBR09, KDS^{+}08, KLC02, KKK08, LS04, LT08, Lok06, LICA09, MT07, MPST06, MA09, MMC05, MDFG08, MSA08, ONZ09, Pon05, Pon06, PQLN04, RRD06, SMG08, SRGL08, SM06a, SSM^{+}09, WDCL08, YWLG02, ZK04a, Zha09, ZL04]. Web-application [Pon03]. Web-based [BM05, CHZY03, FCL^{+}00, HYC02, ONZ09]. Web-crawling [YWLG02]. weight [DDF^{+}05, PIGÖ08]. weighted [SH07]. weighting [KY08, LXG09]. weights [AHGSS05, WZG09]. well [LRvV03, BM07]. Well-formed [BM07]. we’ve [Mea09]. Wheel [HAHH06]. where [SvV08]. Who [JMP07]. wide [HYC04, LY09]. wide-area [HYC04, LY09]. Wikis [RR09]. Win [Sai07, FHT07]. Win-Win [Sai07]. Windows [AS01, CPiLH09, LCH^{+}04]. winner [GSB^{+}07]. WinWin [GSB^{+}07]. wired [LT09]. wired/wireless [LT09]. Wireless [FIBRGCLN05, BLM^{+}08, CC08b, CK00b, CBK02, HHL06, HCC05, KPSK09, LLK04, LG05a, LT09, Lin07, LHP^{+}09, LHP^{+}10, MBM^{+}09, PJ09, SC07, SC08, SKKL07, ZP05, ZK04a, ZCT^{+}09]. within [BKB^{+}07, CLW05, SL03, ZG07]. without [CC01, FM08, LL06]. won’t [HKvVvdV07]. work [GPHS08, IS03a, RR00]. Workflow [CCC05, ACDG02, DCAC09, DZW^{+}09, jHjW08, LH01a, LYC04, LWL04, SK01, SOC^{+}03, ZyCkP01]. workflows [CLW05, ZWX^{+}08]. workload [CBZ00, CHL04, TSSD09]. workload-dependent [TSSD09]. Workshop [CBVD07, DHKV06]. workspace [CCF^{+}04]. workspaces [JF04]. workstations [Boz00, WZJ01]. World [LC06b, Gho01, Iso01, LJS05]. Worm [CNLV07]. Worm-IT [CNLV07]. wormhole [CNLV07]. wormhole-based [CNLV07]. Worst [BCF^{+}05]. wrappers [HL00a]. wrappers/monitors [HL00a]. wrapping [CFFT08]. writing [LNC01]. WS [JRB^{+}06]. WS-Security [JRB^{+}06]. WSDL [JSBR09]. WTA [SOC^{+}03]. Wu [BCW05].

X.509 [RMC05]. Xia [CJT04, Sha05]. XML [BHN02, CDS02, CCCT06, CH07a, CL04a, CLC08a, CM05, CK02a, CK02b, EFG^{+}08, HL09, KY09, LTC01, LWC06, MLC09, NGM08, PK02c, PWLH06, PILO06, SVMAM04, TH02, YSK06, YC08b]. XML-based [CCCT06, CLC08a, NGM08, YSK06]. XML/EDI [LTC01]. XMobile [VA08].

XSL [LDN04].

Y2K [Gla00j, Gla00m, Gla00n]. Yang [SCL07, WL05]. years [FHT07, PTRW04, SM07]. Yen [LLLZ06a, LLLZ06b].

Z [GHKR04]. Zhuge [LC06b, Zha08].

References

Ahmed:2007:MBU

[AA07] Seffah Ahmed and Gaffar Ashraf. Model-based user interface engineering with design patterns. The Journal of Systems and Software, 80 (8):1408–1422, August 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Abi-Antoun:2007:CSR

[AAAC07] Marwan Abi-Antoun, Jonathan Aldrich, and Wesley Coelho. A case study in re-engineering to enforce architectural control flow and data sharing. The Journal of Systems and Software, 80(2): 240–264, February 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Andrews:2002:ICB

[AACL02] David Andrews, Paul Austin, Peter Costello, and David LeVan. Interprocess communications in the AN/BSY-2 distributed computer system: a case study. The Journal of Systems and Software, 61(3): 233–242, April 1, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Al-Ayyoub:2002:ASN

[AADAD02] Abdel-Elah Al-Ayyoub, Khaled Day, and Abdullah Al-Dhelaan. On the attractiveness of the star network. The Journal of Systems and Software, 63(2): 151–158, August 15, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Al-Ayyoub:2000:HSR

[AAM00] Abdel-Elah Al-Ayyoub and Fawaz A. Masoud. Heuristic search revisited. The Journal of Systems and Software, 55(2):103–113, December 27, 2000. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/54/24/25/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/54/
24/25/article.pdf.

Ababneh:2006:EFL

[Aba06] Ismail Ababneh. An efficient free-list submesh allocation scheme for twodimensional mesh-connected multicomputers. The Journal of Systems and Software, 79 (8):1168–1179, August 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ababneh:2008:ABN

[Aba08] Ismail Ababneh. Availabilitybased noncontiguous processor allocation policies for 2D mesh-connected multicomputers. The Journal of Systems and Software, 81(7): 1081–1092, July 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Aversano:2006:TDB

[ABCT06] Lerina Aversano, Thierry Bodhuin, Gerardo Canfora, and Maria Tortorella. Technology-driven business evolution. The Journal of Systems and Software, 79(3): 314–338, March 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anderson:2002:EDM

[ABG02] Bonnie Brinton Anderson, Akhilesh Bajaj, and Wilpen Gorr. An estimation of the decision models of senior IS managers when evaluating the external quality of organizational software. The Journal of Systems and Software, 61(1):59– 75, March 1, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Avritzer:2007:ESP

[ABW07] Alberto Avritzer, André Bondi, and Elaine J. Weyuker. Ensuring system performance for cluster and single server systems. The Journal of Systems and Software, 80 (4):441–454, April 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Antoniol:2001:OOD

[ACDF01] G. Antoniol, G. Casazza, M. Di Penta, and R. Fiutem. Object-oriented design patterns recovery. The Journal of Systems and Software, 59(2):181–196, November 15, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        29/11/71/28/33/abstract.
                                        html.

Aversano:2002:BPR

[ACDG02] Lerina Aversano, Gerardo Canfora, Andrea De Lucia, and Pierpaolo Gallucci. Business process reengineering and workflow automation: a technology transfer experience. The Journal of Systems and Software, 63(1): 29–44, July 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

AAkerholm:2007:SAC

[ÅCF^{+}07] Mikael Åkerholm, Jan Carlson, Johan Fredriksson, Hans Hansson, John Håkansson, Anders Möller, Paul Pettersson, and Massimo Tivoli. The SAVE approach to component-based development of vehicular systems. The Journal of Systems and Software, 80(5):655–667, May 2007. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Aroba:2008:SSC

[ACGS^{+}08] Javier Aroba, Juan J.

Cuadrado-Gallego, Miguel-Ángel Sicilia, Isabel Ramos, and Elena García-Barriocanal. Segmented software cost estimation models based on fuzzy clustering. The Journal of Systems and Software, 81(11):1944–1950, November 2008. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ahmed:2007:ISP

[ACS07] Faheem Ahmed, Luiz Fernando Capretz, and Shahbaz Ali Sheikh. Institutionalization of software product line: an empirical investigation of key organizational factors. The Journal of Systems and Software, 80 (6):836–849, June 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Andronikos:2008:CR

[ACT^{+}08] Theodore Andronikos, Florina M. Ciorba, Panayiotis Theodoropoulos, Dimitrios Kamenopoulos, and George Papakonstantinou. Cronus. The Journal of Systems and Software, 81 (8):1389–1405, August 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ajila:2007:EUC

[AD07] Samuel A. Ajila and Razvan T. Dumitrescu. Experimental use of code delta, code churn, and rate of change to understand software product line evolution. The Journal of Systems and Software, 80(1): 74–91, January 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Adams:2009:UAO

[ADZ^{+}09] Bram Adams, Kris De Schutter, Andy Zaidman, Serge Demeyer, Herman Tromp, and Wolfgang De Meuter. Using aspect orientation in legacy environments for reverse engineering using dynamic analysis—An industrial experience report. The Journal of Systems and Software, 82(4):668–684, April 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Angelov:2008:CRA

[AG08] Samuil Angelov and Paul Grefen. An e-contracting reference architecture. The Journal of Systems and Software, 81(11):1816–1844, November 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Al-Hajri:2005:MSF

[AHGSS05] Mohammed Abdullah Al-Hajri, Abdul Azim Abdul Ghani, Md Nasir Sulaiman, and Mohd Hasan Selamat. Modification of standard Function Point complexity weights system. The Journal of Systems and Software, 74(2):195–206, January 15, 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Al-Jumaily:2008:ODA

[AJCM08] Harith T. Al-Jumaily, Dolores Cuadra, and Paloma Martínez. OCL2Trigger: Deriving active mechanisms for relational databases using Model-Driven Architecture. The Journal of Systems and Software, 81(12):2299– 2314, December 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ajila:2008:ESM

[AK08] Samuel A. Ajila and Ali B.

Kaba. Evolution support mechanisms for software product line process. The Journal of Systems and Software, 81(10):1784–1801, October 2008. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ahn:2004:CAC

[AKP04] Woo Hyun Ahn, Woo Jin Kim, and Daeyeon Park. Content-aware cooperative caching for cluster-based Web servers. The Journal of Systems and Software, 69 (1–2):75–86, January 1, 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Alshayeb:2005:ESS

[AL05] Mohammad Alshayeb and Wei Li. An empirical study of system design instability metric and design evolution in an agile software process. The Journal of Systems and Software, 74(3): 269–274, February 1, 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Alexander:2005:IFU

[Ale05] Perry Alexander. Integrating formalism into undergraduate software engineering. The Journal of Systems and Software, 74(2): 147–154, January 15, 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ahamed:2009:DIM

[ALT^{+}09] Sheikh Iqbal Ahamed, Haifeng Li, Nilothpal Talukder, Mehrab Monjur, and Chowdhury Sharif Hasan. Design and implementation of S-MARKS: a secure middleware for pervasive computing applications. The Journal of Systems and Software, 82(10):1657–1677, October 2009. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Alzamil:2008:ARC

[Alz08] Zakarya A. Alzamil. Application of redundant computation in program debugging. The Journal of Systems and Software, 81(11):2024– 2033, November 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Amin:2004:ABD

[AM04] Kaizar A. Amin and Armin R.

Mikler. Agent-based distance vector routing: a resource efficient and scalable approach to routing in large communication networks. The Journal of Systems and Software, 71 (3):215–227, May 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Arnedo-Moreno:2009:SSJ

[AMHJ09] Joan Arnedo-Moreno and Jordí Herrera-Joancomartí. A survey on security in JXTA applications. The Journal of Systems and Software, 82(9): 1513–1525, September 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Amland:2000:RBT

[Aml00] Ståle Amland. Risk-based testing: Risk analysis fundamentals and metrics for software testing including a financial application case study. The Journal of Systems and Software, 53 (3):287–295, September 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/52/29/
33/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/52/29/abstract.
html.

Autili:2008:SDC

[AMNT08] Marco Autili, Leonardo Mostarda, Alfredo Navarra, and Massimo Tivoli. Synthesis of decentralized and concurrent adaptors for correctly assembling distributed component-based systems. The Journal of Systems and Software, 81(12):2210– 2236, December 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Akkanen:2001:CSE

[AN01] Jyrki Akkanen and Jukka K.

Nurminen. Case study of the evolution of routing algorithms in a network planning tool. The Journal of Systems and Software, 58(3):181–198, September 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/68/34/27/abstract.
html.

Anagnostopoulos:2007:SCI

[ANH07] C. B. Anagnostopoulos, Y. Ntarladimas, and S. Hadjiefthymiades. Situational computing: an innovative architecture with imprecise reasoning. The Journal of Systems and Software, 80(12):1993–2014, December 2007. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2001:Ca

[Ano01a] Anonymous. Contents. The Journal of Systems and Software, 56(2):iii, March 1, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/57/30/24/abstract.
html;   http://www.elsevier.

                                        nl/gej-ng/10/29/11/57/
                                        30/24/article.pdf.

Anonymous:2001:Cb

[Ano01b] Anonymous. Contents. The Journal of Systems and Software, 57(2):iii, June 15, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        29/11/64/29/25/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/29/11/64/
                                        29/25/article.pdf.

Anonymous:2001:CVa

[Ano01c] Anonymous. Contents volume 57. The Journal of Systems and Software, 57(3): 235–236, July 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        29/11/64/32/34/abstract.
                                        html.

Anonymous:2001:CVb

[Ano01d] Anonymous. Contents volume 58. The Journal of Systems and Software, 58(3):281–282, September 15, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        29/11/68/34/35/abstract.
                                        html.

Anonymous:2001:CVc

[Ano01e] Anonymous. Contents volume 59. The Journal of Systems and Software, 59(3):363–364, December 15, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/71/34/40/abstract.
html.

Anonymous:2001:CC

[Ano01f] Anonymous. Controversy Corner. The Journal of Systems and Software, 58(2): 169, September 1, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/68/33/34/abstract.
html.

Anonymous:2001:EC

[Ano01g] Anonymous. Editor’s Corner. The Journal of Systems and Software, 58(1): 63, August 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/68/27/32/abstract.
html.

Anonymous:2002:CPa

[Ano02a] Anonymous. Call for papers. The Journal of Systems and Software, 63(1): 77, July 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2002:CPb

[Ano02b] Anonymous. Call for papers. The Journal of Systems and Software, 63(2): 159, August 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2002:Ca

[Ano02c]  Anonymous.  Contents.

The Journal of Systems and Software, 63(3):259–260, September 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2002:Cb

[Ano02d] Anonymous. Contents. The Journal of Systems and Software, 64(3):257–258, December 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2002:CVa

[Ano02e] Anonymous. Contents of Volume 60. The Journal of Systems and Software, 60(3):263–264, February 15, 2002. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        29/11/72/29/36/abstract.
                                        html.

Anonymous:2002:CVb

[Ano02f] Anonymous. Contents volume 61. The Journal of Systems and Software, 61(3): 243–244, April 1, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2002:CVc

[Ano02g] Anonymous. Contents volume 62. The Journal of Systems and Software, 62(3): 217, June 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2002:EBa

[Ano02h] Anonymous. Editorial board.

The Journal of Systems and Software, 63(1):CO2, July 15, 2002. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2002:EBb

[Ano02i] Anonymous. Editorial board.

The Journal of Systems and Software, 63(2):CO2, August 15, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2002:EBc

[Ano02j] Anonymous. Editorial board.

The Journal of Systems and Software, 63(3):CO2, September 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2002:EBd

[Ano02k] Anonymous. Editorial board. The Journal of Systems and Software, 64(1):CO2, October 15, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2002:EBe

[Ano02l] Anonymous. Editorial board. The Journal of Systems and Software, 64(2):CO2, November 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2002:EBf

[Ano02m] Anonymous. Editorial board. The Journal of Systems and Software, 64(3):CO2, December 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2003:CVa

[Ano03a] Anonymous. Contents volume 65. The Journal of Systems and Software, 65(3): 237–238, March 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2003:CVb

[Ano03b] Anonymous. Contents volume 66. The Journal of Systems and Software, 66(3):

285–286, June 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2003:CVc

[Ano03c] Anonymous. Contents volume 67. The Journal of Systems and Software, 67 (3):225–226, September 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2003:CVd

[Ano03d] Anonymous. Contents volume 68. The Journal of Systems and Software, 68(3): 263–264, December 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2003:EBa

[Ano03e] Anonymous. Editorial board.

The Journal of Systems and Software, 65(1):CO2, January 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2003:EBb

[Ano03f] Anonymous. Editorial board.

The Journal of Systems and Software, 65(2):CO2, February 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2003:EBc

[Ano03g] Anonymous. Editorial board. The Journal of Systems and Software, 65(3):CO2, March 15, 2003. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2003:EBd

[Ano03h] Anonymous. Editorial board. The Journal of Systems and Software, 66(1):CO2, April 15, 2003. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2003:EBe

[Ano03i] Anonymous. Editorial board. The Journal of Systems and Software, 66(2):CO2, May 15, 2003. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2003:EBf

[Ano03j] Anonymous. Editorial board. The Journal of Systems and Software, 66(3):CO2, June 15, 2003. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2003:EBg

[Ano03k] Anonymous. Editorial board. The Journal of Systems and Software, 67(1):CO2, July 15, 2003. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2003:EBh

[Ano03l] Anonymous. Editorial board.

The Journal of Systems and Software, 67(2):CO2, August 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2003:EBi

[Ano03m] Anonymous. Editorial board.

The Journal of Systems and Software, 67(3):CO2, September 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2003:EBj

[Ano03n] Anonymous. Editorial board.

The Journal of Systems and Software, 68(1):CO2, October 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2003:EBk

[Ano03o] Anonymous. Editorial board.

The Journal of Systems and Software, 68(2):CO2, November 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2003:EBl

[Ano03p] Anonymous. Editorial board.

The Journal of Systems and Software, 68(3):CO2, December 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2004:CVa

[Ano04a] Anonymous. Contents volume 69. The Journal of Systems and Software, 69(3): 297–298, January 15, 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2004:CVb

[Ano04b] Anonymous. Contents volume 70. The Journal of Systems and Software, 70(3): 331–332, March 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2004:CVc

[Ano04c] Anonymous. Contents volume 71. The Journal of Systems and Software, 71(3): 297–298, May 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2004:CVd

[Ano04d] Anonymous. Contents volume 72. The Journal of Systems and Software, 72(3): 467–468, August 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2004:CVe

[Ano04e] Anonymous. Contents volume 73. The Journal of Systems and Software, 73(3):563–565, November/ December 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2004:EBa

[Ano04f] Anonymous. Editorial board.

The Journal of Systems and Software, 69(1–2):CO2, January 1, 2004. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2004:EBb

[Ano04g] Anonymous. Editorial board.

The Journal of Systems and Software, 69(3):CO2, January 15, 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2004:EBc

[Ano04h] Anonymous. Editorial board.

The Journal of Systems and Software, 70(1–2):CO2, February 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2004:EBd

[Ano04i] Anonymous. Editorial board.

The Journal of Systems and Software, 70(3):CO2, March 2004. CODEN JSSODM.

ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2004:EBe

[Ano04j] Anonymous. Editorial board. The Journal of Systems and Software, 71(1–2):CO2, April 2004. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2004:EBf

[Ano04k] Anonymous. Editorial board. The Journal of Systems and Software, 71(3):CO2, May 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2004:EBg

[Ano04l] Anonymous. Editorial board. The Journal of Systems and Software, 72(1):CO2, June 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2004:EBh

[Ano04m] Anonymous. Editorial board. The Journal of Systems and Software, 72(2):CO2, July 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2004:EBi

[Ano04n] Anonymous. Editorial board. The Journal of Systems and Software, 72(3):CO2, August 2004. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2004:EBj

[Ano04o] Anonymous. Editorial board.

The Journal of Systems and Software, 73(1):CO2, September 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2004:EBk

[Ano04p] Anonymous. Editorial board.

The Journal of Systems and Software, 73(2):CO2, October 2004. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2004:EBl

[Ano04q] Anonymous. Editorial board.

The Journal of Systems and Software, 73(3):CO2, November/December 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2005:Ca

[Ano05a] Anonymous. Contents. The Journal of Systems and Software, 74(1):iv, January 1, 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2005:Cb

[Ano05b] Anonymous. Contents. The Journal of Systems and Software, 74(2):iv, January 15, 2005. CODEN JSSODM.

ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2005:Cc

[Ano05c] Anonymous. Contents. The Journal of Systems and Software, 75(1–2):iv, February 15, 2005. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2005:Cd

[Ano05d] Anonymous. Contents. The Journal of Systems and Software, 76(1):iv, April 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2005:CVa

[Ano05e] Anonymous. Contents volume 74. The Journal of Systems and Software, 74(3): 337–338, February 1, 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2005:CVb

[Ano05f] Anonymous. Contents volume 75. The Journal of Systems and Software, 75(3): 329–330, March 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2005:CVc

[Ano05g] Anonymous. Contents volume 76. The Journal of Systems and Software, 76(3):

341–342, June 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2005:EBa

[Ano05h] Anonymous. Editorial board.

The Journal of Systems and Software, 74(1):CO2, January 1, 2005. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2005:EBb

[Ano05i] Anonymous. Editorial board.

The Journal of Systems and Software, 74(2):CO2, January 15, 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2005:EBc

[Ano05j] Anonymous. Editorial board.

The Journal of Systems and Software, 74(3):CO2, February 1, 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2005:EBd

[Ano05k] Anonymous. Editorial board.

The Journal of Systems and Software, 75(1–2):CO2, February 15, 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2005:EBe

[Ano05l] Anonymous. Editorial board. The Journal of Systems and Software, 75(3):CO2, March 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2005:EBf

[Ano05m] Anonymous. Editorial board. The Journal of Systems and Software, 76(1):CO2, April 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2005:EBg

[Ano05n] Anonymous. Editorial board. The Journal of Systems and Software, 76(2):CO2, May 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2005:EBh

[Ano05o] Anonymous. Editorial board. The Journal of Systems and Software, 76(3):CO2, June 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2005:EBi

[Ano05p] Anonymous. Editorial board. The Journal of Systems and Software, 77(1):CO2, July 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Anonymous:2005:EBj

[Ano05q] Anonymous. Editorial board.

The Journal of Systems and Software, 77(2):CO2, August 2005. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Abrahao:2009:FEE

[AP09] Silvia Abrahão and Geert Poels. A family of experiments to evaluate a functional size measurement procedure for Web applications. The Journal of Systems and Software, 82(2): 253–269, February 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Arisholm:2000:TFE

[AS00] Erik Arisholm and Dag I. K.

Sjøberg. Towards a framework for empirical assessment of changeability decay. The Journal of Systems and Software, 53(1):3–14, July 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/52/25/
26/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/52/25/abstract.
html.

Ahn:2001:DUG

[AS01] Gail-Joon Ahn and Ravi Sandhu. Decentralized user group assignment in Windows NT. The Journal of Systems and Software, 56(1): 39–49, February 1, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/57/24/27/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/57/
24/27/article.pdf.

Al-Salem:2007:EWA

[ASS07] Lana S. Al-Salem and Ala Abu Samaha. Eliciting Web application requirements — an industrial case study. The Journal of Systems and Software, 80(3): 294–313, March 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Aghdaie:2009:CTF

[AT09] Navid Aghdaie and Yuval Tamir. CoRAL: a transparent fault-tolerant Web service. The Journal of Systems and Software, 82(1): 131–143, January 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Atif:2000:SSS

[Ati00] Yacine Atif. System software support for distributed realtime systems. The Journal of Systems and Software, 53 (3):245–264, September 15, 2000. CODEN JSSODM.

ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/52/29/
30/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/52/29/abstract.
html.

Ahmed:2002:MST

[AV02] Quazi N. Ahmed and Susan V. Vrbsky. Maintaining security and timeliness in real-time database system. The Journal of Systems and Software, 61(1):15– 29, March 1, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Antoniou:2004:SWP

[AV04] G. (Grigoris) Antoniou and Frank Van Harmelen. A Semantic Web primer. Cooperative information systems. MIT Press, Cambridge, MA, USA, 2004. ISBN 0-262-01210-3. xx + 238 pp. LCCN TK5105.88815 .A58 2004.

Antoniou:2008:SWP

[AV08] G. (Grigoris) Antoniou and Frank Van Harmelen. A Semantic Web primer. Cooperative information systems. MIT Press, Cambridge, MA, USA, second edition, 2008. ISBN 0-262-01242-1 (hardcover). xxi + 264 pp. LCCN TK5105.88815 .A58 2008. URL http:/

/www.loc.gov/catdir/toc/
ecip0718/2007020429.html.

Ajila:2007:ESE

[AW07] Samuel A. Ajila and Di Wu. Empirical study of the effects of open source adoption on software development economics. The Journal of Systems and Software, 80(9): 1517–1529, September 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ayres:2004:SPT

[Ayr04] Ronald F. Ayres. A simple and powerful type system for programming languages. The Journal of Systems and Software, 73(2): 301–311, October 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Abreu:2009:PES

[AZGvG09] Rui Abreu, Peter Zoeteweij, Rob Golsteijn, and Arjan J. C. van Gemund. A practical evaluation of spectrum-based fault localization. The Journal of Systems and Software, 82(11): 1780–1792, November 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ahamed:2007:SBT

[AZW07] Sheikh I. Ahamed, Mohammad Zulkernine, and Steve Wolfe. A software-based trust framework for distributed industrial management systems. The Journal of Systems and Software, 80(10):1621–1630, October 2007. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bae:2005:E

[Bae05]  Doo-Hwan Bae.  Edito-

rial. The Journal of Systems and Software, 76(1): 1–2, April 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bae:2006:GE

[Bae06] Doo-Hwan Bae. Guest editorial. The Journal of Systems and Software, 79(7): 889–890, July 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bai:2005:BNB

[Bai05] Cheng-Gang Bai. Bayesian network based software reliability prediction with an operational profile. The Journal of Systems and Software, 77(2):103–112, August 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bannerman:2008:RRM

[Ban08] Paul L. Bannerman. Risk and risk management in software projects: a reassessment. The Journal of Systems and Software, 81(12):

2118–2133, December 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bate:2008:SAU

[Bat08] I. Bate. Systematic approaches to understanding and evaluating design tradeoffs. The Journal of Systems and Software, 81(8):1253– 1271, August 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Burge:2008:SEU

[BB08] Janet E. Burge and David C. Brown. Software Engineering Using RATionale. The Journal of Systems and Software, 81(3):395–413, March 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bracciali:2005:FAC

[BBC05] Andrea Bracciali, Antonio Brogi, and Carlos Canal. A formal approach to component adaptation. The Journal of Systems and Software, 74(1):45–54, January 1, 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bertoni:2008:PSI

[BBC^{+}08] Guido M. Bertoni, Luca Breveglieri, Liqun Chen, Pasqualina Fragneto, Keith A. Harrison, and Gerardo Pelosi.

A pairing SW implementation for Smart-Cards. The Journal of Systems and Software, 81(7):1240–1247, July 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Banavar:2004:TSS

[BBG^{+}04] Guruduth Banavar, Lawrence D.

Bergman, Yves Gaeremynck, Danny Soroker, and Jeremy Sussman. Tooling and system support for authoring multi-device applications. The Journal of Systems and Software, 69(3): 227–242, January 15, 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Benander:2000:EAD

[BBS00] A. C. Benander, B. A. Benander, and Janche Sang. An empirical analysis of debugging performance — differences between iterative and recursive constructs. The Journal of Systems and Software, 54(1): 17–28, September 30, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/53/24/
27/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/53/24/abstract.
html.

Beecher:2009:IED

[BCB09] Karl Beecher, Andrea Capiluppi, and Cornelia Boldyreff. Identifying exogenous drivers and evolutionary stages in FLOSS projects. The Journal of Systems and Software, 82(5):739–750, May 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Belli:2006:ISS

[BCDM06] Fevzi Belli, Kai-Yuan Cai, Raymond DeCarlo, and Aditya Mathur. Introduction to the special section on software cybernetics. The Journal of Systems and Software, 79(11):1483– 1485, November 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Barton:2004:UC

[BCF04] John J. Barton, Renato Cerqueira, and Marcus Fontoura. Ubiquitous computing. The Journal of Systems and Software, 69(3): 207, January 15, 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Brodnik:2005:WCC

[BCF^{+}05] Andrej Brodnik, Svante Carlsson, Michael L. Fredman, Johan Karlsson, and J. Ian Munro. Worst case constant time priority queue. The Journal of Systems and Software, 78(3):249–256, December 2005. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bae:2000:SVR

[BCK00] Hyun Seop Bae, In Sang Chung, and Yong Rae Kwon. Specifying and verifying realtime systems with timing uncertainty. The Journal of Systems and Software, 50(1):85–96, January 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/49/25/
32/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/49/25/abstract.
html.

Bleistein:2006:VSA

[BCV06] Steven J. Bleistein, Karl Cox, and June Verner. Validating strategic alignment of organizational IT requirements using goal modeling and problem diagrams. The Journal of Systems and Software, 79(3): 362–378, March 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bao:2005:RWH

[BCW05] Haiyong Bao, Zhenfu Cao, and Shengbao Wang. Remarks on Wu–Hsu’s threshold signature scheme using self-certified public keys. The Journal of Systems and Software, 78(1):56–59, October 2005. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bourque:2002:FPS

[BDA^{+}02] Pierre Bourque, Robert Dupuis, Alain Abran, James W. Moore, Leonard Tripp, and Sybille Wolff. Fundamental principles of software engineering — a journey. The Journal of Systems and Software, 62(1): 59–70, May 1, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Blundo:2004:HNP

[BDDG04] Carlo Blundo, Paolo D’Arco, Alfredo De Santis, and Clemente Galdi. Hyppocrates: a new proactive password checker. The Journal of Systems and Software, 71(1–2):163–175, April 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bhattacherjee:2001:HAO

[BDGR01] Anol Bhattacherjee, Kurt DeShazer, James H. Gerlach, and Bill Rierden. A hybrid approach to OO development: the SUMMI-Trak project at TCI. The Journal of Systems and Software, 57(2):91–98, June 15, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/64/29/26/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/64/
29/26/article.pdf.

Brukman:2008:SSA

[BDK08] Olga Brukman, Shlomi Dolev, and Elliot K. Kolodner. A self-stabilizing autonomic recoverer for eventual Byzantine software. The Journal of Systems and Software, 81(12):2315–2327, December 2008. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Baldwin:2003:QNA

[BDMK03] Rusty O. Baldwin, Nathaniel J.

Davis IV, Scott F. Midkiff, and John E. Kobza. Queueing network analysis: concepts, terminology, and methods. The Journal of Systems and Software, 66(2): 99–117, May 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Berry:2002:IIR

[Ber02] Daniel M. Berry. The importance of ignorance in requirements engineering: an earlier sighting and a revisitation. The Journal of Systems and Software, 60(1):83– 85, January 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        29/11/72/27/35/abstract.
                                        html.

Berglund:2003:DER

[Ber03]  Erik Berglund.  Design-

ing electronic reference documentation for software component libraries. The Journal of Systems and Software, 68 (1):65–75, October 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Binkley:2009:IDN

[BFLP09] David Binkley, Henry Feild, Dawn Lawrie, and Maurizio Pighin. Increasing diversity: Natural language measures for software fault prediction. The Journal of Systems and Software, 82(11): 1793–1803, November 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Blanco-Fernandez:2008:ESB

[BFPAGS^{+}08] Yolanda Blanco-Fernández, José J. Pazos-Arias, Alberto Gil-Solla, Manuel Ramos-Cabrer, Martín López-Nores, Jorge García-Duque, Ana Fernández-Vilas, and Rebeca P. Díaz-Redondo. Exploiting synergies between semantic reasoning and personalization strategies in intelligent recommender systems: a case study. The Journal of Systems and Software, 81(12):2371–2385, December 2008. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Briand:2004:UMA

[BFV04] Lionel C. Briand, Bernd Freimut, and Ferdinand Vollei. Using multiple adaptive regression splines to support decision making in code inspections. The Journal of Systems and Software, 73(2): 205–217, October 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Brass:2006:SES

[BG06] Stefan Brass and Christian Goldberg. Semantic errors in SQL queries: a quite complete list. The Journal of Systems and Software, 79(5): 630–644, May 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bagheri:2009:BTF

[BG09] Ebrahim Bagheri and Ali A.

Ghorbani. A belief-theoretic framework for the collaborative development and integration of para-consistent conceptual models. The Journal of Systems and Software, 82(4):707–729, April 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bischofs:2006:CED

[BGG^{+}06] Ludger Bischofs, Simon Giesecke, Michael Gottschalk, Wilhelm Hasselbring, Timo Warns, and Stefan Willer. Comparative evaluation of dependability characteristics for peer-to-peer architectural styles by simulation. The Journal of Systems and Software, 79(10):1419– 1432, October 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bernabe:2009:LWT

[BGG09] Gregorio Bernabé, Jose M. García, and José González. A lossy 3D wavelet transform for high-quality compression of medical video. The Journal of Systems and Software, 82(3):526–534, March 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Barber:2003:EDC

[BGH03] K. Suzanne Barber, Tom Graser, and Jim Holt. Evaluating dynamic correctness properties of domain reference architectures. The Journal of Systems and Software, 68(3):217–231, December 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Binkley:2008:ESR

[BGH^{+}08] David Binkley, Nicolas Gold, Mark Harman, Zheng Li, and Kiarash Mahdavi. An empirical study of the relationship between the concepts expressed in source code and dependence. The Journal of Systems and Software, 81(12):2287–2298, December 2008. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Baddoo:2002:MSP

[BH02] Nathan Baddoo and Tracy Hall. Motivators of software process improvement: an analysis of practitioners’ views. The Journal of Systems and Software, 62(2):85– 96, May 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Baddoo:2003:MSP

[BH03] Nathan Baddoo and Tracy Hall. De-motivators for software process improvement: an analysis of practitioners’ views. The Journal of Systems and Software, 66(1):23– 33, April 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Byun:2009:IMU

[BH09] Siwoo Byun and Moonhaeng Hur. An index management using CHCcluster for flash memory databases. The Journal of Systems and Software, 82(5): 825–835, May 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Baousis:2009:AMA

[BHAM09] Vasileios Baousis, Stathes Hadjiefthymiades, George Alyfantis, and Lazaros Merakos. Autonomous mobile agent routing for efficient server resource allocation. The Journal of Systems and Software, 82(5): 891–906, May 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Beecham:2005:UEP

[BHB^{+}05] Sarah Beecham, Tracy Hall, Carol Britton, Michaela Cottee, and Austen Rainer. Using an expert panel to validate a requirements process improvement model. The Journal of Systems and Software, 76(3):251–275, June 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bowden:2000:ESL

[BHL00] P. Bowden, M. Hargreaves, and C. S. Langensiepen. Estimation support by lexical analysis of requirements documents. The Journal of Systems and Software, 51 (2):87–98, April 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/50/25/
26/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/50/25/abstract.
html.

Bi:2002:XAL

[BHN02] Y. Bi, M. E. C. Hull, and P. N. Nicholl. An XML approach for legacy code reuse. The Journal of Systems and Software, 61(2):77– 89, March 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bai:2005:SFP

[BHXN05] C. G. Bai, Q. P. Hu, M. Xie, and S. H. Ng. Software failure prediction based on a Markov Bayesian network model. The Journal of Systems and Software, 74(3): 275–282, February 1, 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Biffl:2003:EDE

[Bif03] S. Biffl. Evaluating defect estimation models with major defects. The Journal of Systems and Software, 65(1):13– 29, January 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bass:2003:LUS

[BJ03] Len Bass and Bonnie E. John. Linking usability to software architecture patterns through general scenarios. The Journal of Systems and Software, 66(3): 187–197, June 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Beck:2006:PMP

[BJK06] Philip Beck, James J. Jiang, and Gary Klein. Prototyping mediators to project performance: Learning and interaction. The Journal of Systems and Software, 79(7): 1025–1035, July 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Barkaoui:2002:GE

[BJM02] K. Barkaoui, M. Jmaiel, and A. Mili. Guest editorial. The Journal of Systems and Software, 60(2):87– 89, February 1, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        29/11/72/28/27/abstract.
                                        html.

Brereton:2007:LAS

[BKB^{+}07] Pearl Brereton, Barbara A. Kitchenham, David Budgen, Mark Turner, and Mohamed Khalil. Lessons from applying the systematic literature review process within the software engineering domain. The Journal of Systems and Software, 80(4): 571–583, April 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Becker:2009:PCM

[BKR09] Steffen Becker, Heiko Koziolek, and Ralf Reussner. The Palladio component model for model-driven performance prediction. The Journal of Systems and Software, 82(1):3–22, January 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Babar:2006:ESG

[BKZ^{+}06] Muhammad Ali Babar, Barbara Kitchenham, Liming Zhu, Ian Gorton, and Ross Jeffery. An empirical study of groupware support for distributed software architecture evaluation process. The Journal of Systems and Software, 79(7): 912–925, July 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bosch:2003:SAX

[BL03] Jan Bosch and Lars Lundberg. Software architecture — engineering quality attributes. The Journal of Systems and Software, 66(3): 183–186, June 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Babar:2009:DDD

[BL09] Muhammad Ali Babar and Patricia Lago. Design decisions and design rationale in software architecture. The Journal of Systems and Software, 82(8):1195– 1197, August 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bengtsson:2004:ALM

[BLBvV04] PerOlof Bengtsson, Nico Lassing, Jan Bosch, and Hans van Vliet. Architecturelevel modifiability analysis (ALMA). The Journal of Systems and Software, 69(1– 2):129–147, January 1, 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Barroso:2002:TUD

[BLL02] A. M. Barroso, J. C. B. Leite, and O. G. Loques. Treating uncertainty in distributed scheduling. The Journal of Systems and Software, 63(2): 129–136, August 15, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Brayner:2008:ANA

[BLM^{+}08] Angelo Brayner, Aretusa Lopes, Diorgens Meira, Ricardo Vasconcelos, and Ronaldo Menezes. An adaptive in-network aggregation operator for query processing in wireless sensor networks. The Journal of Systems and Software, 81(3): 328–342, March 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Briand:2006:AIA

[BLOS06] L. C. Briand, Y. Labiche, L. O’Sullivan, and M. M. Sówka. Automated impact analysis of UML models. The Journal of Systems and Software, 79(3): 339–352, March 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Barrow:2000:IPS

[BM00a] Patrick D. M. Barrow and Pam J. Mayhew. Investigating principles of stakeholder evaluation in a modern IS development approach. The Journal of Systems and Software, 52(2–3):95–103, June 1, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/51/27/
26/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/51/27/abstract.
html.

Burd:2000:UEE

[BM00b] Elizabeth Burd and Malcolm Munro. Using evolution to evaluate reverse engineering technologies: mapping the process of software change. The Journal of Systems and Software, 53(1):43–51, July 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/29/11/52/25/
                                        30/article.pdf;   http:/
                                        /www.elsevier.nl/gej-ng/
                                        10/29/11/52/25/abstract.
                                        html.

Bagert:2005:DUW

[BM05] Donald J. Bagert and Susan A. Mengel. Developing and using a Web-based project process throughout the software engineering curriculum. The Journal of Systems and Software, 74(2): 113–120, January 15, 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bernardi:2007:PEU

[BM07] Simona Bernardi and José Merseguer. Performance evaluation of UML design with Stochastic Well-formed Nets. The Journal of Systems and Software, 80(11):1843– 1865, November 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Barrett:2004:ACB

[BMES04] Christopher L. Barrett, Madhav V. Marathe, D. Charles Engelhart, and Anand Sivasubramaniam. Approximating the connectivity between nodes when simulating largescale mobile ad hoc radio networks. The Journal of Systems and Software, 73(1):63– 74, September 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bani-Mohammad:2009:CEC

[BMOKAM09] S. Bani-Mohammad, M. Ould-Khaoua, I. Ababneh, and Lewis M. Mackenzie. Comparative evaluation of contiguous allocation strategies on 3D mesh multicomputers. The Journal of Systems and Software, 82(2): 307–318, February 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bellini:2009:EOR

[BNR09] P. Bellini, P. Nesi, and D. Rogai. Expressing and organizing real-time specification patterns via temporal logics. The Journal of Systems and Software, 82(2): 183–196, February 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Becker:2005:RST

[BNSG05] L. B. Becker, E. Nett, S. Schemmer, and M. Gergeleit. Robust scheduling in teamrobotics. The Journal of Systems and Software, 77 (1):3–16, July 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Baker:2005:ECG

[BNvdH05] Alex Baker, Emily Oh Navarro, and André van der Hoek. An experimental card game for teaching software engineering processes. The Journal of Systems and Software, 75(1–2):3–16, February 15, 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bass:2008:AAE

[BNW^{+}08] Len Bass, Robert Nord, William Wood, David Zubrow, and Ipek Ozkaya. Analysis of architecture evaluation data. The Journal of Systems and Software, 81(9): 1443–1455, September 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bozyigit:2000:HDD

[Boz00] M. Bozyigit. History-driven dynamic load balancing for recurring applications on networks of workstations. The Journal of Systems and Software, 51(1):61–72, April 1, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/29/11/50/26/

30/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/50/26/abstract.
html.

Baldwin:2006:UPC

[BPM06] Rusty Baldwin, Brian Peterson, and Robert Mills. Using playing cards to estimate interference in frequencyhopping spread spectrum radio networks. The Journal of Systems and Software, 79(12):1782–1788, December 2006. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Balbastre:2009:PSA

[BRC09] Patricia Balbastre, Ismael Ripoll, and Alfons Crespo. Period sensitivity analysis and D–P domain feasibility region in dynamic priority systems. The Journal of Systems and Software, 82(7): 1098–1111, July 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Baloukas:2009:OMD

[BRMA^{+}09] Christos Baloukas, Jose L.

Risco-Martin, David Atienza, Christophe Poucet, Lazaros Papadopoulos, Stylianos Mamagkakis, Dimitrios Soudris, J. Ignacio Hidalgo, Francky Catthoor, and Juan Lanchares. Optimization methodology of dynamic data structures based on genetic algorithms for multimedia embedded systems. The Journal of Systems and Software, 82 (4):590–602, April 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Barioni:2008:AM

[BRTT08] Maria Camila N. Barioni, Humberto L. Razente, Agma J. M. Traina, and Caetano Traina Jr. Accelerating kmedoid. The Journal of Systems and Software, 81(3): 343–355, March 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Barnett:2003:RVN

[BS03] Mike Barnett and Wolfram Schulte. Runtime verification of .NET contracts. The Journal of Systems and Software, 65(3):199–208, March 15, 2003. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Berander:2009:ETW

[BS09] Patrik Berander and Mikael Svahnberg. Evaluating two ways of calculating priorities in requirements hierarchies — an experiment on hierarchical cumulative voting. The Journal of Systems and Software, 82(5): 836–850, May 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Budgen:2003:CTE

[BT03] David Budgen and Mitchell Thomson. CASE tool evaluation: experiences from an empirical study. The Journal of Systems and Software, 67 (2):55–75, August 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Budgen:2005:SCM

[BT05] David Budgen and James E.

Tomayko. The SEI curriculum modules and their influence: Norm Gibbs’ legacy to software engineering education. The Journal of Systems and Software, 75(1–2):55–62, February 15, 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Bertoa:2006:MUS

[BTV06] Manuel F. Bertoa, José M.

Troya, and Antonio Vallecillo. Measuring the usability of software components. The Journal of Systems and Software, 79(3): 427–439, March 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Budgen:2000:EAS

[Bud00] David Budgen. Evaluation and assessment in software engineering. The Journal of Systems and Software, 52 (2–3):93–94, June 1, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/51/27/
25/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/51/27/abstract.
html.

Bruntink:2006:ESC

[BvD06] Magiel Bruntink and Arie van Deursen. An empirical study into class testability. The Journal of Systems and Software, 79(9): 1219–1232, September 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Babar:2007:EMT

[BVN07] Muhammad Ali Babar, June M. Verner, and Phong Thanh Nguyen. Establishing and maintaining trust in software outsourcing relationships: an empirical investigation. The Journal of Systems and Software, 80(9): 1438–1449, September 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Briand:2001:ISB

[BW01] Lionel C. Briand and Jürgen Wüst. Integrating scenariobased and measurementbased software product assessment. The Journal of Systems and Software, 59(1): 3–22, October 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/71/27/28/abstract.
html.

Briand:2000:ERB

[BWDP00] Lionel C. Briand, Jürgen Wüst, John W. Daly, and D. Victor Porter. Exploring the relationships between design measures and software quality in objectoriented systems. The Journal of Systems and Software, 51(3):245–273, May 1, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/50/28/
32/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/50/28/abstract.
html.

Braga:2006:OSM

[BWM06] Regina M. M. Braga, Cláudia M. L. Werner, and Marta Mattoso. Odyssey-search: a multi-agent system for component information search and retrieval. The Journal of Systems and Software, 79(2): 204–215, February 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cam:2000:LRP

[Çam00a] Hasan Çam. List ranking on processor arrays. The Journal of Systems and Software, 55(2):185–192, December 27, 2000. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        29/11/54/24/30/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/29/11/54/
                                        24/30/article.pdf.

Cam:2000:LSP

[Çam00b] Hasan Çam. An on-line scheduling policy for IRIS real-time composite tasks. The Journal of Systems and Software, 52(1):25–32, May 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/29/11/51/25/
                                        28/article.pdf;   http:/
                                        /www.elsevier.nl/gej-ng/
                                        10/29/11/51/25/abstract.
                                        html.

Card:2002:ECS

[Car02] David N. Card. Editor’s corner: Stepping up to the plate. The Journal of Systems and Software, 60(1): 3, January 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        29/11/72/27/28/abstract.
                                        html.

Card:2004:EC

[Car04]  D. Card.  Editor’s cor-

ner. The Journal of Systems and Software, 70(1–2): 1–2, February 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Card:2008:EOE

[Car08]  David N. Card.  Edito-

rial — outgoing Editor-in-Chief. The Journal of Systems and Software, 81 (3):v, March 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Charreteur:2009:MDM

[CBG09] Florence Charreteur, Bernard Botella, and Arnaud Gotlieb. Modelling dynamic memory management in constraintbased testing. The Journal of Systems and Software, 82(11):1755–1766, November 2009. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chung:2002:EBD

[CBK02] Yon Dohn Chung, Su Ho Bang, and Myoung Ho Kim. An efficient broadcast data clustering method for multipoint queries in wireless information systems. The Journal of Systems and Software, 64(3):173–181, December 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Clarke:2008:ACC

[CBKK08] Peter J. Clarke, Djuradj Babich, Tariq M. King, and B. M. Golam Kibria. Analyzing clusters of class characteristics in OO applications. The Journal of Systems and Software, 81(12): 2269–2286, December 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cioch:2000:ISA

[CBS00] Frank A. Cioch, John M. Brabbs, and Larry Sieh. The impact of software architecture reuse on development processes and standards. The Journal of Systems and Software, 50(3): 221–236, March 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/49/28/
29/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/49/28/abstract.
html.

Cox:2007:RIW

[CBVD07] Karl A. Cox, Steven J. Bleistein, June M. Verner, and Alan M. Davis. REBNITA’05 — 1st International Workshop on Requirements Engineering for Business Need and IT Alignment. The Journal of Systems and Software, 80(3):283–284, March 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cao:2000:DES

[CBZ00] Jiannong Cao, Graeme Bennett, and Kang Zhang. Direct execution simulation of load balancing algorithms with real workload distribution. The Journal of Systems and Software, 54(3):227–237, November 1, 2000. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/53/28/30/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/53/
28/30/article.pdf.

Chen:2001:FTG

[CC01] Ching-Wen Chen and Chung-Ping Chung. Fault-tolerant gamma interconnection network without backtracking. The Journal of Systems and Software, 58(1):23– 31, August 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/68/27/29/abstract.
html.

Cai:2002:ASS

[CC02a] Kai-Yuan Cai and Lei Chen.

Analyzing software science data with partial repeatability. The Journal of Systems and Software, 63(3):173–186, September 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chan:2002:SLI

[CC02b] Chi-Kwong Chan and L. M. Cheng. Security of Lin’s image watermarking system. The Journal of Systems and Software, 62(3): 211–215, June 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:2003:PEA

[CC03] Da-Wei Chang and Ruei-Chuan Chang. OS Portal: an economic approach for making an embedded kernel extensible. The Journal of Systems and Software, 67(1):19– 30, July 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chan:2004:IRS

[CC04] Yung-Kuan Chan and Chih-Ya Chen. Image retrieval system based on colorcomplexity and color-spatial features. The Journal of Systems and Software, 71(1– 2):65–70, April 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chou:2005:IFC

[CC05] Shih-Chien Chou and Chin-Yi Chang. An information flow control model for C applications based on access control lists. The Journal of Systems and Software, 78(1): 84–100, October 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chou:2006:MRR

[CC06] Shih-Chien Chou and Yuan-Chien Chen. Managing role relationships in an information flow control model. The Journal of Systems and Software, 79(4):507– 522, April 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:2007:DPS

[CC07] Ching-Pao Chang and Chih-Ping Chu. Defect prevention in software processes: an action-based approach. The Journal of Systems and Software, 80(4): 559–570, April 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cai:2008:ART

[CC08a] Kai-Yuan Cai and David Card. An analysis of research topics in software engineering — 2006. The Journal of Systems and Software, 81(6): 1051–1058, June 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:2008:ABM

[CC08b] Yaw-Chung Chen and Wen-Yen Chen. An agentbased metric for quality of services over wireless networks. The Journal of Systems and Software, 81(10): 1625–1639, October 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chow:2008:SSC

[CC08c] Tsun Chow and Dac-Buu Cao. A survey study of critical success factors in agile software projects. The Journal of Systems and Software, 81(6):961–971, June 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cao:2009:IBU

[CC09a] Feng Cao and Zhenfu Cao. An identity based universal designated verifier signature scheme secure in the standard model. The Journal of Systems and Software, 82(4): 643–649, April 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:2009:HAM

[CC09b] Chung-Yang Chen and Pei-Chi Chen. A holistic approach to managing software change impact. The Journal of Systems and Software, 82(12):2051–2067, December 2009. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cao:2005:WAW

[CCC05] Jiannong Cao, Catherine Chan, and Keith Chan. Workflow analysis for Web publishing using a stageactivity process model. The Journal of Systems and Software, 76(3):221–235, June 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cheung:2006:PNB

[CCC06] K. S. Cheung, T. Y. Cheung, and K. O. Chow. A petrinet-based synthesis methodology for use-case-driven system design. The Journal of Systems and Software, 79(6): 772–790, June 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chan:2006:AGO

[CCCT06] W. K. Chan, M. Y. Cheng, S. C. Cheung, and T. H. Tse. Automatic goal-oriented classification of failure behaviors for testing XMLbased multimedia software applications: an experimental case study. The Journal of Systems and Software, 79 (5):602–612, May 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Clark:2004:SCL

[CCD^{+}04] Justin Clark, Chris Clarke, Stefano De Panfilis, Giampiero Granatella, Paolo Predonzani, Alberto Sillitti, Giancarlo Succi, and Tullio Vernazza. Selecting components in large COTS repositories. The Journal of Systems and Software, 73(2): 323–331, October 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Canfora:2000:DLP

[CCDD00] Gerardo Canfora, Aniello Cimitile, Andrea De Lucia, and Giuseppe A. Di Lucca. Decomposing legacy programs: a first step towards migrating to client-server platforms. The Journal of Systems and Software, 54(2): 99–110, October 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/53/25/
26/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/53/25/abstract.
html.

Chen:2004:PEW

[CCF^{+}04] Shyh-Kwei Chen, Trieu C. Chieu, Shiwa S. Fu, Yew-Huey Liu, Florian Pinel, and Jih-Shyr Yih. A practical experience in workspace separation for developing multiple storefronts on customized commerce engines. The Journal of Systems and Software, 73(3):415–424, November/December 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chalmeta:2001:RAE

[CCG01] Ricardo Chalmeta, Christina Campos, and Reyes Grangel. References architectures for enterprise integration. The Journal of Systems and Software, 57(3):175–191, July 15, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/64/32/27/abstract.
html.

Canfora:2007:EPP

[CCG^{+}07] Gerardo Canfora, Aniello Cimitile, Felix Garcia, Mario Piattini, and Corrado Aaron Visaggio. Evaluating performances of pair designing in industry. The Journal of Systems and Software, 80 (8):1317–1327, August 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:2009:SOC

[CCH09] Chin-Chen Chang, Yung-Chen Chou, and Yi-Pei Hsieh. Search-order coding method with indicatorelimination property. The Journal of Systems and Software, 82(3):516–525, March 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chan:2009:PPC

[CCHT09] W. K. Chan, S. C. Cheung, Jeffrey C. F. Ho, and T. H. Tse. PAT: a pattern classification approach to automatic reference oracles for the testing of mesh simplification programs. The Journal of Systems and Software, 82(3):422–434, March 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:2009:EHR

[CCHW09] Chien-Tsun Chen, Yu Chin Cheng, Chin-Yun Hsieh, and I-Lang Wu. Exception handling refactorings: Directed by goals and driven by bug fixing. The Journal of Systems and Software, 82(2): 333–345, February 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:2002:VRR

[CCK02] Deng-Jyi Chen, Wu-Chi Chen, and Krishna M. Kavi. Visual requirement representation. The Journal of Systems and Software, 61(2): 129–143, March 15, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:2009:APA

[CCKM09] Ya-Shu Chen, Li-Pin Chang, Tei-Wei Kuo, and Aloysius K. Mok. An anomaly prevention approach for realtime task scheduling. The Journal of Systems and Software, 82(1):144–154, January 2009. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chan:2001:EFM

[CCL01] Edward Chan, Daoxu Chen, and Victor C. S. Lee. Effectiveness of the FDDI-M protocol in supporting synchronous traffic. The Journal of Systems and Software, 56(1):51–62, February 1, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/57/24/28/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/57/
24/28/article.pdf.

Chen:2005:CCT

[CCP05] Jeanne Chen, Tung-Shou Chen, and Yu-Mei Pan. A crisscross checking technique for tamper detection in halftone images. The Journal of Systems and Software, 76 (3):287–295, June 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:2001:RGS

[CCSC01] Hsung-Pin Chang, Ray-I Chang, Wei-Kuan Shih, and Ruei-Chuan Chang. Reschedulable-group-SCAN scheme for mixed real-time/ non-real-time disk scheduling in a multimedia system. The Journal of Systems and Software, 59(2):143–152, November 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/71/28/30/abstract.
html.

Chang:2007:GGS

[CCSC07] Hsung-Pin Chang, Ray-I Chang, Wei-Kuan Shih, and Ruei-Chuan Chang. GSR: a global seek-optimizing realtime disk-scheduling algorithm. The Journal of Systems and Software, 80(2): 198–215, February 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cai:2001:NNA

[CCW^{+}01] Kai-Yuan Cai, Lin Cai, Wei-Dong Wang, Zhou-Yi Yu, and David Zhang. On the neural network approach in software reliability modeling. The Journal of Systems and Software, 58(1):47– 62, August 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.

elsevier.com/gej-ng/10/
29/11/68/27/31/abstract.
html.

Chang:2002:PDA

[CCW02a] Weng-Long Chang, Chih-Ping Chu, and Jia-Hwa Wu. A precise dependence analysis for multi-dimensional arrays under specific dependence direction. The Journal of Systems and Software, 63 (2):99–112, August 15, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:2002:SGA

[CCW02b] Weng-Long Chang, Chih-Ping Chu, and Jia-Hwa Wu. A simple and general approach to parallelize loops with arbitrary control flow and uniform data dependence distances. The Journal of Systems and Software, 63 (2):91–98, August 15, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chan:2009:HBR

[CCY^{+}09] Yung-Kuan Chan, Wen-Tang Chen, Shyr-Shen Yu, Yu-An Ho, Chwei-Shyong Tsai, and Yen-Ping Chu. A HDWTbased reversible data hiding method. The Journal of Systems and Software, 82(3): 411–421, March 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:2000:EDS

[CD00] J. Morris Chang and Charles H. Daugherty. An efficient data structure for dynamic memory management. The Journal of Systems and Software, 54(3):219–226, November 1, 2000. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        29/11/53/28/29/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/29/11/53/
                                        28/29/article.pdf.

Cook:2005:DTI

[CD05] Jonathan E. Cook and Zhidian Du. Discovering thread interactions in a concurrent system. The Journal of Systems and Software, 77(3): 285–297, September 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chiang:2007:VMD

[CD07] Tsung-Hsi Chiang and Lan-Rong Dung. Verification method of dataflow algorithms in high-level synthesis. The Journal of Systems and Software, 80(8):1256– 1270, August 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chai:2009:SOA

[CDC09] Yunpeng Chai, Zhihui Du, and Yinong Chen. A stepwise optimization algorithm of clustered streaming media servers. The Journal of Systems and Software, 82 (8):1344–1361, August 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Canfora:2008:FQA

[CDEV08] Gerardo Canfora, Massimiliano Di Penta, Raffaele Esposito, and Maria Luisa Villani. A framework for QoS-aware binding and rebinding of composite Web services. The Journal of Systems and Software, 81(10): 1754–1769, October 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Caporuscio:2007:MBS

[CDI07] Mauro Caporuscio, Antinisca Di Marco, and Paola Inverardi. Model-based system reconfiguration for dynamic performance management. The Journal of Systems and Software, 80(4): 455–473, April 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Costa:2007:ESP

[CdOBT07] Hélio R. Costa, Marcio de O. Barros, and Guilherme H. Travassos. Evaluating software project portfolio risks. The Journal of Systems and Software, 80(1):

16–31, January 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cimato:2005:OOJ

[CDP05] S. Cimato, A. De Santis, and U. Ferraro Petrillo. Overcoming the obfuscation of Java programs by identifier renaming. The Journal of Systems and Software, 78(1): 60–72, October 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chan:2002:AMA

[CDS02] Stephen Chan, Tharam Dillon, and Andrew Siu. Applying a mediator architecture employing XML to retailing inventory control. The Journal of Systems and Software, 60(3):239–248, February 15, 2002. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        29/11/72/29/34/abstract.
                                        html.

Castiglione:2007:TAD

[CDS07] A. Castiglione, A. De Santis, and C. Soriente. Taking advantages of a disadvantage: Digital forensics and steganography using document metadata. The Journal of Systems and Software, 80 (5):750–764, May 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cavanaugh:2007:GEI

[CDW07] Charles Cavanaugh, Frank Drews, and Lonnie Welch. Guest Editor’s introduction. The Journal of Systems and Software, 80(7): 935–937, July 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Corbin:2007:TTK

[CDZ07] Richard D. Corbin, Christopher B. Dunbar, and Qiuming Zhu. A threetier knowledge management scheme for software engineering support and innovation. The Journal of Systems and Software, 80(9): 1494–1505, September 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cobb:2008:WPC

[CE08] Jake Cobb and Hala ElAarag.

Web proxy cache replacement scheme based on backpropagation neural network. The Journal of Systems and Software, 81(9):1539– 1558, September 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chaari:2007:CAM

[CELS07] Tarak Chaari, Dejene Ejigu, Frédérique Laforest, and Vasile-Marian Scuturici. A comprehensive approach to model and use context for adapting applications in pervasive environments. The Journal of Systems and Software, 80(12):1973–1992, December 2007. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:2008:DAD

[CET^{+}08] Dan Chen, Roland Ewald, Georgios K. Theodoropoulos, Robert Minson, Ton Oguara, Michael Lees, Brian Logan, and Adelinde M. Uhrmacher. Data access in distributed simulations of multi-agent systems. The Journal of Systems and Software, 81(12): 2345–2360, December 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cerri:2007:OSO

[CF07] Davide Cerri and Alfonso Fuggetta. Open standards, open formats, and open source. The Journal of Systems and Software, 80(11): 1930–1937, November 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cooper:2008:E

[CF08] Kendra Cooper and Xavier Franch. Editorial. The Journal of Systems and Software, 81(6):841–842, June 2008.

CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Canfora:2008:WAM

[CFFT08] Gerardo Canfora, Anna Rita Fasolino, Gianni Frattolillo, and Porfirio Tramontana. A wrapping approach for migrating legacy system interactive functionalities to Service Oriented Architectures. The Journal of Systems and Software, 81(4): 463–480, April 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cicirelli:2007:EAM

[CFN07] Franco Cicirelli, Angelo Furfaro, and Libero Nigro. Exploiting agents for modelling and simulation of coverage control protocols in large sensor networks. The Journal of Systems and Software, 80(11):1817–1832, November 2007. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chalmeta:2003:AEV

[CG03] Ricardo Chalmeta and Reyes Grangel. ARDIN extension for virtual enterprise integration. The Journal of Systems and Software, 67(3):141–152, September 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Coskun:2005:SCI

[CG05] Erman Coskun and Martha Grabowski. Software complexity and its impacts in embedded intelligent real-time systems. The Journal of Systems and Software, 78 (2):128–145, November 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cuadrado-Gallego:2008:SIS

[CGA08] Juan J. Cuadrado-Gallego and Alain Abran. Special issue on software process and product measurement. The Journal of Systems and Software, 81(5): 601–602, May 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cai:2007:AST

[CGHL07] Kai-Yuan Cai, Bo Gu, Hai Hu, and Yong-Chao Li. Adaptive software testing with fixed-memory feedback. The Journal of Systems and Software, 80(8):1328– 1348, August 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cesario:2004:OBH

[CGL^{+}04] W. Cesário, L. Gauthier, D. Lyonnard, G. Nicolescu, and A. A. Jerraya. Object-based hardware/software component interconnection model for interface design in system-on-a-chip circuits. The Journal of Systems and Software, 70(3):229–244, March 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cuadrado-Gallego:2008:CBI

[CGMPAP08] Juan J. Cuadrado-Gallego, Fernando Machado-Piriz, and Javier Aroba-Páez. On the conversion between IF-PUG and COSMIC software functional size units: a theoretical and empirical study. The Journal of Systems and Software, 81(5): 661–672, May 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Canfora:2005:FEV

[CGP^{+}05] G. Canfora, F. García, M. Piattini, F. Ruiz, and C. A. Visaggio. A family of experiments to validate metrics for software process models. The Journal of Systems and Software, 77(2): 113–129, August 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Capozucca:2009:FDI

[CGP^{+}09] Alfredo Capozucca, Nicolas Guelfi, Patrizio Pelliccione, Alexander Romanovsky, and Avelino F. Zorzo. Frameworks for designing and implementing dependable systems using Coordinated Atomic Actions: a comparative study. The Journal of Systems and Software, 82(2): 207–228, February 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cuadrado-Gallego:2006:ESP

[CGSGR06] Juan J. Cuadrado-Gallego, Miguel-Ángel Sicilia, Miguel Garre, and Daniel Rodríguez. An empirical study of process-related attributes in segmented software costestimation relationships. The Journal of Systems and Software, 79(3):353–361, March 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Colvin:2008:TBT

[CGW08] Robert Colvin, Lars Grunske, and Kirsten Winter. Timed behavior trees for failure mode and effects analysis of time-critical systems. The Journal of Systems and Software, 81(12):2163–2182, December 2008. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:2005:AAT

[CH05] Ing-Chau Chang and Sheng-Wen Hsieh. ATF: an Adaptive Three-layer Framework for inter-stream synchronization of SMIL multimedia presentations. The Journal of Systems and Software, 75(3): 283–303, March 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:2007:DIA

[CH07a] Tao-Ku Chang and Gwan-Hwan Hwang. The design and implementation of an application program interface for securing XML documents. The Journal of Systems and Software, 80 (8):1362–1374, August 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chiu:2007:AAB

[CH07b] Nan-Hsing Chiu and Sun-Jen Huang. The adjusted analogy-based software effort estimation based on similarity distances. The Journal of Systems and Software, 80 (4):628–640, April 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:2009:EAI

[CH09] Jie-Cherng Chen and Sun-Jen Huang. An empirical analysis of the impact of software development problem factors on software maintainability. The Journal of Systems and Software, 82(6): 981–992, June 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chalmeta:2006:MCR

[Cha06] Ricardo Chalmeta. Methodology for customer relationship management. The Journal of Systems and Software, 79(7):1015–1024, July 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:2009:I

[Cha09] Carl K. Chang. Introduction. The Journal of Systems and Software, 82(10):1565– 1567, October 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:2001:NEA

[CHC01] Chin-Chen Chang, Min-Shian Hwang, and Tung-Shou Chen. A new encryption algorithm for image cryptosystems. The Journal of Systems and Software, 58(2):83–91, September 1, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        29/11/68/33/27/abstract.
                                        html.

Chu:2004:IWB

[CHL04] Slo-Li Chu, Tsung-Chuan Huang, and Lan-Chi Lee. Improving workload balance and code optimization on processor-in-memory systems. The Journal of Systems and Software, 71(1– 2):71–82, April 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chou:2005:DPD

[CHL05] Shih-Chien Chou, Wei-Chuan Hsu, and Wei-Kuang Lo. DPE/PAC: decentralized process engine with product access control. The Journal of Systems and Software, 76 (3):207–219, June 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chung:2008:BAE

[CHL^{+}08] Yu Fang Chung, Kuo Hsuan Huang, Hsiu Hui Lee, Feipei Lai, and Tzer Shyong Chen. Bidder-anonymous English auction scheme with privacy and public verifiability. The Journal of Systems and Software, 81(1): 113–119, January 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chou:2004:ERB

[Cho04a]  Shih-Chien Chou.  Em-

bedding role-based access control model in objectoriented systems to protect privacy. The Journal of Systems and Software, 71(1–2): 143–161, April 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chou:2004:PFA

[Cho04b] Shih-Chien Chou. Providing flexible access control to an information flow control model. The Journal of Systems and Software, 73(3):425–439, November/December 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chou:2005:ABI

[Cho05] Shih-Chien Chou. An agentbased inter-application information flow control model. The Journal of Systems and Software, 75(1–2):179–187, February 15, 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Crnkovic:2007:GE

[CHS^{+}07] Ivica Crnkovic, George T. Heineman, Heinz W. Schmidt, Judith Stafford, and Kurt Wallnau. Guest editorial. The Journal of Systems and Software, 80(5): 641–642, May 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chow:2005:GAS

[CHY^{+}05] Sherman S. M. Chow, Lucas C. K. Hui, S. M. Yiu, K. P. Chow, and Richard W. C. Lui. A generic anti-spyware solution by access control list at kernel level. The Journal of Systems and Software, 75(1–2):227– 234, February 15, 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cheung:2003:SIT

[CHZY03] B. Cheung, L. Hui, J. Zhang, and S. M. Yiu. SmartTutor: An intelligent tutoring system in Web-based adult education. The Journal of Systems and Software, 68(1):11– 25, October 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chien:2003:HAP

[CJ03] Hung-Yu Chien and Jinn-Ke Jan. A hybrid authentication protocol for large mobile network. The Journal of Systems and Software, 67(2): 123–130, August 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:2005:IES

[CJ05] Wen-Kui Chang and Shuen-Lin Jeng. Impartial evaluation in software reliability practice. The Journal of Systems and Software, 76 (2):99–110, May 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:2009:CSN

[CJ09] Yaw-Chung Chen and Wen-Kang Jia. Challenge and solutions of NAT traversal for ubiquitous and pervasive applications on the Internet. The Journal of Systems and Software, 82(10):1620– 1626, October 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cai:2008:ESA

[CJHB08] Kai-Yuan Cai, Chang-Hai Jiang, Hai Hu, and Cheng-Gang Bai. An experimental study of adaptive testing for software reliability assessment. The Journal of Systems and Software, 81 (8):1406–1429, August 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:2009:RSR

[CJKC09] Houn-Gee Chen, James J. Jiang, Gary Klein, and Jengchung V. Chen. Reducing software requirement perception gaps through coordination mechanisms. The Journal of Systems and Software, 82(4):650–655, April 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chien:2001:MRL

[CJT01] Hung-Yu Chien, Jinn-Ke Jan, and Yuh-Min Tseng.

A modified remote login authentication scheme based on geometric approach. The Journal of Systems and Software, 55(3):287–290, January 15, 2001. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/54/27/29/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/54/
27/29/article.pdf.

Chien:2004:SIS

[CJT04] Hung-Yu Chien, Jinn-Ke Jan, and Yuh-Min Tseng. Solving the invalid signerverified signature problem and comments on Xia–You group signature. The Journal of Systems and Software, 73(3):369–373, November/December 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:2004:RSA

[CJZ04] Changgui Chen, Weijia Jia, and Wanlei Zhou. A reactive system architecture for building fault-tolerant distributed applications. The Journal of Systems and Software, 72(3): 401–415, August 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:2000:IRS

[CK00a] Pei-Min Chen and Fong-Chou Kuo. An information retrieval system based on a user profile. The Journal of Systems and Software, 54 (1):3–8, September 30, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/53/24/
25/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/53/24/abstract.
html.

Chung:2000:IRS

[CK00b] Yon Dohn Chung and Myoung Ho Kim. An index replication scheme for wireless data broadcasting. The Journal of Systems and Software, 51(3):191–199, May 1, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/50/28/
28/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/50/28/abstract.
html.

Chung:2002:TPO

[CK02a] Tae-Sun Chung and Hyoung-Joo Kim. A two phase optimization technique for XML queries with multiple regular path expressions. The Journal of Systems and Software, 64(3):183–193, December 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chung:2002:XQP

[CK02b] Tae-Sun Chung and Hyoung-Joo Kim. XML query processing using document type definitions. The Journal of Systems and Software, 64(3): 195–205, December 15, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:2008:DTC

[CKL08] Tsong Yueh Chen, Fei-Ching Kuo, and Huai Liu. Distributing test cases more evenly in adaptive random testing. The Journal of Systems and Software, 81(12): 2146–2162, December 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:2009:ART

[CKL09] Tsong Yueh Chen, Fei-Ching Kuo, and Huai Liu. Adaptive random testing based on distribution metrics. The Journal of Systems and Software, 82(9): 1419–1433, September 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:2006:SPT

[CKM06] Tsong Yueh Chen, Fei-Ching Kuo, and Robert Merkel. On the statistical properties of testing effectiveness measures. The Journal of Systems and Software, 79(5):

591–601, May 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Crnkovic:2002:CCB

[CL02] Ivica Crnkovic and Magnus Larsson. Challenges of component-based development. The Journal of Systems and Software, 61(3): 201–212, April 1, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:2004:CSI

[CL04a] Cheng-Wei Chen and Jenq Kuen Lee. Case study: an infrastructure for C/ATLAS environments with objectoriented design and XML representation. The Journal of Systems and Software, 71 (1–2):83–95, April 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chunlin:2004:AFS

[CL04b] Li Chunlin and Li Layuan. Agent framework to support the computational grid. The Journal of Systems and Software, 70(1–2):177– 187, February 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:2006:DEO

[CL06a] Chin-Chen Chang and Tzu-Chuen Lu. A difference expansion oriented data hiding scheme for restoring the original host images. The Journal of Systems and Software, 79(12):1754–1766, December 2006. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:2006:RID

[CL06b] Chin-Chen Chang and Tzu-Chuen Lu. Reversible indexdomain information hiding scheme based on side-match vector quantization. The Journal of Systems and Software, 79(8):1120–1129, August 2006. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:2008:AWM

[CL08] Chin-Chen Chang and Pei-Yu Lin. Adaptive watermark mechanism for rightful ownership protection. The Journal of Systems and Software, 81(7):1118–1129, July 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chirinos:2005:CDM

[CLB05] L. Chirinos, F. Losavio, and J. Bøegh. Characterizing a data model for software measurement. The Journal of Systems and Software, 74(2): 207–226, January 15, 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:2003:DGI

[CLC03] Ching-Wen Chen, Neng-Pin Lu, and Chung-Ping Chung. 3-disjoint gamma interconnection networks. The Journal of Systems and Software, 66(2):129–134, May 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:2008:XBA

[CLC08a] Bo Chen, David D. Linz, and Harry H. Cheng. XMLbased agent communication, migration and computation in mobile agent systems. The Journal of Systems and Software, 81(8):1364– 1376, August 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:2008:RCE

[CLC08b] Tzung-Her Chen, Wei-Bin Lee, and Hsing-Bai Chen. A round- and computationefficient three-party authenticated key exchange protocol. The Journal of Systems and Software, 81(9): 1581–1590, September 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:2004:PII

[CLCY04] Yue-Shan Chang, Kai-Chih Liang, Ming-Chun Cheng, and Shyan-Ming Yuan. Prototyping an integrated information gathering system on CORBA. The Journal of Systems and Software, 72(2): 281–294, July 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chiang:2008:DIE

[CLG08] Mei-Ling Chiang, Yu-Chen Lin, and Lian-Feng Guo. Design and implementation of an efficient Web cluster with content-based request distribution and file caching. The Journal of Systems and Software, 81(11): 2044–2058, November 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:2005:PPC

[CLGL05] Shiping Chen, Yan Liu, Ian Gorton, and Anna Liu. Performance prediction of component-based applications. The Journal of Systems and Software, 74(1):35– 43, January 1, 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:2007:TEM

[CLH07] Chin-Chen Chang, Yu-Chiang Li, and Wen-Hung Huang. TFRP: an efficient microaggregation algorithm for statistical disclosure control. The Journal of Systems and Software, 80(11):1866– 1878, November 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chan:2005:OOS

[CLL05] Ben C. B. Chan, John C. F. Lau, and John C. S. Lui. OPERA: an opensource extensible router architecture for adding new network services and protocols. The Journal of Systems and Software, 78(1): 24–36, October 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:2001:SAB

[CLSa01] J. Morris Chang, Woo Hyong Lee, and Witawas Srisaan. A study of the allocation behavior of C++ programs. The Journal of Systems and Software, 57(2): 107–118, June 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        29/11/64/29/28/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/29/11/64/
                                        29/28/article.pdf.

Chou:2005:PIL

[CLW05] Shih-Chien Chou, An-Feng Liu, and Chien-Jung Wu. Preventing information leakage within workflows that execute among competing organizations. The Journal of Systems and Software, 75(1–2):109–123, February 15, 2005. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cao:2004:DIR

[CLX^{+}04] J. Cao, Y. Liu, Li Xie, B. Mao, and K. Zhang. The design and implementation of a runtime system for graph-oriented parallel and distributed programming. The Journal of Systems and Software, 72(3): 389–399, August 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Choi:2005:LML

[CM05] Yonggoo Choi and Songchun Moon. Lightweight multigranularity locking for transaction management in XML database systems. The Journal of Systems and Software, 78(1):37–46, October 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Choi:2004:CMS

[CMC04] Yong-Jin Choi, Jun-Ki Min, and Chin-Wan Chung. A cost model for spatio-temporal queries using the TPRtree. The Journal of Systems and Software, 73(1): 101–112, September 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cucurull:2009:FMA

[CMNA^{+}09] J. Cucurull, R. Martí, G. Navarro-Arribas, S. Robles, and J. Borrell. Full mobile agent interoperability in an IEEE-FIPA context. The Journal of Systems and Software, 82(12):1927– 1940, December 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ciminiera:2004:IIS

[CMS04] L. Ciminiera, P. Maggi, and R. Sisto. Implementing innovative services supporting user and terminal mobility: the SCARAB architecture. The Journal of Systems and Software, 72(3): 367–376, August 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Castello:2002:VTS

[CMT02] R. Castelló, R. Mili, and I. G. Tollis. ViSta: a tool suite for the visualization of behavioral requirements. The Journal of Systems and Software, 62(3): 141–159, June 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Counsell:2000:UFC

[CN00] S. Counsell and P. Newson.

Use of friends in C++ software: an empirical investigation. The Journal of Systems and Software, 53 (1):15–21, July 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/52/25/
27/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/52/25/abstract.
html.

Campos:2004:PCC

[CN04] Alvaro E. Campos and Juan E. Navarro. A pagecoherent, causally consistent protocol for distributed shared memory. The Journal of Systems and Software, 72(3):305–319, August 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Correia:2007:WIW

[CNLV07] Miguel Correia, Nuno Ferreira Neves, Lau Cheuk Lung, and Paulo Veríssimo. Worm-IT — a wormholebased intrusion-tolerant group communication system. The Journal of Systems and Software, 80(2):178–197, February 2007. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Coleman:2008:ISP

[CO08] Gerry Coleman and Rory O’Connor. Investigating software process in practice: a grounded theory perspective. The Journal of Systems and Software, 81(5): 772–784, May 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cowling:2005:RMS

[Cow05] A. J. Cowling. The role of modelling in the software engineering curriculum. The Journal of Systems and Software, 75(1–2): 41–53, February 15, 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cox:2007:PEE

[CP07] Karl Cox and Keith T. Phalp. Practical experience of eliciting classes from use case descriptions. The Journal of Systems and Software, 80(8):1286–1304, August 2007. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:2009:UPF

[CP09] Matthew Chang and Chung Keung Poon. Using phrases as features in email classification. The Journal of Systems and Software, 82(6): 1036–1045, June 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Choi:2009:SAB

[CPiLH09] Seokwoo Choi, Heewan Park, Hyun il Lim, and Taisook Han. A static API birthmark for Windows binary executables. The Journal of Systems and Software, 82(5): 862–873, May 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cho:2004:CBL

[CPL^{+}04] Hyeon-Gyu Cho, Yong-Mun Park, Chan-Seob Lee, Dong-Hyuk Kim, Joo-Hyun Baek, and Eui-In Choi. A clientbased logging technique using backward analysis of log in client/server environment. The Journal of Systems and Software, 72(3): 455–466, August 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Campo:2005:CPT

[CPT05] Marcelo R. Campo, J. Andrés Díaz Pace, and Federico U. Trilnik. “Computer, please, tell me what I have to do… ”: an approach to agent-aided application composition. The Journal of Systems and Software, 74(1):55– 64, January 1, 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cassez:2006:STT

[CR06] Franck Cassez and Olivier H. Roux. Structural translation from Time Petri Nets to Timed Automata. The Journal of Systems and Software, 79(10):1456–1468, October 2006. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Choi:2001:MSS

[CS01] S. James Choi and Walt Scacchi. Modeling and simulating software acquisition process architectures. The Journal of Systems and Software, 59(3):343–354, December 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        29/11/71/34/38/abstract.
                                        html.

Chung:2004:AAG

[CS04] Lawrence Chung and Nary Subramanian. Adaptable architecture generation for embedded systems. The Journal of Systems and Software, 71 (3):271–295, May 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:2002:DDM

[CSaLG02] J. Morris Chang, Witawas Srisa-an, Chia-Tien Dan Lo, and Edward F. Gehringer. DMMX: Dynamic memory management extensions. The Journal of Systems and Software, 63(3):187–199, September 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:2005:ARC

[CSGL05] Shu-Ching Chen, Mei-Ling Shyu, Irina Gray, and Hongli Luo. An adaptive ratecontrol streaming mechanism with optimal buffer utilization. The Journal of Systems and Software, 75(3): 271–282, March 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Carrington:2005:IUC

[CSNS05] David Carrington, Paul Strooper, Sharron Newby, and Terry Stevenson. An industry/university collaboration to upgrade software engineering knowledge and skills in industry. The Journal of Systems and Software, 75(1–2):29–39, February 15, 2005. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Crnkovic:2003:GE

[CSSW03] Ivica Crnkovic, Heinz Schmidt, Judith Stafford, and Kurt Wallnau. Guest editorial. The Journal of Systems and Software, 65(3): 169–171, March 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Crnkovic:2005:ACB

[CSSW05] Ivica Crnkovic, Heinz W. Schmidt, Judith Stafford, and Kurt Wallnau. Automated Component-Based Software Engineering. The Journal of Systems and Software, 74(1):1–3, January 1, 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chin:2000:THP

[CT00] Charles Changli Chin and Shang-Rong Tsai. TOFF-2: a high-performance faulttolerant file service. The Journal of Systems and Software, 53(2):173–182, August 31, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/52/28/
30/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/52/28/abstract.
html.

Calzarossa:2008:CEN

[CT08] Maria Carla Calzarossa and Daniele Tessera. Characterization of the evolution of a news Web site. The Journal of Systems and Software, 81(12):2336–2344, December 2008. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cabot:2009:IIC

[CT09] Jordi Cabot and Ernest Teniente. Incremental integrity checking of UML/OCL conceptual schemas. The Journal of Systems and Software, 82(9):1459–1478, September 2009. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chu:2008:EAM

[CTL08] Chun-Jung Chu, Vincent S.

Tseng, and Tyne Liang. An efficient algorithm for mining temporal high utility itemsets from data streams. The Journal of Systems and Software, 81(7):1105–1117, July 2008. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:2001:PSS

[CTY01] T. Y. Chen, T. H. Tse, and Y. T. Yu. Proportional sampling strategy: a compendium and some insights. The Journal of Systems and Software, 58(1):65– 81, August 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/68/27/33/abstract.
html.

Cortellessa:2009:SIS

[CUY09] Vittorio Cortellessa, Sebastian Uchitel, and Daniel Yankelevich. Special issue: Software performance — modeling and analysis. The Journal of Systems and Software, 82(1): 1–2, January 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chechik:2002:FMC

[CW02] Marsha Chechik and Andre Wong. Formal modeling in a commercial setting: a case study. The Journal of Systems and Software, 60(1):59– 82, January 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        29/11/72/27/34/abstract.
                                        html.

Chung:2009:ADB

[CW09] Yu-Fang Chung and Zhen-Yu Wu. Approach to designing bribery-free and coercionfree electronic voting scheme. The Journal of Systems and Software, 82(12):2081– 2090, December 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:2004:ARA

[CWC04] Ing-Ray Chen, Ding-Chau Wang, and Chih-Ping Chu.

Analyzing reconfigurable algorithms for managing replicated data. The Journal of Systems and Software, 72(3): 417–430, August 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:2000:ELD

[CWH00] Yuh-Shihng Chang, Tzong-Chen Wu, and Shih-Chan Huang. ElGamal-like digital signature and multisignature schemes using self-certified public keys. The Journal of Systems and Software, 50(2): 99–105, February 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/49/27/
26/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/49/27/abstract.
html.

Chang:2009:RDB

[CWP09] Chuan-Yu Chang, Hung-Jen Wang, and Sheng-Wen Pan. A robust DWT-based copyright verification scheme with Fuzzy ART. The Journal of Systems and Software, 82(11):1906–1915, November 2009. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:2000:OBV

[CY00] Her-Kun Chang and Shyan-Ming Yuan. Optimal binary vote assignment for replicated data. The Journal of Systems and Software, 53(1):73–82, July 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/29/11/52/25/
                                        32/article.pdf;   http:/
                                        /www.elsevier.nl/gej-ng/
                                        10/29/11/52/25/abstract.
                                        html.

Chan:2004:AOT

[CY04] Jien-Tsai Chan and Wuu Yang. Advanced obfuscation techniques for Java bytecode. The Journal of Systems and Software, 71(1– 2):1–10, April 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chan:2004:TJ

[CYH04] Jien-Tsai Chan, Wuu Yang, and Jing-Wei Huang. Traps in Java. The Journal of Systems and Software, 72(1): 33–47, June 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cornelissen:2008:ETA

[CZH^{+}08] Bas Cornelissen, Andy Zaidman, Danny Holten, Leon Moonen, Arie van Deursen, and Jarke J. van Wijk. Execution trace analysis through massive sequence and circular bundle views. The Journal of Systems and Software, 81(12):2252–2268, December 2008. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:2007:IBR

[CZL07] Xiaofeng Chen, Fangguo Zhang, and Shengli Liu. IDbased restrictive partially blind signatures and applications. The Journal of Systems and Software, 80(2): 164–171, February 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Diaz:2007:TEW

[DA07] Paloma Díaz and Ignacio Aedo. Towards efficient Web engineering approaches through flexible process models. The Journal of Systems and Software, 80(8):1375– 1389, August 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Dong:2007:CPB

[DACY07] Jing Dong, Paulo S. C.

Alencar, Donald D. Cowan, and Sheng Yang. Composing pattern-based components and verifying correctness. The Journal of Systems and Software, 80(11):1755– 1769, November 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Darwish:2002:CCP

[Dar02] Nasir Darwish. COPS: cooperative problem solving using DCOM. The Journal of Systems and Software, 63(2): 79–90, August 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Donzelli:2006:PFE

[DB06] Paolo Donzelli and Victor Basili. A practical framework for eliciting and modeling system dependability requirements: Experience from the NASA high dependability computing project. The Journal of Systems and Software, 79(1):107–119, January 2006. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Deeprasertkul:2005:ADC

[DBO05] Prattana Deeprasertkul, Pattarasinee Bhattarakosol, and Fergus O’Brien. Automatic detection and correction of programming faults for software applications. The Journal of Systems and Software, 78(2):101–110, November 2005. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tronto:2008:IAN

[dBTdSS08] Iris Fabiana de Barcelos Tronto, José Demísio Simões da Silva, and Nilson Sant’Anna. An investigation of artificial neural networks based prediction systems in software project management. The Journal of Systems and Software, 81(3):356– 367, March 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

deBruin:2003:QDS

[dBvV03] Hans de Bruin and Hans van Vliet. Quality-driven software architecture composition. The Journal of Systems and Software, 66(3): 269–284, June 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

deBoer:2008:AKD

[dBvV08] Remco C. de Boer and Hans van Vliet. Architectural knowledge discovery with latent semantic analysis: Constructing a reading guide for software product audits. The Journal of Systems and Software, 81(9): 1456–1469, September 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

deBoer:2009:SBR

[dBvV09] Remco C. de Boer and Hans van Vliet. On the similarity between requirements and architecture. The Journal of Systems and Software, 82(3):

544–550, March 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Duan:2009:EAT

[DC09] Lihua Duan and Jessica Chen. Exploring alternatives for transition verification. The Journal of Systems and Software, 82(9): 1388–1402, September 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Dai:2009:LQB

[DCAC09] W. Dai, D. Covvey, P. Alencar, and D. Cowan. Lightweight query-based analysis of workflow process dependencies. The Journal of Systems and Software, 82(6):915–931, June 2009. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Dow:2002:CMA

[DCH02] Chyi-Ren Dow, Jong-Shin Chen, and Min-Chang Hsieh. Checkpointing MPI applications on symmetric multiprocessor machines using SMPCkpt. The Journal of Systems and Software, 63(2): 137–150, August 15, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Drehmer:2001:NES

[DD01] David E. Drehmer and Sasa M. Dekleva. A note on the evolution of software engineering practices. The Journal of Systems and Software, 57(1): 1–7, April 27, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.elsevier.nl/gej%
2Dng/10/29/11/64/25/25/
abstract.html;   http:
//www.elsevier.nl/gej%
2Dng/10/29/11/64/25/25/
article.pdf.

Danicic:2005:CLW

[DDF^{+}05] Sebastian Danicic, Mohammed Daoudi, Chris Fox, Mark Harman, Robert M. Hierons, John R. Howroyd, Lahcen Ourabya, and Martin Ward. ConSUS: a light-weight program conditioner. The Journal of Systems and Software, 77(3): 241–262, September 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

DeLucia:2009:DPR

[DDGR09] Andrea De Lucia, Vincenzo Deufemia, Carmine Gravino, and Michele Risi. Design pattern recovery through visual language parsing and source code analysis. The Journal of Systems and Software, 82(7): 1177–1193, July 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

DelRosso:2008:SPT

[Del08] Christian Del Rosso. Software performance tuning of software product family architectures: Two case studies in the real-time embedded systems domain. The Journal of Systems and Software, 81 (1):1–19, January 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Deubler:2001:EMV

[Deu01]  Hanns-Helmuth  Deubler.

Employing multiple views to separate large-scale software systems. The Journal of Systems and Software, 56(2): 105–113, March 1, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        29/11/57/30/25/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/29/11/57/
                                        30/25/article.pdf.

DiFelice:2000:SRS

[DF00] Paolino Di Felice and Goffredo Fonzi. Selecting reusable stand-alone routines: a proposal and a case study. The Journal of Systems and Software, 54(3):171–178, November 1, 2000. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/53/28/25/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/53/
28/25/article.pdf.

Dieste:2003:CMC

[DGJ^{+}03] Oscar Dieste, Marcela Genero, Natalia Juristo, José L. Maté, and Ana M. Moreno. A conceptual model completely independent of the implementation paradigm. The Journal of Systems and Software, 68(3):183–198, December 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Diaz:2008:RAC

[DGL^{+}08] M. Díaz, D. Garrido, L. Llopis, F. Rus, and J. M. Troya. UM-RTCOM: an analyzable component model for real-time distributed systems. The Journal of Systems and Software, 81(5): 709–726, May 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Davis:2002:ICA

[DGP02] L. Davis, R. F. Gamble, and J. Payton. The impact of component architectures on interoperability. The Journal of Systems and Software, 61(1):31–45, March 1, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Dasarathy:2007:ANQ

[DGV^{+}07] Balakrishnan Dasarathy, Shrirang Gadgil, Ravi Vaidyanathan, Arnie Neidhardt, Brian Coan, Kirthika Parmeswaran, Allen McIntosh, and Frederick Porter. Adaptive network QoS in layer-3/layer-2 networks as a middleware service for mission-critical applications. The Journal of Systems and Software, 80(7): 972–983, July 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lucca:2008:GEI

[DGV08] Giuseppe A. Di Lucca, Nicolas Gold, and Giuseppe Visaggio. Guest Editor’s introduction: 10th Conference on Software Maintenance and Reengineering. The Journal of Systems and Software, 81 (4):461–462, April 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Dalton:2009:NSA

[DH09] Andrew R. Dalton and Jason O. Hallstrom. nAIT: a source analysis and instrumentation framework for nesC. The Journal of Systems and Software, 82(7): 1057–1072, July 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Dick:2005:SPM

[DHJ05] A. J. J. Dick, M. E. C. Hull, and K. Jackson. Specifying process and measuring progress in terms of information state. The Journal of Systems and Software, 76(3): 311–322, June 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Dean:2006:SPF

[DHKV06] Tom Dean, Mark Harman, Rainer Koschke, and Michael Van De Vanter. Selected papers from the fourth Source Code Analysis and Manipulation (SCAM 2004) Workshop. The Journal of Systems and Software, 79(9): 1217–1218, September 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Deng:2006:OOC

[DHL06] Yan-Xiang Deng, Chao-Jang Hwang, and Jiang-Lung Liu. An object-oriented cryptosystem based on two-level reconfigurable computing architecture. The Journal of Systems and Software, 79(4): 466–479, April 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Donzelli:2001:DSS

[DI01a] Paolo Donzelli and Giuseppe Iazeolla. A dynamic simulator of software processes to test process assumptions. The Journal of Systems and Software, 56(1):81– 90, February 1, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/57/24/31/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/57/
24/31/article.pdf.

Donzelli:2001:HSM

[DI01b] Paolo Donzelli and Giuseppe Iazeolla. Hybrid simulation modelling of the software process. The Journal of Systems and Software, 59(3):227–235, December 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/71/34/28/abstract.
html.

DAmbrogio:2005:MDD

[DI05] Andrea D’Ambrogio and Giuseppe Iazeolla. Metadatadriven design of integrated environments for software performance validation. The Journal of Systems and Software, 76(2):127–146, May 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Dietrich:2008:USN

[DJW08] Jens Dietrich, Nathan Jones, and Jevon Wright. Using social networking and Semantic Web technology in software engineering — Use cases, patterns, and a case study. The Journal of Systems and Software, 81(12): 2183–2193, December 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Dowling:2001:PBS

[DK01] Andy Dowling and John G.

Keating. A proxy-based security architecture for Internet applications in an extranet environment. The Journal of Systems and Software, 58(2):107–118, September 1, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/68/33/29/abstract.
html.

Doller:2008:MMD

[DK08] Mario Döller and Harald Kosch. The MPEG-7 Multimedia Database System (MPEG-7 MMDB). The Journal of Systems and Software, 81(9):1559–1580, September 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

deLemos:2004:AFB

[dL04] Rogério de Lemos. Analysing failure behaviours in component interaction. The Journal of Systems and Software, 71(1–2):97–115, April 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Damm:2006:RIC

[DL06] Lars-Ola Damm and Lars Lundberg. Results from introducing component-level test automation and Test-Driven Development. The Journal of Systems and Software, 79(7):1001–1014, July 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Duran-Limon:2004:QMS

[DLB04] Hector A. Duran-Limon and Gordon S. Blair. QoS management specification support for multimedia middleware. The Journal of Systems and Software, 72 (1):1–23, June 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

deLemos:2006:ADS

[dLGR06] Rogério de Lemos, Cristina Gacek, and Alexander Romanovsky. Architecting dependable systems. The Journal of Systems and Software, 79(10):1359–1360, October 2006. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

delaRiva:2006:AGA

[dlRT06] Claudio de la Riva and Javier Tuya. Automatic generation of assumptions for modular verification of software specifications. The Journal of Systems and Software, 79(9): 1324–1340, September 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Damm:2008:MSR

[DLW08] Lars-Ola Damm, Lars Lundberg, and Claes Wohlin. A model for software rework reduction through a combination of anomaly metrics. The Journal of Systems and Software, 81(11):1968– 1982, November 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Santis:2007:NRN

[DM07] Alfredo De Santis and Barbara Masucci. New results on non-perfect sharing of multiple secrets. The Journal of Systems and Software, 80(2): 216–223, February 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Dabrowski:2007:UFR

[DMQ07] C. Dabrowski, K. Mills, and S. Quirolgico. Understanding failure response in service discovery systems. The Journal of Systems and Software, 80(6):896–917, June 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Penta:2005:LIS

[DNAM05] M. Di Penta, M. Neteler, G. Antoniol, and E. Merlo. A language-independent software renovation framework. The Journal of Systems and Software, 77(3):225– 240, September 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Barros:2004:SRS

[dOBWT04] Márcio de Oliveira Barros, Cláudia Maria Lima Werner, and Guilherme Horta Travassos. Supporting risks in software project management. The Journal of Systems and Software, 70(1–2): 21–35, February 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

deOliveira:2004:DOS

[dOZR^{+}04] Káthia Marçal de Oliveira, Fábio Zlot, Ana Regina Rocha, Guilherme Horta Travassos, Cátia Galotta, and Crediné Silva de Menezes. Domain-oriented software development environment. The Journal of Systems and Software, 72(2):145–161, July 2004. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Drakatos:2007:CAC

[DPMD07] Stylianos Drakatos, Niki Pissinou, Kia Makki, and Christos Douligeris. A context-aware cache structure for mobile computing environments. The Journal of Systems and Software, 80(7):1102–1119, July 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lucia:2003:AMP

[DPS03] Andrea De Lucia, Eugenio Pompella, and Silvio Stefanucci. Assessing the maintenance processes of a software organization: an empirical analysis of a large industrial project. The Journal of Systems and Software, 65(2): 87–103, February 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Dietrich:2006:CAE

[DPSU06] Suzanne W. Dietrich, Rohini Patil, Amy Sundermier, and Susan D. Urban. Component adaptation for event-based application integration using active rules. The Journal of Systems and Software, 79(12): 1725–1734, December 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Dunsmore:2000:RCS

[DRW00] A. Dunsmore, M. Roper, and M. Wood. The role of comprehension in software inspection. The Journal of Systems and Software, 52(2– 3):121–129, June 1, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/51/27/
29/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/51/27/abstract.
html.

Ding:2004:EJP

[DS04] Haibiao Ding and Mansur H. Samadzadeh. Extraction of Java program fingerprints for software authorship identification. The Journal of Systems and Software, 72(1): 49–57, June 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Deligiannis:2004:CEI

[DSA^{+}04] Ignatios Deligiannis, Ioannis Stamelos, Lefteris Angelis, Manos Roumeliotis, and Martin Shepperd. A controlled experiment investigation of an object-oriented design heuristic for maintainability. The Journal of Systems and Software, 72(2): 129–143, July 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Deelstra:2005:PDS

[DSB05] Sybren Deelstra, Marco Sinnema, and Jan Bosch. Product derivation in software product families: a case study. The Journal of Systems and Software, 74(2): 173–194, January 15, 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Deng:2008:CCV

[DSC^{+}08] Yi Deng, S. Masoud Sadjadi, Peter J. Clarke, Vagelis Hristidis, Raju Rangaswami, and Yingbo Wang. CVM — a communication virtual machine. The Journal of Systems and Software, 81(10): 1640–1662, October 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Deligiannis:2003:EIO

[DSRS03] Ignatios Deligiannis, Martin Shepperd, Manos Roumeliotis, and Ioannis Stamelos. An empirical investigation of an object-oriented design heuristic for maintainability. The Journal of Systems and Software, 65(2): 127–139, February 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Soares:2008:RTS

[dSSJV08] Michel dos Santos Soares, Stéphane Julia, and Jos Vrancken. Real-time scheduling of batch systems using Petri nets and linear logic. The Journal of Systems and Software, 81(11):1983– 1996, November 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

DeBardeleben:2009:BPS

[DSSL09] Nathan DeBardeleben, Ron Sass, Daniel Stanzione, Jr., and Walter B. Ligon III. Building problem-solving environments with the Arches framework. The Journal of Systems and Software, 82(7): 1137–1151, July 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Demestichas:2004:SPO

[DST^{+}04] P. Demestichas, V. Stavroulaki, V. Tountopoulos, M. Theologou, and N. Mitrou. Service provider oriented management systems over open cellular network infrastructures. The Journal of Systems and Software, 69(1–2): 1–14, January 1, 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

DiModica:2009:DSM

[DTV09] Giuseppe Di Modica, Orazio Tomarchio, and Lorenzo Vita. Dynamic SLAs management in service oriented environments. The Journal of Systems and Software, 82 (5):759–771, May 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Dai:2003:OTR

[DXPY03] Y. S. Dai, M. Xie, K. L.

Poh, and B. Yang. Optimal testing-resource allocation with genetic algorithm for modular software systems. The Journal of Systems and Software, 66(1):47– 55, April 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Depeng:2003:CCR

[DY03] Dang Depeng and Liu Yunsheng. Concurrency control in real-time broadcast environments. The Journal of Systems and Software, 68(2): 137–144, November 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Davis:2000:MPS

[DZ00] Alan M. Davis and Ann S.

Zweig. The missing piece of software development. The Journal of Systems and Software, 53(3):205– 206, September 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/29/11/52/29/
                                        25/article.pdf;   http:/
                                        /www.elsevier.nl/gej-ng/
                                        10/29/11/52/29/abstract.
                                        html.

Davis:2005:CSS

[DZ05] Alexander Davis and Du Zhang. A comparative study of SOAP and DCOM. The Journal of Systems and Software, 76(2):157–169, May 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Deppe:2004:RPR

[DZRH04] M. Deppe, M. Zanella, M. Robrecht, and W. Hardt. Rapid prototyping of realtime control laws for complex mechatronic systems: a case study. The Journal of Systems and Software, 70(3): 263–274, March 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Duan:2009:CET

[DZW^{+}09] Hua Duan, Qingtian Zeng, Huaiqing Wang, Sherry X. Sun, and Dongming Xu. Classification and evaluation of timed running schemas for workflow based on process mining. The Journal of Systems and Software, 82(3): 400–410, March 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

ElEmam:2000:VII

[EB00] Khaled El Emam and Andreas Birk. Validating the ISO/IEC 15504 measures of software development process capability. The Journal of Systems and Software, 51 (2):119–149, April 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/50/25/
29/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/50/25/abstract.
html.

Eriksson:2009:MRS

[EBB09] Magnus Eriksson, Jürgen Börstler, and Kjell Borg. Managing requirements specifications for product lines — an approach and industry case study. The Journal of Systems and Software, 82(3): 435–447, March 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ebert:2007:ISP

[Ebe07] Christof Ebert. The impacts of software product management. The Journal of Systems and Software, 80(6): 850–861, June 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

ElEmam:2001:CCB

[EBGR01] Khaled El Emam, Saida Benlarbi, Nishith Goel, and Shesh N. Rai. Comparing case-based reasoning classifiers for predicting high risk software components. The Journal of Systems and Software, 55(3):301–320, January 15, 2001. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/54/27/31/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/54/
27/31/article.pdf.

Emam:2004:ASS

[EC04] Khaled El Emam and Anita D. Carleton. Applications of statistics in software engineering. The Journal of Systems and Software, 73(2): 181–182, October 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Eeckhout:2004:HAS

[ED04] Lieven Eeckhout and Koen De Bosschere. How accurate should early design stage power/performance tools be? A case study with statistical simulation. The Journal of Systems and Software, 73(1): 45–62, September 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Eeckhout:2006:YSW

[ED06] Lieven Eeckhout and Koen De Bosschere. Yet shorter warmup by combining nostate-loss and MRRL for sampled LRU cache simulation. The Journal of Systems and Software, 79(5): 645–652, May 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Elish:2008:PDP

[EE08] Karim O. Elish and Mahmoud O. Elish. Predicting defect-prone software modules using support vector machines. The Journal of Systems and Software, 81(5): 649–660, May 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ekelhart:2008:XSC

[EFG^{+}08] Andreas Ekelhart, Stefan Fenz, Gernot Goluch, Markus Steinkellner, and Edgar Weippl. XML security — a comparative literature review. The Journal of Systems and Software, 81(10): 1715–1724, October 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

ElEmam:2000:EES

[EG00] Khaled El Emam and Iñigo Garro. Estimating the extent of standards use: the case of ISO/IEC 15504. The Journal of Systems and Software, 53(2):137–143, August 31, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/29/11/52/28/
                                        27/article.pdf;   http:/
                                        /www.elsevier.nl/gej-ng/
                                        10/29/11/52/28/abstract.
                                        html.

Ernst:2004:FBH

[EHKH04] M. Ernst, B. Henhapl, S. Klupsch, and S. Huss. FPGA based hardware acceleration for elliptic curve public key cryptosystems. The Journal of Systems and Software, 70(3):299– 313, March 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

ElEmam:2001:EEI

[EJ01] Khaled El Emam and Ho-Won Jung. An empirical evaluation of the ISO/ IEC 15504 assessment model. The Journal of Systems and Software, 59(1):23–41, October 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        29/11/71/27/29/abstract.
                                        html.

Eracar:2000:ASA

[EK00] Yönet A. Eracar and Mieczysław M. Kokar. An architecture for software that adapts to changes in requirements. The Journal of Systems and Software, 50(3):209–219, March 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/49/28/
28/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/49/28/abstract.
html.

Eisenbarth:2005:SOT

[EKV05] Thomas Eisenbarth, Rainer Koschke, and Gunther Vogel. Static object trace extraction for programs with pointers. The Journal of Systems and Software, 77(3): 263–284, September 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Engel:2007:MST

[EL07] Avner Engel and Mark Last.

Modeling software testing costs and risks using fuzzy logic paradigm. The Journal of Systems and Software, 80 (6):817–835, June 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

ElEmam:2000:ASE

[ELH00] Khaled El Emam, Oliver Laitenberger, and Thomas Harbich. The application of subjective estimates of effectiveness to controlling software inspections. The Journal of Systems and Software, 54(2):119–136, October 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/53/25/
28/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/53/25/abstract.
html.

Engel:2006:MSC

[ELK06] Jacob Engel, Daniel Lacks, and Taskin Kocak. Modelling and simulation of off-chip communication architectures for high-speed packet processors. The Journal of Systems and Software, 79(12):1701– 1714, December 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

ElEmam:2001:PFC

[EMM01] Khaled El Emam, Walcelio Melo, and Javam C. Machado. The prediction of faulty classes using object-oriented design metrics. The Journal of Systems and Software, 56(1):63– 75, February 1, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/57/24/29/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/57/
24/29/article.pdf.

Ehrich:2006:E

[EST06] Hans-Dieter Ehrich, Klaus-Dieter Schewe, and T. H. Tse. Editorial. The Journal of Systems and Software, 79 (5):575–576, May 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Edwards:2006:AFL

[ESW06] Dennis Edwards, Sharon Simmons, and Norman Wilde. An approach to feature location in distributed systems. The Journal of Systems and Software, 79(1): 57–68, January 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Feitelson:2007:FGA

[FAB^{+}07] Dror G. Feitelson, Tokunbo O. S. Adeshiyan, Daniel Balasubramanian, Yoav Etsion, Gabor Madl, Esteban P. Osses, Sameer Singh, Karlkim Suwanmongkol, Minhui Xie, and Stephen R. Schach. Fine-grain analysis of common coupling and its application to a Linux case study. The Journal of Systems and Software, 80 (8):1239–1255, August 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fairley:2007:ICS

[Fai07] Richard E. (Dick) Fairley. The influence of CO-COMO on software engineering education and training. The Journal of Systems and Software, 80(8):1201– 1208, August 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Folmer:2004:AUS

[FB04] Eelke Folmer and Jan Bosch. Architecting for usability: a survey. The Journal of Systems and Software, 70(1–2): 61–78, February 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Feredj:2009:MDP

[FBM09] Mohamed Feredj, Frédéric Boulanger, and Aimé Mokhoo Mbobi. A model of domainpolymorph component for heterogeneous system design. The Journal of Systems and Software, 82(1): 112–120, January 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fontoura:2000:UVD

[FCL^{+}00] Marcus Fontoura, Sérgio Crespo, Carlos José Lucena, Paulo S. C. Alencar, and Donald D. Cowan. Using viewpoints to derive objectoriented frameworks: a case study in the Web-based education domain. The Journal of Systems and Software, 54(3):239–257, November 1, 2000. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/53/28/31/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/53/
28/31/article.pdf.

Ferreira:2009:UER

[FCSM09] Susan Ferreira, James Collofello, Dan Shunk, and Gerald Mackulak. Understanding the effects of requirements volatility in software engineering by using analytical modeling and software process simulation. The Journal of Systems and Software, 82(10):1568–1577, October 2009. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Filho:2004:FIA

[FdOdL04] Ivan Mathias Filho, Toacy C.

de Oliveira, and Carlos J. P. de Lucena. A framework instantiation approach based on the Features Model. The Journal of Systems and Software, 73(2):333–349, October 2004. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Filho:2006:SEF

[FdSBR06] Fernando Castor Filho, Patrick Henrique da S. Brito, and Cecília Mary F. Rubira. Specification of exception flow in software architectures. The Journal of Systems and Software, 79(10): 1397–1418, October 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Forte:2008:UOW

[FdSdP08] Marcos Forte, Wanderley Lopes de Souza, and Antonio Francisco do Prado. Using ontologies and Web services for content adaptation in Ubiquitous Computing. The Journal of Systems and Software, 81(3): 368–381, March 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ferneley:2000:CCF

[Fer00] Elaine Ferneley. Coupling and control flow measures in practice. The Journal of Systems and Software, 51 (2):99–109, April 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/29/11/50/25/
                                        27/article.pdf;   http:/
                                        /www.elsevier.nl/gej-ng/
                                        10/29/11/50/25/abstract.
                                        html.

Fan:2009:FAR

[FHHL09] Chun-I Fan, Shi-Yuan Huang, Pei-Hsiu Ho, and Chin-Laung Lei. Fair anonymous rewarding based on electronic cash. The Journal of Systems and Software, 82(7): 1168–1176, July 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fornaro:2007:RYS

[FHT07] Robert J. Fornaro, Margaret R. Heil, and Alan L. Tharp. Reflections on 10 years of sponsored senior design projects: Students win– clients win! The Journal of Systems and Software, 80 (8):1209–1216, August 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fernandez-Iglesias:2005:WPT

[FIBRGCLN05] Manuel J. Fernández-Iglesias, Juan C. Burguillo-Rial, Francisco J. González-Castaño, and Martín Llamas-Nistal. Wireless protocol testing and validation supported by formal methods. A hands-on report. The Journal of Systems and Software, 75(1–2):139– 154, February 15, 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fernandez-Iglesias:2002:AFD

[FIGCLN^{+}02] Manuel J. Fernández-Iglesias, Francisco J. González-Castaño, Martín Llamas-Nistal, Jose M. Pousada-Carballo, and Javier Vales-Alonso. On the application of formal description techniques to the design of interception systems for GSM mobile terminals. The Journal of Systems and Software, 60(1):51– 58, January 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/72/27/33/abstract.
html.

Fichman:2001:ICS

[FK01] Robert G. Fichman and Chris F. Kemerer. Incentive compatibility and systematic software reuse. The Journal of Systems and Software, 57(1):45–60, April 27, 2001. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.elsevier.nl/gej%
2Dng/10/29/11/64/25/29/
abstract.html;   http:
//www.elsevier.nl/gej%
2Dng/10/29/11/64/25/29/
article.pdf.

Feng:2005:BDS

[FL05] Qian Feng and Robyn R. Lutz. Bi-directional safety analysis of product lines. The Journal of Systems and Software, 78(2):111–127, November 2005. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fung:2009:MEF

[FL09] Kam Hay Fung and Graham Cedric Low. Methodology evaluation framework for dynamic evolution in composition-based distributed applications. The Journal of Systems and Software, 82(12):1950–1965, December 2009. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fontoura:2001:UUF

[FLA^{+}01] Marcus Fontoura, Carlos J.

Lucena, Alexandre Andreatta, Sérgio E. Carvalho, and Celso C. Ribeiro. Using UML-F to enhance framework development: a case study in the local search heuristics domain. The Journal of Systems and Software, 57(3):201–206, July 15, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/64/32/29/abstract.
html.

Ferrari:2008:SAR

[FM08] Remo Ferrari and Nazim H.

Madhavji. Software architecting without requirements knowledge and experience: What are the repercussions? The Journal of Systems and Software, 81(9): 1470–1490, September 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Farooq:2009:AEQ

[FMP09] Umar Farooq, Shikharesh Majumdar, and Eric W. Parsons. Achieving efficiency, quality of service and robustness in multi-organizational Grids. The Journal of Systems and Software, 82(1): 23–38, January 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Frantzeskou:2008:ESH

[FMSG08] Georgia Frantzeskou, Stephen MacDonell, Efstathios Stamatatos, and Stefanos Gritzalis. Examining the significance of high-level programming features in source code author classification. The Journal of Systems and Software, 81(3):447– 460, March 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fioravanti:2000:MTA

[FN00] F. Fioravanti and P. Nesi. A method and tool for assessing object-oriented projects and metrics management. The Journal of Systems and Software, 53(2):111–136, August 31, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/29/11/52/28/
                                        26/article.pdf;   http:/
                                        /www.elsevier.nl/gej-ng/
                                        10/29/11/52/28/abstract.
                                        html.

Frakes:2004:CSR

[Fra04] William B. Frakes. A case study of a reusable component collection in the information retrieval domain. The Journal of Systems and Software, 72(2): 265–270, July 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Frailey:2007:ETB

[Fra07] Dennis J. Frailey. Experience teaching Barry Boehm’s techniques in industrial and academic settings. The Journal of Systems and Software, 80(8):1217–1221, August 2007. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Filho:2009:IRC

[FRR09] Fernando Castor Filho, Alexander Romanovsky, and Cecília Mary F. Rubira. Improving reliability of cooperative concurrent systems with exception flow analysis. The Journal of Systems and Software, 82(5): 874–890, May 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Frakes:2001:ISR

[FS01] William B. Frakes and Giancarlo Succi. An industrial study of reuse, quality, and productivity. The Journal of Systems and Software, 57(2): 99–106, June 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        29/11/64/29/27/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/29/11/64/
                                        29/27/article.pdf.

Freeman:2005:TDN

[FS05] Gordon L. Freeman, Jr. and Stephen R. Schach. The task-dependent nature of the maintenance of objectoriented programs. The Journal of Systems and Software, 76(2):195–206, May 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Farrugia:2006:ESP

[FS06] Emerson Farrugia and Robert Simon. An efficient and secure protocol for sensor network time synchronization. The Journal of Systems and Software, 79(2): 147–162, February 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fuggetta:2003:OSS

[Fug03]  Alfonso Fuggetta.  Open

source software — an evaluation. The Journal of Systems and Software, 66(1):77– 90, April 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Frankl:2000:TSD

[FW00] Phyllis G. Frankl and Elaine J.

Weyuker. Testing software to detect and reduce risk. The Journal of Systems and Software, 53(3):275– 286, September 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/52/29/
32/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/52/29/abstract.
html.

Fraser:2009:IUM

[FWA09] Gordon Fraser, Franz Wotawa, and Paul Ammann. Issues in using model checkers for test case generation. The Journal of Systems and Software, 82(9): 1403–1418, September 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Feng:2005:NMS

[FWTC05] Jen-Bang Feng, Hsien-Chu Wu, Chwei-Shyong Tsai, and Yen-Ping Chu. A new multisecret images sharing scheme using Lagrange’s interpolation. The Journal of Systems and Software, 76(3): 327–339, June 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Fan:2004:BBS

[FY04] Chin-Feng Fan and Yuan-Chang Yu. BBN-based software project risk management. The Journal of Systems and Software, 73(2): 193–203, October 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Guan:2007:DTP

[GAWW07] L. Guan, I. U. Awan, M. E. Woodward, and Xingang Wang. Discrete-time performance analysis of a congestion control mechanism based on RED under multiclass bursty and correlated traffic. The Journal of Systems and Software, 80(10): 1716–1725, October 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Garousi:2008:TAS

[GBL08] Vahid Garousi, Lionel C. Briand, and Yvan Labiche. Traffic-aware stress testing of distributed real-time systems based on UML models using genetic algorithms. The Journal of Systems and Software, 81(2):161– 185, February 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:2001:ASS

[GC01] Robert L. Glass and T. Y. Chen. An assessment of Systems and Software Engineering scholars and institutions (1996–2000). The Journal of Systems and Software, 59(1):107–113, October 15, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/71/27/35/abstract.
html.

Glass:2002:ASS

[GC02] R. L. Glass and T. Y.

Chen. An assessment of systems and software engineering scholars and institutions (1997–2001). The Journal of Systems and Software, 64 (1):79–86, October 15, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:2003:ASS

[GC03] R. L. Glass and T. Y.

Chen. An assessment of systems and software engineering scholars and institutions (1998–2002). The Journal of Systems and Software, 68 (1):77–84, October 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:2005:ASS

[GC05] Robert L. Glass and T. Y.

Chen. An assessment of systems and software engineering scholars and institutions (1999–2003). The Journal of Systems and Software, 76 (1):91–97, April 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Giusto:2004:RDE

[GD04] Paolo Giusto and Thilo Demmeler. Rapid design exploration of safety-critical distributed automotive applications via virtual integration platforms. The Journal of Systems and Software, 70(3): 245–262, March 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Grundy:2005:DSC

[GDH05] John Grundy, Guoliang Ding, and John Hosking. Deployed software component testing using dynamic validation agents. The Journal of Systems and Software, 74(1): 5–14, January 1, 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Giguette:2002:DRF

[GH02] Ray Giguette and Johnette Hassell. Designing a resourceful fault-tolerance system. The Journal of Systems and Software, 62(1): 47–57, May 1, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Goumopoulos:2004:ETG

[GH04] Christos Goumopoulos and Efthymios Housos. Efficient trip generation with a rule modeling system for crew scheduling problems. The Journal of Systems and Software, 69(1–2):43– 56, January 1, 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Grundy:2008:SIB

[GH08] John Grundy and Jun Han.

Special issue: Best papers from ASWEC 2007. The Journal of Systems and Software, 81(12):2105, December 2008. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gelenbe:2005:SAA

[GHK05] Erol Gelenbe, Khaled Hussain, and Varol Kaptan. Simulating autonomous agents in augmented reality. The Journal of Systems and Software, 74(3):255–268, February 1, 2005. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gruer:2004:HFS

[GHKR04] Juan Pablo Gruer, V. Hilaire, A. Koukam, and P. Rovarini. Heterogeneous formal specification based on Object-Z and statecharts: semantics and verification. The Journal of Systems and Software, 70 (1–2):95–105, February 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ghosh:2001:UCR

[Gho01] Sumit Ghosh. Understanding complex, real-world systems through asynchronous, distributed decision-making algorithms. The Journal of Systems and Software, 58(2):153–167, September 1, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/68/33/33/abstract.
html.

Grimstad:2007:IEJ

[GJ07] Stein Grimstad and Magne Jørgensen. Inconsistency of expert judgment-based estimates of software development effort. The Journal of Systems and Software, 80(11):1770–1777, November 2007. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Grunske:2008:QRB

[GJ08] Lars Grunske and David Joyce. Quantitative riskbased security prediction for component-based systems with explicitly modeled attack profiles. The Journal of Systems and Software, 81 (8):1327–1345, August 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ghosh:2008:BFI

[GK08] Sudipto Ghosh and John L.

Kelly. Bytecode fault injection for Java software. The Journal of Systems and Software, 81(11):2034– 2043, November 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Guan:2005:AMC

[GL05] Steven Guan and Richard Lai. Adaptive multimedia computing. The Journal of Systems and Software, 75(3): 235–236, March 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Glass:1999:ASS

[Gla99] Robert L. Glass. An assessment of systems and software engineering scholars and institutions (1994– 1998). The Journal of Systems and Software, 49(1): 81–86, December 15, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/46/25/
32/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/46/25/abstract.

html. See corrigendum [Gla00d].

Glass:2000:LFA

[Gla00a] R. L. Glass. A letter from the frustrated author of a journal paper. The Journal of Systems and Software, 54 (1):1, September 30, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/29/11/53/24/
                                        24/article.pdf;   http:/
                                        /www.elsevier.nl/gej-ng/
                                        10/29/11/53/24/abstract.
                                        html.

Glass:2000:ASL

[Gla00b] Robert L. Glass. Academics, and the scarlet letter ‘A’. The Journal of Systems and Software, 51(3):155, May 1, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/29/11/50/28/
                                        25/article.pdf;   http:/
                                        /www.elsevier.nl/gej-ng/
                                        10/29/11/50/28/abstract.
                                        html.

Glass:2000:ASS

[Gla00c] Robert L. Glass. An assessment of systems and software engineering scholars and institutions (1995– 1999). The Journal of Systems and Software, 54(1): 77–82, September 30, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/53/24/
32/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/53/24/abstract.
html.

Glass:2000:CAS

[Gla00d] Robert L. Glass. Corrigendum to: an assessment of systems and software engineering scholars and institutions (1994–1998) [The Journal of Systems and Software 49 (1) (1999) 81–86]. The Journal of Systems and Software, 51(3):275, May 1, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/50/28/
33/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/50/28/abstract.

html. See [Gla99].

Glass:2000:EOE

[Gla00e] Robert L. Glass. The end of the “outsourcing era”. The Journal of Systems and Software, 53(2):95–97, August 31, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/52/28/
24/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/52/28/abstract.
html.

Glass:2000:GBS

[Gla00f] Robert L. Glass. A goodbye of sorts. The Journal of Systems and Software, 55(2): 101, December 27, 2000. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        29/11/54/24/24/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/29/11/54/
                                        24/24/article.pdf.

Glass:2000:MFS

[Gla00g]  Robert L. Glass.  The

“maintenance-first” software era. The Journal of Systems and Software, 50(3): 171–174, March 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/29/11/49/28/
                                        25/article.pdf;   http:/
                                        /www.elsevier.nl/gej-ng/
                                        10/29/11/49/28/abstract.
                                        html.

Glass:2000:D

[Gla00h] Robert L. Glass. On design. The Journal of Systems and Software, 52(1):1–2, May 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/29/11/51/25/
                                        25/article.pdf;   http:/
                                        /www.elsevier.nl/gej-ng/
                                        10/29/11/51/25/abstract.
                                        html.

Glass:2000:SFR

[Gla00i]  Robert L. Glass.  The

“software-first” revolution in computer hardware design. The Journal of Systems and Software, 51(1):1–5, April 1, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/50/26/
25/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/50/26/abstract.
html.

Glass:2000:SMY

[Gla00j] Robert L. Glass. Software maintenance, Y2K and other software non-crises. The Journal of Systems and Software, 54(2):83–85, October 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/53/25/
24/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/53/25/abstract.
html.

Glass:2000:SRS

[Gla00k] Robert L. Glass. Software requirements success predictors — behavioral factors beat technical ones. The Journal of Systems and Software, 51(2):85–86, April 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/

                                        gej-ng/10/29/11/50/25/
                                        25/article.pdf;   http:/
                                        /www.elsevier.nl/gej-ng/
                                        10/29/11/50/25/abstract.
                                        html.

Glass:2000:TAS

[Gla00l]  Robert L. Glass.  Talk

about a software crisis — not! The Journal of Systems and Software, 55(1):1– 2, November 5, 2000. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        29/11/54/28/24/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/29/11/54/
                                        28/24/article.pdf.

Glass:2000:YBS

[Gla00m] Robert L. Glass. Y2K, and believing in software practice. The Journal of Systems and Software, 54(3):169–170, November 1, 2000. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        29/11/53/28/24/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/29/11/53/
                                        28/24/article.pdf.

Glass:2000:YBU

[Gla00n] Robert L. Glass. Y2K behind us: smooth sailing and blue skies? The Journal of Systems and Software, 50(1):1–3, January 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/49/25/
25/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/49/25/abstract.
html.

Glass:2002:ECF

[Gla02]  Robert L. Glass.  Edi-

tor’s corner: a final goodbye. The Journal of Systems and Software, 60(1): 1, January 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/72/27/27/abstract.
html.

Ghosh:2000:FRP

[GLJ00] Sumit Ghosh, Tony S. Lee, and Seong-Soon Joo. A frame of reference for the performance evaluation of asynchronous, distributed decision-making algorithms. The Journal of Systems and Software, 55(1):45–56, November 5, 2000. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/54/28/28/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/54/
28/28/article.pdf.

Gates:2002:FCB

[GM02] Ann Q. Gates and Oscar Mondragon. FasTLInC: a constraint-based tracing approach. The Journal of Systems and Software, 63 (3):241–258, September 15, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Garrigues:2009:PMA

[GMB^{+}09] Carles Garrigues, Nikos Migas, William Buchanan, Sergi Robles, and Joan Borrell. Protecting mobile agents from external replay attacks. The Journal of Systems and Software, 82(2): 197–206, February 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gannod:2005:ASS

[GML05] Gerald C. Gannod, Sudhakiran V. Mudiam, and Timothy E. Lindquist. Automated support for servicebased software development and integration. The Journal of Systems and Software, 74(1):65–71, January 1, 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Galindo:2008:ICB

[GMR08] David Galindo, Paz Morillo, and Carla Ràfols. Improved certificate-based encryption in the standard model. The Journal of Systems and Software, 81(7): 1218–1226, July 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Grassi:2007:FGB

[GMS07] Vincenzo Grassi, Raffaela Mirandola, and Antonino Sabetta. Filling the gap between design and performance/reliability models of component-based systems: a model-driven approach. The Journal of Systems and Software, 80(4): 528–558, April 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gokhale:2009:MBP

[Gok09] Swapna S. Gokhale. Modelbased performance analysis using block coverage measurements. The Journal of Systems and Software, 82(1): 121–130, January 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gondra:2008:AML

[Gon08]  Iker Gondra.  Applying

machine learning to software fault-proneness prediction. The Journal of Systems and Software, 81(2): 186–195, February 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gertphol:2005:MFR

[GP05] Sethavidh Gertphol and Viktor K. Prasanna. MIP formulation for robust resource allocation in dynamic realtime systems. The Journal of Systems and Software, 77 (1):55–65, July 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gonzalez-Perez:2007:MSD

[GPHS07] Cesar Gonzalez-Perez and Brian Henderson-Sellers. Modelling software development methodologies: a conceptual foundation. The Journal of Systems and Software, 80(11):1778–1796, November 2007. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gonzalez-Perez:2008:WPP

[GPHS08] Cesar Gonzalez-Perez and Brian Henderson-Sellers. A work product pool approach to methodology specification and enactment. The Journal of Systems and Software, 81 (8):1288–1305, August 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Goncalves:2008:RED

[GPM08] Gilles Goncalves and Aneta Poniszewska-Maranda. Role engineering: From design to evolution of security schemes. The Journal of Systems and Software, 81(8):1306– 1326, August 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Griman:2006:FAA

[GPML06] A. Grimán, M. Pérez, L. Mendoza, and F. Losavio. Feature analysis for architectural evaluation methods. The Journal of Systems and Software, 79(6): 871–888, June 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Germain:2005:EBP

[GR05] Éric Germain and Pierre N.

Robillard. Engineering-based processes and agile methodologies for software development: a comparative case study. The Journal of Systems and Software, 75(1–2): 17–27, February 15, 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Garcia:2001:CSE

[GRRX01] Alessandro F. Garcia, Cecília M. F. Rubira, Alexander Romanovsky, and Jie Xu. A comparative study of exception handling mechanisms for building dependable objectoriented software. The Journal of Systems and Software, 59(2):197–222, November 15, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/

                                        29/11/71/28/34/abstract.
                                        html.

Grunske:2007:EQP

[Gru07] Lars Grunske. Early quality prediction of componentbased systems — a generic framework. The Journal of Systems and Software, 80(5): 678–686, May 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gui:2007:RRS

[GS07] Gui Gui and Paul D. Scott. Ranking reusability of software components using coupling metrics. The Journal of Systems and Software, 80(9): 1450–1459, September 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Grunbacher:2007:MES

[GSB^{+}07] Paul Grünbacher, Norbert Seyff, Robert O. Briggs, Hoh Peter In, Hasan Kitapci, and Daniel Port. Making every student a winner: The WinWin approach in software engineering education. The Journal of Systems and Software, 80 (8):1191–1200, August 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gavalas:2009:MAP

[GTA09] Damianos Gavalas, George E. Tsekouras, and Christos Anagnostopoulos. A mobile agent platform for distributed network and systems management. The Journal of Systems and Software, 82(2):355–371, February 2009. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gursaran:2001:VRV

[Gur01] Gursaran. Viewpoint representation validation: a case study on two metrics from the Chidamber and Kemerer suite. The Journal of Systems and Software, 59(1):83– 97, October 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/71/27/33/abstract.
html.

Gruhn:2001:APL

[GW01] Volker Gruhn and Ursula Wellen. Analysing a process landscape by simulation. The Journal of Systems and Software, 59(3):333–342, December 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/71/34/37/abstract.
html.

Gu:2007:CRD

[GWDE07] Dazhang Gu, Lonnie Welch, Frank Drews, and Klaus Ecker. Characterizing robustness in dynamic realtime systems. The Journal of Systems and Software, 80(7): 1005–1014, July 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Graaf:2008:MDM

[GWvD08] Bas Graaf, Sven Weber, and Arie van Deursen. Modeldriven migration of supervisory machine control architectures. The Journal of Systems and Software, 81(4): 517–535, April 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hassapis:2003:MVC

[HA03] George Hassapis and Dimitra Ananidou. Modeling and verification of a class of real-time systems by the use of High Level Petri Nets. The Journal of Systems and Software, 68(2):153–165, November 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Helms:2006:FSW

[HAHH06] James W. Helms, James D. Arthur, Deborah Hix, and H. Rex Hartson. A field study of the Wheel — a usability engineering process model. The Journal of Systems and Software, 79(6): 841–858, June 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Habra:2008:FDV

[HALS08] Naji Habra, Alain Abran, Miguel Lopez, and Asma Sellami. A framework for the design and verification of software measurement methods. The Journal of Systems and Software, 81(5): 633–648, May 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Harrison:2000:ESS

[Har00] Rachel Harrison. Empirical studies of software development and evolution. The Journal of Systems and Software, 53(1):1, July 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/52/25/
25/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/52/25/abstract.
html.

Harrison:2004:FMM

[Har04] Warren Harrison. A flexible method for maintaining software metrics data: a universal metrics repository. The Journal of Systems and Software, 72(2): 225–234, July 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hazzan:2002:RPP

[Haz02] Orit Hazzan. The reflective practitioner perspective in software engineering education. The Journal of Systems and Software, 63(3):161–171, September 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Harman:2003:APS

[HBD03] Mark Harman, David Binkley, and Sebastian Danicic. Amorphous program slicing. The Journal of Systems and Software, 68(1):45– 64, October 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hedin:2005:TEP

[HBM05] Görel Hedin, Lars Bendix, and Boris Magnusson. Teaching extreme programming to large groups of students. The Journal of Systems and Software, 74(2): 133–146, January 15, 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hochstein:2008:PSC

[HBVG08] Lorin Hochstein, Victor R. Basili, Uzi Vishkin, and John Gilbert. A pilot study to compare programming effort for two parallel programming models. The Journal of Systems and Software, 81(11):

1920–1930, November 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Horng:2001:MVC

[HC01a] Jorng-Tzong Horng and Chi-Wei Chen. A mechanism for view consistency in a data warehousing system. The Journal of Systems and Software, 56(1):23– 37, February 1, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/57/24/26/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/57/
24/26/article.pdf.

Huang:2001:PBR

[HC01b] Yin-Fu Huang and Bo-Wei Chao. A priority-based resource allocation strategy in distributed computing networks. The Journal of Systems and Software, 58(3):221–233, September 15, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/68/34/30/abstract.
html.

Huang:2004:AFC

[HC04a] Chung-Ming Huang and Chi-Kuang Chang. An adaptive flow control with the re-transmission policy over the server–proxy–client networking environment. The Journal of Systems and Software, 69(1–2):15–27, January 1, 2004. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Huang:2004:NDE

[HC04b] Hui-Feng Huang and Chin-Chen Chang. A new design of efficient partially blind signature scheme. The Journal of Systems and Software, 73(3):397–403, November/December 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hasan:2006:THM

[HC06] Yusuf Hasan and J. Morris Chang. A tunable hybrid memory allocator. The Journal of Systems and Software, 79(8):1051–1063, August 2006. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hwang:2005:IWN

[HCC05] Liang-Yi Hwang, Mei-Ling Chiang, and Ruei-Chuan Chang. Implementation of wireless network environments supporting inter access point protocol and dual packet filtering. The Journal of Systems and Software, 76 (3):297–309, June 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hsueh:2008:QAE

[HCC08] Nien-Lin Hsueh, Peng-Hua Chu, and William Chu. A quantitative approach for evaluating the quality of design patterns. The Journal of Systems and Software, 81 (8):1430–1439, August 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hladik:2008:SRT

[HCDJ08] Pierre-Emmanuel Hladik, Hadrien Cambazard, Anne-Marie Déplanche, and Narendra Jussien. Solving a real-time allocation problem with constraint programming. The Journal of Systems and Software, 81(1): 132–149, January 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hwang:2008:DTD

[HCKY08] Jae-Hyun Hwang, Jin-Hee Choi, Se-Won Kim, and Chuck Yoo. DR-TCP: Downloadable and reconfigurable TCP. The Journal of Systems and Software, 81(1): 83–99, January 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Harrison:2000:EAE

[HCN00] R. Harrison, S. Counsell, and R. Nithi. Experimental assessment of the effect of inheritance on the maintainability of object-oriented systems. The Journal of Systems and Software, 52(2– 3):173–179, June 1, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/51/27/
35/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/51/27/abstract.
html.

Hansel:2004:DPV

[HCS04] David Hansel, Rance Cleaveland, and Scott A. Smolka. Distributed prototyping from validated specifications. The Journal of Systems and Software, 70(3):275–298, March 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hong:2009:RDH

[HCS09] Wien Hong, Tung-Shou Chen, and Chih-Wei Shiu. Reversible data hiding for high quality images using modification of prediction errors. The Journal of Systems and Software, 82(11):1833– 1842, November 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chu:2004:RSA

[hCSW^{+}04] Hao hua Chu, Henry Song, Candy Wong, Shoji Kurakake, and Masaji Katagiri. Roam, a seamless application framework. The Journal of Systems and Software, 69(3): 209–226, January 15, 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Henderson:2005:EIS

[HCWN05] Peter Henderson, Stephen Crouch, Robert John Walters, and Qinglai Ni. Effects of introducing survival behaviours into automated negotiators specified in an environmental and behavioural framework. The Journal of Systems and Software, 76(1): 65–76, April 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hansson:2006:HAI

[HDGZ06] Christina Hansson, Yvonne Dittrich, Björn Gustafsson, and Stefan Zarnak. How agile are industrial software development practices? The Journal of Systems and Software, 79(9):1295– 1311, September 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Huang:2000:SRB

[HDLK00] P. W. Huang, S. K. Dai, P. L.

Lin, and R. T. Kuo. Similarity retrieval based on group bounding and angle sequence matching in shape database systems. The Journal of Systems and Software, 54(1): 9–16, September 30, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/53/24/
26/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/53/24/abstract.
html.

Hanssen:2008:PFI

[HF08] Geir K. Hanssen and Tor E. Fægri. Process fusion: an industrial case study on agile software product line engineering. The Journal of Systems and Software, 81(6): 843–854, June 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Houston:2001:BCF

[HFC^{+}01] Dan X. Houston, Susan Ferreira, James S. Collofello, Douglas C. Montgomery, Gerald T. Mackulak, and Dan L. Shunk. Behavioral characterization: finding and using the influential factors in software process simulation models. The Journal of Systems and Software, 59(3):259–270, December 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.

elsevier.com/gej-ng/10/
29/11/71/34/31/abstract.
html.

Hug:2009:MBI

[HFRHS09] Charlotte Hug, Agnès Front, Dominique Rieu, and Brian Henderson-Sellers. A method to build information systems engineering process metamodels. The Journal of Systems and Software, 82(10): 1730–1742, October 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hepner:2006:PCA

[HGK^{+}06] M. Hepner, R. Gamble, M. Kelkar, L. Davis, and D. Flagg. Patterns of conflict among software components. The Journal of Systems and Software, 79(4): 537–551, April 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Huang:2000:PRT

[HH00] Tsung-Chuan Huang and Po-Hsueh Hsu. A practical runtime technique for exploiting loop-level parallelism. The Journal of Systems and Software, 54(3):259–271, November 1, 2000. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/53/28/32/abstract.
html;   http://www.elsevier.

                                        nl/gej-ng/10/29/11/53/
                                        28/32/article.pdf.

Huang:2005:PDP

[HH05] Yin-Fu Huang and Hung-Ming Ho. Pinned demand paging based on the access frequency of video files in video servers. The Journal of Systems and Software, 78 (3):223–233, December 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hsu:2006:ISU

[HH06] Ching-Sheng Hsu and Young-Chang Hou. An image size unconstrained ownership identification scheme for gray-level and color ownership statements based on sampling methods. The Journal of Systems and Software, 79(8):1130–1140, August 2006. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Han:2007:EAR

[HH07] Wen-Ming Han and Sun-Jen Huang. An empirical analysis of risk components and performance on software projects. The Journal of Systems and Software, 80(1): 42–50, January 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hazzan:2008:WHC

[HH08a] Orit Hazzan and Irit Hadar.

Why and how can humanrelated measures support software development processes? The Journal of Systems and Software, 81(7): 1248–1252, July 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hu:2008:AIB

[HH08b] Xiaoming Hu and Shangteng Huang. Analysis of IDbased restrictive partially blind signatures and applications. The Journal of Systems and Software, 81(11): 1951–1954, November 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Huang:2006:RPL

[HHL06] Chung-Ming Huang, Tz-Heng Hsu, and Yi-Wei Lin. REDUP: a packet loss recovery scheme for real-time audio streaming over wireless IP networks. The Journal of Systems and Software, 79(1): 29–42, January 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Henderson:2001:TES

[HHW01] Peter Henderson, Yvonne Margaret Howard, and Robert John Walters. A tool for evaluation of the software development process. The Journal of Systems and Software, 59(3):355–362, December 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/71/34/39/abstract.
html.

Host:2000:ECR

[HJ00] Martin Höst and Conny Johansson. Evaluation of code review methods through interviews and experimentation. The Journal of Systems and Software, 52(2– 3):113–120, June 1, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/51/27/
28/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/51/27/abstract.
html.

Haring:2007:E

[HJP07] Günter Haring, Carlos Juiz, and Dorina C. Petriu. Editorial. The Journal of Systems and Software, 80(4): 439–440, April 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Horgan:2009:UAQ

[HK09] Gerard Horgan and Souheil Khaddaj. Use of an adaptable quality model approach in a production support environment. The Journal of Systems and Software, 82(4): 730–738, April 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hofmeister:2007:GMS

[HKN^{+}07] Christine Hofmeister, Philippe Kruchten, Robert L. Nord, Henk Obbink, Alexander Ran, and Pierre America. A general model of software architecture design derived from five industrial approaches. The Journal of Systems and Software, 80(1): 106–126, January 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hoorn:2007:RCF

[HKvVvdV07] Johan F. Hoorn, Elly A.

Konijn, Hans van Vliet, and Gerrit van der Veer. Requirements change: Fears dictate the must haves; desires the won’t haves. The Journal of Systems and Software, 80(3): 328–355, March 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Huang:2000:TIM

[HKW00] Chung-Ming Huang, Cheng-Yi Kuo, and Chian Wang. TVIS: an interactive multimedia communication engine and its applications. The Journal of Systems and Software, 51(3):201–216, May 1, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/50/28/
29/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/50/28/abstract.
html.

Huang:2001:SFA

[HKY01] Chung-Ming Huang, Hsu-Yang Kung, and Jiong-Liang Yang. Synchronization and flow adaptation schemes for reliable multiple-stream transmission in multimedia presentations. The Journal of Systems and Software, 56(2):133–151, March 1, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/57/30/27/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/57/
30/27/article.pdf.

Horng:2000:MDW

[HL00a] Jorng-Tzong Horng and Jye Lu. Modularized design for wrappers/monitors in data warehouse systems. The Journal of Systems and Software, 54(3):185–199, November 1, 2000. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/53/28/27/abstract.

html;   http://www.elsevier.
nl/gej-ng/10/29/11/53/
28/27/article.pdf.

Huang:2000:IID

[HL00b] Chung-Ming Huang and Pei-Chuan Liu. IDRS: an interactive digital radio station over Internet. The Journal of Systems and Software, 51 (3):229–243, May 1, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/50/28/
31/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/50/28/abstract.
html.

Haggander:2001:SPM

[HL01] Daniel Häggander and Lars Lundberg. A simple process for migrating server applications to SMP:s. The Journal of Systems and Software, 57(1):31–43, April 27, 2001. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.elsevier.nl/gej%
2Dng/10/29/11/64/25/28/
abstract.html;   http:
//www.elsevier.nl/gej%
2Dng/10/29/11/64/25/28/
article.pdf.

Huang:2002:PSM

[HL02] Yin-Fu Huang and Jung-Hau Lin. A placement strategy of multimedia objects in multimedia information systems. The Journal of Systems and Software, 62(3): 181–193, June 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Huang:2006:ORA

[HL06a] Chin-Yu Huang and Jung-Hua Lo. Optimal resource allocation for cost and reliability of modular software systems in the testing phase. The Journal of Systems and Software, 79(5): 653–664, May 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hung:2006:EIC

[HL06b] Jen-Jou Hung and Yungho Leu. Efficient index caching for data dissemination in mobile computing environments. The Journal of Systems and Software, 79(1):93– 106, January 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Haw:2009:EPS

[HL09] Su-Cheng Haw and Chien-Sing Lee. Extending path summary and region encoding for efficient structural query processing in native XML databases. The Journal of Systems and Software, 82(6):1025–1035, June 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hsiung:2009:MVR

[HLC^{+}09] Pao-Ann Hsiung, Shang-Wei Lin, Yean-Ru Chen, Chun-Hsian Huang, Chihhsiong Shih, and William C. Chu. Modeling and verification of real-time embedded systems with urgency. The Journal of Systems and Software, 82(10): 1627–1641, October 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Huang:2001:OSU

[HLL01a] P. W. Huang, P. L. Lin, and H. Y. Lin. Optimizing storage utilization in Rtree dynamic index structure for spatial databases. The Journal of Systems and Software, 55(3):291–299, January 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/54/27/30/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/54/
27/30/article.pdf.

Hwang:2001:SMP

[HLL01b] Min-Shiang Hwang, Iuon-Chang Lin, and Li-Hua Li. A simple micropayment scheme. The Journal of Systems and Software, 55(3):221–229, January 15, 2001. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        29/11/54/27/24/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/29/11/54/
                                        27/24/article.pdf.

Harman:2009:TPM

[HLM^{+}09] Mark Harman, Zheng Li, Phil McMinn, Jeff Offutt, and John Clark. TAIC PART 2007 and Mutation 2007 special issue editorial. The Journal of Systems and Software, 82(11):1753– 1754, November 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

He:2007:OPC

[HLMB07] Zhen He, Richard Lai, Alonso Marquez, and Stephen Blackburn. Opportunistic prioritised clustering framework for improving OODBMS performance. The Journal of Systems and Software, 80(3):371–387, March 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Huang:2009:MIR

[HLT09] Chung-Ming Huang, Jian-Wei Li, and I-Ting Tseng. Multimedia Internet Rekeying for secure session mobility in ubiquitous mobile networks. The Journal of Systems and Software, 82(9): 1526–1539, September 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

He:2008:PRM

[HLW08] Zhen He, Byung Suk Lee, and X. Sean Wang. Proactive and reactive multidimensional histogram maintenance for selectivity estimation. The Journal of Systems and Software, 81(3): 414–430, March 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Horng:2004:PED

[HLWC04] Jong-Tzong Horng, Feng-Mao Lin, Li-Cheng Wu, and Chia-Hui Chang. Performance evaluation of a database of repetitive elements in complete genomes. The Journal of Systems and Software, 73(3):491–506, November/December 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Huang:2006:LAM

[HLYL06] Chung-Ming Huang, Chao-Hsien Lee, Yuan-Tse Yu, and Chung-Yi Lai. Locationaware multimedia proxy handoff over the IPv6 mobile network environment. The Journal of Systems and Software, 79(8):1037– 1050, August 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hall:2000:SEC

[HM00] Gregory A. Hall and John C. Munson. Software evolution: code delta and code churn. The Journal of Systems and Software, 54(2): 111–118, October 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/29/11/53/25/
                                        27/article.pdf;   http:/
                                        /www.elsevier.nl/gej-ng/
                                        10/29/11/53/25/abstract.
                                        html.

Hierons:2009:MTP

[HM09] Robert M. Hierons and Mercedes G. Merayo. Mutation testing from probabilistic and stochastic finite state machines. The Journal of Systems and Software, 82(11):1804–1818, November 2009. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Houston:2001:SSR

[HMC01] Dan X. Houston, Gerald T. Mackulak, and James S. Collofello. Stochastic simulation of risk factor potential effects for software development risk management. The Journal of Systems and Software, 59(3):247–257, December 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/71/34/30/abstract.
html.

Hissam:2003:EPA

[HMSW03] Scott Hissam, Gabriel Moreno, Judith Stafford, and Kurt Wallnau. Enabling predictable assembly. The Journal of Systems and Software, 65(3):185–198, March 15, 2003. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hall:2001:TFB

[HOR01] Benny Hall, George Orr, and Thomas E. Reeves. A technique for function block counting. The Journal of Systems and Software, 57(3): 217–220, July 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/64/32/31/abstract.
html.

Ha:2007:EST

[HPT07] Phuong Hoai Ha, Marina Papatriantafilou, and Philippas Tsigas. Efficient self-tuning spin-locks using competitive analysis. The Journal of Systems and Software, 80(7): 1077–1090, July 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Harn:2009:DDB

[HRL09] Lein Harn, Jian Ren, and Changlu Lin. Design of DL-based certificateless digital signatures. The Journal of Systems and Software, 82 (5):789–793, May 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Host:2001:EBM

[HRN^{+}01] Martin Höst, Björn Regnell, Johan Natt och Dag, Josef Nedstam, and Christian Nyberg. Exploring bottlenecks in market-driven requirements management processes with discrete event simulation. The Journal of Systems and Software, 59(3):323–332, December 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        29/11/71/34/36/abstract.
                                        html.

Hericko:2006:FRF

[HRŽ06] Marjan Heričko, Ivan Rozman, and Aleš Živkovič. A formal representation of functional size measurement methods. The Journal of Systems and Software, 79(9): 1341–1358, September 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hoffman:2003:ADE

[HS03] Daniel Hoffman and Paul Strooper. API documentation with executable examples. The Journal of Systems and Software, 66(2): 143–156, May 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

He:2007:FCB

[HSM^{+}07] Tian He, John A. Stankovic, Michael Marley, Chenyang Lu, Ying Lu, Tarek Abdelzaher, Sang Son, and Gang Tao. Feedback control-based dynamic resource management in distributed real-time systems. The Journal of Systems and Software, 80(7): 997–1004, July 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hoffbeck:2001:IMP

[HSR01] Joseph P. Hoffbeck, Mansoor Sarwar, and Eric J. Rix. Interfacing MAT-LAB with a parallel virtual processor for matrix algorithms. The Journal of Systems and Software, 56(1):77– 80, February 1, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/57/24/30/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/57/
24/30/article.pdf.

Huang:2009:CBS

[HTH09] Shi-Ming Huang, Chih-Fong Tsai, and Po-Chun Huang. Component-based software version management based on a Component-Interface Dependency Matrix. The Journal of Systems and Software, 82(3):382–399, March 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hashimoto:2000:NAF

[HTK00] Koji Hashimoto, Tatsuhiro Tsuchiya, and Tohru Kikuno. A new approach to faulttolerant scheduling using task duplication in multiprocessor systems. The Journal of Systems and Software, 53(2):159–171, August 31, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/29/11/52/28/
                                        29/article.pdf;   http:/
                                        /www.elsevier.nl/gej-ng/
                                        10/29/11/52/28/abstract.
                                        html.

Huang:2005:CRO

[Hua05a]  Chin-Yu Huang.  Cost-

reliability-optimal release policy for software reliability models incorporating improvements in testing efficiency. The Journal of Systems and Software, 77(2): 139–155, August 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Huang:2005:PAS

[Hua05b]  Chin-Yu Huang.  Per-

formance analysis of software reliability growth models with testing-effort and change-point. The Journal of Systems and Software, 76 (2):181–194, May 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Huston:2001:EDP

[Hus01]  Brian Huston.  The ef-

fects of design pattern application on metric scores. The Journal of Systems and Software, 58(3):261–269, September 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/68/34/33/abstract.
html.

Helmer:2003:LAI

[HWH^{+}03] Guy Helmer, Johnny S. K.

Wong, Vasant Honavar, Les Miller, and Yanxin Wang. Lightweight agents for intrusion detection. The Journal of Systems and Software, 67(2):109–122, August 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Helmer:2002:ADC

[HWHM02] Guy Helmer, Johnny S. K. Wong, Vasant Honavar, and Les Miller. Automated discovery of concise predictive rules for intrusion detection. The Journal of Systems and Software, 60(3):165–175, February 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/72/29/27/abstract.
html.

Helmer:2001:AID

[HWM01] Guy Helmer, Johnny Wong, and Subhasri Madaka. Anomalous intrusion detection system for hostile Java applets. The Journal of Systems and Software, 55(3):273–286, January 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/54/27/28/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/54/
27/28/article.pdf.

Hwang:2004:TAD

[HWML04] Jae-Joon Hwang, Kyu-Young Whang, Yang-Sae Moon, and Byung-Suk Lee. A top-down approach for density-based clustering using multidimensional indexes. The Journal of Systems and Software, 73(1):169–180, September 2004. CODEN JSSODM.

ISSN 0164-1212 (print), 1873-1228 (electronic).

Hsu:2001:NNT

[HWW01] Chien-Lung Hsu, Tzong-Sun Wu, and Tzong-Chen Wu. New nonrepudiable threshold proxy signature scheme with known signers. The Journal of Systems and Software, 58(2):119–124, September 1, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/68/33/30/abstract.
html.

Huang:2000:DDA

[HY00] Tsung-Chuan Huang and Cheng-Ming Yang. Data dependence analysis for array references. The Journal of Systems and Software, 52(1):55–65, May 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/51/25/
30/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/51/25/abstract.
html.

Huang:2001:NLA

[HY01] Tsung-Chuan Huang and Cheng-Ming Yang. Nonlinear array data dependence test. The Journal of Systems and Software, 57(2): 145–154, June 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        29/11/64/29/30/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/29/11/64/
                                        29/30/article.pdf.

Hwang:2003:CAL

[HY03] Min-Shiang Hwang and Wei-Pang Yang. Controlling access in large partially ordered hierarchies using cryptographic keys. The Journal of Systems and Software, 67 (2):99–107, August 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hou:2002:OCI

[HYC02] Ting-Chao Hou, Chorng-Horng Yang, and Kim-Joan Chen. Optimizing controllability of an interactive videoconferencing system with Web-based control interfaces. The Journal of Systems and Software, 62(2): 97–109, May 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hou:2004:AMS

[HYC04] Ting-Chao Hou, Chorng-Horng Yang, and Kim-Joan Chen. Adaptive message scheduling for supporting causal ordering in wide-area group communications. The Journal of Systems and Software, 72(3):

321–333, August 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hwang:2004:MID

[HYJL04] Seong Oun Hwang, Ki Song Yoon, Kyung Pyo Jun, and Kwang Hyung Lee. Modeling and implementation of digital rights. The Journal of Systems and Software, 73(3):533–549, November/December 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

He:2004:FAS

[HYS^{+}04] Xudong He, Huiqun Yu, Tianjun Shi, Junhua Ding, and Yi Deng. Formally analyzing software architectural specifications using SAM. The Journal of Systems and Software, 71(1– 2):11–29, April 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hussein:2007:IDA

[HZ07] Mohammed Hussein and Mohammad Zulkernine. Intrusion detection aware component-based systems: a specification-based framework. The Journal of Systems and Software, 80(5): 700–710, May 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Huang:2005:PPR

[HZCD05] Heyuan Huang, Shensheng Zhang, Jian Cao, and Yonghong Duan. A practical pattern recovery approach based on both structural and behavioral analysis. The Journal of Systems and Software, 75(1–2):69–87, February 15, 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Immich:2003:PAF

[IBP03] Patricia K. Immich, Ravi S. Bhagavatula, and Ravi Pendse. Performance analysis of five interprocess communication mechanisms across UNIX operating systems. The Journal of Systems and Software, 68(1):27– 43, October 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Iyengar:2003:TEA

[IJC03] Arun Iyengar, Shudong Jin, and Jim Challenger. Techniques for efficiently allocating persistent storage. The Journal of Systems and Software, 68(2):85–102, November 15, 2003. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Itzkovitz:2000:DAS

[INS00] Ayal Itzkovitz, Nitzan Niv, and Assaf Schuster. Dynamic adaptation of sharing granularity in DSM systems. The Journal of Systems and Software, 55(1):19– 32, November 5, 2000. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/54/28/26/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/54/
28/26/article.pdf.

Ilavarasan:2003:SWR

[IS03a] P. Vigneswara Ilavarasan and Arun Kumar Sharma. Is software work routinized?: Some empirical observations from Indian software industry. The Journal of Systems and Software, 66(1):1– 6, April 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Iosif:2003:TLP

[IS03b] Radu Iosif and Riccardo Sisto. Temporal logic properties of Java objects. The Journal of Systems and Software, 68(3):243–251, December 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Isoda:2001:OOR

[Iso01]  Sadahiro Isoda.  Object-

oriented real-world modeling revisited. The Journal of Systems and Software, 59(2):153–162, November 15, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        29/11/71/28/31/abstract.
                                        html.

Inverardi:2003:DFS

[IT03] Paola Inverardi and Massimo Tivoli. Deadlockfree software architectures for COM/DCOM applications. The Journal of Systems and Software, 65(3): 173–183, March 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Israr:2007:ITA

[IWF07] Tauseef Israr, Murray Woodside, and Greg Franks. Interaction tree algorithms to extract effective architecture and layered performance models from traces. The Journal of Systems and Software, 80(4):474–492, April 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jansen:2009:ESA

[JAvdV09] Anton Jansen, Paris Avgeriou, and Jan Salvador van der Ven. Enriching software architecture documentation. The Journal of Systems and Software, 82 (8):1232–1248, August 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jansen:2008:DAF

[JBA08] Anton Jansen, Jan Bosch, and Paris Avgeriou. Documenting after the fact: Recovering architectural design decisions. The Journal of Systems and Software, 81(4): 536–557, April 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jeon:2002:PME

[JC02] Gwangil Jeon and Yookun Cho. A partitioning method for efficient system-level diagnosis. The Journal of Systems and Software, 63(1):1– 16, July 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jiang:2005:HFT

[JCC05] Ying-Jie Jiang, Da-Wei Chang, and Ruei-Chuan Chang. HANet: a framework toward ultimately reliable network services. The Journal of Systems and Software, 78(3):270–281, December 2005. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jo:2004:UEA

[JCYC04] Jang-Wu Jo, Byeong-Mo Chang, Kwangkeun Yi, and Kwang-Moo Choe. An uncaught exception analysis for Java. The Journal of Systems and Software, 72(1): 59–69, June 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jaoua:2002:GCF

[JE02a] Ali Jaoua and Samir Elloumi. Galois connection, formal concepts and Galois lattice in real relations: application in a real classifier. The Journal of Systems and Software, 60(2):149–163, February 1, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/72/28/33/abstract.
html.

Jarraya:2002:IDI

[JE02b] Mohamed Jarraya and Didier El Baz. Implementation of distributed iterative algorithm for optimal control problems on several parallel architectures. The Journal of Systems and Software, 60(2):141–148, February 1, 2002. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/72/28/32/abstract.
html.

Johanson:2004:ETC

[JF04] Brad Johanson and Armando Fox. Extending tuplespaces for coordination in interactive workspaces. The Journal of Systems and Software, 69(3):243–266, January 15, 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jimenez:2008:PAI

[JFC08] Ernesto Jiménez, Antonio Fernández, and Vicent Cholvi. A parametrized algorithm that implements sequential, causal, and cache memory consistencies. The Journal of Systems and Software, 81(1):120–131, January 2008. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jorgensen:2007:CSE

[JFG07] Magne Jørgensen, Bjørn Faugli, and Tanja Gruschke. Characteristics of software engineers with optimistic predictions. The Journal of Systems and Software, 80(9): 1472–1482, September 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jeffrey:2008:ETC

[JG08] Dennis Jeffrey and Neelam Gupta. Experiments with test case prioritization using relevant slices. The Journal of Systems and Software, 81 (2):196–221, February 2008.

CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jung:2001:RBI

[JH01] Ho-Won Jung and Robin Hunter. The relationship between ISO/IEC 15504 process capability levels, ISO 9001 certification and organization size: an empirical study. The Journal of Systems and Software, 59(1):43– 55, October 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        29/11/71/27/30/abstract.
                                        html.

Hsu:2008:IAR

[jHjW08] Hwai jung Hsu and Feng jian Wang. An incremental analysis for resource conflicts to workflow specifications. The Journal of Systems and Software, 81(10): 1770–1783, October 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jeon:2009:DPS

[JHSB09] Sang-Uk Jeon, Jang-Eui Hong, In-Gwon Song, and Doo-Hwan Bae. Developing platform specific model for MPSoC architecture from UML-based embedded software models. The Journal of Systems and Software, 82(10):1695–1708, October 2009. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jorgensen:2003:SEE

[JIS03] Magne Jørgensen, Ulf Indahl, and Dag Sjøberg. Software effort estimation by analogy and “regression toward the mean”. The Journal of Systems and Software, 68(3): 253–262, December 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jalote:2006:ATH

[JJ06] Pankaj Jalote and Gourav Jain. Assigning tasks in a 24-h software development model. The Journal of Systems and Software, 79(7): 904–911, July 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jeong:2002:CCF

[JJP02] Bongju Jeong, Ho-Sang Jung, and Nam-Kyu Park. A computerized causal forecasting system using genetic algorithms in supply chain management. The Journal of Systems and Software, 60(3):223–237, February 15, 2002. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/

29/11/72/29/33/abstract.
html.

Jiang:2000:SDR

[JK00] James Jiang and Gary Klein. Software development risks to project effectiveness. The Journal of Systems and Software, 52(1):3–10, May 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/51/25/
26/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/51/25/abstract.
html.

Jiang:2002:PDS

[JKD02] James J. Jiang, Gary Klein, and Richard Discenza. Perception differences of software success: provider and user views of system metrics. The Journal of Systems and Software, 63(1):17– 27, July 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jiang:2009:RRU

[JKWL09] James J. Jiang, Gary Klein, Shelly P. J. Wu, and T. P. Liang. The relation of requirements uncertainty and stakeholder perception gaps to project management performance. The Journal of Systems and Software, 82(5): 801–808, May 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Juang:2004:FBT

[JL04] Wen-Shenq Juang and Horng-Twu Liaw. Fair blind threshold signatures in wallet with observers. The Journal of Systems and Software, 72 (1):25–31, June 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jung:2004:MCR

[JLC04] Ho-Won Jung, YiKyong Lim, and Chang-Shin Chung. Modeling change requests due to faults in a largescale telecommunication system. The Journal of Systems and Software, 72(2): 235–247, July 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jeon:2009:HEE

[JLYK09] Joo Hyuk Jeon, Ki Yong Lee, Jae Soo Yoo, and Myoung Ho Kim. HIPaG: an energy-efficient in-network join for distributed condition tables in sensor networks. The Journal of Systems and Software, 82(7): 1073–1086, July 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jalote:2007:WWH

[JMP07] Pankaj Jalote, Rajesh Munshi, and Todd Probsting. The When–Who–How analysis of defects for improving the quality control process. The Journal of Systems and Software, 80(4): 584–589, April 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Juristo:2007:AIU

[JMSS07] Natalia Juristo, Ana M. Moreno, and Maria-Isabel Sanchez-Segura. Analysing the impact of usability on software design. The Journal of Systems and Software, 80(9):1506–1516, September 2007. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jorgensen:2004:RSE

[Jør04] M. Jørgensen. A review of studies on expert estimation of software development effort. The Journal of Systems and Software, 70(1–2): 37–60, February 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jarzabek:2003:HVR

[JOZ03] Stan Jarzabek, Wai Chun Ong, and Hongyu Zhang. Handling variant requirements in domain modeling. The Journal of Systems and Software, 68(3):171–182, December 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jung:2000:ESC

[JPK00] Ho-Won Jung, Marjan Pivka, and Jong-Yoon Kim. An empirical study of complexity metrics in Cobol programs. The Journal of Systems and Software, 51(2): 111–118, April 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/50/25/
28/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/50/25/abstract.
html.

Jalote:2004:TPM

[JPKP04] Pankaj Jalote, Aveejeet Palit, Priya Kurien, and V. T. Peethamber. Timeboxing: a process model for iterative software development. The Journal of Systems and Software, 70(1–2): 117–127, February 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Janzen:2009:ENG

[JR09] David S. Janzen and Jungwoo Ryoo. Engaging the net generation with evidencebased software engineering through a community-driven Web database. The Journal of Systems and Software, 82 (4):563–570, April 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Juric:2006:CPW

[JRB^{+}06] Matjaz B. Juric, Ivan Rozman, Bostjan Brumen, Matjaz Colnaric, and Marjan Hericko. Comparison of performance of Web services, WS-Security, RMI, and RMI–SSL. The Journal of Systems and Software, 79 (5):689–700, May 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Juric:2009:WUE

[JSBR09] Matjaz B. Juric, Ana Sasa, Bostjan Brumen, and Ivan Rozman. WSDL and UDDI extensions for version support in Web services. The Journal of Systems and Software, 82(8):1326–1343, August 2009. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jorgensen:2004:BST

[JTM04] Magne Jørgensen, Karl Halvor Teigen, and Kjetil Moløkken. Better sure than safe? Overconfidence in judgement based software development effort prediction intervals. The Journal of Systems and Software, 70(1–2):79– 93, February 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jun:2000:MGL

[Jun00] Woochun Jun. A multigranularity locking-based concurrency control in objectoriented database systems. The Journal of Systems and Software, 54(3):201–217, November 1, 2000. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/53/28/28/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/53/
28/28/article.pdf.

Jeng:2006:EKM

[JW06] Fuh-Gwo Jeng and Chung-Ming Wang. An efficient keymanagement scheme for hierarchical access control based on elliptic curve cryptosystem. The Journal of Systems and Software, 79(8):1161– 1167, August 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jiang:2007:MAS

[JX07] Letian Jiang and Guozhi Xu. Modeling and analysis of software aging and software failure. The Journal of Systems and Software, 80(4): 590–595, April 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jeske:2005:SSA

[JZ05] Daniel R. Jeske and Xuemei Zhang. Some successful approaches to software reliability modeling in industry. The Journal of Systems and Software, 74(1):85– 99, January 1, 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jeske:2007:AVO

[JZ07] Daniel R. Jeske and Qi Zhang. Assessing the validity of one-part software reliability models using likelihood ratio and early detection tests. The Journal of Systems and Software, 80(6): 805–816, June 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jayaputera:2007:ERT

[JZL07] G. T. Jayaputera, A. Zaslavsky, and S. W. Loke. Enabling run-time composition and support for heterogeneous pervasive multi-agent systems. The Journal of Systems and Software, 80(12): 2039–2062, December 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Karatza:2000:CAR

[Kar00] Helen D. Karatza. Cache affinity and resequencing in a shared-memory multiprocessing system. The Journal of Systems and Software, 51(1):7–18, April 1, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/50/26/
26/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/50/26/abstract.
html.

Karatza:2001:JSH

[Kar01]  Helen D. Karatza.  Job

scheduling in heterogeneous distributed systems. The Journal of Systems and Software, 56(3):203–212, March 15, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/57/29/24/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/57/
29/24/article.pdf.

Karatza:2004:CS

[Kar04a] Helen D. Karatza. Computer systems. The Journal of Systems and Software, 71 (3):199–200, May 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Karatza:2004:PMA

[Kar04b] Helen D. Karatza. Performance modeling and analysis of computer systems and networks. The Journal of Systems and Software, 73(1): 1–2, September 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kitchenham:2007:ISS

[KB07] Barbara Kitchenham and Pearl Brereton. Introduction to special section on Evaluation and Assessment in Software Engineering EASE06. The Journal of Systems and Software, 80(9):1423– 1424, September 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kitchenham:2005:ISE

[KBBW05] Barbara Kitchenham, David Budgen, Pearl Brereton, and Philip Woodall. An investigation of software engineering curricula. The Journal of Systems and Software, 74(3): 325–335, February 1, 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kang:2009:NIB

[KBD09] Baoyuan Kang, Colin Boyd, and Ed Dawson. A novel identity-based strong designated verifier signature scheme. The Journal of Systems and Software, 82(2):

270–273, February 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kim:2007:CCM

[KBH07] In-Gyu Kim, Doo-Hwan Bae, and Jang-Eui Hong. A component composition model providing dynamic, flexible, and hierarchical composition of components for supporting software evolution. The Journal of Systems and Software, 80(11):1797– 1816, November 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kazman:2006:ECS

[KBK06] Rick Kazman, Len Bass, and Mark Klein. The essential components of software architecture design and analysis. The Journal of Systems and Software, 79(8):1207– 1216, August 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kaiser:2005:CRT

[KBM05] Jörg Kaiser, Cristiano Brudna, and Carlos Mitidieri. COS-MIC: a real-time event-based middleware for the CANbus. The Journal of Systems and Software, 77(1): 27–36, July 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kieu:2009:HSI

[KC09] The Duc Kieu and Chin-Chen Chang. A high stegoimage quality steganographic scheme with reversibility and high payload using multiple embedding strategy. The Journal of Systems and Software, 82(10):1743–1752, October 2009. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kim:2005:RFU

[KCB05] Deok-Hwan Kim, Chin-Wan Chung, and Kobus Barnard. Relevance feedback using adaptive clustering for image similarity retrieval. The Journal of Systems and Software, 78(1): 9–23, October 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kim:2001:JSG

[KCS01] Taehyoun Kim, Naehyuck Chang, and Heonshik Shin. Joint scheduling of garbage collector and hard realtime tasks for embedded applications. The Journal of Systems and Software, 58(3):247–260, September 15, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/68/34/32/abstract.
html.

Kouskouras:2008:FSE

[KCS08] Konstantinos G. Kouskouras, Alexander Chatzigeorgiou, and George Stephanides. Facilitating software extension with design patterns and Aspect-Oriented Programming. The Journal of Systems and Software, 81(10): 1725–1737, October 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kusmierek:2005:SVD

[KD05] Ewa Kusmierek and David H. C. Du. Streaming video delivery over Internet with adaptive end-to-end QoS. The Journal of Systems and Software, 75(3): 237–252, March 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Koriem:2004:NPN

[KDEK04] Samir M. Koriem, T. E. Dabbous, and W. S. El-Kilani. A new Petri net modeling technique for the performance analysis of discrete event dynamic systems. The Journal of Systems and Software, 72(3):335–348, August 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Karam:2008:PLA

[KDS^{+}08] Marcel Karam, Sergiu Dascalu, Haidar Safa, Rami Santina, and Zeina Koteich. A product-line architecture for Web service-based visual composition of Web applications. The Journal of Systems and Software, 81(6): 855–867, June 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Koriem:2004:NDB

[KEK04] Samir M. Koriem and W. S. El-Kilani. A new diskbased technique for solving the largeness problem of stochastic modeling formalisms. The Journal of Systems and Software, 72(3): 349–365, August 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kelly:2009:DFA

[Kel09] Diane Kelly. Determining factors that affect long-term evolution in scientific application software. The Journal of Systems and Software, 82 (5):851–861, May 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kendall:2002:SEM

[KFS^{+}02] H. Richard Kendall, Vincent W. Freeh, Paul W. Schermerhorn, Robert J. Minerick, and Peter W. Rijks. Streaming extensibility in the Modify-on-Access file system. The Journal of Systems and Software, 60(1):21– 36, January 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/72/27/31/abstract.
html.

Khurum:2009:SRD

[KG09] Mahvish Khurum and Tony Gorschek. A systematic review of domain analysis solutions for product lines. The Journal of Systems and Software, 82(12):1982– 2003, December 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kawaguchi:2006:MAC

[KGMI06] Shinji Kawaguchi, Pankaj K.

Garg, Makoto Matsushita, and Katsuro Inoue. MUD-ABlue: an automatic categorization system for Open Source repositories. The Journal of Systems and Software, 79(7):939–953, July 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Klosch:2002:TAL

[KGT02] René R. Klösch, Paul W.

Glaser, and Robert J. Truschnegg. A testing approach for large system portfolios in industrial environments. The Journal of Systems and Software, 62(1): 11–20, May 1, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kotini:2006:VRH

[KH06] Isabella Kotini and George Hassapis. Verification of rectangular hybrid automata models. The Journal of Systems and Software, 79(10): 1433–1443, October 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kim:2007:ICI

[Kim07a] Sangkyun Kim. IT compliance of industrial information systems: Technology management and industrial engineering perspective. The Journal of Systems and Software, 80(10):1590– 1593, October 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kim:2007:MSE

[Kim07b] Sangkyun Kim. Methodology of security engineering for industrial security management systems. The Journal of Systems and Software, 80(10):1589, October 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Klein:2001:SCI

[KJ01] Gary Klein and James J. Jiang. Seeking consonance in information systems. The Journal of Systems and Software, 56(2): 195–202, March 1, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/57/30/31/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/57/
30/31/article.pdf.

Kirk:2004:ITB

[KJ04] Steven R. Kirk and Samantha Jenkins. Information theory-based software metrics and obfuscation. The Journal of Systems and Software, 72(2):179–186, July 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kwon:2007:CDI

[KJLK07] Sungho Kwon, Sangsoo Jang, Jaeill Lee, and Sangkyun Kim. Common defects in information security management system of Korean companies. The Journal of Systems and Software, 80(10): 1631–1638, October 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kesseler:2006:THC

[KK06] Ernst Kesseler and Ed G.

Knapen. Towards humancentred design: Two case studies. The Journal of Systems and Software, 79(3): 301–313, March 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kania:2007:LSP

[KK07a] Dariusz Kania and Józef Kulisz. Logic synthesis for PAL-based CPLD-s based on two-stage decomposition. The Journal of Systems and Software, 80(7): 1129–1141, July 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kocaballi:2007:GBM

[KK07b] A. Baki Kocaballı and Altan Koçyiḡit. Granular best match algorithm for contextaware computing systems. The Journal of Systems and Software, 80(12):2015– 2024, December 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ko:2008:QSO

[KKK08] Jong Myoung Ko, Chang Ouk Kim, and Ick-Hyun Kwon. Quality-of-service oriented Web service composition algorithm and planning architecture. The Journal of Systems and Software, 81(11): 2079–2090, November 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kim:2009:QDA

[KKLP09] Suntae Kim, Dae-Kyoo Kim, Lunjin Lu, and Sooyong Park. Quality-driven architecture development using architectural tactics. The Journal of Systems and Software, 82(8):1211–1231, August 2009. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kim:2006:GSB

[KKP06] Jintae Kim, Minseong Kim, and Sooyong Park. Goal and scenario based domain requirements analysis environment. The Journal of Systems and Software, 79(7): 926–938, July 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Koru:2007:ICC

[KL07] A. Güneş Koru and Hongfang Liu. Identifying and characterizing change-prone classes in two large-scale open-source products. The Journal of Systems and Software, 80(1):63–73, January 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kim:2002:HID

[KLC02] Deok-Hwan Kim, Seok-Lyong Lee, and Chin-Wan Chung. Heterogeneous image database selection on the Web. The Journal of Systems and Software, 64(2):131–149, November 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kuz:2007:CCM

[KLGH07] Ihor Kuz, Yan Liu, Ian Gorton, and Gernot Heiser. CAmkES: a component model for secure microkernelbased embedded systems. The Journal of Systems and Software, 80(5):687–699, May 2007. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kraft:2006:IES

[KLMC06] Nicholas A. Kraft, Errol L. Lloyd, Brian A. Malloy, and Peter J. Clarke. The implementation of an extensible system for comparison and visualization of class ordering methodologies. The Journal of Systems and Software, 79 (8):1092–1109, August 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Keil:2008:ICR

[KLMZ08] Mark Keil, Lei Li, Lars Mathiassen, and Guangzhi Zheng. The influence of checklists and roles on software practitioner risk perception and decision-making. The Journal of Systems and Software, 81(6):908–919, June 2008. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kim:2007:SSP

[KLNS07] Chulyun Kim, Jong-Hwa Lim, Raymond T. Ng, and Kyuseok Shim. SQUIRE: Sequential pattern mining with quantities. The Journal of Systems and Software, 80(10):1726–1745, October 2007. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kahen:2001:SDM

[KLRW01] G. Kahen, M. M. Lehman, J. F. Ramil, and P. Wernick. System dynamics modelling of software evolution processes for policy investigation: Approach and example. The Journal of Systems and Software, 59(3):271–281, December 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/71/34/32/abstract.
html.

Karahasanovic:2007:CSD

[KLT07] Amela Karahasanović, Annette Kristin Levine, and Richard Thomas. Comprehension strategies and difficulties in maintaining objectoriented systems: an explorative study. The Journal of Systems and Software, 80(9): 1541–1559, September 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kim:2003:SAS

[KLY03] Daeyoung Kim, Yann-Hang Lee, and Mohamed Younis. Software architecture supporting integrated real-time systems. The Journal of Systems and Software, 65(1):71– 86, January 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kouvatsos:2004:BSH

[KM04] D. D. Kouvatsos and I. M. Mkwawa. Broadcasting schemes for hypercubes with background traffic. The Journal of Systems and Software, 73(1):3–14, September 2004. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kokune:2007:FSM

[KMKY07] Atsushi Kokune, Masuhiro Mizuno, Kyoichi Kadoya, and Shuichiro Yamamoto. FBCM: Strategy modeling method for the validation of software requirements. The Journal of Systems and Software, 80(3): 314–327, March 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Krishna:2009:EAD

[KMOS09] P. Venkata Krishna, Sudip Misra, Mohammad S. Obaidat, and V. Saritha. An efficient approach for distributed dynamic channel allocation with queues for real-time and non-real-time traffic in cellular networks. The Journal of Systems and Software, 82(7):1112–1124, July 2009. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Knoke:2005:E

[KMR05] Pete Knoke, Ana M. Moreno, and Michael Ryan. Editorial. The Journal of Systems and Software, 75(1–2): 1–2, February 15, 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kim:2004:SOS

[KMS04] Namgyu Kim, Songchun Moon, and Yonglak Sohn. Secure one snapshot protocol for concurrency control in real-time stock trading systems. The Journal of Systems and Software, 73(3):441–454, November/December 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kalogeraki:2008:RMU

[KMSMD08] Vana Kalogeraki, P. M.

Melliar-Smith, L. E. Moser, and Y. Drougas. Resource management using multiple feedback loops in soft realtime distributed object systems. The Journal of Systems and Software, 81(7): 1144–1162, July 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Karanikolas:2009:CLS

[KNYS09] Nikitas N. Karanikolas, Maria Nitsiou, Emmanuel J. Yannakoudakis, and Christos Skourlas. CUDL language semantics: Updating FDB data. The Journal of Systems and Software, 82(6): 947–962, June 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kurian:2007:MER

[KP07] Susan Kurian and Michael J. Pont. The maintenance and evolution of resourceconstrained embedded systems created using design patterns. The Journal of Systems and Software, 80(1): 32–41, January 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kostoulas:2007:APT

[KPG^{+}07] Dionysios Kostoulas, Dimitrios Psaltoulis, Indranil Gupta, Kenneth P. Birman, and Alan J. Demers. Active and passive techniques for group size estimation in large-scale and dynamic distributed systems. The Journal of Systems and Software, 80(10):1639–1658, October 2007. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kitchenham:2002:ESM

[KPME02] Barbara Kitchenham, Shari Lawrence Pfleeger, Beth McColl, and Suzanne Eagan. An empirical study of maintenance and development estimation accuracy. The Journal of Systems and Software, 64 (1):57–77, October 15, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). See erratum [KPME05].

Kitchenham:2005:ESE

[KPME05] Barbara Kitchenham, Shari Lawrence Pfleeger, Beth McColl, and Suzanne Eagan. Erratum to “An empirical study of maintenance and development estimation accuracy” [The Journal of Systems and Software 64 (2002) 57– 77]. The Journal of Systems and Software, 74(2): 227, January 15, 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). See [KPME02].

Khajenoori:2004:KCA

[KPS^{+}04] Soheil Khajenoori, Lorenz Prem, Karen Stevens, Ban Seng Keng, and Nader Kameli. Knowledge centered assessment pattern: an effective tool for assessing safety concerns in software architecture. The Journal of Systems and Software, 73(2): 313–322, October 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kim:2008:DFD

[KPS08] Jintae Kim, Sooyong Park, and Vijayan Sugumaran. DRAMA: a framework for domain requirements analysis and modeling architectures in software product lines. The Journal of Systems and Software, 81(1): 37–55, January 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ko:2009:EVR

[KPSK09] Young Il Ko, Chang-Sup Park, In Chul Song, and Myoung Ho Kim. An efficient void resolution method for geographic routing in wireless sensor networks. The Journal of Systems and Software, 82(6):963–973, June 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kim:2009:DRH

[KPT09] Youngmann Kim, E. K. Park, and Sungwoo Tak. Dynamically reconfigurable hardware–software architecture for partitioning networking functions on the SoC platform. The Journal of Systems and Software, 82(10):1588–1599, October 2009. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kapitsaki:2009:CAS

[KPTV09] Georgia M. Kapitsaki, George N.

Prezerakos, Nikolaos D. Tselikas, and Iakovos S. Venieris. Context-aware service engineering: a survey. The Journal of Systems and Software, 82(8):1285– 1297, August 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kiran:2008:SRP

[KR08] N. Raj Kiran and V. Ravi.

Software reliability prediction by soft computing techniques. The Journal of Systems and Software, 81(4): 576–583, April 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kim:2000:NRC

[KRC00] Dong Ho Kim, Keun Ho Ryu, and Young So Cho. A new routing control technique using active temporal data management. The Journal of Systems and Software, 51(1):37–48, April 1, 2000.

CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/29/11/50/26/
                                        28/article.pdf;   http:/
                                        /www.elsevier.nl/gej-ng/
                                        10/29/11/50/26/abstract.
                                        html.

Kumar:2008:SDC

[KRCK08] K. Vinay Kumar, V. Ravi, Mahil Carr, and N. Raj Kiran. Software development cost estimation using wavelet neural networks. The Journal of Systems and Software, 81(11):1853– 1867, November 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kralj:2005:ISF

[KRHŽ05] Tomaž Kralj, Ivan Rozman, Marjan Heričko, and Aleš Živkovič. Improved standard FPA method — resolving problems with upper boundaries in the rating complexity process. The Journal of Systems and Software, 77(2): 81–90, August 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Krinke:2006:ECP

[Kri06]  Jens Krinke.  Effects of

context on program slicing. The Journal of Systems and Software, 79(9): 1249–1260, September 2006.

CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kim:2000:SDM

[KRK00] Dong Ho Kim, Keun Ho Ryu, and Hong Soo Kim. A spatiotemporal database model and query language. The Journal of Systems and Software, 55(2):129–149, December 27, 2000. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/54/24/27/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/54/
24/27/article.pdf.

Kim:2002:DIS

[KRP02] Dong Ho Kim, Keun Ho Ryu, and Chee Hang Park. Design and implementation of spatiotemporal database query processing system. The Journal of Systems and Software, 60(1):37–49, January 15, 2002. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/72/27/32/abstract.
html.

Kruchten:2008:WDS

[Kru08] Philippe Kruchten. What do software architects really do? The Journal of Systems and Software, 81(12):2413– 2416, December 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kelly:2004:TDS

[KS04] Diane Kelly and Terry Shepard. Task-directed software inspection. The Journal of Systems and Software, 73(2): 361–368, October 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Khonsari:2004:ATF

[KSAOK04] A. Khonsari, H. Sarbazi-Azad, and M. Ould-Khaoua. Analysis of true fully adaptive routing with softwarebased deadlock recovery. The Journal of Systems and Software, 71(3):259–270, May 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Khoshgoftaar:2005:ROS

[KSH05] Taghi M. Khoshgoftaar, Naeem Seliya, and Angela Herzberg. Resource-oriented software quality classification models. The Journal of Systems and Software, 76(2): 111–126, May 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kraemer:2009:TSR

[KSH09] Frank Alexander Kraemer, Vidar Slåtten, and Peter Herrmann. Tool support for the rapid composition, analysis and implementation of reactive services. The Journal of Systems and Software, 82(12):2068–2080, December 2009. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Koo:2003:MFR

[KSS03] Seo Ryong Koo, Han Seong Son, and Poong Hyun Seong. A method of formal requirement analysis for NPP I&C systems based on UML modeling with software cost reduction. The Journal of Systems and Software, 67 (3):213–224, September 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Koru:2003:ECC

[KT03] A. Güne Koru and Jeff Tian. An empirical comparison and characterization of high defect and high complexity modules. The Journal of Systems and Software, 67(3):153–163, September 15, 2003. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kobayashi:2001:MMD

[KTK01] Noritaka Kobayashi, Tatsuhiro Tsuchiya, and Tohru Kikuno. Minimizing the mean delay of quorum-based mutual exclusion schemes. The Journal of Systems and Software, 58(1):1–9, August 15, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/68/27/27/abstract.
html.

Kasai:2007:SPS

[KUK07] Hiroyuki Kasai, Wataru Uchida, and Shoji Kurakake. A service provisioning system for distributed personalization with private data protection. The Journal of Systems and Software, 80(12):2025– 2038, December 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kuo:2000:KKC

[Kuo00] Feng Yang Kuo. To Kanji or not to Kanji: a cognitive perspective. The Journal of Systems and Software, 52(1):79–91, May 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/51/25/
32/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/51/25/abstract.
html.

Kudikyala:2005:SRU

[KV05] Udai Kumar Kudikyala and Rayford B. Vaughn. Software requirement understanding using Pathfinder networks: discovering and evaluating mental models. The Journal of Systems and Software, 74 (1):101–108, January 1, 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Koning:2006:MDI

[KvV06] Henk Koning and Hans van Vliet. A method for defining IEEE Std 1471 viewpoints. The Journal of Systems and Software, 79(1): 120–131, January 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kommareddy:2000:NBD

[KW00] Manjula Kommareddy and Johnny Wong. Non-blocking distributed transaction processing system. The Journal of Systems and Software, 54(1):65–76, September 30, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/53/24/
31/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/53/24/abstract.
html.

Keil:2000:IRP

[KWT^{+}00] Mark Keil, Linda Wallace, Dan Turk, Gayle Dixon-Randall, and Urban Nulden. An investigation of risk perception and risk propensity on the decision to continue a software development project. The Journal of Systems and Software, 53(2): 145–157, August 31, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/29/11/52/28/
                                        28/article.pdf;   http:/
                                        /www.elsevier.nl/gej-ng/
                                        10/29/11/52/28/abstract.
                                        html.

Kim:2008:NRF

[KY08] Deok-Hwan Kim and Seung-Hoon Yu. A new region filtering and region weighting approach to relevance feedback in content-based image retrieval. The Journal of Systems and Software, 81(9): 1525–1538, September 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kim:2009:DTB

[KY09] Seung Min Kim and Suk I. Yoo. DOM tree browsing of a very large XML document: Design and implementation. The Journal of Systems and Software, 82(11): 1843–1858, November 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kim:2003:DPC

[KYP^{+}03] Myungchul Kim, Sangjo Yoo, Jinhee Park, Sungwon Kang, Soon J. Hyun, and Hyuckjae Lee. A dynamic protocol conformance test method. The Journal of Systems and Software, 67(1): 31–43, July 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kim:2006:SBR

[KYPW06] Sang-Wook Kim, Jeehee Yoon, Sanghyun Park, and Jung-Im Won. Shapebased retrieval in time-series databases. The Journal of Systems and Software, 79(2): 191–203, February 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kong:2009:SBS

[KZDX09] Jun Kong, Kang Zhang, Jing Dong, and Dianxiang Xu. Specifying behavioral semantics of UML diagrams through graph transformations. The Journal of Systems and Software, 82(2): 292–306, February 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lai:2002:SCP

[Lai02] R. Lai. A survey of communication protocol testing. The Journal of Systems and Software, 62(1): 21–46, May 1, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Laitenberger:2000:ECR

[LASE00] Oliver Laitenberger, Colin Atkinson, Maud Schlich, and Khaled El Emam. An experimental comparison of reading techniques for defect detection in UML design documents. The Journal of Systems and Software, 53(2): 183–204, August 31, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/52/28/
31/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/52/28/abstract.
html.

Lardieri:2007:MLR

[LBS^{+}07] Patrick Lardieri, Jaiganesh Balasubramanian, Douglas C. Schmidt, Gautam Thaker, Aniruddha Gokhale, and Thomas Damiano. A multilayered resource management framework for dynamic resource management in enterprise DRE systems. The Journal of Systems and Software, 80(7):984–996, July 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lassing:2002:EAA

[LBvVB02] Nico Lassing, PerOlof Bengtsson, Hans van Vliet, and Jan Bosch. Experiences with ALMA: Architecture-Level Modifiability Analysis. The Journal of Systems and Software, 61(1):47– 57, March 1, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:2000:BFS

[LC00] Chin-Feng Lee and Chin-Chen Chang. Breadth-first search oriented symbolic picture representation for spatial match retrieval. The Journal of Systems and Software, 52(1):11–23, May 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/51/25/
27/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/51/25/abstract.
html.

Lee:2002:PVC

[LC02] Wei-Bin Lee and Tung-Her Chen. A public verifiable copy protection technique for still images. The Journal of Systems and Software, 62(3): 195–204, June 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:2005:TLS

[LC05] Yi-Hsuan Lee and Cheng Chen. A two-level scheduling method: an effective parallelizing technique for uniform nested loops on a DSP multiprocessor. The Journal of Systems and Software, 75(1–2):155–170, February 15, 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:2006:HLK

[LC06a] Yueh-Feng Lee and Ruei-Chuan Chang. Hotswapping Linux kernel modules. The Journal of Systems and Software, 79(2): 163–175, February 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Liu:2006:BRH

[LC06b] Ling Liu and Xue Chen. Book review: H. Zhuge, The Knowledge Grid, World Scientific Publishing Co., Singapore, 2004. The Journal of Systems and Software, 79 (5):743–745, May 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). See [Zhu04b].

Lee:2007:EEC

[LC07] Yi-Hsuan Lee and Cheng Chen. An effective and efficient code generation algorithm for uniform loops on non-orthogonal DSP architecture. The Journal of Systems and Software, 80(3): 410–428, March 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Liu:2008:RBM

[LC08] Shaoying Liu and Yuting Chen. A relation-based method combining functional and structural testing for test case generation. The Journal of Systems and Software, 81 (2):234–248, February 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lai:2009:MBD

[LC09] Jung-Yu Lai and Wei-Heng Chen. Measuring e-business dependability: The employee perspective. The Journal of Systems and Software, 82(6): 1046–1055, June 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lim:2002:MBA

[LCCH02] Andrew Lim, Oon Wee Chong, and Chi Chi-Hung. A matching-based algorithm for page access sequencing in join processing. The Journal of Systems and Software, 60(1):11–19, January 15, 2002. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/72/27/30/abstract.
html.

Liu:2006:MTO

[LCF^{+}06] Yan Liu, Bojan Cukic, Edgar Fuller, Sampath Yerramalla, and Srikanth Gururajan.

Monitoring techniques for an online neuro-adaptive controller. The Journal of Systems and Software, 79(11): 1527–1540, November 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lemos:2008:PPS

[LCF08] Melissa Lemos, Marco A. Casanova, and Antonio L. Furtado. Process pipeline scheduling. The Journal of Systems and Software, 81(3): 307–327, March 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Liang:2004:NSS

[LCH^{+}04] Deron Liang, P. Emerald Chung, Yennun Huang, Chandra Kintala, Woei-Jyh Lee, Timothy K. Tsai, and Chung-Yih Wang. NT-SwiFT: software implemented fault tolerance on Windows NT. The Journal of Systems and Software, 71(1– 2):127–141, April 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:2004:DEC

[LCL04] Choongseok Lee, Byounggu Choi, and Heeseok Lee. A development environment for customer-oriented Internet business: eBizBench. The Journal of Systems and Software, 72(2):163–178, July 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lin:2007:RAN

[LCLL07] Yi-Neng Lin, Yao-Chung Chang, Ying-Dar Lin, and Yuan-Chen Lai. Resource allocation in network processors for network intrusion prevention systems. The Journal of Systems and Software, 80(7):1030–1036, July 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lin:2008:DEI

[LCLL08] Ying-Dar Lin, Szu-Hao Chen, Po-Ching Lin, and Yuan-Cheng Lai. Designing and evaluating interleaving decompressing and virus scanning in a stream-based mail proxy. The Journal of Systems and Software, 81(9): 1517–1524, September 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:2006:MAR

[LcLsW06] Anthony J. T. Lee, Wan chuen Lin, and Chun sheng Wang. Mining association rules with multi-dimensional constraints. The Journal of Systems and Software, 79(1): 79–92, January 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Losavio:2004:IQS

[LCM^{+}04] F. Losavio, L. Chirinos, A. Matteo, N. Lévy, and A. Ramdane-Cherif. ISO quality standards for measuring architectures. The Journal of Systems and Software, 72(2):209–223, July 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lam:2000:ABT

[LCY00] Kam-Yiu Lam, Edward Chan, and Joe Chun-Hung Yuen. Approaches for broadcasting temporal data in mobile computing systems. The Journal of Systems and Software, 51(3):175–189, May 1, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/50/28/
27/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/50/28/abstract.
html.

Laitenberger:2000:ELC

[LD00] Oliver Laitenberger and Jean-Marc DeBaud. An encompassing life cycle centric survey of software inspection. The Journal of Systems and Software, 50(1): 5–31, January 2000. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/49/25/

26/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/49/25/abstract.
html.

Liu:2007:SAS

[LDL07] Jing Liu, Josh Dehlinger, and Robyn Lutz. Safety analysis of software product lines using state-based modeling. The Journal of Systems and Software, 80(11): 1879–1892, November 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Luk:2004:SMX

[LDN04] Robert W. P. Luk, Tharam S.

Dillon, and Vincent T. Y. Ng. Supporting metasearch with XSL. The Journal of Systems and Software, 73(1): 159–168, September 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lucredio:2008:SRB

[LdSBA^{+}08] Daniel Lucrédio, Kellyton dos Santos Brito, Alexandre Alvaro, Vinicius Cardoso Garcia, Eduardo Santana de Almeida, Renata Pontin de Mattos Fortes, and Silvio Lemos Meira. Software reuse: The Brazilian industry scenario. The Journal of Systems and Software, 81(6): 996–1013, June 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Leach:2008:SCL

[Lea08] Ronald J. Leach. Setting checkpoints in legacy code to improve fault-tolerance. The Journal of Systems and Software, 81(6):920–928, June 2008. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:2007:EAR

[Lee07]  Gun Ho Lee.  Efficient

approach for restructuring multiple inheritance hierarchies. The Journal of Systems and Software, 80(9): 1560–1568, September 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lethbridge:2000:PET

[Let00] Timothy C. Lethbridge. Priorities for the education and training of software engineers. The Journal of Systems and Software, 53 (1):53–71, July 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/29/11/52/25/
                                        31/article.pdf;   http:/
                                        /www.elsevier.nl/gej-ng/
                                        10/29/11/52/25/abstract.
                                        html.

Lutz:2003:ASP

[LG03] Robyn R. Lutz and Gerald C. Gannod. Analysis of a software product line architecture: an experience report. The Journal of Systems and Software, 66(3): 253–267, June 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lei:2005:AVT

[LG05a] Zhijun Lei and Nicolas D.

Georganas. Adaptive video transcoding and streaming over wireless channels. The Journal of Systems and Software, 75(3):253–270, March 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Liu:2005:SLM

[LG05b] Xin Liu and Steve Goddard.

Scheduling legacy multimedia applications. The Journal of Systems and Software, 75(3):319–328, March 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Liu:2008:AEC

[LGL08] Yan Liu, Ian Gorton, and Vinh Kah Lee. The architecture of an event correlation service for adaptive middleware-based applications. The Journal of Systems and Software, 81(12): 2134–2145, December 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Li:2009:DFA

[LGW09] Wei Li, Dawu Gu, and Yi Wang. Differential fault analysis on the contracting UFN structure, with application to SMS4 and MacGuffin. The Journal of Systems and Software, 82(2): 346–354, February 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Leung:2001:HSP

[LH01a] Karl R. P. H. Leung and Lucas C. K. Hui. Handling signature purposes in workflow systems. The Journal of Systems and Software, 55(3):245–259, January 15, 2001. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/54/27/26/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/54/
27/26/article.pdf.

Lu:2001:DEM

[LH01b] Eric Jui-Lin Lu and Rong-Ji Hwang. A distributed EDI model. The Journal of Systems and Software, 56(1):1– 7, February 1, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/57/24/24/abstract.
html;   http://www.elsevier.

nl/gej-ng/10/29/11/57/
24/24/article.pdf.

Lee:2004:AME

[LH04] S. K. Lee and C. S. Hwang.

Architecture modeling and evaluation for design of agent-based system. The Journal of Systems and Software, 72(2):195–208, July 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lo:2006:IFD

[LH06] Jung-Hua Lo and Chin-Yu Huang. An integration of fault detection and correction processes in software reliability analysis. The Journal of Systems and Software, 79(9):1312–1323, September 2006. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lin:2008:EMP

[LH08] Chu-Ti Lin and Chin-Yu Huang. Enhancing and measuring the predictive capabilities of testing-effort dependent software reliability models. The Journal of Systems and Software, 81(6): 1025–1038, June 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lo:2005:RAS

[LHC^{+}05] Jung-Hua Lo, Chin-Yu Huang, Ing-Yi Chen, Sy-Yen Kuo, and Michael R. Lyu. Reliability assessment and sensitivity analysis of software reliability growth modeling based on software module structure. The Journal of Systems and Software, 76 (1):3–13, April 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lan:2005:FDB

[LHLY05] Kwok Cheung Lan, Kei Shiu Ho, Robert Wing Pong Luk, and Daniel So Yeung. FNDS: a dialogue-based system for accessing digested financial news. The Journal of Systems and Software, 78(2): 180–193, November 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Loden:2009:WSS

[LHP^{+}09] Philip Loden, Qi Han, Lisa Porta, Tissa Illangasekare, and Anura P. Jayasumana. A wireless sensor system for validation of real-time automatic calibration of groundwater transport models. The Journal of Systems and Software, 82(11):1859– 1868, November 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). See corrigendum [LHP^{+}10].

Loden:2010:CWS

[LHP^{+}10] Philip Loden, Qi Han, Lisa Porta, Tissa Illangasekare, and Anura P. Jayasumana. Corrigendum to “A wireless sensor system for validation of real-time automatic calibration of groundwater transport models” [J. Syst. Software 82 (2009) 1859–1868]. The Journal of Systems and Software, 83 (4):710, April 2010. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). See [LHP^{+}09].

Lu:2006:FES

[LHSK06] Wan-Chen Lu, Jen-Wei Hsieh, Wei-Kuan Shih, and Tei-Wei Kuo. A faster exact schedulability analysis for fixed-priority scheduling. The Journal of Systems and Software, 79(12):1744– 1753, December 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lopez:2009:VCA

[LICA09] Claudia López, Pablo Inostroza, Luiz Marcio Cysneiros, and Hernán Astudillo. Visualization and comparison of architecture rationale with Semantic Web technologies. The Journal of Systems and Software, 82 (8):1198–1210, August 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lin:2000:RTI

[Lin00] Phen-Lan Lin. Robust transparent image watermarking system with spatial mechanisms. The Journal of Systems and Software, 50(2): 107–116, February 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/29/11/49/27/
                                        27/article.pdf;   http:/
                                        /www.elsevier.nl/gej-ng/
                                        10/29/11/49/27/abstract.
                                        html.

Lin:2001:DWM

[Lin01] Phen-Lan Lin. Digital watermarking models for resolving rightful ownership and authenticating legitimate customer. The Journal of Systems and Software, 55(3):261–271, January 15, 2001. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        29/11/54/27/27/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/29/11/54/
                                        27/27/article.pdf.

Lin:2007:PFT

[Lin07]  Jenn-Wei Lin.  Provid-

ing fault-tolerant authentication and authorization in wireless mobile IP networks. The Journal of Systems and Software, 80(2): 149–163, February 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lankes:2005:DPC

[LJB05] Stefan Lankes, Andreas Jabs, and Thomas Bemmerl. Design and performance of a CAN-based connectionoriented protocol for Real-Time CORBA. The Journal of Systems and Software, 77(1):37–45, July 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lim:2005:EII

[LJS05] Joa Sang Lim, Seung Ryul Jeong, and Stephen R. Schach. An empirical investigation of the impact of the object-oriented paradigm on the maintainability of realworld mission-critical software. The Journal of Systems and Software, 77(2): 131–138, August 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:2001:FCB

[LK01] Dong-Ho Lee and Hyoung-Joo Kim. A fast contentbased indexing and retrieval technique by the shape information in large image database. The Journal of Systems and Software, 56 (2):165–182, March 1, 2001. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/29/11/57/30/
                                        29/00007152.pdf;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/29/11/57/30/29/abstract.
                                        html.

Leem:2002:IIM

[LK02] Choon Seong Leem and Sangkyun Kim. Introduction to an integrated methodology for development and implementation of enterprise information systems. The Journal of Systems and Software, 60(3):249–261, February 15, 2002. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        29/11/72/29/35/abstract.
                                        html.

Lim:2004:RTB

[LK04] Sung-Hwa Lim and J.-H. Kim. Real-time broadcast algorithm for mobile computing. The Journal of Systems and Software, 69(1–2):173– 181, January 1, 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lacks:2009:DRS

[LK09] Daniel Lacks and Taskin Kocak. Developing reusable simulation core code for networking: The grid resource discovery example. The Journal of Systems and Software, 82(1):89–100, January 2009. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Liu:2006:IEW

[LKB06] Xiaoqing (Frank) Liu, Gautam Kane, and Monu Bambroo. An intelligent early warning system for software quality improvement and project management. The Journal of Systems and Software, 79(11):1552– 1564, November 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:2008:SAF

[LKH^{+}08] Changhoon Lee, Jongsung Kim, Seokhie Hong, Jaechul Sung, and Sangjin Lee. Security analysis of the fullround DDO-64 block cipher. The Journal of Systems and Software, 81(12):2328– 2335, December 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:2009:SWT

[LKH09] Sangho Lee, Jong Kim, and Sung Je Hong. Security weakness of Tseng’s fault-tolerant conference-key agreement protocol. The Journal of Systems and Software, 82(7):1163–1167, July 2009. CODEN JSSODM.

ISSN 0164-1212 (print), 1873-1228 (electronic).

Lim:2001:SAW

[LKJL01] Shin-Young Lim, Jeong-Ho Ko, Eun-Ah Jun, and Gang-Soo Lee. Specification and analysis of n-way key recovery system by Extended Cryptographic Timed Petri Net. The Journal of Systems and Software, 58(2):93– 106, September 1, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/68/33/28/abstract.
html.

Lam:2002:SRI

[LKL02] Kam-Yiu Lam, Tei-Wei Kuo, and Tony S. H. Lee. Strategies for resolving inter-class data conflicts in mixed realtime database systems. The Journal of Systems and Software, 61(1):1–14, March 1, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:2004:MDQ

[LKL04] Sangjun Lee, Dongseop Kwon, and Sukho Lee. Minimum distance queries for time series data. The Journal of Systems and Software, 69(1–2):105–113, January 1, 2004. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lo:2005:MDS

[LKL05] Shi-Wu Lo, Tei-Wei Kuo, and Kam-Yiu Lam. Multidisk scheduling for timeconstrained requests in RAID-0 devices. The Journal of Systems and Software, 76(3): 237–250, June 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:2009:DIM

[LKW^{+}09] Ki Yong Lee, Hyojun Kim, Kyoung-Gu Woo, Yon Dohn Chung, and Myoung Ho Kim. Design and implementation of MLC NAND flashbased DBMS for mobile devices. The Journal of Systems and Software, 82(9): 1447–1458, September 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:2000:CFT

[LL00] Victor C. S. Lee and K.-W.

Lam. Conflict free transaction scheduling using serialization graph for realtime databases. The Journal of Systems and Software, 55(1):57–65, November 5, 2000. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/54/28/29/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/54/
28/29/article.pdf.

Lundell:2004:CPC

[LL04] Björn Lundell and Brian Lings. Changing perceptions of CASE technology. The Journal of Systems and Software, 72(2): 271–280, July 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:2006:ISC

[LL06] Wei-Bin Lee and Kuan-Chieh Liao. Improved selfcertified group-oriented cryptosystem without a combiner. The Journal of Systems and Software, 79(4): 502–506, April 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Li:2007:SPI

[LL07] Shujun Li and Kwok-Tung Lo. Security problems with improper implementations of improved FEA-M. The Journal of Systems and Software, 80(5):791–794, May 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lai:2009:IKF

[LL09] Chin-Hui Lai and Duen-Ren Liu. Integrating knowledge flow mining and collaborative filtering to support document recommendation. The Journal of Systems and Software, 82(12):2023– 2037, December 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:2009:MTI

[LLC^{+}09] Yong Joon Lee, Jun Wook Lee, Duck Jin Chai, Bu Hyun Hwang, and Keun Ho Ryu. Mining temporal interval relational rules from temporal data. The Journal of Systems and Software, 82(1): 155–167, January 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Li:2008:CRR

[LLCL08] Shujun Li, Chengqing Li, Guanrong Chen, and Kwok-Tung Lo. Cryptanalysis of the RCES/RSES image encryption scheme. The Journal of Systems and Software, 81(7):1130–1143, July 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Liu:2008:DRK

[LLH08] Duen-Ren Liu, Chin-Hui Lai, and Chiu-Wen Huang. Document recommendation for knowledge sharing in personal folder environments. The Journal of Systems and Software, 81(8):1377– 1388, August 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:2004:EVM

[LLK04] Victor C. S. Lee, Kwok Wa Lam, and Tei-Wei Kuo. Efficient validation of mobile transactions in wireless environments. The Journal of Systems and Software, 69(1– 2):183–193, January 1, 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:2005:DIE

[LLK05] Hyun-Joo Lee, Sang-Won Lee, and Hyoung-Joo Kim. Design and implementation of an extended relationship semantics in an ODMGcompliant OODBMS. The Journal of Systems and Software, 76(2):171–180, May 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Li:2004:LMC

[LLKL04] Guo-Hui Li, Kam-Yiu Lam, Tei-Wei Kuo, and Shi-Wu Lo. Location management in cellular mobile computing systems with dynamic hierarchical location databases. The Journal of Systems and Software, 69(1–2):159–171, January 1, 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lam:2000:PDA

[LLL00] Kam-Yiu Lam, Gary C. K. Law, and Victor C. S. Lee.

Priority and deadline assignment to triggered transactions in distributed real-time active databases. The Journal of Systems and Software, 51(1):49–60, April 1, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/50/26/
29/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/50/26/abstract.
html.

Li:2006:EFA

[LLL06] Jung-Shian Li, Chu-Fen Li, and Cheng-Pong Lo. Effective fair allocation using smart market label auction with CSLF and CR-CSFQ. The Journal of Systems and Software, 79(5): 665–673, May 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Li:2006:ESY

[LLLZ06a] Chengqing Li, Shujun Li, Der-Chyuan Lou, and Dan Zhang. Erratum to “On the security of the Yen– Guo’s domino signal encryption algorithm (DSEA)” [The Journal of Systems and Software 79 (2006) 253– 258]. The Journal of Systems and Software, 79(12): 1789, December 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). See [LLLZ06b].

Li:2006:SYG

[LLLZ06b] Chengqing Li, Shujun Li, Der-Chyuan Lou, and Dan Zhang. On the security of the Yen–Guo’s domino signal encryption algorithm (DSEA). The Journal of Systems and Software, 79 (2):253–258, February 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). See erratum [LLLZ06a].

Lee:2009:MFP

[LLT^{+}09] Anthony J. T. Lee, Ying-Ho Liu, Hsin-Mu Tsai, Hsiu-Hui Lin, and Huei-Wen Wu. Mining frequent patterns in image databases with 9D-SPA representation. The Journal of Systems and Software, 82 (4):603–618, April 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Liu:2009:RTN

[LLV^{+}09] Jason Liu, Yue Li, Nathanael Van Vorst, Scott Mann, and Keith Hellman. A realtime network simulation infrastructure based on Open-VPN. The Journal of Systems and Software, 82(3): 473–485, March 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:2007:SSC

[LLY07] Patrick P. C. Lee, John C. S. Lui, and David K. Y. Yau.

SEAL: a secure communication library for building dynamic group key agreement applications. The Journal of Systems and Software, 80(3): 356–370, March 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lutz:2003:OAC

[LM03] Robyn R. Lutz and Ines Carmen Mikulski. Operational anomalies as a cause of safety-critical requirements evolution. The Journal of Systems and Software, 65(2): 155–161, February 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lago:2009:SAT

[LMvV09] Patricia Lago, Henry Muccini, and Hans van Vliet. A scoped approach to traceability management. The Journal of Systems and Software, 82(1):168–182, January 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lopez-Martin:2008:PAC

[LMYMGT08] Cuauhtémoc López-Martín, Cornelio Yáñez-Márquez, and Agustín Gutiérrez-Tornés. Predictive accuracy comparison of fuzzy models for software development effort of small programs. The Journal of Systems and Software, 81(6):

949–960, June 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:2001:IAD

[LNC01] Byong G. Lee, N. Hari Narayanan, and Kai H. Chang. An integrated approach to distributed version management and role-based access control in computer supported collaborative writing. The Journal of Systems and Software, 59(2):119–134, November 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/71/28/28/abstract.
html.

Lopez-Nores:2006:FSA

[LNPAGD^{+}06] Martín López-Nores, José Pazos-Arias, Jorge García-Duque, Yolanda Blanco-Fernández, Manuel Ramos-Cabrer, Alberto Gil-Solla, Ana Fernández-Vilas, and Rebeca Díaz-Redondo. Formal specification applied to multiuser distributed services: Experiences in collaborative t-learning. The Journal of Systems and Software, 79(8):1141–1155, August 2006. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Leung:2006:AEP

[LNY06] Karl R. P. H. Leung, Joseph K.-Y. Ng, and W. L. Yeung. Accessing embedded program in untestable mobile environment: Experience of a trustworthiness approach. The Journal of Systems and Software, 79(7): 962–968, July 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lukaszuk:2004:ADH

[LO04] Jack Lukaszuk and Ratko Orlandic. On accessing data in high-dimensional spaces: a comparative study of three space partitioning strategies. The Journal of Systems and Software, 73(1): 147–157, September 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Loke:2006:DPI

[Lok06] Seng W. Loke. Declarative programming of integrated peer-to-peer and Web based systems: the case of Prolog. The Journal of Systems and Software, 79(4): 523–536, April 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Loo:2005:DMS

[Loo05] Alfred Loo. Distributed multiple selection algorithm for peer-to-peer systems. The Journal of Systems and Software, 78(3):234–248, December 2005. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lopez:2003:AEF

[Lop03] Marta Lopez. Application of an evaluation framework for analyzing the architecture tradeoff analysis method SM. The Journal of Systems and Software, 68(3):233–241, December 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

LeTraon:2003:DDA

[LORB03] Yves Le Traon, Farid Ouabdesselam, Chantal Robach, and Benoit Baudry. From diagnosis to diagnosability: axiomatization, measurement and application. The Journal of Systems and Software, 65 (1):31–50, January 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lederer:2000:SMC

[LP00] Albert L. Lederer and Jayesh Prasad. Software management and cost estimating error. The Journal of Systems and Software, 50(1):33–42, January 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/29/11/49/25/
                                        27/article.pdf;   http:/

/www.elsevier.nl/gej-ng/
10/29/11/49/25/abstract.
html.

Lim:2005:EEC

[LP05] Sang Seok Lim and Kyu Ho Park. ECEM: an event correlation based event manager for an I/O-intensive application. The Journal of Systems and Software, 74(3): 229–242, February 1, 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lefevre:2007:SII

[LP07] Laurent Lefevre and Jean-Marc Pierson. Special issue: International Conference on Pervasive Services (ICPS 2006). The Journal of Systems and Software, 80(12): 1939–1940, December 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:2009:GEM

[LPJP09] Hyoung-Gon Lee, Namkyu Park, Han-Il Jeong, and Jinwoo Park. Grid enabled MRP process improvement under distributed database environment. The Journal of Systems and Software, 82(7): 1087–1097, July 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:2004:TMP

[LPR04] Jun Wook Lee, Ok Hyun Paek, and Keun Ho Ryu.

Temporal moving pattern mining for location-based service. The Journal of Systems and Software, 73(3):481–490, November/ December 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Leszak:2002:CED

[LPS02] Marek Leszak, Dewayne E. Perry, and Dieter Stoll. Classification and evaluation of defects in a project retrospective. The Journal of Systems and Software, 61(3): 173–187, April 1, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Li:2004:PQN

[LR04] Peng Li and Binoy Ravindran. Proactive QoS negotiation in asynchronous real-time distributed systems. The Journal of Systems and Software, 73(1):75– 88, September 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Li:2007:RMR

[LRS^{+}07] Huan Li, Krithi Ramamritham, Prashant Shenoy, Roderic A. Grupen, and John D. Sweeney. Resource management for real-time tasks in mobile robotics. The Journal of Systems and Software, 80(7):962–971, July 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lassing:2003:HWC

[LRvV03] Nico Lassing, Daan Rijsenbrij, and Hans van Vliet. How well can we predict changes at architecture design time? The Journal of Systems and Software, 65(2): 141–153, February 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:2004:CBM

[LS04] Seung C. Lee and Ashraf I.

Shirani. A component based methodology for Web application development. The Journal of Systems and Software, 71(1–2):177–187, April 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Leung:2005:MBE

[LS05a] C. M. Leung and J. A. Schormans. Measurement-based end to end latency performance prediction for SLA verification. The Journal of Systems and Software, 74(3): 243–254, February 1, 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Liu:2005:HAP

[LS05b] Duen-Ren Liu and Ya-Yueh Shih. Hybrid approaches to product recommendation based on customer lifetime value and purchase preferences. The Journal of Systems and Software, 77(2): 181–191, August 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Li:2007:ESB

[LS07] Wei Li and Raed Shatnawi. An empirical study of the bad smells and class error probability in the postrelease object-oriented system evolution. The Journal of Systems and Software, 80(7):1120–1128, July 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lo:2001:SPR

[LSaC01] Chia-Tien Dan Lo, Witawas Srisa-an, and J. Morris Chang. A study of page replacement performance in garbage collection heap. The Journal of Systems and Software, 58(3):235–245, September 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/68/34/31/abstract.
html.

Lo:2004:DAQ

[LSaC04] Chia-Tien Dan Lo, Witawas Srisa-an, and J. Morris Chang. The design and analysis of a quantitative simulator for dynamic memory management. The Journal of Systems and Software, 72(3): 443–453, August 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lin:2004:CCR

[LSC04] Kelvin Lin, Jean Jyh-Jiun Shann, and Chung-Ping Chung. Code compression by register operand dependency. The Journal of Systems and Software, 72(3): 295–304, August 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lian:2009:FCD

[LSH09] Jiexin Lian, Sol M. Shatz, and Xudong He. Flexible coordinator design for modeling resource sharing in multiagent systems. The Journal of Systems and Software, 82(10):1709–1729, October 2009. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Liu:2006:PAS

[LSV^{+}06] Xiaoqing (Frank) Liu, Yan Sun, Chandra Sekhar Veera, Yuji Kyoya, and Kunio Noguchi. Priority assessment of software process requirements from multiple perspectives. The Journal of Systems and Software, 79(11):1649– 1660, November 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:2007:RES

[LSZ^{+}07] Ken C. K. Lee, Josh Schiffman, Baihua Zheng, Wang-Chien Lee, and Hong Va Leong. Round-eye: a system for tracking nearest surrounders in moving object environments. The Journal of Systems and Software, 80(12):2063–2076, December 2007. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lin:2004:SIS

[LT04] Chang-Chou Lin and Wen-Hsiang Tsai. Secret image sharing with steganography and authentication. The Journal of Systems and Software, 73(3):405–414, November/December 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lai:2007:TVR

[LT07] Richard Lai and Tony Tsang. Timed verification of the reliable adaptive multicast protocol. The Journal of Systems and Software, 80(2): 224–239, February 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Liu:2008:TIV

[LT08] Hui Liu and Hee Beng Kuan Tan. Testing input validation in Web applications through automated model recovery. The Journal of Systems and Software, 81(2): 222–233, February 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Leu:2009:PDP

[LT09] Jenq-Shiou Leu and Cheng-Wei Tsai. Practical design of a proxy agent to facilitate adaptive video streaming service across wired/wireless networks. The Journal of Systems and Software, 82(11):1916–1925, November 2009. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lu:2001:ESX

[LTC01] Eric Jui-Lin Lu, Ru-Hui Tsai, and Shihyu Chou. An empirical study of XML/ EDI. The Journal of Systems and Software, 58(3):271–279, September 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/68/34/34/abstract.
html.

Lai:2006:MAM

[LTK^{+}06] Richard Lai, Tony Tsang, Yong Soo Kim, Dae Hoon Hwang, Chul Soo Lee, and Yoon Kyung Han. A method for analysing multimedia protocol performance. The Journal of Systems and Software, 79(4):454–465, April 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lam:2006:ASL

[LU06] Kam-Yiu Lam and Özgür Ulusoy. Adaptive schemes for location update generation in execution locationdependent continuous queries. The Journal of Systems and Software, 79(4):441– 453, April 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Liang:2000:DST

[LUS^{+}00] Tyng-Yeu Liang, Jyh-Chang Ueng, Ce-Kuen Shieh, Deh-Yuan Chuang, and Jun-Qi Lee. Distinguishing sharing types to minimize communication in software distributed shared memory systems. The Journal of Systems and Software, 55(1):73– 85, November 5, 2000. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/54/28/31/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/54/
28/31/article.pdf.

Lutz:2000:EPF

[Lut00] Robyn R. Lutz. Extending the product family approach to support safe reuse. The Journal of Systems and Software, 53(3):207– 217, September 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/52/29/
26/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/52/29/abstract.
html.

Lemos:2007:CDF

[LVMM07] Otávio Augusto Lazzarini Lemos, Auri Marcelo Rizzo Vincenzi, José Carlos Maldonado, and Paulo Cesar Masiero. Control and data flow structural testing criteria for aspect-oriented programs. The Journal of Systems and Software, 80(6): 862–882, June 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Leem:2002:GES

[LW02] Choon Seong Leem and Cho-Li Wang. Guest editorial: Special section on industrial information systems: progresses and perspectives in Pacific Rim. The Journal of Systems and Software, 60(3): 183, February 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        29/11/72/29/29/abstract.
                                        html.

Liu:2007:AAS

[LW07] Shaoying Liu and Hao Wang. An automated approach to specification animation for validation. The Journal of Systems and Software, 80 (8):1271–1285, August 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lu:2006:ESX

[LWC06] Eric Jui-Lin Lu, Bo-Chan Wu, and Po-Yun Chuang. An empirical study of XML data management in business information systems. The Journal of Systems and Software, 79(7):984– 1000, July 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lee:2005:ARM

[LWHS05] Min-Jae Lee, Kyu-Young Whang, Wook-Shin Han, and Il-Yeol Song. Adaptive row major order: a new space filling curve for efficient spatial join processing in the transform space. The Journal of Systems and Software, 78 (3):257–269, December 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Liu:2004:RBA

[LWL04] Duen-Ren Liu, Mei-Yu Wu, and Shu-Teng Lee. Rolebased authorizations for workflow systems in support of task-based separation of duty. The Journal of Systems and Software, 73(3):375–387, November/December 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lu:2009:ILD

[LWL09] Zhe-Ming Lu, Jun-Xiang Wang, and Bei-Bei Liu. An improved lossless data hiding scheme based on image VQ-index residual value coding. The Journal of Systems and Software, 82(6): 1016–1024, June 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Leung:2003:GTC

[LWN03] Karl R. P. H. Leung, Wai Wong, and Joseph Kee-Yin Ng. Generating test cases from class vectors. The Journal of Systems and Software, 66(1):35–46, April 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Laird:2003:DII

[LWS^{+}03] Shawn P. Laird, Johnny S. K.

Wong, William J. Schaller, Bradley J. Erickson, and Piet C. de Groen. Design and implementation of an Internet-based medical image viewing system. The Journal of Systems and Software, 66(2):167–181, May 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Li:2009:SPS

[LXG09] Y. F. Li, M. Xie, and T. N. Goh. A study of project selection and feature weighting for analogy based software cost estimation. The Journal of Systems and Software, 82 (2):241–252, February 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lung:2006:PRU

[LXZS06] Chung-Horng Lung, Xia Xu, Marzia Zaman, and Anand Srinivasan. Program restructuring using clustering techniques. The Journal of Systems and Software, 79(9): 1261–1279, September 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lou:2001:SDE

[LY01] Der-Chyuan Lou and Te-Lung Yin. Spatial database with each picture selfcontained multiscape and access control in a hierarchy. The Journal of Systems and Software, 56(2): 153–163, March 1, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/57/30/28/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/57/
30/28/article.pdf.

Lin:2009:FTD

[LY09] Jenn-Wei Lin and Ming-Feng Yang. Fault-tolerant design for wide-area Mobile IPv6 networks. The Journal of Systems and Software, 82(9): 1434–1446, September 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Li:2004:RCA

[LYC04] Hongchen Li, Yun Yang, and T. Y. Chen. Resource constraints analysis of workflow specifications. The Journal of Systems and Software, 73 (2):271–285, October 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Liu:2009:OSF

[LYX09] Xingchang Liu, Xiaofan Yang, and Min Xiang. Onestep t-fault diagnosis for hypermesh optical interconnection multiprocessor systems. The Journal of Systems and Software, 82(9): 1491–1496, September 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Liu:2006:SBP

[LZ06] Jie Liu and Hai Zhuge. A semantic-based P2P resource organization model R-Chord. The Journal of Systems and Software, 79(11): 1619–1631, November 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Liao:2007:CGA

[LZ07] Jianxin Liao and Xiaomin Zhu. Comments on Grigoris Antoniou and Frank Van Harmelen, “A Semantic Web Primer”, The MIT Press, Cambridge, Massachusetts, London, England (2004). The Journal of Systems and Software, 80(9): 1586–1587, September 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). See [AV04].

Lin:2007:DEP

[LZG07] Yongjing Lin, Youtao Zhang, and Rajiv Gupta. The design and evaluation of path matching schemes on compressed control flow traces. The Journal of Systems and Software, 80(3):396– 409, March 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Li:2006:SRS

[LZL^{+}06] Ge Li, Lu Zhang, Yan Li, Bing Xie, and Weizhong Shao. Shortening retrieval sequences in browsing-based component retrieval using information entropy. The Journal of Systems and Software, 79(2):216–230, February 2006. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Lung:2004:ACT

[LZN04] Chung-Horng Lung, Marzia Zaman, and Amit Nandi. Applications of clustering techniques to software partitioning, recovery and restructuring. The Journal of Systems and Software, 73(2): 227–244, October 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Li:2009:RCD

[LZXS09] Yan Li, Lu Zhang, Bing Xie, and Jiasu Sun. Refining component description by leveraging user query logs. The Journal of Systems and Software, 82(5): 751–758, May 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mittas:2008:CCP

[MA08] Nikolaos Mittas and Lefteris Angelis. Comparing cost prediction models by resampling techniques. The Journal of Systems and Software, 81(5): 616–632, May 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mathur:2009:ORC

[MA09] Vipul Mathur and Varsha Apte. An overhead and resource contention aware analytical model for overloaded Web servers. The Journal of Systems and Software, 82(1): 39–55, January 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Maqbool:2006:ASC

[MB06] O. Maqbool and H. A. Babri. Automated software clustering: an insight using cluster labels. The Journal of Systems and Software, 79(11): 1632–1648, November 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mansour:2001:ECR

[MBB01] Nashat Mansour, Rami Bahsoon, and Ghinwa Baradhi. Empirical comparison of regression test selection algorithms. The Journal of Systems and Software, 57 (1):79–90, April 27, 2001. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.elsevier.nl/gej%
                                        2Dng/10/29/11/64/25/32/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej%

2Dng/10/29/11/64/25/32/
article.pdf.

Marsh:2009:SPL

[MBM^{+}09] David W. Marsh, Rusty O.

Baldwin, Barry E. Mullins, Robert F. Mills, and Michael R. Grimaila. A security policy language for wireless sensor networks. The Journal of Systems and Software, 82(1): 101–111, January 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ma:2001:DRE

[MC01] Yung-Cheng Ma and Chung-Ping Chung. A dominance relation enhanced branchand-bound task allocation. The Journal of Systems and Software, 58(2):125–134, September 1, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/68/33/31/abstract.
html.

Min:2004:DEP

[MC04] Jun-Ki Min and Chin-Wan Chung. DIPS: an efficient pointer swizzling strategy for incremental uncaching environments. The Journal of Systems and Software, 71(1– 2):189–198, April 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

McBride:2008:MPM

[McB08] Tom McBride. The mechanisms of project management of software development. The Journal of Systems and Software, 81(12): 2386–2395, December 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ma:2002:PFP

[MCC02] Yung-Cheng Ma, Tien-Fu Chen, and Chung-Ping Chung. Posting file partitioning and parallel information retrieval. The Journal of Systems and Software, 63(2): 113–127, August 15, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ma:2003:VSD

[MCCC03] Yung-Cheng Ma, Jih-Ching Chiu, Tien-Fu Chen, and Chung-Ping Chung. Variablesize data item placement for load and storage balancing. The Journal of Systems and Software, 66(2): 157–166, May 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

McDonald:2002:SPM

[McD02] James McDonald. Software project management audits — update and experience report. The Journal of Systems and Software, 64(3):247–255, December 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mendes:2008:CCV

[MDFG08] Emilia Mendes, Sergio Di Martino, Filomena Ferrucci, and Carmine Gravino. Crosscompany vs. single-company Web effort models using the Tukutuku database: an extended study. The Journal of Systems and Software, 81 (5):673–690, May 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Miller:2006:CTA

[MDMC06] Scott D. Miller, Raymond A.

DeCarlo, Aditya P. Mathur, and João W. Cangussu. A control-theoretic approach to the management of the software system test phase. The Journal of Systems and Software, 79(11):1486– 1503, November 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Muccini:2006:SAB

[MDR06] Henry Muccini, Marcio Dias, and Debra J. Richardson. Software architecture-based regression testing. The Journal of Systems and Software, 79(10):1379–1396, October 2006. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mead:2009:SEE

[Mea09] Nancy R. Mead. Software engineering education: How far we’ve come and how far we have to go. The Journal of Systems and Software, 82 (4):571–575, April 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Manimaran:2005:PDR

[MEH05] G. Manimaran, Klaus Ecker, and Eui-Nam Huh. Parallel and distributed realtime systems. The Journal of Systems and Software, 77(1):1–2, July 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Misic:2004:SAA

[MG04] Mark M. Misic and David K. Graf. Systems analyst activities and skills in the new millennium. The Journal of Systems and Software, 71(1– 2):31–36, April 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Medvidovic:2003:BMA

[MGEB03] Nenad Medvidovic, Paul Grünbacher, Alexander Egyed, and Barry W. Boehm. Bridging models across the software lifecycle. The Journal of Systems and Software, 68(3):

199–215, December 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mokhtar:2007:CCB

[MGI07] Sonia Ben Mokhtar, Nikolaos Georgantas, and Valérie Issarny. COCOA: COnversationbased service COmposition in pervAsive computing environments with QoS support. The Journal of Systems and Software, 80(12):1941– 1955, December 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Munch:2004:SPC

[MH04] Jürgen Münch and Jens Heidrich. Software project control centers: concepts and approaches. The Journal of Systems and Software, 70(1– 2):3–19, February 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

McCann:2000:KAI

[MHC00] J. A. McCann, P. Howlett, and J. S. Crane. Kendra: adaptive Internet system. The Journal of Systems and Software, 55(1):3–17, November 5, 2000. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/54/28/25/abstract.
html;   http://www.elsevier.

nl/gej-ng/10/29/11/54/
28/25/article.pdf.

Mikler:2001:AAC

[MHW01] Armin R. Mikler, Vasant Honavar, and Johnny S. K. Wong. Autonomous agents for coordinated distributed parameterized heuristic routing in large dynamic communication networks. The Journal of Systems and Software, 56(3):231–246, March 15, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/57/29/26/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/57/
29/26/article.pdf.

Miller:2000:AMA

[Mil00a]  James Miller.  Applying

meta-analytical procedures to software engineering experiments. The Journal of Systems and Software, 54(1): 29–39, September 30, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/53/24/
28/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/53/24/abstract.
html.

Millet:2000:IQA

[Mil00b]  Ido Millet.  Improving

the quality of the analysis phase. The Journal of Systems and Software, 51 (2):151–153, April 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/50/25/
30/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/50/25/abstract.
html.

Miller:2002:ISI

[Mil02] J. Miller. On the independence of software inspectors. The Journal of Systems and Software, 60(1):5– 10, January 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/72/27/29/abstract.
html.

Miller:2004:SST

[Mil04] James Miller. Statistical significance testing — a panacea for software technology experiments? The Journal of Systems and Software, 73(2): 183–192, October 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Kim:2001:SSC

[mJKME01] Gyeung min J. Kim, Rohae Myung, and Ali F. Emdad. A simulation study on coordination strategies: decision cycle-time perspective. The Journal of Systems and Software, 57(1):61– 72, April 27, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.elsevier.nl/gej%
2Dng/10/29/11/64/25/30/
abstract.html;   http:
//www.elsevier.nl/gej%
2Dng/10/29/11/64/25/30/
article.pdf.

Mishra:2000:NTI

[MK00] Shivakant Mishra and Sudha M. Kuntur. Newsmonger: a technique to improve the performance of atomic broadcast protocols. The Journal of Systems and Software, 55(2):167–183, December 27, 2000. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/54/24/29/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/54/
24/29/article.pdf.

Mavromoustakis:2006:EPE

[MK06] Constandinos X. Mavromoustakis and Helen D. Karatza. On the efficiency and performance evaluation of the bandwidth clustering scheme for adaptive and reliable resource allocation. The Journal of Systems and Software, 79(8):1064– 1078, August 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mavromoustakis:2008:USC

[MK08] Constandinos X. Mavromoustakis and Helen D. Karatza. Under storage constraints of epidemic backup node selection using HyMIS architecture for data replication in mobile peer-to-peer networks. The Journal of Systems and Software, 81(1): 100–112, January 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Misra:2009:ISI

[MKK09] Subhas Chandra Misra, Vinod Kumar, and Uma Kumar. Identifying some important success factors in adopting agile software development practices. The Journal of Systems and Software, 82(11):1869–1890, November 2009. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mair:2000:IML

[MKL^{+}00] Carolyn Mair, Gada Kadoda, Martin Lefley, Keith Phalp, Chris Schofield, Martin Shepperd, and Steve Webster. An investigation of machine learning based prediction systems. The Journal of Systems and Software, 53(1):23–29, July 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/29/11/52/25/
                                        28/article.pdf;   http:/
                                        /www.elsevier.nl/gej-ng/
                                        10/29/11/52/25/abstract.
                                        html.

Mohapatra:2006:DDS

[MKM^{+}06] Durga P. Mohapatra, Rajeev Kumar, Rajib Mall, D. S. Kumar, and Mayank Bhasin. Distributed dynamic slicing of Java programs. The Journal of Systems and Software, 79(12): 1661–1678, December 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Malik:2005:MSC

[MKMS05] Kamran Malik, Osama Khan, Tahir Mobashir, and Mansoor Sarwar. Migratable sockets in cluster computing. The Journal of Systems and Software, 75(1–2):171– 177, February 15, 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Medvidovic:2006:UPI

[MKNS06] Nenad Medvidovic, René Krikhaar, Robert Nord, and Judith Stafford. Understanding the past, improving the present, and mapping out the future of software architecture. The Journal of Systems and Software, 79(12):1790– 1791, December 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mahaney:2003:ISP

[ML03] Robert C. Mahaney and Albert L. Lederer. Information systems project management: an agency theory interpretation. The Journal of Systems and Software, 68(1): 1–9, October 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Montesi:2008:SEA

[ML08] Michela Montesi and Patricia Lago. Software engineering article types: an analysis of the literature. The Journal of Systems and Software, 81(10):1694– 1714, October 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Marew:2009:TBA

[MLB09] Tegegne Marew, Joon-Sang Lee, and Doo-Hwan Bae. Tactics based approach for integrating non-functional requirements in object-oriented analysis and design. The Journal of Systems and Software, 82(10):1642–1656, October 2009. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Min:2009:EXE

[MLC09] Jun-Ki Min, Jihyun Lee, and Chin-Wan Chung. An efficient XML encoding and labeling method for query processing and updating on dynamic XML data. The Journal of Systems and Software, 82(3):503–515, March 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Miller:2000:EIA

[MM00a] J. Miller and F. Macdonald. An empirical incremental approach to tool evaluation and improvement. The Journal of Systems and Software, 51(1):19–35, April 1, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/50/26/
27/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/50/26/abstract.
html.

Mustafa:2000:CCB

[MM00b] Yousif Mustafa and Olugnenga Mejabi. Changing class behaviors at run-time in MRP systems. The Journal of Systems and Software, 55(1):67–71, November 5, 2000. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/54/28/30/abstract.

html;   http://www.elsevier.
nl/gej-ng/10/29/11/54/
28/30/article.pdf.

Middleton:2001:MPI

[MM01a] Peter Middleton and Barry McCollum. Management of process improvement by prescription. The Journal of Systems and Software, 57 (1):9–19, April 27, 2001. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.elsevier.nl/gej%
2Dng/10/29/11/64/25/26/
abstract.html;   http:
//www.elsevier.nl/gej%
2Dng/10/29/11/64/25/26/
article.pdf.

Mustafa:2001:HGC

[MM01b] Suleiman Hussein Mustafa and Fawaz Ahmed Masoud. How good is the critical factor concept in estimating the average number of character comparisons per item in string sorting? The Journal of Systems and Software, 57(3): 221–226, July 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/64/32/32/abstract.
html.

Mund:2006:EID

[MM06] G. B. Mund and Rajib Mall. An efficient interprocedural dynamic slicing method. The Journal of Systems and Software, 79(6): 791–806, June 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mendes:2005:IWS

[MMC05] Emilia Mendes, Nile Mosley, and Steve Counsell. Investigating Web size metrics for early Web cost estimation. The Journal of Systems and Software, 77(2): 157–172, August 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Madria:2000:OSN

[MMCB00] Sanjay Kumar Madria, S. N. Maheshwari, B. Chandra, and Bharat Bhargava. An open and safe nested transaction model: concurrency and recovery. The Journal of Systems and Software, 55(2):151–165, December 27, 2000. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        29/11/54/24/28/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/29/11/54/
                                        24/28/article.pdf.

Manimaran:2000:DTE

[MMM00] G. Manimaran, Anand Manikutty, and C. Siva Ram Murthy. DHARMA: a tool for evaluating dynamic scheduling algorithms for real-time multiprocessor systems. The Journal of Systems and Software, 50(2):131–149, February 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/49/27/
29/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/49/27/abstract.
html.

Meso:2006:KME

[MMTL06] Peter Meso, Gregory Madey, Marvin D. Troutt, and Jens Liegle. The knowledge management efficacy of matching information systems development methodologies with application characteristics— an experimental study. The Journal of Systems and Software, 79(1):15–28, January 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Masiero:2008:E

[MN08] Paulo C. Masiero and Mario A. Nascimento. Editorial. The Journal of Systems and Software, 81(3): 305–306, March 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Molokken-Ostvold:2008:UPP

[MØHB08] Kjetil Moløkken-Østvold, Nils Christian Haugen, and Hans Christian Benestad.

Using planning poker for combining expert estimates in software projects. The Journal of Systems and Software, 81(12):2106–2117, December 2008. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Moynihan:2000:CRU

[Moy00]  Tony Moynihan.  Cop-

ing with ‘requirementsuncertainty’: the theoriesof-action of experienced IS/ software project managers. The Journal of Systems and Software, 53(2):99–109, August 31, 2000. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/52/28/
25/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/52/28/abstract.
html.

Mokhtar:2008:EES

[MPG^{+}08] Sonia Ben Mokhtar, Davy Preuveneers, Nikolaos Georgantas, Valérie Issarny, and Yolande Berbers. EASY: Efficient semAntic Service discoverY in pervasive computing environments with QoS and context support. The Journal of Systems and Software, 81(5):785–808, May 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Makris:2006:EAD

[MPST06] Christos Makris, Yannis Panagis, Evangelos Sakkopoulos, and Athanasios Tsakalidis. Efficient and adaptive discovery techniques of Web Services handling large data sets. The Journal of Systems and Software, 79(4): 480–495, April 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Misic:2000:RBL

[MR00a] Mark M. Misic and Nancy L.

Russo. Reading between the lines: an examination of systems analysis and design texts. The Journal of Systems and Software, 50(1):65–73, January 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/49/25/
30/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/49/25/abstract.
html.

Morasca:2000:HAA

[MR00b] Sandro Morasca and Günther Ruhe. A hybrid approach to analyze empirical software engineering data and its application to predict module fault-proneness in maintenance. The Journal of Systems and Software, 53 (3):225–237, September 15, 2000. CODEN JSSODM.

ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/29/11/52/29/
                                        28/article.pdf;   http:/
                                        /www.elsevier.nl/gej-ng/
                                        10/29/11/52/29/abstract.
                                        html.

Martin:2001:AHP

[MR01] Robert Martin and David Raffo. Application of a hybrid process simulation model to a software development project. The Journal of Systems and Software, 59(3):237–246, December 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        29/11/71/34/29/abstract.
                                        html.

MacDonell:2003:CTO

[MS03] Stephen G. MacDonell and Martin J. Shepperd. Combining techniques to optimize effort predictions in software project management. The Journal of Systems and Software, 66(2): 91–98, May 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mendonca:2008:CSS

[MSA08] Nabor C. Mendonça, José Airton F. Silva, and Ricardo O. Anido. Clientside selection of replicated Web services: an empirical assessment. The Journal of Systems and Software, 81 (8):1346–1363, August 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Morisio:2002:CBS

[MSB^{+}02] M. Morisio, C. B. Seaman, V. R. Basili, A. T. Parra, S. E. Kraft, and S. E. Condon. COTS-based software development: Processes and open issues. The Journal of Systems and Software, 61(3): 189–199, April 1, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Shen:2005:NIW

[mSgFtL05] Rui min Shen, Yong gang Fu, and Hong tao Lu. A novel image watermarking scheme based on support vector regression. The Journal of Systems and Software, 78(1): 1–8, October 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ma:2007:WEC

[MT07] Li Ma and Jeff Tian. Web error classification and analysis for reliability improvement. The Journal of Systems and Software, 80(6): 795–804, June 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Muller:2005:TCE

[Mül05] Matthias M. Müller. Two controlled experiments concerning the comparison of pair programming to peer review. The Journal of Systems and Software, 78 (2):166–179, November 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Muller:2007:DPP

[Mül07] Matthias M. Müller. Do programmer pairs make different mistakes than solo programmers? The Journal of Systems and Software, 80(9): 1460–1471, September 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Murrill:2008:EPO

[Mur08] Branson W. Murrill. An empirical, path-oriented approach to software analysis and testing. The Journal of Systems and Software, 81(2): 249–261, February 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mustafa:2003:MDS

[Mus03] Suleiman H. Mustafa. A morphology-driven string matching approach to Arabic text searching. The Journal of Systems and Software, 67 (2):77–87, August 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Manvi:2005:ABA

[MV05] S. S. Manvi and P. Venkataram.

An agent based adaptive bandwidth allocation scheme for multimedia applications. The Journal of Systems and Software, 75(3):305– 318, March 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Manvi:2006:ABS

[MV06] S. S. Manvi and P. Venkataram.

An agent based synchronization scheme for multimedia applications. The Journal of Systems and Software, 79 (5):701–713, May 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

McGrew:2009:DVC

[MV09] Robert Wesley McGrew and Rayford B. Vaughn. Discovering vulnerabilities in control system human–machine interface software. The Journal of Systems and Software, 82(4):583–589, April 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mesbah:2008:CPB

[MvD08] Ali Mesbah and Arie van Deursen. A componentand push-based architectural style for AJAX. The Journal of Systems and Software, 81(12):2194–2209, December 2008. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Mathur:2008:MGE

[MW08] Aditya Mathur and Johnny Wong. Message from the Guest Editors. The Journal of Systems and Software, 81(10):1623–1624, October 2008. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ma:2006:QAC

[MYZC06] Hui Ma, I.-Ling Yen, Jia Zhou, and Kendra Cooper. QoS analysis for componentbased embedded software: Model and methodology. The Journal of Systems and Software, 79(6):859–870, June 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Naedele:2001:AME

[Nae01] Martin Naedele. An approach to modeling and evaluation of functional and timing specifications of real-time systems. The Journal of Systems and Software, 57(2): 155–174, June 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/

29/11/64/29/31/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/64/
29/31/article.pdf.

Nuseibeh:2001:MIR

[NER01] Bashar Nuseibeh, Steve Easterbrook, and Alessandra Russo. Making inconsistency respectable in software development. The Journal of Systems and Software, 58(2):171–180, September 1, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/68/33/35/abstract.
html.

Noh:2008:BTD

[NG08] Seo-Young Noh and Shashi K.

Gadia. Benchmarking temporal database models with interval-based and temporal element-based timestamping. The Journal of Systems and Software, 81(11):1931– 1943, November 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ng:2002:ECB

[NGC02] Celeste See Pui Ng, Guy G.

Gable, and Taizan Chan. An ERP-client benefit-oriented maintenance taxonomy. The Journal of Systems and Software, 64(2):87–109, November 15, 2002. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Noh:2008:XBM

[NGM08] Seo-Young Noh, Shashi K. Gadia, and Shihe Ma. An XML-based methodology for parametric temporal database model implementation. The Journal of Systems and Software, 81(6): 929–948, June 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Needham:2007:SFT

[NJ07] D. M. Needham and S. A. Jones. A software fault tree key node metric. The Journal of Systems and Software, 80(9):1530–1540, September 2007. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Nou:2009:AQC

[NKJT09] Ramon Nou, Samuel Kounev, Ferran Julià, and Jordi Torres. Autonomic QoS control in enterprise Grid environments using online simulation. The Journal of Systems and Software, 82(3): 486–502, March 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Nam:2005:DBG

[NLKW05] Junghyun Nam, Jinwoo Lee, Seungjoo Kim, and Dongho Won. DDH-based group key agreement in a mobile environment. The Journal of Systems and Software, 78(1): 73–83, October 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Na:2004:UPS

[NLSK04] Kwan-Sik Na, Xiaotong Li, James T. Simpson, and Ki-Yoon Kim. Uncertainty profile and software project performance: a cross-national comparison. The Journal of Systems and Software, 70(1– 2):155–163, February 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Nazareth:2004:ACE

[NR04] Derek L. Nazareth and Marcus A. Rothenberger. Assessing the cost-effectiveness of software reuse: a model for planned reuse. The Journal of Systems and Software, 73(2):245–255, October 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Noor:2008:APL

[NRG08] Muhammad A. Noor, Rick Rabiser, and Paul Grünbacher. Agile product line planning: a collaborative approach and a case study. The Journal of Systems and Software, 81 (6):868–882, June 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ng:2000:PET

[NsL00] Joseph Kee-Yin Ng and Victor Chung sing Lee. Performance evaluation of transmission schemes for realtime traffic in a highspeed timed-token MAC network. The Journal of Systems and Software, 54(1): 41–60, September 30, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/53/24/
29/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/53/24/abstract.
html.

Na:2007:SDR

[NSL^{+}07] Kwan-Sik Na, James T. Simpson, Xiaotong Li, Tushar Singh, and Ki-Yoon Kim. Software development risk and project performance measurement: Evidence in Korea. The Journal of Systems and Software, 80(4): 596–605, April 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Niazi:2005:FAD

[NWZ05a] Mahmood Niazi, David Wilson, and Didar Zowghi. A framework for assisting the design of effective software process improvement implementation strategies. The Journal of Systems and Software, 78(2):204–222, November 2005. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Niazi:2005:MMI

[NWZ05b] Mahmood Niazi, David Wilson, and Didar Zowghi. A maturity model for the implementation of software process improvement: an empirical study. The Journal of Systems and Software, 74(2): 155–172, January 15, 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ng:2000:MSV

[NXS00] Joseph Kee-Yin Ng, Shuhua Xiong, and Hong Shen. A multi-server video-on-demand system with arbitraryrate playback support. The Journal of Systems and Software, 51(3):217–227, May 1, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/50/28/
30/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/50/28/abstract.
html.

Oliveira:2007:RLF

[OAdLC07] Toacy C. Oliveira, Paulo S. C. Alencar, Carlos J. P. de Lucena, and Donald D. Cowan. RDL: a language for framework instantiation representation. The Journal of Systems and Software, 80(11):1902–1929, November 2007. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ormandjieva:2008:EQM

[OAZ08] O. Ormandjieva, V. S. Alagar, and M. Zheng. Early quality monitoring in the development of real-time reactive systems. The Journal of Systems and Software, 81(10):1738–1753, October 2008. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

OBrien:2008:AST

[O’B08] Wayne O’Brien. Avoiding semantic and temporal gaps in developing software intensive systems. The Journal of Systems and Software, 81(11): 1997–2013, November 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ortin:2004:DAA

[OC04] Francisco Ortin and Juan Manuel Cueva. Dynamic adaptation of application aspects. The Journal of Systems and Software, 71(3): 229–243, May 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Otero:2005:ECD

[OD05] Mari Carmen Otero and José Javier Dolado. An empirical comparison of the dynamic modeling in OML and UML. The Journal of Systems and Software, 77(2): 91–102, August 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Omari:2007:EPM

[OFWP07] Tariq Omari, Greg Franks, Murray Woodside, and Amy Pan. Efficient performance models for layered server systems with replicated servers and parallel behaviour. The Journal of Systems and Software, 80(4): 510–527, April 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Opdahl:2001:GOM

[OHS01] Andreas L. Opdahl and Brian Henderson-Sellers. Grounding the OML metamodel in ontology. The Journal of Systems and Software, 57(2): 119–143, June 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/64/29/29/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/64/
29/29/article.pdf.

Oi:2008:LVA

[Oi08] Hitoshi Oi. Local variable access behavior of a hardwaretranslation based Java virtual machine. The Journal of Systems and Software, 81(11):2059–2068, November 2008. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Oluyomi:2008:DTA

[OKS08] Ayodele Oluyomi, Shanika Karunasekera, and Leon Sterling. Description templates for agent-oriented patterns. The Journal of Systems and Software, 81(1): 20–36, January 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ozogul:2009:ROA

[ÖKT09] C. Okan Özogul, E. Ertugrul Karsak, and Ethem Tolga. A real options approach for evaluation and justification of a hospital information system. The Journal of Systems and Software, 82(12):2091– 2102, December 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Owei:2002:ACB

[ONR02] Vesper Owei, Shamkant B. Navathe, and Hyeun-Suk Rhee. An abbreviated concept-based query language and its exploratory evaluation. The Journal of Systems and Software, 63(1): 45–67, July 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Oztekin:2009:UAM

[ONZ09] Asil Oztekin, Alexander Nikov, and Selim Zaim. UWIS: an assessment methodology for usability of Webbased information systems. The Journal of Systems and Software, 82(12):2038– 2050, December 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

OKeeffe:2008:SBR

[OÓ08] Mark O’Keeffe and Mel Ó Cinnéide. Search-based refactoring for software maintenance. The Journal of Systems and Software, 81 (4):502–516, April 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ohishi:2009:GSR

[OOD09] Koji Ohishi, Hiroyuki Okamura, and Tadashi Dohi. Gompertz software reliability model: Estimation algorithm and empirical validation. The Journal of Systems and Software, 82(3): 535–543, March 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Orr:2000:FPC

[OR00] George Orr and Thomas E. Reeves. Function point counting: one program’s experience. The Journal of Systems and Software, 53 (3):239–244, September 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/52/29/
29/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/52/29/abstract.
html.

Obaidat:2009:NES

[OS09] Mohammad S. Obaidat and Mukund Sundararajan. New enhancements to the SOCKS communication network security protocol: Schemes and performance evaluation. The Journal of Systems and Software, 82(12):1941– 1949, December 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Oh:2004:GAB

[OW04] Jaewon Oh and Chisu Wu. Genetic-algorithm-based real-time task scheduling with multiple goals. The Journal of Systems and Software, 71(3):245–258, May 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ozmen:2009:EBA

[Özm09] Ahmet Özmen. An entropybased algorithm for data elimination in time-driven software instrumentation. The Journal of Systems and Software, 82(5):907–913, May 2009. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Park:2000:SRS

[Par00] Young Park. Software retrieval by samples using concept analysis. The Journal of Systems and Software, 54(3):179–183, November 1, 2000. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/53/28/26/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/53/
28/26/article.pdf.

Pighin:2000:FEI

[PB00] Maurizio Pighin and Giorgio Brajnik. A formative evaluation of information retrieval techniques applied to software catalogues. The Journal of Systems and Software, 52 (2–3):131–138, June 1, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/51/27/
30/article.pdf;   http:/

                                        /www.elsevier.nl/gej-ng/
                                        10/29/11/51/27/abstract.
                                        html.

Pont:2004:DES

[PB04] Michael J. Pont and Mark P. Banner. Designing embedded systems using patterns: a case study. The Journal of Systems and Software, 71 (3):201–213, May 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pean:2001:ONE

[PC01] Der-Lin Pean and Cheng Chen. ODCHP: a new effective mechanism to maximize parallelism of nested loops with non-uniform dependences. The Journal of Systems and Software, 56(3):279–297, March 15, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        29/11/57/29/29/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/29/11/57/
                                        29/29/article.pdf.

Park:2002:EQP

[PC02] Chang-Won Park and Chin-Wan Chung. An effective query pruning technique for multiple regular path expressions. The Journal of Systems and Software, 64(3): 219–233, December 15, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Park:2004:FAH

[PC04] Moonju Park and Yookun Cho. Feasibility analysis of hard real-time periodic tasks. The Journal of Systems and Software, 73(1):89– 100, September 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Park:2002:HAE

[PCC02] Ho-Hyun Park, Hyung-Ju Cho, and Chin-Wan Chung. Heuristic approach for early separated filter and refinement strategy in spatial query optimization. The Journal of Systems and Software, 62(3):161–179, June 15, 2002. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pepper:2002:PLS

[PCDG02] P. Pepper, M. Cebulla, K. Didrich, and W. Grieskamp. From program languages to software languages. The Journal of Systems and Software, 60(2):91–101, February 1, 2002. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/72/28/28/abstract.
html.

Pereira:2008:WDS

[PCV^{+}08] Javier Pereira, Narciso Cerpa, June Verner, Mario Rivas, and J. Drew Procaccino. What do software practitioners really think about project success: a cross-cultural comparison. The Journal of Systems and Software, 81(6): 897–907, June 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Parrish:2001:CFC

[PDC01] Allen Parrish, Brandon Dixon, and David Cordes. A conceptual foundation for component-based software deployment. The Journal of Systems and Software, 57(3): 193–200, July 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/64/32/28/abstract.
html.

Pfleeger:2000:RBW

[Pfl00] Shari Lawrence Pfleeger. Risky business: what we have yet to learn about risk management. The Journal of Systems and Software, 53 (3):265–273, September 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/52/29/
31/article.pdf;   http:/
/www.elsevier.nl/gej-ng/

10/29/11/52/29/abstract.
html.

Porwal:2004:EEW

[PG04] Rabins Porwal and Gursaran. An experimental evaluation of weak-branch criterion for class testing. The Journal of Systems and Software, 70(1–2):209– 224, February 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Papadopoulos:2005:ECD

[PG05] Yiannis Papadopoulos and Christian Grante. Evolving car designs using modelbased automated safety analysis and optimisation techniques. The Journal of Systems and Software, 76(1): 77–89, April 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Paulish:2008:E

[PGTS08] Dan Paulish, Ian Gorton, Jeff Tyree, and Dilip Soni. Editorial. The Journal of Systems and Software, 81(9): 1441–1442, September 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pai:2006:SRF

[PH06] Ping-Feng Pai and Wei-Chiang Hong. Software reliability forecasting by support vector machines with simulated annealing algorithms. The Journal of Systems and Software, 79(6): 747–755, June 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Philippi:2007:CSS

[PH07] Stephan Philippi and Hermann Josef Hill. Communication support for systems engineering — process modelling and animation with APRIL. The Journal of Systems and Software, 80 (8):1305–1316, August 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Philippi:2004:FBM

[Phi04] Stephan Philippi. Formally based modeling and inheritance of behaviour in objectoriented systems. The Journal of Systems and Software, 70(1–2):129–142, February 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Philippi:2005:MDG

[Phi05] Stephan Philippi. Model driven generation and testing of object-relational mappings. The Journal of Systems and Software, 77(2): 193–207, August 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Philippi:2006:ACG

[Phi06] Stephan Philippi. Automatic code generation from highlevel Petri-Nets for model driven systems engineering. The Journal of Systems and Software, 79(10):1444– 1455, October 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Park:2008:UVF

[PHN08] JinKyu Park, EenJun Hwang, and Yunyoung Nam. Utilizing venation features for efficient leaf image retrieval. The Journal of Systems and Software, 81(1):71–82, January 2008. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pettersson:2008:PGL

[PIGÖ08] F. Pettersson, M. Ivarsson, T. Gorschek, and P. Öhman. A practitioner’s guide to light weight software process assessment and improvement planning. The Journal of Systems and Software, 81(6): 972–995, June 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Penna:2006:XES

[PILO06] Giuseppe Della Penna, Benedetto Intrigila, Anna Rita Laurenzi, and Sergio Orefice. An XML environment for scenario based requirements engineering. The Journal of Systems and Software, 79(3): 379–403, March 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Park:2009:EEM

[PJ09] Sunggeun Park and Sungwon Jung. An energy-efficient mobile transaction processing method using random back-off in wireless broadcast environments. The Journal of Systems and Software, 82(12):2012–2022, December 2009. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Park:2001:OSA

[PK01a] Hee-Jun Park and Byung Kook Kim. An optimal scheduling algorithm for minimizing the computing period of cyclic synchronous tasks on multiprocessors. The Journal of Systems and Software, 56(3):213–229, March 15, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/57/29/25/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/57/
29/25/article.pdf.

Post:2001:DMS

[PK01b] Gerald Post and Albert Kagan. Database management systems: design considerations and attribute facilities. The Journal of Systems and Software, 56(2): 183–193, March 1, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/57/30/30/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/57/
30/30/article.pdf.

Park:2002:AAI

[PK02a] Kiejin Park and Sungsoo Kim. Availability analysis and improvement of Active/Standby cluster systems using software rejuvenation. The Journal of Systems and Software, 61(2): 121–128, March 15, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Park:2002:CPM

[PK02b] Kiejin Park and Sungsoo Kim. A capacity planning model of unreliable multimedia service systems. The Journal of Systems and Software, 63(1):69–76, July 15, 2002. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Park:2002:SEX

[PK02c] Sangwon Park and Hyoung-Joo Kim. SigDAQ: an enhanced XML query optimization technique. The Journal of Systems and Software, 61 (2):91–103, March 15, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Punter:2009:SET

[PKB09] Teade Punter, René L. Krikhaar, and Reinder J. Bril. Software engineering technology innovation — Turning research results into industrial success. The Journal of Systems and Software, 82(6):993–1003, June 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Park:2003:UBC

[PKL03] Chang-Sup Park, Myoung Ho Kim, and Yoon-Joon Lee. Usability-based caching of query results in OLAP systems. The Journal of Systems and Software, 68(2): 103–119, November 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pfahl:2001:CMI

[PKR01] Dietmar Pfahl, Marco Klemm, and Günther Ruhe. A CBT module with integrated simulation component for software project management education and training. The Journal of Systems and Software, 59(3):283–298, December 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/71/34/33/abstract.
html.

Paes:2009:EDH

[PLCC09] Rodrigo Paes, Carlos Lucena, Gustavo Carvalho, and Don Cowan. An eventdriven high level model for the specification of laws in open multi-agent systems. The Journal of Systems and Software, 82(4): 629–642, April 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Poon:2005:PSI

[PLF05] Wing-Fai Poon, Kwok-Tung Lo, and J. Feng. Performance study on implementation of VCR functionality in staggered broadcast video-on-demand systems. The Journal of Systems and Software, 75(1–2): 95–107, February 15, 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Peng:2007:MEO

[PLM07] Yi Peng, Fu Li, and Ali Mili.

Modeling the evolution of operating systems: An empirical study. The Journal of Systems and Software, 80(1): 1–15, January 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pretschner:2004:MBT

[PLP04] Alexander Pretschner, Heiko Lötzbeyer, and Jan Philipps. Model based testing in incremental system development. The Journal of Systems and Software, 70(3): 315–329, March 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pulkkinen:2007:MIS

[PNL07] Mirja Pulkkinen, Anton Naumenko, and Kari Luostarinen. Managing information security in a business network of machinery maintenance services business — Enterprise architecture as a coordination tool. The Journal of Systems and Software, 80(10):1607– 1620, October 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Purhonen:2004:VDS

[PNM04] Anu Purhonen, Eila Niemelä, and Mari Matinlassi. Viewpoints of DSP software and service architectures. The Journal of Systems and Software, 69(1–2):57–73, January 1, 2004. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pons:2003:WAC

[Pon03] Alexander P. Pons. Webapplication centric object prefetching. The Journal of Systems and Software, 67(3):193–200, September 15, 2003. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pons:2005:IPC

[Pon05] Alexander P. Pons. Improving the performance of client Web object retrieval. The Journal of Systems and Software, 74(3):303–311, February 1, 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pons:2006:SPO

[Pon06] Alexander P. Pons. Semantic prefetching objects of slower Web site pages. The Journal of Systems and Software, 79(12):1715– 1724, December 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Prowell:2004:CSR

[PP04] S. J. Prowell and J. H. Poore.

Computing system reliability using Markov chain usage models. The Journal of Systems and Software, 73(2): 219–225, October 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Priyadarshini:2004:PDS

[PQLN04] Pallavi Priyadarshini, Fengqiong Qin, Ee-Peng Lim, and Wee-Keong Ng. Parameter driven synthetic Web database generation. The Journal of Systems and Software, 69(1–2): 29–42, January 1, 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Phalp:2000:QAS

[PS00] Keith Phalp and Martin Shepperd. Quantitative analysis of static models of processes. The Journal of Systems and Software, 52(2– 3):105–112, June 1, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/29/11/51/27/
                                        27/article.pdf;   http:/
                                        /www.elsevier.nl/gej-ng/
                                        10/29/11/51/27/abstract.
                                        html.

Pedrycz:2005:GGC

[PS05] Witold Pedrycz and Giancarlo Succi. Genetic granular classifiers in modeling software quality. The Journal of Systems and Software, 76 (3):277–285, June 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Park:2009:FFA

[PS09] Hyoungmin Park and Kyuseok Shim. FAST: Flash-aware external sorting for mobile database systems. The Journal of Systems and Software, 82(8):1298–1312, August 2009. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pustina:2009:PAP

[PSG^{+}09] Lukas Pustina, Simon Schwarzer, Michael Gerharz, Peter Martini, and Volker Deichmann. A practical approach for performancedriven UML modelling of handheld devices — a case study. The Journal of Systems and Software, 82(1): 75–88, January 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Park:2006:ADD

[PSH06] KwangJin Park, Moon-Bae Song, and Chong-Sun Hwang. Adaptive data dissemination schemes for location-aware mobile services. The Journal of Systems and Software, 79(5): 674–688, May 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Park:2005:AIM

[PSK05] Hyun Kyoo Park, Jin Hyun Son, and Myoung Ho Kim. Adaptive index management for future location-based queries. The Journal of Systems and Software, 74(3): 313–324, February 1, 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pedrycz:2001:USO

[PSMB01] Witold Pedrycz, Giancarlo Succi, Petr Musílek, and Xiao Bai. Using selforganizing maps to analyze object-oriented software measures. The Journal of Systems and Software, 59(1):65–82, October 15, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        29/11/71/27/32/abstract.
                                        html.

Pelliccione:2008:AAC

[PTBP08] Patrizio Pelliccione, Massimo Tivoli, Antonio Bucchiarone, and Andrea Polini. An architectural approach to the correct and automatic assembly of evolving component-based systems. The Journal of Systems and Software, 81(12): 2237–2251, December 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Poo:2000:EOC

[PTK00] Danny C. C. Poo, Tech-Kang Toh, and Christopher S. G. Khoo. Enhancing online catalog searches with an electronic referencer. The Journal of Systems and Software, 55(2):203–219, December 27, 2000. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/54/24/32/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/54/
24/32/article.pdf.

Papadimitriou:2008:RCR

[PTM08] Panagiotis Papadimitriou, Vassilis Tsaoussidis, and Lefteris Mamatas. A receivercentric rate control scheme for layered video streams in the Internet. The Journal of Systems and Software, 81(12):2396–2412, December 2008. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Petersson:2004:CRS

[PTRW04] Håkan Petersson, Thomas Thelin, Per Runeson, and Claes Wohlin. Capture– recapture in software inspections after 10 years research

— theory, evaluation and application. The Journal of Systems and Software, 72(2): 249–264, July 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Prechelt:2003:CEI

[PUPT03] Lutz Prechelt, Barbara Unger, Michael Philippsen, and Walter Tichy. A controlled experiment on inheritance depth as a cost factor for code maintenance. The Journal of Systems and Software, 65(2):115–126, February 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Procaccino:2006:SPM

[PV06] J. Drew Procaccino and June M. Verner. Software project managers and project success: an exploratory study. The Journal of Systems and Software, 79(11): 1541–1551, November 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Procaccino:2005:WDS

[PVSG05] J. Drew Procaccino, June M. Verner, Katherine M. Shelfer, and David Gefen. What do software practitioners really think about project success: an exploratory study. The Journal of Systems and Software, 78(2):194– 203, November 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Preiss:2003:TCM

[PW03] Otto Preiss and Alain Wegmann. Towards a composition model problem based on IEC61850. The Journal of Systems and Software, 65 (3):227–236, March 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Petersen:2009:CIA

[PW09] Kai Petersen and Claes Wohlin. A comparison of issues and advantages in agile and incremental development between state of the art and an industrial case. The Journal of Systems and Software, 82(9): 1479–1490, September 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Pean:2001:DSM

[PWCC01] Der-Lin Pean, Chao-Chin Wu, Huey-Ting Chua, and Cheng Chen. Design of a scalable multiprocessor architecture and its simulation. The Journal of Systems and Software, 58(2):135–152, September 1, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/68/33/32/abstract.
html.

Park:2006:EEL

[PWLH06] Young-Ho Park, Kyu-Young Whang, Byung Suk Lee, and Wook-Shin Han. Efficient evaluation of linear path expressions on large-scale heterogeneous XML documents using information retrieval techniques. The Journal of Systems and Software, 79(2): 180–190, February 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Qumer:2008:FSE

[QHS08] A. Qumer and B. Henderson-Sellers. A framework to support the evaluation, adoption and improvement of agile methods in practice. The Journal of Systems and Software, 81(11):1899– 1919, November 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Quante:2008:DOP

[QK08] Jochen Quante and Rainer Koschke. Dynamic object process graphs. The Journal of Systems and Software, 81 (4):481–501, April 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Qin:2003:HPD

[QL03] Biao Qin and Yunsheng Liu. High performance distributed real-time commit protocol. The Journal of Systems and Software, 68(2): 145–152, November 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Radenski:2004:AFC

[Rad04] Atanas Radenski. Anomalyfree component adaptation with class overriding. The Journal of Systems and Software, 71(1–2):37–48, April 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ravindran:2003:LDA

[Rav03] Binoy Ravindran. LMR, DTA: adaptive communication algorithms for asynchronous real-time distributed systems using token-ring networks. The Journal of Systems and Software, 65(1):51– 70, January 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Roblet:2002:FDD

[RDD02] Etienne Roblet, Khalil Drira, and Michel Diaz. Formal design and development of a Corba-based application for cooperative HTML group editing support. The Journal of Systems and Software, 60(2):113–127, February 1, 2002. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/72/28/30/abstract.
html.

Rout:2007:SRD

[REF^{+}07] Terence P. Rout, Khaled El Emam, Mario Fusani, Dennis Goldenson, and Ho-Won Jung. SPICE in retrospect: Developing a standard for process assessment. The Journal of Systems and Software, 80(9):1483– 1493, September 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Reifer:2000:CF

[Rei00]  Donald J. Reifer.  The

CMMI: it’s formidable. The Journal of Systems and Software, 50(2):97–98, February 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/49/27/
25/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/49/27/abstract.
html.

Reynolds:2007:MRU

[Rey07] Peter Reynolds. Managing requirements for a US$1bn IT-based business transformation: New approaches and challenges. The Journal of Systems and Software, 80(3): 285–293, March 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Radenski:2008:JGC

[RFZ08] Atanas Radenski, Jeff Furlong, and Vladimir Zanev. The Java 5 generics compromise orthogonality to keep compatibility. The Journal of Systems and Software, 81(11):2069–2078, November 2008. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Robles:2006:BSC

[RGBM06] Gregorio Robles, Jesus M.

Gonzalez-Barahona, and Juan Julian Merelo. Beyond source code: The importance of other artifacts in software development (a case study). The Journal of Systems and Software, 79(9): 1233–1248, September 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ramesh:2004:RCS

[RGV04] V. Ramesh, Robert L. Glass, and Iris Vessey. Research in computer science: an empirical study. The Journal of Systems and Software, 70(1– 2):165–176, February 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rainer:2002:KSF

[RH02] Austen Rainer and Tracy Hall. Key success factors for implementing software process improvement: a maturity-based analysis. The Journal of Systems and Software, 62(2):71–84, May 15, 2002. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rainer:2003:QQA

[RH03] Austen Rainer and Tracy Hall. A quantitative and qualitative analysis of factors affecting software processes. The Journal of Systems and Software, 66(1):7– 21, April 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rho:2006:FQA

[RH06] Seungmin Rho and Eenjun Hwang. FMF: Query adaptive melody retrieval system. The Journal of Systems and Software, 79(1): 43–56, January 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rho:2008:MNM

[RjHHK08] Seungmin Rho, Byeong jun Han, Eenjun Hwang, and Minkoo Kim. MUSEMBLE: a novel music retrieval system with automatic voice query transcription and reformulation. The Journal of Systems and Software, 81(7): 1065–1080, July 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Raffo:2000:EAS

[RK00] David M. Raffo and Marc I. Kellner. Empirical analysis in software process simulation modeling. The Journal of Systems and Software, 53(1):31–41, July 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/52/25/
29/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/52/25/abstract.
html.

Roeller:2006:RAA

[RLvV06] Ronny Roeller, Patricia Lago, and Hans van Vliet. Recovering architectural assumptions. The Journal of Systems and Software, 79(4): 552–573, April 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Raghunathan:2005:SAC

[RMC05] Subhashini Raghunathan, Armin R. Mikler, and Cliff Cozzolino. Secure agent computation: X.509 Proxy Certificates in a multi-lingual agent framework. The Journal of Systems and Software, 75(1–2):125–137, February 15, 2005. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rombach:2008:TDS

[RMO^{+}08] Dieter Rombach, Jürgen Münch, Alexis Ocampo, Watts S. Humphrey, and Dan Burton. Teaching disciplined software development. The Journal of Systems and Software, 81(5): 747–763, May 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ryan:2009:DTM

[RO09] Sharon Ryan and Rory V. O’Connor. Development of a team measure for tacit knowledge in software development teams. The Journal of Systems and Software, 82 (2):229–240, February 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Robillard:2000:TCW

[RR00] Pierre N. Robillard and Martin P. Robillard. Types of collaborative work in software engineering. The Journal of Systems and Software, 53 (3):219–224, September 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/52/29/
27/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/52/29/abstract.
html.

Ras:2009:UWS

[RR09] Eric Ras and Jörg Rech. Using Wikis to support the Net Generation in improving knowledge acquisition in capstone projects. The Journal of Systems and Software, 82 (4):553–562, April 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rosenfeld:2007:ABC

[RRC07] Shalom N. Rosenfeld, Ioana Rus, and Michel Cukier. Archetypal behavior in computer security. The Journal of Systems and Software, 80(10):1594–1606, October 2007. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rafla:2006:IIU

[RRD06] Tamer Rafla, Pierre N. Robillard, and Michel Desmarais. Investigating the impact of usability on software architecture through scenarios: a case study on Web systems. The Journal of Systems and Software, 79(3): 415–426, March 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ruiz:2001:SMS

[RRT01] Mercedes Ruiz, Isabel Ramos, and Miguel Toro. A simplified model of software project dynamics. The Journal of Systems and Software, 59(3):299–309, December 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.

elsevier.com/gej-ng/10/
29/11/71/34/34/abstract.
html.

Regnell:2000:TIU

[RRW00] Björn Regnell, Per Runeson, and Claes Wohlin. Towards integration of use case modelling and usage-based testing. The Journal of Systems and Software, 50(2): 117–130, February 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/49/27/
28/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/49/27/abstract.
html.

RodriguesdeCarvalho:2000:MIF

[RS00] Sergio Eduardo Rodrigues de Carvalho and Julio Cesar Sampaio do Prado Leite. Module interconnection features in object-oriented development tools. The Journal of Systems and Software, 50(1):57–64, January 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/49/25/
29/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/49/25/abstract.
html.

Ryoo:2006:AHA

[RS06] Jungwoo Ryoo and Hossein Saiedian. AVDL: a highly adaptable architecture view description language. The Journal of Systems and Software, 79 (8):1180–1206, August 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rajlich:2000:PCS

[RSM00] Václav Rajlich, Waseed Abdul Syed, and James Martinez. Perceptions of contribution in software teams. The Journal of Systems and Software, 54(1):61–63, September 30, 2000. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/53/24/
30/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/53/24/abstract.
html.

Reussner:2003:RPC

[RSP03] Ralf H. Reussner, Heinz W.

Schmidt, and Iman H. Poernomo. Reliability prediction for component-based software architectures. The Journal of Systems and Software, 66(3):241–252, June 15, 2003. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ramanujan:2000:EII

[RSS00] Sam Ramanujan, Richard W.

Scamell, and Jaymeen R.

Shah. An experimental investigation of the impact of individual, program, and organizational characteristics on software maintenance effort. The Journal of Systems and Software, 54(2): 137–157, October 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/53/25/
29/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/53/25/abstract.
html.

Riva:2007:DAS

[RT07] Oriana Riva and Santtu Toivonen. The DYNAMOS approach to support contextaware service provisioning in mobile environments. The Journal of Systems and Software, 80(12):1956–1972, December 2007. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Raptopoulou:2006:PTI

[RVM06] Katerina Raptopoulou, Michael Vassilakopoulos, and Yannis Manolopoulos. On pasttime indexing of moving objects. The Journal of Systems and Software, 79(8):1079– 1091, August 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rogers:2000:SRA

[RW00] P. Rogers and A. J. Wellings.

State restoration in Ada 95: a portable approach to supporting software fault tolerance. The Journal of Systems and Software, 50(3): 237–255, March 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/49/28/
30/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/49/28/abstract.
html.

Raffo:2001:GES

[RW01] David Raffo and Paul Wernick. Guest editorial: Software process simulation modelling. The Journal of Systems and Software, 59(3):223–225, December 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/71/34/27/abstract.
html.

Rim:2001:HCO

[RwJK01] Hwakyung Rim, Ju wook Jang, and Sungchun Kim. Hiding communication overheads in dynamic load balancing for multicomputers. The Journal of Systems and Software, 58(1):11–22, August 15, 2001. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        29/11/68/27/28/abstract.
                                        html.

Srinivasan:2005:FSD

[SA05] Anand Srinivasan and James H. Anderson. Fair scheduling of dynamic task systems on multiprocessors. The Journal of Systems and Software, 77(1):67–80, July 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sentas:2006:CMD

[SA06] Panagiotis Sentas and Lefteris Angelis. Categorical missing data imputation for software cost estimation by multinomial logistic regression. The Journal of Systems and Software, 79(3): 404–414, March 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sofokleous:2008:AET

[SA08] Anastasis A. Sofokleous and Andreas S. Andreou. Automatic, evolutionary test data generation for dynamic software testing. The Journal of Systems and Software, 81(11):1883–1898, November 2008. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Saiedian:2002:BPS

[Sai02] Hossein Saiedian. Best practices in software engineering. The Journal of Systems and Software, 61(3): 171–172, April 1, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Saiedian:2007:RIC

[Sai07] Hossein Saiedian. Reflections on the influences of the COCOMO, spiral and the win-win models on software project and risk management. The Journal of Systems and Software, 80 (8):1183–1184, August 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Saiedian:2009:SPI

[Sai09] Hossein Saiedian. Selected papers from the 2008 IEEE Conference on Software Engineering Education and Training (CSEET08). The Journal of Systems and Software, 82 (4):551–552, April 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Salmeron:2002:EDF

[Sal02] Jose L. Salmeron. EIS data:

findings from an evolutionary study. The Journal of Systems and Software, 64(2): 111–114, November 15, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Schreck:2000:BGM

[SC00] T. Schreck and Z. Chen. Branch grafting method for R-tree implementation. The Journal of Systems and Software, 53(1):83–93, July 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/52/25/
33/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/52/25/abstract.
html.

Subramanian:2001:ESC

[SC01] Girish Subramanian and William Corbin. An empirical study of certain objectoriented software metrics. The Journal of Systems and Software, 59(1):57–63, October 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/71/27/31/abstract.
html.

Shen:2007:SDI

[SC07] Jun-Hong Shen and Ye-In Chang. A skewed distributed indexing for skewed access patterns on the wireless broadcast. The Journal of Systems and Software, 80 (5):711–723, May 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Shen:2008:ENI

[SC08] Jun-Hong Shen and Ye-In Chang. An efficient nonuniform index in the wireless broadcast environments. The Journal of Systems and Software, 81(11): 2091–2103, November 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sun:2009:DGI

[SC09] Wen-Chen Sun and Yi-Ming Chen. On the design of a global intrusion tolerance network architecture against the Internet catastrophes. The Journal of Systems and Software, 82(8):1313– 1325, August 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Serrano:2002:RLS

[SCdO02] Miguel A. Serrano, Doris L.

Carver, and Carlos Montes de Oca. Reengineering legacy systems for distributed environments. The Journal of Systems and Software, 64 (1):37–55, October 15, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sardinha:2006:CSL

[SCdS^{+}06] José Alberto R. P. Sardinha, Ricardo Choren, Viviane Torres da Silva, Ruy Milidiú, and Carlos J. P. de Lucena. A combined specification language and development framework for agentbased application engineering. The Journal of Systems and Software, 79(11):1565– 1577, November 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Schmidt:2003:TCX

[Sch03] H. Schmidt. Trustworthy components — compositionality and prediction. The Journal of Systems and Software, 65(3):215–225, March 15, 2003. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sun:2005:SKA

[SCH05] Hung-Min Sun, Bing-Chang Chen, and Tzonelih Hwang. Secure key agreement protocols for three-party against guessing attacks. The Journal of Systems and Software, 75(1–2):63–68, February 15, 2005. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Shao:2007:IYA

[SCL07] Jun Shao, Zhenfu Cao, and Rongxing Lu. Improvement of Yang et al.’s threshold proxy signature scheme. The Journal of Systems and Software, 80(2):172–177, February 2007. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

St-Denis:2002:DRS

[SD02] Richard St-Denis. Designing reactive systems: integration of abstraction techniques into a synthesis procedure. The Journal of Systems and Software, 60(2):103–112, February 1, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/72/28/29/abstract.
html.

Sinnema:2008:IVC

[SD08] Marco Sinnema and Sybren Deelstra. Industrial validation of COVAMOF. The Journal of Systems and Software, 81(4):584–600, April 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Subramonian:2007:DPC

[SDG^{+}07] Venkita Subramonian, Gan Deng, Christopher Gill, Jaiganesh Balasubramanian, Liang-Jui Shen, William Otte, Douglas C. Schmidt, Aniruddha Gokhale, and Nanbor Wang. The design and performance of component middleware for QoS-enabled deployment and configuration of DRE systems. The Journal of Systems and Software, 80(5): 668–677, May 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Skotiniotis:2002:EIM

[SeMC02] Therapon Skotiniotis and Ji en Morris Chang. Estimating internal memory fragmentation for Java programs. The Journal of Systems and Software, 64(3): 235–246, December 15, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Seffah:2004:MDM

[SFJ04] Ahmed Seffah, Peter Forbrig, and Homa Javahery. Multidevices “multiple” user interfaces: development models and research opportunities. The Journal of Systems and Software, 73(2): 287–300, October 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Siemers:2005:RET

[SFSE05] Christian Siemers, Rainer Falsett, Reinhard Seyer, and Klaus Ecker. Reliable eventtriggered systems for mechatronic applications. The Journal of Systems and Software, 77(1):17–26, July 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Stallinger:2001:SDM

[SG01] Friedrich Stallinger and Paul Grünbacher. System dynamics modelling and simulation of collaborative requirements engineering. The Journal of Systems and Software, 59(3):311–321, December 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/71/34/35/abstract.
html.

Seceleanu:2006:DAS

[SG06] Tiberiu Seceleanu and David Garlan. Developing adaptive systems with synchronized architectures. The Journal of Systems and Software, 79(11):1514–1526, November 2006. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Su:2007:NNB

[SH07] Yu-Shen Su and Chin-Yu Huang. Neural-network-based approaches for software reliability estimation using dynamic weighted combinational models. The Journal of Systems and Software, 80(4):606–615, April 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Shama:2001:DCC

[Sha01] Avi Shama. Dot-coms’ coma.

The Journal of Systems and Software, 56(1):101–104, February 1, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        29/11/57/24/33/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/29/11/57/
                                        24/33/article.pdf.

Shao:2005:CXY

[Sha05] Zuhua Shao. Cryptanalysis of Xia–You group signature scheme. The Journal of Systems and Software, 75(1– 2):89–94, February 15, 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Shao:2007:SCS

[Sha07] Zuhua Shao. Self-certified signature scheme from pairings. The Journal of Systems and Software, 80(3): 388–395, March 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Shao:2009:IIB

[Sha09] Zuhua Shao. Improvement of identity-based proxy multisignature scheme. The Journal of Systems and Software, 82(5):794–800, May 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sheetz:2002:IDO

[She02] Steven D. Sheetz. Identifying the difficulties of objectoriented development. The Journal of Systems and Software, 64(1):23–36, October 15, 2002. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Senger:2007:EIC

[SHS^{+}07] Hermes Senger, Eduardo R.

Hruschka, Fabrício A. B. Silva, Liria M. Sato, Calebe P. Bianchini, and Bruno F. Jerosch. Exploiting idle cycles to execute data mining applications on clusters of PCs. The Journal of Systems and Software, 80(5): 778–790, May 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sun:2005:SSP

[SHT05] Hung-Min Sun, Bin-Tsan Hsieh, and Shin-Mu Tseng. On the security of some proxy blind signature schemes. The Journal of Systems and Software, 74(3):297–302, February 1, 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Shu:2003:ARB

[Shu03] LihChyun Shu. On avoiding remote blocking via real-time concurrency control protocols. The Journal of Systems and Software, 68(2):121–136, November 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Smidts:2002:PRS

[SHW02] Carol Smidts, Xin Huang, and James C. Widmaier. Producing reliable software: an experiment. The Journal of Systems and Software, 61(3):213–224, April 1, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sheetz:2009:UDM

[SHW09] Steven D. Sheetz, David Henderson, and Linda Wallace. Understanding developer and manager perceptions of function points and source lines of code. The Journal of Systems and Software, 82(9):1540– 1549, September 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Shyur:2003:SSR

[Shy03] Huan-Jyh Shyur. A stochastic software reliability model with imperfect-debugging and change-point. The Journal of Systems and Software, 66(2):135–141, May 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Schneider:2005:EPH

[SJ05] Jean-Guy Schneider and Lorraine Johnston. eXtreme Programming — helpful or harmful in educating undergraduates? The Journal of Systems and Software, 74(2): 121–132, January 15, 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Subramanian:2007:SQP

[SJK07] Girish H. Subramanian, James J. Jiang, and Gary Klein. Software quality and IS project performance improvements from software development process maturity and IS implementation strategies. The Journal of Systems and Software, 80(4): 616–627, April 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Son:2001:IPT

[SK01] Jin Hyun Son and Myoung Ho Kim. Improving the performance of time-constrained workflow processing. The Journal of Systems and Software, 58(3):211–219, September 15, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/68/34/29/abstract.
html.

Spanoudakis:2002:DSI

[SK02] George Spanoudakis and Hyoseob Kim. Diagnosis of the significance of inconsistencies in object-oriented designs: a framework and its experimental evaluation. The Journal of Systems and Software, 64(1):3–22, October 15, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Seo:2003:ISP

[SK03] Kyung-Hee Seo and Sung-Chun Kim. Improving system performance in contiguous processor allocation for mesh-connected parallel systems. The Journal of Systems and Software, 67(1):45– 54, July 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Son:2004:AVP

[SK04] Jin Hyun Son and Myoung Ho Kim. An adaptable vertical partitioning method in distributed systems. The Journal of Systems and Software, 73(3):551–561, November/December 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Soman:2007:ASG

[SK07] Sunil Soman and Chandra Krintz. Application-specific garbage collection. The Journal of Systems and Software, 80(7):1037–1056, July 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Skianis:2007:ESI

[SKKL07] C. Skianis, G. Kormentzas, K. Kontovasilis, and G. Lisa. Efficiency study of the information flow mechanism enabling interworking of heterogeneous wireless systems. The Journal of Systems and Software, 80(10):1746– 1753, October 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Staron:2006:EAU

[SKW06] Mirosław Staron, Ludwik Kuzniarz, and Claes Wohlin. Empirical assessment of using stereotypes to improve comprehension of UML models: a set of experiments. The Journal of Systems and Software, 79(5):727–742, May 2006. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sangpachatanaruk:2004:DAR

[SKZ^{+}04] Chatree Sangpachatanaruk, Sherif M. Khattab, Taieb Znati, Rami Melhem, and Daniel Mossé. Design and analysis of a replicated elusive server scheme for mitigating denial of service attacks. The Journal of Systems and Software, 73(1):15– 29, September 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Suh:2001:MBC

[SL01] Woojong Suh and Heeseok Lee. A methodology for building content-oriented hypermedia systems. The Journal of Systems and Software, 56(2):115–131, March 1, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        29/11/57/30/26/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/29/11/57/
                                        30/26/article.pdf.

Shin:2002:RSI

[SL02] Kitae Shin and Choon Seong Leem. A reference system for Internet based interenterprise electronic commerce. The Journal of Systems and Software, 60(3): 195–209, February 15, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/29/11/72/29/
                                        31/abstract.html.

Seruca:2003:TSA

[SL03] Isabel Seruca and Pericles Loucopoulos. Towards a systematic approach to the capture of patterns within a business domain. The Journal of Systems and Software, 67(1):1–18, July 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Spinellis:2007:FSV

[SL07] Diomidis Spinellis and Panagiotis Louridas. A framework for the static verification of API. The Journal of Systems and Software, 80(7): 1156–1168, July 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Shatnawi:2008:ESM

[SL08] Raed Shatnawi and Wei Li.

The effectiveness of software metrics in identifying error-prone classes in postrelease software evolution process. The Journal of Systems and Software, 81(11): 1868–1882, November 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Shiu:2000:ASS

[SLC00] R-Ming Shiu, Neng-Pin Lu, and Chung-Ping Chung. Applying stack simulation for branch target buffers. The Journal of Systems and Software, 52(1):67–78, May 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/

gej-ng/10/29/11/51/25/
31/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/51/25/abstract.
html.

Santos:2008:ISS

[SLS08] Rodrigo Santos, Giuseppe Lipari, and Jorge Santos. Improving the schedulability of soft real-time open dynamic systems: The inheritor is actually a debtor. The Journal of Systems and Software, 81(7):1093–1104, July 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Schollmeyer:2000:ERT

[SM00] Martina Schollmeyer and Bruce McMillin. Efficient run-time assurance in distributed systems through selection of executable assertions. The Journal of Systems and Software, 52 (1):33–54, May 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/51/25/
29/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/51/25/abstract.
html.

Striegel:2003:DCB

[SM03] A. Striegel and G. Manimaran. Dynamic class-based queue management for scalable media servers. The Journal of Systems and Software, 66(2):119–128, May 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Shi:2006:PEP

[SM06a] Weisong Shi and Yonggen Mao. Performance evaluation of peer-to-peer Web caching systems. The Journal of Systems and Software, 79(5): 714–726, May 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sidiropoulos:2006:GCG

[SM06b] Antonis Sidiropoulos and Yannis Manolopoulos. Generalized comparison of graphbased ranking algorithms for publications and authors. The Journal of Systems and Software, 79(12): 1679–1700, December 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Svahnberg:2007:SYE

[SM07] Mikael Svahnberg and Frans Mårtensson. Six years of evaluating software architectures in student projects. The Journal of Systems and Software, 80(11):1893– 1901, November 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sohn:2008:SAS

[SM08] So Young Sohn and Min Seok Mok. A strategic analysis for successful open source software utilization based on a structural equation model. The Journal of Systems and Software, 81(6): 1014–1024, June 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sadou:2009:DBA

[SM09] Salah Sadou and Hafedh Mili. A delegation-based approach for the unanticipated dynamic evolution of distributed objects. The Journal of Systems and Software, 82(6):932–946, June 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Shokoufandeh:2005:SMH

[SMDM05] Ali Shokoufandeh, Spiros Mancoridis, Trip Denton, and Matthew Maycock. Spectral and meta-heuristic algorithms for software clustering. The Journal of Systems and Software, 77(3): 213–223, September 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sadat-Mohtasham:2008:LHL

[SMG08] S. Hossein Sadat-Mohtasham and Ali A. Ghorbani. A language for high-level description of adaptive Web systems. The Journal of Systems and Software, 81(7): 1196–1217, July 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Seffah:2008:RUI

[SMHMA08] Ahmed Seffah, Taleb Mohamed, Halima Habieb-Mammar, and Alain Abran. Reconciling usability and interactive system architecture using patterns. The Journal of Systems and Software, 81(11):1845–1852, November 2008. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sarkar:2009:DAL

[SMR09] Santonu Sarkar, Girish Maskeri, and Shubha Ramachandran. Discovery of architectural layers and measurement of layering violations in source code. The Journal of Systems and Software, 82(11):1891–1905, November 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Staples:2007:EUS

[SN07] Mark Staples and Mahmood Niazi. Experiences using systematic review guidelines. The Journal of Systems and Software, 80(9): 1425–1437, September 2007.

CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sangwan:2008:ISA

[SNBH08] Raghvinder Sangwan, Colin Neill, Matthew Bass, and Zakaria El Houda. Integrating a software architecturecentric method into objectoriented analysis and design. The Journal of Systems and Software, 81(5): 727–746, May 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Staples:2007:ESW

[SNJ^{+}07] Mark Staples, Mahmood Niazi, Ross Jeffery, Alan Abrahams, Paul Byatt, and Russell Murphy. An exploratory study of why organizations do not adopt CMMI. The Journal of Systems and Software, 80(6): 883–895, June 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Schmidt:2003:PPD

[SO03] Douglas C. Schmidt and Carlos O’Ryan. Patterns and performance of distributed real-time and embedded publisher/subscriber architectures. The Journal of Systems and Software, 66(3): 213–223, June 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Son:2003:GWE

[SOC^{+}03] Jin Hyun Son, Seok Kyun Oh, Kyung Hoon Choi, Yoon Joon Lee, and Myoung Ho Kim. GM-WTA: An efficient workflow task allocation method in a distributed execution environment. The Journal of Systems and Software, 67 (3):165–179, September 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Short:2008:AHI

[SP08] Michael Short and Michael J.

Pont. Assessment of highintegrity embedded automotive control systems using hardware in the loop simulation. The Journal of Systems and Software, 81(7): 1163–1183, July 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Souliou:2006:CFI

[SPDT06] Dora Souliou, Aris Pagourtzis, Nikolaos Drosinos, and Panayiotis Tsanakas. Computing frequent itemsets in parallel using partial support trees. The Journal of Systems and Software, 79(12): 1735–1743, December 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Spinellis:2001:NDP

[Spi01] Diomidis Spinellis. Notable design patterns for domainspecific languages. The Journal of Systems and Software, 56(1):91–99, February 1, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/57/24/32/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/57/
24/32/article.pdf.

Succi:2003:PAM

[SPSM03] Giancarlo Succi, Witold Pedrycz, Milorad Stefanovic, and James Miller. Practical assessment of the models for identification of defect-prone classes in object-oriented commercial systems using design metrics. The Journal of Systems and Software, 65(1): 1–12, January 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sutcliffe:2006:CRA

[SPZ06] Alistair Sutcliffe, George Papamargaritis, and Liping Zhao. Comparing requirements analysis methods for developing reusable component libraries. The Journal of Systems and Software, 79(2): 273–289, February 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Shimizu:2009:PIM

[SRDLCP09] Shuichi Shimizu, Raju Rangaswami, Hector A. Duran-Limon, and Manuel Corona-Perez. Platform-independent modeling and prediction of application resource usage characteristics. The Journal of Systems and Software, 82(12):2117–2127, December 2009. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Santos:2008:WSB

[SRGL08] Rodrygo L. T. Santos, Pablo A. Roberto, Marcos André Gonçalves, and Alberto H. F. Laender. A Web services-based framework for building componentized digital libraries. The Journal of Systems and Software, 81 (5):809–822, May 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sridhar:2007:S

[Sri07] Nigamanth Sridhar. Serfs.

The Journal of Systems and Software, 80(5):736–749, May 2007. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sohn:2004:QES

[SS04] SeDo Sohn and PoongHyun Seong. Quantitative evaluation of safety critical software testability based on fault tree analysis and entropy. The Journal of Systems and Software, 73(2): 351–360, October 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Song:2007:NIM

[SS07] Qinbao Song and Martin Shepperd. A new imputation method for small software project data sets. The Journal of Systems and Software, 80(1):51–62, January 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sowe:2008:UKS

[SSA08] Sulayman K. Sowe, Ioannis Stamelos, and Lefteris Angelis. Understanding knowledge sharing activities in free/open source software projects: an empirical study. The Journal of Systems and Software, 81(3): 431–446, March 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Song:2008:CNI

[SSCL08] Qinbao Song, Martin Shepperd, Xiangru Chen, and Jun Liu. Can k-NN imputation improve the performance of C4.5 with small software project data sets? A comparative evaluation. The Journal of Systems and Software, 81(12):2361– 2370, December 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sanchez-Segura:2004:VRS

[SSCM^{+}04] María I. Sánchez-Segura, Juan J. Cuadrado, Ana-María Moreno, Antonio de Amescua, Angélica de Antonio, and Oscar Marbán. Virtual reality systems estimation vs. traditional systems estimation. The Journal of Systems and Software, 72(2):187–194, July 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sioutas:2009:DWS

[SSM^{+}09] S. Sioutas, E. Sakkopoulos, Ch. Makris, B. Vassiliadis, A. Tsakalidis, and P. Triantafillou. Dynamic Web Service discovery architecture based on a novel peer based overlay network. The Journal of Systems and Software, 82(5): 809–824, May 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Santos:2005:LUB

[SSO05] Rodrigo M. Santos, Jorge Santos, and Javier D. Orozco. A least upper bound on the fault tolerance of realtime systems. The Journal of Systems and Software, 78(1): 47–55, October 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sheetz:2001:ICC

[ST01] Steven D. Sheetz and David P. Tegarden. Illustrating the cognitive consequences of object-oriented systems development. The Journal of Systems and Software, 59(2):163–179, November 15, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        29/11/71/28/32/abstract.
                                        html.

Sharma:2007:QSP

[ST07] Vibhu Saujanya Sharma and Kishor S. Trivedi. Quantifying software performance, reliability and security: an architecture-based approach. The Journal of Systems and Software, 80(4):493– 509, April 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Stamelos:2002:LKC

[Sta02] I. Stamelos. Linguistic kleptomania in computer science. The Journal of Systems and Software, 64(2):171–172, November 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Stamelos:2003:DAS

[Sta03] Ioannis Stamelos. Detecting associative shift faults in predicate testing. The Journal of Systems and Software, 66(1):57–63, April 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Stankovic:2009:SDP

[Sta09] Nenad Stankovic. Single development project. The Journal of Systems and Software, 82(4):576–582, April 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Santos:2004:NMR

[SUSO04] R. M. Santos, J. Urriza, J. Santos, and J. Orozco. New methods for redistributing slack time in realtime systems: applications and comparative evaluations. The Journal of Systems and Software, 69(1–2):115–128, January 1, 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sutcliffe:2000:DAS

[Sut00]  Alistair Sutcliffe.  Do-

main analysis for software reuse. The Journal of Systems and Software, 50(3): 175–199, March 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/49/28/
26/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/49/28/abstract.
html.

Sipani:2004:DHP

[SVMAM04] Sudhanshu Sipani, Kunal Verma, John A. Miller, and Boanerges Aleman-Meza. Designing a highperformance database engine for the ‘Db4XML’ native XML database system. The Journal of Systems and Software, 69(1–2):87– 104, January 1, 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Schalken:2008:MWI

[SvV08] Joost Schalken and Hans van Vliet. Measuring where it matters: Determining starting points for metrics collection. The Journal of Systems and Software, 81(5): 603–615, May 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Saiedian:2005:NCS

[SW05] Hossein Saiedian and Bruce W. Weide. The new context for software engineering education and training. The Journal of Systems and Software, 74(2): 109–111, January 15, 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Smith:2009:SST

[SW09] Ben H. Smith and Laurie Williams. Should software testers use mutation analysis to augment a test set? The Journal of Systems and Software, 82(11):1819– 1832, November 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sun:2009:TDS

[SWH^{+}09] Hung-Min Sun, Mu-En Wu, M. Jason Hinek, Cheng-Ta Yang, and Vincent S. Tseng. Trading decryption for speeding encryption in Rebalanced-RSA. The Journal of Systems and Software, 82(9):1503– 1512, September 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Shu:2002:VCC

[SY02] LihChyun Shu and Michal Young. Versioning concurrency control for hard realtime systems. The Journal of Systems and Software, 63(3):201–218, September 15, 2002. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sanden:2006:DSB

[SZ06] Bo Sandén and Janusz Zalewski. Designing state-based systems with entity-life modeling. The Journal of Systems and Software, 79(1): 69–78, January 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Spanoudakis:2004:RBG

[SZPMK04] George Spanoudakis, Andrea Zisman, Elena Pérez-Miñana, and Paul Krause. Rule-based generation of requirements traceability relations. The Journal of Systems and Software, 72(2): 105–127, July 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Shi:2006:AES

[SZZ06] Xiaoqing Shi, Jingzhu Zhao, and Ouyang Zhiyun. Assessment of eco-security in the Knowledge Grid e-science environment. The Journal of Systems and Software, 79(2): 246–252, February 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tabary:2002:SET

[TA02] Dimitri Tabary and Mourad Abed. A software environment task object-oriented design (ETOOD). The Journal of Systems and Software, 60(2):129–140, February 1, 2002. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/72/28/31/abstract.
html.

Tang:2000:IFM

[Tan00] Shung-Ming Tang. An impact factor model of Intranet adoption: an exploratory and empirical research. The Journal of Systems and Software, 51(3):157–173, May 1, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/50/28/
26/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/50/28/abstract.
html.

Tandler:2004:BAM

[Tan04] Peter Tandler. The BEACH application model and software framework for synchronous collaboration in ubiquitous computing environments. The Journal of Systems and Software, 69(3): 267–296, January 15, 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tempero:2000:SMI

[TB00] Ewan Tempero and Robert Biddle. Simulating multiple inheritance in Java. The Journal of Systems and Software, 55(1):87–100, November 5, 2000. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        29/11/54/28/32/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/29/11/54/
                                        28/32/article.pdf.

Trinidad:2008:AEA

[TBD^{+}08] P. Trinidad, D. Benavides, A. Durán, A. Ruiz-Cortés, and M. Toro. Automated error analysis for the agilization of feature modeling. The Journal of Systems and Software, 81(6): 883–896, June 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tang:2006:SAD

[TBGH06] Antony Tang, Muhammad Ali Babar, Ian Gorton, and Jun Han. A survey of architecture design rationale. The Journal of Systems and Software, 79(12):1792– 1804, December 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tseng:2006:ERL

[TC06] Hsien-Wen Tseng and Chin-Chen Chang. Error resilient locally adaptive data compression. The Journal of Systems and Software, 79 (8):1156–1160, August 2006.

CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tsai:2002:SMS

[TCC02] Chwei-Shyong Tsai, Chin-Chen Chang, and Tung-Shou Chen. Sharing multiple secrets in digital images. The Journal of Systems and Software, 64(2):163–170, November 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tse:2006:ASS

[TCG06] T. H. Tse, T. Y. Chen, and Robert L. Glass. An assessment of systems and software engineering scholars and institutions (2000– 2004). The Journal of Systems and Software, 79(6): 816–819, June 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ton:2004:SHC

[TCSC04] Lee-Ren Ton, Lung-Chung Chang, Jyh-Jiun Shann, and Chung-Ping Chung. A software/hardware cooperated stack operations folding model for Java processors. The Journal of Systems and Software, 72(3): 377–387, August 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tselikas:2007:DSP

[TDK^{+}07] Nikolaos D. Tselikas, Nikolaos L. Dellas, Eleftherios A. Koutsoloukas, Sofia H. Kapellaki, George N. Prezerakos, and Iakovos S. Venieris. Distributed service provision using open APIs-based middleware: “OSA/Parlay vs. JAIN” performance evaluation study. The Journal of Systems and Software, 80(5): 765–777, May 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tzifa:2002:IAC

[TDL^{+}02] E. C. Tzifa, P. P. Demestichas, N. I. Liossis, N. D. Polydorou, and M. E. Anagnostou. Introduction of accounting capabilities in future service architectures. The Journal of Systems and Software, 64(2):115–129, November 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Togay:2008:SCO

[TDT08] Cengiz Togay, Ali H. Dogru, and John Urcun Tanik. Systematic component-oriented development with axiomatic design. The Journal of Systems and Software, 81(11): 1803–1815, November 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tseng:2002:ALE

[TH02] Frank S. C. Tseng and Wen-Jong Hwung. An automatic load/extract scheme for XML documents through object-relational repositories. The Journal of Systems and Software, 64(3):207–218, December 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tikir:2005:EOC

[TH05] Mustafa M. Tikir and Jeffrey K. Hollingsworth. Efficient online computation of statement coverage. The Journal of Systems and Software, 78(2):146–165, November 2005. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tomaszewski:2007:SMV

[THGL07] Piotr Tomaszewski, Jim Håkansson, Håkan Grahn, and Lars Lundberg. Statistical models vs. expert estimation for fault prediction in modified code — an industrial case study. The Journal of Systems and Software, 80 (8):1227–1238, August 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Thomasian:2006:SMR

[Tho06] Alexander Thomasian. Shortcut method for reliability comparisons in RAID.

The Journal of Systems and Software, 79(11):1599– 1605, November 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Thornberg:2006:PSG

[THP^{+}06] Benny Thörnberg, Qubo Hu, Martin Palkovic, Mattias O’Nils, and Per Gunnar Kjeldsberg. Polyhedral space generation and memory estimation from interface and memory models of real-time video systems. The Journal of Systems and Software, 79 (2):231–245, February 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tang:2007:RBA

[TJH07] Antony Tang, Yan Jin, and Jun Han. A rationale-based architecture model for design traceability and reasoning. The Journal of Systems and Software, 80(6): 918–934, June 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tsagias:2000:EBO

[TK00] Manolis Tsagias and Barbara Kitchenham. An evaluation of the business object approach to software development. The Journal of Systems and Software, 52(2– 3):149–156, June 1, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/51/27/
32/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/51/27/abstract.
html.

Tripathi:2002:DAS

[TKA^{+}02] Anand R. Tripathi, Neeran M.

Karnik, Tanvir Ahmed, Ram D. Singh, Arvind Prakash, Vineet Kakani, Manish K. Vora, and Mukta Pathak. Design of the Ajanta system for mobile agent programming. The Journal of Systems and Software, 62(2): 123–140, May 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tahvildari:2003:QDS

[TKM03] Ladan Tahvildari, Kostas Kontogiannis, and John Mylopoulos. Quality-driven software re-engineering. The Journal of Systems and Software, 66(3):225–239, June 15, 2003. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tseng:2007:EES

[TL07] Vincent S. Tseng and Kawuu W.

Lin. Energy efficient strategies for object tracking in sensor networks: a data mining approach. The Journal of Systems and Software, 80(10):1678–1698, October 2007. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tsai:2009:EKB

[TL09a] Jeffrey J. P. Tsai and Alan Liu. Experience on knowledge-based software engineering: a logic-based requirements language and its industrial applications. The Journal of Systems and Software, 82(10):1578– 1587, October 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tseng:2009:EER

[TL09b] Vincent S. Tseng and Eric Hsueh-Chan Lu. Energyefficient real-time object tracking in multi-level sensor networks by mining and predicting movement patterns. The Journal of Systems and Software, 82(4): 697–706, April 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tan:2007:VIT

[TLW07] Zhenyu Tan, William Leal, and Lonnie Welch. Verification of instrumentation techniques for resource management of real-time systems. The Journal of Systems and Software, 80(7): 1015–1022, July 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tang:2006:CHA

[TM06] Qiang Tang and Chris J.

Mitchell. Cryptanalysis of a hybrid authentication protocol for large mobile networks. The Journal of Systems and Software, 79(4): 496–501, April 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tubaishat:2002:PEL

[TMB02] Malik Ayed Tubaishat, Sanjay Kumar Madria, and Bharat Bhargava. Performance evaluation of linear hash structure model in a nested transaction environment. The Journal of Systems and Software, 63(3):219–239, September 15, 2002. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tselikis:2007:EMI

[TMD07] Christos Tselikis, Sarandis Mitropoulos, and Christos Douligeris. An evaluation of the middleware’s impact on the performance of object oriented distributed systems. The Journal of Systems and Software, 80(7): 1169–1181, July 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tian:2005:LPS

[TN05] Liang Tian and Afzel Noore. On-line prediction of software reliability using an evolutionary connectionist model. The Journal of Systems and Software, 77(2): 173–180, August 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tian:2001:EIC

[TNAA01] Jeff Tian, Anthony Nguyen, Curt Allen, and Ravi Appan. Experience with identifying and characterizing problem-prone modules in telecommunication software systems. The Journal of Systems and Software, 57(3): 207–215, July 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/64/32/30/abstract.
html.

Tang:2007:UBB

[TNJH07] Antony Tang, Ann Nicholson, Yan Jin, and Jun Han. Using Bayesian belief networks for change impact analysis in architecture design. The Journal of Systems and Software, 80(1): 127–148, January 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tiakas:2009:SST

[TPN^{+}09] E. Tiakas, A. N. Papadopoulos, A. Nanopoulos, Y. Manolopoulos, Dragan Stojanovic, and Slobodanka Djordjevic-Kajan. Searching for similar trajectories in spatial networks. The Journal of Systems and Software, 82 (5):772–788, May 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Thelin:2004:ASI

[TPRW04] Thomas Thelin, Håkan Petersson, Per Runeson, and Claes Wohlin. Applying sampling to improve software inspections. The Journal of Systems and Software, 73(2): 257–269, October 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Thwin:2005:ANN

[TQ05] Mie Mie Thet Thwin and Tong-Seng Quah. Application of neural networks for software quality prediction using object-oriented metrics. The Journal of Systems and Software, 76(2): 147–156, May 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Thelin:2000:REF

[TR00] Thomas Thelin and Per Runeson. Robust estimations of fault content with capturerecapture and detection profile estimators. The Journal of Systems and Software, 52 (2–3):139–148, June 1, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/51/27/
31/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/51/27/abstract.
html.

Tekinerdogan:2008:SAR

[TSA08] Bedir Tekinerdogan, Hasan Sozer, and Mehmet Aksit. Software architecture reliability analysis using failure scenarios. The Journal of Systems and Software, 81(4): 558–575, April 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tseng:2007:CEF

[Tse07] Yuh-Min Tseng. A communicationefficient and fault-tolerant conference-key agreement protocol with forward secrecy. The Journal of Systems and Software, 80(7): 1091–1101, July 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tsetsos:2006:SFE

[TSPH06] Vassileios Tsetsos, Odysseas Sekkas, Ioannis Priggouris, and Stathes Hadjiefthymiades. A scheduling framework for enterprise services. The Journal of Systems and Software, 79(2): 259–272, February 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tsakalozos:2009:ADS

[TSSD09] Konstantinos Tsakalozos, Vassilis Stoumpos, Kostas Saidis, and Alex Delis. Adaptive disk scheduling with workload-dependent anticipation intervals. The Journal of Systems and Software, 82 (2):274–291, February 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Thiry:2009:FMS

[TT09] Laurent Thiry and Bernard Thirion. Functional metamodels for systems and software. The Journal of Systems and Software, 82(7): 1125–1136, July 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tsai:2004:NAM

[TTC04] Cheng-Fa Tsai, Chun-Wei Tsai, and Chi-Pin Chen. A novel algorithm for multimedia multicast routing in a large scale network. The Journal of Systems and Software, 72(3):431–441, August 2004. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tsai:2004:AND

[TTWY04] Cheng-Fa Tsai, Chun-Wei Tsai, Han-Chang Wu, and Tzer Yang. ACODF: a novel data clustering approach for data mining in large databases. The Journal of Systems and Software, 73(1): 133–145, September 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Terzi:2004:SAA

[TVA04] Evimaria Terzi, Athena Vakali, and Lefteris Angelis. A simulated annealing approach for multimedia data placement. The Journal of Systems and Software, 73(3):467–480, November/December 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tsai:2007:NDH

[TW07] Yuan-Yu Tsai and Chung-Ming Wang. A novel data hiding scheme for color images using a BSP tree. The Journal of Systems and Software, 80(3):429–437, March 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tolfo:2008:IOC

[TW08a] Cristiano Tolfo and Raul Sidnei Wazlawick. The influence of organizational culture on the adoption of extreme programming. The Journal of Systems and Software, 81(11):1955–1967, November 2008. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tse:2008:E

[TW08b] T. H. Tse and W. Eric Wong.

Editorial. The Journal of Systems and Software, 81(2): 159–160, February 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tong:2004:VCP

[TYH04] Sheau-Ru Tong, Yuan-Tse Yu, and Chung-Ming Huang. A video caching policy for providing differentiated service grades and maximizing system revenue in hierarchical video servers. The Journal of Systems and Software, 73(3):515–531, November/December 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Unel:2004:EQO

[ÜDUG04] G. Ünel, M. E. Dönderler, Ö. Ulusoy, and U. Güdükbay. An efficient query optimization strategy for spatiotemporal queries in video databases. The Journal of Systems and Software, 73(1): 113–131, September 2004.

CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Utke:2006:ERI

[ULN06] Jean Utke, Andrew Lyons, and Uwe Naumann. Efficient reversal of the intraprocedural flow of control in adjoint computations. The Journal of Systems and Software, 79(9): 1280–1294, September 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Uzoka:2009:EAB

[UN09] Faith-Michael E. Uzoka and Tshepo Ndzinge. Empirical analysis of biometric technology adoption and acceptance in Botswana. The Journal of Systems and Software, 82(9): 1550–1564, September 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ueng:2001:PER

[USLC01] Jyh-Chang Ueng, Ce-Kuen Shieh, Tyng-Yue Liang, and Jyh-Biau Chang. Proteus: an efficient runtime reconfigurable distributed shared memory system. The Journal of Systems and Software, 56(3):247–260, March 15, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/

29/11/57/29/27/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/57/
29/27/article.pdf.

Umar:2009:RSO

[UZ09] Amjad Umar and Adalberto Zordan. Reengineering for service oriented architectures: a strategic decision model for integration versus migration. The Journal of Systems and Software, 82(3): 448–462, March 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Viana:2008:XMU

[VA08] Windson Viana and Rossana M. C. Andrade. XMobile: a MB-UID environment for semi-automatic generation of adaptive applications for mobile devices. The Journal of Systems and Software, 81(3): 382–394, March 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Vilas:2004:ISS

[VAS^{+}04] Ana Fernández Vilas, José J. Pazos Arias, Alberto Gil Solla, Rebeca P. Díaz Redondo, Jorge García Duque, and A. Belén Barragáns Martínez. Incremental specification with SCTL/MUS-T: a case study. The Journal of Systems and Software, 70(1– 2):189–208, February 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Vaughn:2007:LEP

[Vau07] Rayford B. Vaughn. Leadership by example: a perspective on the influence of Barry Boehm. The Journal of Systems and Software, 80 (8):1222–1226, August 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

vanDeursen:2005:SRE

[vDB05] Arie van Deursen and Elizabeth Burd. Software reverse engineering. The Journal of Systems and Software, 77(3): 209–211, September 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

vanderStok:2007:HRA

[vdSJK^{+}07] Peter van der Stok, Dmitri Jarnikov, Sergei Kozlov, Michael van Hartskamp, and Johan Lukkien. Hierarchical resource allocation for robust in-home video streaming. The Journal of Systems and Software, 80(7): 951–961, July 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Vegas:2003:BPS

[VE03] S. Vegas and M. Estayno. Best papers on software engineering from the SEKE’01 Conference. The Journal of Systems and Software, 68(3): 167–169, December 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Verner:2001:DVS

[VEM^{+}01] J. M. Verner, W. M. Evanco, K. W. McCain, G. W. Hislop, and V. J. Cole. The determinants of visibility of software engineering researchers. The Journal of Systems and Software, 59(1):99–106, October 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/71/27/34/abstract.
html.

vanGurp:2002:DEP

[vGB02] Jilles van Gurp and Jan Bosch. Design erosion: problems and causes. The Journal of Systems and Software, 61 (2):105–119, March 15, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Vaughn:2002:ESI

[VHF02] Rayford B. Vaughn, Jr., Ronda Henning, and Kevin Fox. An empirical study of industrial security-engineering practices. The Journal of Systems and Software, 61(3): 225–232, April 1, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Vegas:2006:PEI

[VJB06] Sira Vegas, Natalia Juristo, and Victor Basili. Packaging experiences for improving testing technique selection. The Journal of Systems and Software, 79(11):1606– 1618, November 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

VanHulse:2008:CEE

[VK08] Jason Van Hulse and Taghi M. Khoshgoftaar. A comprehensive empirical evaluation of missing value imputation in noisy software measurement data. The Journal of Systems and Software, 81(5): 691–708, May 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Vakali:2000:DPS

[VM00] Athena Vakali and Yannis Manolopoulos. Data placement schemes in replicated mirrored disk systems. The Journal of Systems and Software, 55(2):115–128, December 27, 2000. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej-ng/10/
                                        29/11/54/24/26/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/29/11/54/
                                        24/26/article.pdf.

Valerdi:2007:ICM

[VM07] Ricardo Valerdi and Ray Madachy. Impact and contributions of MBASE on software engineering graduate courses. The Journal of Systems and Software, 80 (8):1185–1190, August 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Vandecruys:2008:MSR

[VMB^{+}08] Olivier Vandecruys, David Martens, Bart Baesens, Christophe Mues, Manu De Backer, and Raf Haesen. Mining software repositories for comprehensible software fault prediction models. The Journal of Systems and Software, 81(5): 823–839, May 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Vergilio:2006:CBS

[VMJS06] Silvia Regina Vergilio, José Carlos Maldonado, Mario Jino, and Inali Wisniewski Soares. Constraint based structural testing criteria. The Journal of Systems and Software, 79 (6):756–771, June 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Voas:2000:DCS

[VP00] Jeffrey Voas and Jeffery Payne. Dependability certification of software components. The Journal of Systems and Software, 52(2– 3):165–172, June 1, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/51/27/
34/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/51/27/abstract.
html.

Valente:2007:COA

[VP07] Marco Tulio Valente and Rodrigo Palhares. Collocation optimizations in an aspect-oriented middleware system. The Journal of Systems and Software, 80(10): 1659–1666, October 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

vanVliet:2008:ET

[vV08] Hans van Vliet. Editorial — Taking over. The Journal of Systems and Software, 81(3): vii–viii, March 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Walters:2005:CMB

[Wal05] Robert John Walters. Checking of models built using a graphically based formal modelling language. The Journal of Systems and Software, 76(1):55–64, April 2005. CODEN JSSODM.

ISSN 0164-1212 (print), 1873-1228 (electronic).

Wilde:2003:CML

[WBP^{+}03] Norman Wilde, Michelle Buckellew, Henry Page, Vaclav Rajlich, and LaTreva Pounds. A comparison of methods for locating features in legacy software. The Journal of Systems and Software, 65(2):105–114, February 15, 2003. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wang:2006:TAG

[WBW^{+}06] Yanxin Wang, Smruti Ranjan Behera, Johnny Wong, Guy Helmer, Vasant Honavar, Les Miller, Robyn Lutz, and Mark Slagell. Towards the automatic generation of mobile agents for distributed intrusion detection system. The Journal of Systems and Software, 79 (1):1–14, January 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wu:2002:DRT

[WC02] Hsien-Chu Wu and Chin-Chen Chang. Detection and restoration of tampered JPEG compressed images. The Journal of Systems and Software, 64(2):151–161, November 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wang:2007:IBP

[WC07] Qin Wang and Zhenfu Cao. Identity based proxy multisignature. The Journal of Systems and Software, 80(7): 1023–1029, July 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wang:2003:DRT

[WCH03] Ching-Te Wang, Tung-Shou Chen, and Shao-Hau He. Detecting and restoring the tampered images based on iteration-free fractal compression. The Journal of Systems and Software, 67(2): 131–140, August 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wang:2007:FOR

[WCLK07] Xiaorui Wang, Yingming Chen, Chenyang Lu, and Xenofon Koutsoukos. FC-ORB: a robust distributed real-time embedded middleware with end-to-end utilization control. The Journal of Systems and Software, 80 (7):938–950, July 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wang:2009:RIH

[WCLL09] Zhi-Hui Wang, Chin-Chen Chang, Chia-Chen Lin, and Ming-Chu Li. A reversible information hiding scheme using left-right and up-down Chinese character representation. The Journal of Systems and Software, 82 (8):1362–1369, August 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wicks:2007:NRA

[WD07] M. N. Wicks and R. G.

Dewar. A new research agenda for tool integration. The Journal of Systems and Software, 80(9): 1569–1585, September 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wang:2008:VBA

[WDCL08] Xiaoying Wang, Zhihui Du, Yinong Chen, and Sanli Li. Virtualization-based autonomic resource management for multi-tier Web applications in shared data center. The Journal of Systems and Software, 81(9):1591– 1608, September 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wuyts:2005:DCA

[WDN05] Roel Wuyts, Stéphane Ducasse, and Oscar Nierstrasz. A data-centric approach to composing embedded, realtime software components. The Journal of Systems and Software, 74(1):25–34, January 1, 2005. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

White:2009:SHO

[WDS09] Jules White, Brian Dougherty, and Douglas C. Schmidt. Selecting highly optimal architectural feature sets with Filtered Cartesian Flattening. The Journal of Systems and Software, 82(8):1268– 1284, August 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wenger:2003:FPL

[Wen03] James Wenger. The future of programming languages: evidence to support a midwest university information systems curriculum. The Journal of Systems and Software, 65(2):163–167, February 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Westland:2002:CES

[Wes02] J. Christopher Westland. The cost of errors in software development: evidence from industry. The Journal of Systems and Software, 62 (1):1–9, May 1, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Weyuker:2001:GEC

[Wey01] Elaine J. Weyuker. Guest Editor’s corner: Transitioning from academia to industrial research. The Journal of Systems and Software, 59(1): 1–2, October 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/71/27/27/abstract.
html.

Wang:2007:HKM

[WF07] Nen-Chung Wang and Shian-Zhang Fang. A hierarchical key management scheme for secure group communications in mobile ad hoc networks. The Journal of Systems and Software, 80(10): 1667–1677, October 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wang:2009:MDM

[WFWL09] Juan Wang, Dan Feng, Fang Wang, and Chengtao Lu. MHS: a distributed metadata management strategy. The Journal of Systems and Software, 82(12):2004– 2011, December 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wong:2005:SDD

[WG05] W. Eric Wong and Swapna Gokhale. Static and dynamic distance metrics for feature-based code analysis. The Journal of Systems and Software, 74(3): 283–295, February 1, 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wang:2002:IPC

[WGC02] Tianhan Wang, Sheng-Uei Guan, and Tai Khoon Chan. Integrity protection for Code-on-Demand mobile agents in e-commerce. The Journal of Systems and Software, 60(3):211–221, February 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/72/29/32/abstract.
html.

Wong:2000:QCB

[WGH00] W. Eric Wong, Swapna S. Gokhale, and Joseph R. Horgan. Quantifying the closeness between program components and features. The Journal of Systems and Software, 54(2):87–98, October 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/53/25/
25/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/53/25/abstract.
html.

Wang:2009:NAS

[WGW^{+}09] Xingwei Wang, Lei Guo, Xuetao Wei, Lan Pang, Tengfei Wu, Juan Du, and Xuekui Wang. A new algorithm with segment protection and load balancing for single-link failure in multicasting survivable networks. The Journal of Systems and Software, 82(3): 377–381, March 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wang:2008:MLS

[WGY^{+}08] Xingwei Wang, Lei Guo, Fei Yang, Tengfei Wu, and Wei Ji. Multi-layer survivable routing mechanism in GMPLS based optical networks. The Journal of Systems and Software, 81(11): 2014–2023, November 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wu:2002:CAE

[WH02] Tzong-Sun Wu and Chien-Lung Hsu. Convertible authenticated encryption scheme. The Journal of Systems and Software, 62(3): 205–209, June 15, 2002. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wu:2003:TSS

[WH03] Tzong-Sun Wu and Chien-Lung Hsu. Threshold signature scheme using selfcertified public keys. The Journal of Systems and Software, 67(2):89–97, August 15, 2003. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wilson:2001:FEP

[WHB01] David N. Wilson, Tracy Hall, and Nathan Baddoo. A framework for evaluation and prediction of software process improvement success. The Journal of Systems and Software, 59(2):135–142, November 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        29/11/71/28/29/abstract.
                                        html.

Wuu:2007:BIP

[WHC07] Lih-Chyau Wuu, Chi-Hsiang Hung, and Sout-Fong Chen. Building intrusion pattern miner for Snort network intrusion detection system. The Journal of Systems and Software, 80(10):1699– 1715, October 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wu:2001:DMS

[WHG01] Tzong-Chen Wu, Chih-Chan Huang, and D.-J. Guan. Delegated multisignature scheme with document decomposition. The Journal of Systems and Software, 55(3):321–328, January 15, 2001. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/54/27/32/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/54/
27/32/article.pdf.

Wu:2008:RPG

[WHHT08] Tzong-Chen Wu, Thsia-Tzu Huang, Chien-Lung Hsu, and Kuo-Yu Tsai. Recursive protocol for group-oriented authentication with key distribution. The Journal of Systems and Software, 81(7): 1227–1239, July 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wong:2001:SMA

[WHN^{+}01] Johnny Wong, Guy Helmer, Venkatraman Naganathan, Sriniwas Polavarapu, Vasant Honavar, and Les Miller. SMART mobile agent facility. The Journal of Systems and Software, 56(1):9– 22, February 1, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/57/24/25/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/57/
24/25/article.pdf.

Wen:2006:TSA

[WHYT06] Jyh-Horng Wen, Kuo-Ting Huang, Cheng-Ying Yang, and Tzu-Chen Tsai. Timeslotsharing algorithm with a dynamic grouping for WDM broadcast-and-select star networks. The Journal of Systems and Software, 79 (8):1110–1119, August 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wijnstra:2003:PSQ

[Wij03] Jan Gerben Wijnstra. From problem to solution with quality attributes and design aspects. The Journal of Systems and Software, 66(3): 199–211, June 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wile:2003:RCP

[Wil03] David S. Wile. Revealing component properties through architectural styles. The Journal of Systems and Software, 65(3):209– 214, March 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wang:2009:EFD

[WJT09] Shuenn-Jyi Wang, Dah-Lih Jeng, and Meng-Tsai Tsai. Early fire detection method in video for vessels. The Journal of Systems and Software, 82(4):

656–667, April 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wilkie:2000:CMC

[WK00] F. G. Wilkie and B. A.

Kitchenham. Coupling measures and change ripples in C++ application software. The Journal of Systems and Software, 52(2– 3):157–164, June 1, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/51/27/
33/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/51/27/abstract.
html.

Waszniowski:2009:CSD

[WKH09] Libor Waszniowski, Jan Krákora, and Zdeněk Hanzálek. Case study on distributed and fault tolerant system modeling based on timed automata. The Journal of Systems and Software, 82(10): 1678–1694, October 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wills:2004:RSP

[WKL04] Linda M. Wills, Fabrice Kordon, and Luqi. Rapid system prototyping. The Journal of Systems and Software, 70(3): 225–227, March 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wang:2005:CHY

[WL05] Shyh-Yih Wang and Chi-Sung Laih. Cryptanalysis of Hwang–Yang scheme for controlling access in large partially ordered hierarchies. The Journal of Systems and Software, 75(1–2):189–192, February 15, 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wu:2009:ESC

[WL09] Tzong-Sun Wu and Han-Yu Lin. Efficient self-certified proxy CAE scheme and its variants. The Journal of Systems and Software, 82(6): 974–980, June 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wu:2007:NIC

[WLC07] Ming-Ni Wu, Chia-Chen Lin, and Chin-Chen Chang. Novel image copy detection with rotating tolerance. The Journal of Systems and Software, 80(7):1057–1069, July 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wu:2008:ETB

[WLC08] Ming-Ni Wu, Chia-Chen Lin, and Chin-Chen Chang. An embedding technique based upon block prediction. The Journal of Systems and Software, 81(9): 1505–1516, September 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wu:2009:HCR

[WLT^{+}09] Hsien-Chu Wu, Chih-Chiang Lee, Chwei-Shyong Tsai, Yen-Ping Chu, and Hung-Ruei Chen. A high capacity reversible data hiding scheme with edge prediction and difference expansion. The Journal of Systems and Software, 82(12):1966– 1973, December 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wong:2009:E

[WM09] W. Eric Wong and Aditya Mathur. Editorial. The Journal of Systems and Software, 82(9):1375, September 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wu:2008:AEI

[WOH08] Liang-Chuan Wu, Chorng-Shyong Ong, and Yao-Wen Hsu. Active ERP implementation management: a Real Options perspective. The Journal of Systems and Software, 81(6): 1039–1050, June 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wang:2006:ABS

[WPC06] Wen-Li Wang, Dai Pan, and Mei-Hwa Chen. Architecturebased software reliability modeling. The Journal of Systems and Software, 79(1): 132–146, January 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Woodside:2009:PAS

[WPP^{+}09] Murray Woodside, Dorina C. Petriu, Dorin B. Petriu, Jing Xu, Tauseef Israr, Geri Georg, Robert France, James M. Bieman, Siv Hilde Houmb, and Jan Jürjens. Performance analysis of security aspects by weaving scenarios extracted from UML models. The Journal of Systems and Software, 82(1): 56–74, January 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wong:2006:EPD

[WQ06] W. Eric Wong and Yu Qi. Effective program debugging based on execution slices and inter-block data dependency. The Journal of Systems and Software, 79(7): 891–903, July 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wang:2008:CAF

[WSJK08] Eric T. G. Wang, Sheng-Pao Shih, James J. Jiang, and Gary Klein. The consistency among facilitating factors and ERP implementation success: a holistic view of fit. The Journal of Systems and Software, 81(9): 1609–1621, September 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wong:2005:SDS

[WSQM05] W. Eric Wong, Tatiana Sugeta, Yu Qi, and Jose C. Maldonado. Smart debugging software architectural design in SDL. The Journal of Systems and Software, 76(1): 15–28, April 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wang:2001:SPC

[WT01] Chen Wang and Yong Meng Teo. Supporting parallel computing on a distributed object architecture. The Journal of Systems and Software, 56(3):261–278, March 15, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/57/29/28/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/57/
29/28/article.pdf.

Wong:2008:ASS

[WTG^{+}08] W. Eric Wong, T. H. Tse, Robert L. Glass, Victor R. Basili, and T. Y. Chen. An assessment of systems and software engineering scholars and institutions (2001– 2005). The Journal of Systems and Software, 81(6): 1059–1062, June 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wong:2009:ASS

[WTG^{+}09] W. Eric Wong, T. H. Tse, Robert L. Glass, Victor R. Basili, and T. Y. Chen. An assessment of systems and software engineering scholars and institutions (2002– 2006). The Journal of Systems and Software, 82 (8):1370–1373, August 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wynn:2000:ECP

[WW00] Allen Wynn and Jie Wu. The effect of compression on performance in a demand paging operating system. The Journal of Systems and Software, 50(2): 151–170, February 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/49/27/
30/article.pdf;   http:/
/www.elsevier.nl/gej-ng/

10/29/11/49/27/abstract.
html.

Wang:2009:DAA

[WW09] Ching-Huey Wang and Feng-Jian Wang. Detecting artifact anomalies in business process specifications with a formal model. The Journal of Systems and Software, 82(10):1600–1619, October 2009. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wappler:2009:ETS

[WWB09] Stefan Wappler, Joachim Wegener, and André Baresel. Evolutionary testing of software with function-assigned flags. The Journal of Systems and Software, 82(11):1767– 1779, November 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wang:2000:HLS

[WWC00] Yi-Min Wang, Hsiao-Hsi Wang, and Ruei-Chuan Chang. Hierarchical loop scheduling for clustered NUMA machines. The Journal of Systems and Software, 55(1):33–44, November 5, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/54/28/
27/00007102.pdf;   http:
//www.elsevier.nl/gej-

                                        ng/10/29/11/54/28/27/abstract.
                                        html.

Wang:2008:HQS

[WWTH08] Chung-Ming Wang, Nan-I Wu, Chwei-Shyong Tsai, and Min-Shiang Hwang. A high quality steganographic method with pixel-value differencing and modulus function. The Journal of Systems and Software, 81(1): 150–158, January 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wang:2004:DLF

[WY04] S. C. Wang and K. Q. Yan. Dual link fault diagnosis agreement. The Journal of Systems and Software, 71(1–2):117–125, April 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wu:2006:ASA

[WYL06] Lin-Chuan Wu, Yi-Shiung Yeh, and Tsann-Shyong Liu. Analysis of Sun et al.’s linkability attack on some proxy blind signature schemes. The Journal of Systems and Software, 79(2):176–179, February 2006. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wynn:2001:GEC

[Wyn01] Eleanor Wynn. Guest Editor’s corner: Some informal thoughts about reviewing as a social process. The Journal of Systems and Software, 59(2):115–117, November 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/71/28/27/abstract.
html.

Wang:2009:CAW

[WZG09] Wei Wang, Xiangliang Zhang, and Sylvain Gombault. Constructing attribute weights from computer audit data for effective intrusion detection. The Journal of Systems and Software, 82(12):1974– 1981, December 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wang:2001:DIA

[WZJ01] Li Wang, Wanlei Zhou, and Weijia Jia. The design and implementation of an active replication scheme for distributing services in a cluster of workstations. The Journal of Systems and Software, 58(3):199–210, September 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/68/34/28/abstract.
html.

Xia:2000:CCM

[Xia00] Franck Xia. On the concept of coupling, its modeling and measurement. The Journal of Systems and Software, 50(1):75–84, January 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/29/11/49/25/
                                        31/article.pdf;   http:/
                                        /www.elsevier.nl/gej-ng/
                                        10/29/11/49/25/abstract.
                                        html.

Xue:2007:ISE

[XNP07] Jingling Xue, Phung Hua Nguyen, and John Potter. Interprocedural sideeffect analysis for incomplete object-oriented software modules. The Journal of Systems and Software, 80(1):92–105, January 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Xu:2006:RPE

[XSS06] Hong Xu, Pete Sawyer, and Ian Sommerville. Requirement process establishment and improvement from the viewpoint of cybernetics. The Journal of Systems and Software, 79(11):1504– 1513, November 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Xia:2002:GSS

[XY02] Shundong Xia and Jinyuan You. A group signature scheme with strong separability. The Journal of Systems and Software, 60(3):177–182, February 15, 2002. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/72/29/28/abstract.
html.

Xie:2007:CBH

[XYS07] Fei Xie, Guowu Yang, and Xiaoyu Song. Componentbased hardware/software co-verification for building trustworthy embedded systems. The Journal of Systems and Software, 80(5): 643–654, May 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Xu:2006:APN

[XZAR06] Lihua Xu, Hadar Ziv, Thomas A. Alspaugh, and Debra J. Richardson. An architectural pattern for nonfunctional dependability requirements. The Journal of Systems and Software, 79(10):1370–1378, October 2006. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yeh:2008:EII

[YC08a] Wei-Horng Yeh and Ye-In Chang. An efficient iconic indexing strategy for image rotation and reflection in image databases. The Journal of Systems and Software, 81(7): 1184–1195, July 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yun:2008:DIB

[YC08b] Jung-Hee Yun and Chin-Wan Chung. Dynamic intervalbased labeling scheme for efficient XML query and update processing. The Journal of Systems and Software, 81(1):56–70, January 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yang:2009:ETP

[YC09] Jen-Ho Yang and Chin-Chen Chang. An efficient threeparty authenticated key exchange protocol using elliptic curve cryptography for mobile-commerce environments. The Journal of Systems and Software, 82(9): 1497–1502, September 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yang:2007:IIS

[YCYW07] Ching-Nung Yang, Tse-Shih Chen, Kun Hsuan Yu, and Chung-Chun Wang. Improvements of image sharing with steganography and authentication. The Journal of Systems and Software, 80(7): 1070–1076, July 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yeung:2000:ATJ

[Yeu00] W. L. Yeung. Automated translation of JSD into CSP

— a case study in methods integration. The Journal of Systems and Software, 55(2):193–202, December 27, 2000. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/54/24/31/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/54/
24/31/article.pdf.

Yau:2008:SDA

[YGH^{+}08] S. S. Yau, H. Gong, D. Huang, W. Gao, and L. Zhu. Specification, decomposition and agent synthesis for situationaware service-based systems. The Journal of Systems and Software, 81(10):1663– 1680, October 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yun:2003:MAR

[YHHR03] Hyunyoon Yun, Danshim Ha, Buhyun Hwang, and Keun Ho Ryu. Mining association rules on significant rare data using relative support. The Journal of Systems and Software, 67(3):181–191, September 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yang:2009:QAT

[YHZ^{+}09] Jie Yang, Gang Huang, Wenhui Zhu, Xiaofeng Cui, and Hong Mei. Quality attribute tradeoff through adaptive architectures at runtime. The Journal of Systems and Software, 82(2):319–332, February 2009. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yoo:2005:FSR

[YKC^{+}05] Junbeom Yoo, Taihyo Kim, Sungdeok Cha, Jang-Soo Lee, and Han Seong Son. A formal software requirements specification method for digital nuclear plant protection systems. The Journal of Systems and Software, 74(1):73– 83, January 1, 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yu:2006:CMD

[YL06] Yuen Tak Yu and Man Fai Lau. A comparison of MC/DC, MUMCUT and several other coverage criteria for logical decisions. The Journal of Systems and Software, 79(5):577–590, May 2006. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yin:2009:NRF

[YL09] Peng-Yeng Yin and Chin-Wen Liu. A new relevance feedback technique for iconic image retrieval based on spatial relationships. The Journal of Systems and Software, 82(4):685–696, April 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yu:2006:AGT

[YLC06] Yuen Tak Yu, Man Fai Lau, and Tsong Yueh Chen. Automatic generation of test cases from Boolean specifications using the MUMCUT strategy. The Journal of Systems and Software, 79(6): 820–840, June 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yeh:2008:EER

[YLC08] Dowming Yeh, Yuwen Li, and William Chu. Extracting entity-relationship diagram from a table-based legacy database. The Journal of Systems and Software, 81 (5):764–771, May 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yoo:2009:RTT

[Yoo09]  Myungryun Yoo.  Real-

time task scheduling by multiobjective genetic algorithm. The Journal of Systems and Software, 82(4):

619–628, April 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yu:2009:EAE

[YR09] Liguo Yu and Srini Ramaswamy. An empirical approach to evaluating dependency locality in hierarchically structured software systems. The Journal of Systems and Software, 82(3): 463–472, March 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yoo:2002:EAS

[YS02] Chang Sik Yoo and Poong Hyun Seong. Experimental analysis of specification language diversity impact on NPP software diversity. The Journal of Systems and Software, 62 (2):111–122, May 15, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yeh:2004:PBU

[YS04] Her-Tyan Yeh and Hung-Min Sun. Password-based user authentication and key distribution protocols for client– server applications. The Journal of Systems and Software, 72(1):97–103, June 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yu:2006:MKO

[YSC^{+}06] Liguo Yu, Stephen R.

Schach, Kai Chen, Gillian Z. Heller, and Jeff Offutt. Maintainability of the kernels of open-source operating systems: a comparison of Linux with FreeBSD, NetBSD, and OpenBSD. The Journal of Systems and Software, 79(6): 807–815, June 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yoo:2006:ESR

[YSK06] Sanghyun Yoo, Jin Hyun Son, and Myoung Ho Kim. An efficient subscription routing algorithm for scalable XML-based publish/subscribe systems. The Journal of Systems and Software, 79(12): 1767–1781, December 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yoo:2009:SPS

[YSK09] Sanghyun Yoo, Jin Hyun Son, and Myoung Ho Kim. A scalable publish/subscribe system for large mobile ad hoc networks. The Journal of Systems and Software, 82(7): 1152–1162, July 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yang:2004:ENT

[YTH04] Cheng-Ying Yang, Shiang-Feng Tzeng, and Min-Shiang Hwang. On the efficiency of nonrepudiable threshold proxy signature scheme with known signers. The Journal of Systems and Software, 73(3):507–514, November/December 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yan:2002:ADE

[YWLG02] Hongfei Yan, Jianyong Wang, Xiaoming Li, and Lin Guo. Architectural design and evaluation of an efficient Web-crawling system. The Journal of Systems and Software, 60(3):185–193, February 15, 2002. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/72/29/30/abstract.
html.

Yang:2007:SMA

[YWT07] Chun-Chuan Yang, Yung-Chi Wang, and Chih-Wen Tien. Synchronization modeling and its application for SMIL2.0 presentations. The Journal of Systems and Software, 80(7):1142–1155, July 2007. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yang:2004:DIJ

[YY04] Chun-Chuan Yang and Yi-Zheng Yang. Design and implementation of the justin-time retrieving policy for schedule-based distributed multimedia presentations. The Journal of Systems and Software, 71(1–2):49– 63, April 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yoo:2006:UMI

[YYL^{+}06] Chanwoo Yoo, Junho Yoon, Byungjeong Lee, Chongwon Lee, Jinyoung Lee, Seunghun Hyun, and Chisu Wu. A unified model for the implementation of both ISO 9001:2000 and CMMI by ISO-certified organizations. The Journal of Systems and Software, 79 (7):954–961, July 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yin:2007:TAM

[YYWW07] Peng-Yeng Yin, Shiuh-Sheng Yu, Pei-Pei Wang, and Yi-Te Wang. Task allocation for maximizing reliability of a distributed system using hybrid particle swarm optimization. The Journal of Systems and Software, 80(5): 724–735, May 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yau:2005:MSS

[YZ05] Stephen S. Yau and Xinyu Zhang. A middleware service for secure group communication in mobile ad hoc networks. The Journal of Systems and Software, 76(1): 29–43, April 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yan:2008:BST

[YZ08] Jun Yan and Jian Zhang. A backtracking search tool for constructing combinatorial test suites. The Journal of Systems and Software, 81(10):1681–1693, October 2008. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zheng:2008:AGT

[ZAO08] Mao Zheng, Vasu Alagar, and Olga Ormandjieva. Automated generation of test suites from formal specifications of real-time reactive systems. The Journal of Systems and Software, 81(2): 286–304, February 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhu:2007:MCB

[ZBLG07] Liming Zhu, Ngoc Bao Bui, Yan Liu, and Ian Gorton. MDABench: Customized benchmark generation using MDA. The Journal of Systems and Software, 80(2): 265–282, February 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhang:2005:CHC

[ZC05] Fangguo Zhang and Xiaofeng Chen. Cryptanalysis of Huang–Chang partially blind signature scheme. The Journal of Systems and Software, 76(3):323–325, June 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhang:2006:UTL

[ZC06] Ji Zhang and Betty H. C.

Cheng. Using temporal logic to specify adaptive program semantics. The Journal of Systems and Software, 79(10):1361–1369, October 2006. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhan:2008:SBF

[ZC08] Yuan Zhan and John A.

Clark. A search-based framework for automatic testing of MATLAB/Simulink. The Journal of Systems and Software, 81(2):262–285, February 2008. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhang:2009:RPC

[ZCT^{+}09] Zhenyu Zhang, W. K. Chan, T. H. Tse, Heng Lu, and Lijun Mei. Resource prioritization of code optimization techniques for program synthesis of wireless sensor network applications. The Journal of Systems and Software, 82(9):1376– 1387, September 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhao:2003:QAU

[ZE03] Luyin Zhao and Sebastian Elbaum. Quality assurance under the open source development model. The Journal of Systems and Software, 66(1):65–75, April 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zelkowitz:2009:UEM

[Zel09] Marvin V. Zelkowitz. An update to experimental models for validating computer technology. The Journal of Systems and Software, 82(3): 373–376, March 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zaki:2000:SCA

[ZERO00] Mohammed Zaki, Abdallah El-Ramsisi, and Rostom Omran. A soft computing approach for recognition of occluded shapes. The Journal of Systems and Software, 51(1):73–83, April 1, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/50/26/
31/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/50/26/abstract.
html.

Zaki:2004:EEM

[ZEY04] M. Zaki, M. Y. El Nahas, and M. Youssef. EMBOT: An enhanced motion-based object tracker. The Journal of Systems and Software, 69(1– 2):149–158, January 1, 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhang:2000:LMP

[ZG00] Kang Zhang and Narasimhaiah Gorla. Locality metrics and program physical structures. The Journal of Systems and Software, 54(2): 159–166, October 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/53/25/
30/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/53/25/abstract.
html.

Zhuge:2007:VKS

[ZG07] Hai Zhuge and Weiyu Guo.

Virtual knowledge service market—For effective knowledge flow within knowledge grid. The Journal of Systems and Software, 80(11):1833– 1842, November 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhu:2007:PMT

[ZGH^{+}07] Nianping Zhu, John Grundy, John Hosking, Na Liu, Shuping Cao, and Akhil Mehra. Pounamu: a meta-tool for exploratory domain-specific visual language tool development. The Journal of Systems and Software, 80 (8):1390–1407, August 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zimmermann:2005:TME

[ZH05] Armin Zimmermann and Günter Hommel. Towards modeling and evaluation of ETCS real-time communication and operation. The Journal of Systems and Software, 77(1):47–54, July 2005. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhang:2008:HZW

[Zha08] Yanchun Zhang. H. Zhuge, The Web Resource Space Model, Springer (2007). The Journal of Systems and Software, 81(6):1063–1064, June 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhao:2009:DIB

[Zha09]  Wenbing Zhao.  Design

and implementation of a Byzantine fault tolerance framework for Web services. The Journal of Systems and Software, 82(6): 1004–1015, June 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zaki:2001:LDS

[ZHS01] M. Zaki, Hany Harb, and T. S. Sobh. A learning database system to observe malfunctions and to support network planning. The Journal of Systems and Software, 58(1):33–46, August 15, 2001. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/68/27/30/abstract.
html.

Zhuge:2000:POR

[Zhu00] Hai Zhuge. A problemoriented and rule-based component repository. The Journal of Systems and Software, 50(3):201–208, March 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/49/28/
27/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/49/28/abstract.
html.

Zhuge:2003:IMM

[Zhu03] Hai Zhuge. An inexact model matching approach and its applications. The Journal of Systems and Software, 67(3):201–212, September 15, 2003. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhuge:2004:FRS

[Zhu04a] Hai Zhuge. Fuzzy resource space model and platform. The Journal of Systems and Software, 73(3):389–396, November/December 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhuge:2004:KG

[Zhu04b] Hai Zhuge. The Knowledge Grid. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2004. ISBN 981-256-140-4. xiii + 264 pp. LCCN TK5105.88815 .Z48 2004.

Zhuge:2004:RSM

[Zhu04c] Hai Zhuge. Resource space model, its design method and applications. The Journal of Systems and Software, 72 (1):71–81, June 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhuge:2004:RIU

[Zhu04d] Hai Zhuge. Retrieve images by understanding semantic links and clustering image fragments. The Journal of Systems and Software, 73(3):455–466, November/December 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhuge:2006:SCN

[Zhu06] Hai Zhuge. Semantic component networking: Toward the synergy of static reuse and dynamic clustering of resources in the knowledge grid. The Journal of Systems and Software, 79(10):1469– 1482, October 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhu:2002:SRV

[ZJDB02] Hong Zhu, Lingzi Jin, Dan Diaper, and Ganghong Bai. Software requirements validation via task analysis. The Journal of Systems and Software, 61(2):145–169, March 15, 2002. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zerfiridis:2004:BFW

[ZK04a] Konstantinos G. Zerfiridis and Helen D. Karatza. Brute force Web search for wireless devices using mobile agents. The Journal of Systems and Software, 69(1–2): 195–206, January 1, 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zerfiridis:2004:FDU

[ZK04b] Konstantinos G. Zerfiridis and Helen D. Karatza. File distribution using a peer-to-peer network — a simulation study. The Journal of Systems and Software, 73(1):31– 44, September 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zikos:2009:CCE

[ZK09] Stylianos Zikos and Helen D. Karatza. Communication cost effective scheduling policies of nonclairvoyant jobs with load balancing in a grid. The Journal of Systems and Software, 82(12):2103– 2116, December 2009. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zimmermann:2009:MAD

[ZKL^{+}09] Olaf Zimmermann, Jana Koehler, Frank Leymann, Ronny Polley, and Nelly Schuster. Managing architectural decision models with dependency relations, integrity constraints, and production rules. The Journal of Systems and Software, 82 (8):1249–1267, August 2009. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhuge:2004:FRW

[ZL04] Hai Zhuge and Jie Liu. Flexible retrieval of Web Services. The Journal of Systems and Software, 70(1–2): 107–116, February 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhuge:2006:AGD

[ZL06] Hai Zhuge and Xiangfeng Luo. Automatic generation of document semantics for the e-science Knowledge Grid. The Journal of Systems and Software, 79(7): 969–983, July 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhou:2007:POO

[ZL07] Yuming Zhou and Hareton Leung. Predicting objectoriented software maintainability using multivariate adaptive regression splines. The Journal of Systems and Software, 80(8):1349– 1361, August 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhao:2006:SRG

[ZLCY06] Jing Zhao, Hong-Wei Liu, Gang Cui, and Xiao-Zong Yang. Software reliability growth model with changepoint and environmental function. The Journal of Systems and Software, 79(11): 1578–1587, November 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhao:2006:ABD

[ZM06] Jane Zhao and Hui Ma. ASM-based design of data warehouses and on-line analytical processing systems. The Journal of Systems and Software, 79(5):613–629, May 2006. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhao:2008:PLD

[ZMAV08] Liping Zhao, Linda Macaulay, Jonathan Adams, and Paul Verschueren. A pattern language for designing ebusiness architecture. The Journal of Systems and Software, 81(8):1272–1287, August 2008. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhu:2005:FSA

[ZMN05] Feng Zhu, Matt Mutka, and Lionel Ni. Facilitating secure ad hoc service discovery in public environments. The Journal of Systems and Software, 76(1): 45–54, April 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhang:2000:AFA

[ZP00] Xuemei Zhang and Hoang Pham. An analysis of factors affecting software reliability. The Journal of Systems and Software, 50 (1):43–56, January 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/29/11/49/25/
28/article.pdf;   http:/
/www.elsevier.nl/gej-ng/
10/29/11/49/25/abstract.
html.

Zeadally:2005:JSW

[ZP05] Sherali Zeadally and Junhua Pan. J2EE support for wireless services. The Journal of Systems and Software, 75(1–2):215–226, February 15, 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhang:2006:SFF

[ZP06] Xuemei Zhang and Hoang Pham. Software field failure rate prediction before software deployment. The Journal of Systems and Software, 79(3):291–300, March 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zendler:2001:ECC

[ZPEL01] Andreas Zendler, Till Pfeiffer, Michael Eicks, and Franz Lehner. Experimental comparison of coarse-grained concepts in UML, OML, and TOS. The Journal of Systems and Software, 57 (1):21–30, April 27, 2001. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.elsevier.nl/gej%
                                        2Dng/10/29/11/64/25/27/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej%
                                        2Dng/10/29/11/64/25/27/
                                        article.pdf.

Zhang:2006:IUC

[ZQZ^{+}06] Lu Zhang, Tao Qin, Zhiying Zhou, Dan Hao, and Jiasu Sun. Identifying use cases in source code. The Journal of Systems and Software, 79(11):1588–1598, November 2006. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zelkowitz:2004:DEP

[ZR04] Marvin V. Zelkowitz and Ioana Rus. Defect evolution in a product line environment. The Journal of Systems and Software, 70(1–2): 143–154, February 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhuge:2001:CCC

[ZS01] Hai Zhuge and Xiaoqing Shi. Communication cost of cognitive co-operation for distributed team development. The Journal of Systems and Software, 57(3): 227–233, July 15, 2001. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
29/11/64/32/33/abstract.
html.

Zhao:2005:ESL

[ZS05a] Weibin Zhao and Henning Schulzrinne. Enhancing Service Location Protocol for efficiency, scalability and advanced discovery. The Journal of Systems and Software, 75(1–2):193–204, February 15, 2005. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zulkernine:2005:TAM

[ZS05b] Mohammad Zulkernine and Rudolph Seviora. Towards automatic monitoring of component-based software systems. The Journal of Systems and Software, 74(1):15– 24, January 1, 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhang:2004:UAF

[ZSM04] N. Zhang, Q. Shi, and M. Merabti. A unified approach to a fair document exchange system. The Journal of Systems and Software, 72 (1):83–96, June 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhang:2005:RPE

[ZSM05] N. Zhang, Q. Shi, and M. Merabti. Revocation of privacy-enhanced publickey certificates. The Journal of Systems and Software, 75(1–2):205–214, February 15, 2005. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhang:2001:EAE

[ZSP01] Xuemei Zhang, Mi-Young Shin, and Hoang Pham. Exploratory analysis of environmental factors for enhancing the software reliability assessment. The Journal of Systems and Software, 57 (1):73–78, April 27, 2001. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.elsevier.nl/gej%
2Dng/10/29/11/64/25/31/
abstract.html;   http:
//www.elsevier.nl/gej%
2Dng/10/29/11/64/25/31/
article.pdf.

Zeng:2008:CDR

[ZWX^{+}08] Qingtian Zeng, Huaiqing Wang, Dongming Xu, Hua Duan, and Yanbo Han. Conflict detection and resolution for workflows constrained by resources and non-determined durations. The Journal of Systems and Software, 81(9):1491– 1504, September 2008. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Zhuge:2001:TWP

[ZyCkP01] Hai Zhuge, To yat Cheung, and Hung keng Pung. A timed workflow process model. The Journal of Systems and Software, 55(3):231–243, January 15, 2001. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/10/
29/11/54/27/25/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/54/
27/25/article.pdf.